A Complete Bibliography of Publications in Mathematics and Computers in Simulation: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 1.00

Title word cross-reference

(1 + 1) [651]. (2n + 1) [1801]. 2 + 1 [1673]. ^{+} [156]. ^{++} [156]. ^{2} [245, 704]. _{2} [784]. _{∞} [422, 420]. a [1226]. AX − XB = C [478]. b [1226]. C^{2} [1395]. C^{r} [1396]. χ [51]. D [1497]. ϵ [575]. F [784]. G^{1} [1175]. G^{2} [1175]. H [1135, 422, 420, 539]. H_{∞} [1497, 1024]. I [1612]. m [733, 537, 538]. mK(m, n) [684]. N [383, 1890, 782, 1530]. p [497, 1659]. q [1905]. r [228]. s [537, 538]. sech^{2} [1807]. t [537, 538]. θ [1690]. V [1612]. φ^{4} [293]. W [73].

-adaptive [228]. -algorithms [51]. -component [782]. -conditioned [539]. -control [420]. -dimensional [1673]. -discretized [1905].

-fold-decimation-based [733]. -function [73]. -invariant [1226]. -Laplacian [497]. -nets [537, 538]. -Queen [383]. -scheme [1690]. -Shell [575]. -stability [1659, 1497]. -surfaces [1396]. -theorem [1135]. -type [1807].

1 [391, 448, 457, 1047, 918, 1021, 1361, 1581, 125]. 10 [580, 703, 1213]. 11 [1085, 1094]. 12 [824]. 14th [627, 659]. 15 [76, 143, 68, 281, 151, 402, 438, 850, 1233, 1392]. 15th [930, 952, 929, 951]. 16 [928]. 17 [572, 766]. 19 [315]. 1999 [36]. 1D [1891, 1910, 803].

2 [1300]. 2-dimensional [35]. 20 [910]. 2000 [49, 76, 143, 31, 38, 21, 58, 68, 44, 127, 113, 82]. 2001 [269, 281, 308, 315, 151, 244, 224, 459, 525, 627, 659]. 2002 [334, 325, 391, 402, 448, 438]. 2003 [572, 457, 580, 626, 1049, 929, 951]. 2004 [737, 824, 690, 658, 776, 703, 766, 850]. 2005 [910, 1047, 899, 889, 918, 928, 1038, 1021, 1030]. 2006 [1085, 1133, 1075, 1094]. 2007 [1283, 1361, 1223, 1233, 1213, 1300, 1291]. 2008 [1428, 1503, 1581, 1392, 1456, 1537, 1526]. 2009 [1670, 1799, 1813, 1714, 1900, 1885]. 22 [737]. 24 [626]. 25 [690]. 250 [153]. 26 [334]. 27 [658, 1291]. 29 [224]. 2D [752, 1237, 235, 318, 1891, 621, 79]. 2D/3D [752].

3 [308, 899, 889, 1133]. 30 [58, 127, 113, 82, 1283]. 3D [1192, 1816, 771, 752, 1141, 467, 473, 1679, 924, 815, 318, 488, 1334, 1868, 1160, 117, 304]. 3D-cycles [771, 1868].

4 [244, 776, 1428]. 4H [172]. 4H-SiC [172]. 4th [778].

5 [1743, 1774].

6 [1223]. 6-pulse [584]. 625 [1099].

7 [325, 1075]. 7-triangle [1728]. 71 [1209, 1219]. 78 [1931].

8 [1038, 1030]. 80 [1932, 1929, 1933]. 8th [1096].

A-EBDF [769]. a-priori [793]. A-splines [1175]. Abelian [194]. ablation [743]. Ablowitz [1907]. absolute [1364, 345]. absorption [552, 1186]. ac [606]. academic [1473]. accelerate [1422]. Accelerated [1730, 1789, 838]. accelerated-time [838]. accelerating [1877]. acceleration [1913]. acceptable [716]. accidental [63]. accommodation [1881]. account [110]. accumulation [674, 1169]. Accuracy [121, 369, 1319]. Accurate [584, 1609, 601]. accurately [1624]. ACD [1735]. acetylcholine [361]. achieving [716]. acidophilus [718]. Acoustic [59, 798, 759, 1553, 504, 711, 487]. acoustics [741]. acquire [9]. acslXtreme [1750]. act [92]. action [871, 1849]. activated [246]. active [606, 1546, 740, 1105, 1107]. active/passive [606]. activities [942]. activity [960]. ACW [716]. ACW-gradient [716]. Adams [1630]. Adaptation [103, 565, 1818]. Adapting [661, 107]. Adaptive [336, 1672, 1287, 615, 1509, 6, 1538, 566, 1518, 759, 1628, 228, 1506, 1294, 1823, 769, 9, 419, 900, 1110, 546, 1725, 1510, 417, 901, 343, 15, 1663, 115]. adaptivity [732, 1388]. additives [876]. adjoint [1411]. adjusting [1478]. admixtures [864]. ADRs [1464]. adsorbent [1825]. adsorption [1825]. Advanced [594]. advances [1673]. advancing [1307]. advantages [1066]. advection [995, 1247, 1563]. aerodynamic [1509, 1724]. aerospacecraft [102]. aesthetic [1434]. affine [1528]. against [1515]. Age [328, 1080, 1209, 1219, 1515, 674]. age-structured [1080, 1209, 1219]. agent [1057, 630, 362, 411, 1485, 1066]. agent-based [1057, 630, 1485]. agents [1365]. agglomeration [1758]. aggregated [1148]. aggregates [1141]. aggregating [1548]. aggregation [1461, 1662, 1373]. Agincourt [652]. agricultural [937, 1640, 712, 355]. agro [415]. agro-food [415]. aid [27]. aided [25]. aiming [256]. air [757, 872, 114, 63, 265, 107]. Aircraft [1683, 1546]. airlift [70]. airport [838]. algal [1487]. Algebra [283, 22, 825, 287, 87, 430, 155, 34, 833, 23, 326, 327, 1715, 834, 1717]. algebra-based [326]. Algebraic [1353, 1632, 89, 785, 232, 1718, 1564]. algebras [54]. algorithm [1394, 1752, 1518, 759, 533, 1628, 1819, 280, 1791, 1879, 1624, 416, 475, 1218, 786, 1439, 158, 331, 1826, 1522, 900, 318, 382, 1681, 1510, 597, 171, 77, 901, 758, 1341, 1834, 528, 1017, 14, 1869, 1487, 131, 1039, 1617, 109, 384, 1412, 716]. Algorithmic [1620]. algorithms [532, 425, 1088, 11, 1623, 155, 617, 1801, 1877, 1323, 221, 279, 618, 51, 414, 611, 422, 385, 575, 738, 550, 1691, 452, 767, 1378, 892, 1303]. alignment [473]. allocation [1879, 1128, 1203]. Almost [1207, 1499, 466]. almost-unidirectional [466]. along [1154]. alpha [1208]. alternative [1735, 359, 576, 1661]. Alternatives [98]. aluminium [156]. Alzheimer [1892]. American [535, 139]. among [1552, 1521]. amplifier [1616]. amplifiers [262]. amplitude [975, 1903]. anaesthesia [846]. analog [84]. Analogies [1734]. analyses [124]. analysi [2]. Analysis [504, 1016, 329, 1803, 957, 618, 604, 900, 830, 443, 1099, 1410, 1795, 1738, 1549, 696, 668, 1101, 1097, 1708, 1363, 312, 768, 814, 1569, 1778, 844, 1090, 934, 1401, 290, 1225, 978, 1146, 935, 936, 40, 695, 1031, 1532, 979, 787, 1889, 1204, 1621, 1575, 1061, 1377, 373, 1464, 1285, 1880, 386, 1116, 168, 575, 1750, 273, 487, 361, 1475, 1767, 963, 1599, 1520, 176, 729, 1868, 190, 317, 360, 1378, 179, 645, 15, 1718, 374, 1890, 1896, 1529, 1860]. analytic [1932, 1923, 1933, 530, 567]. Analytical [73, 441, 1448, 603, 295, 71, 727, 1597, 1697]. analyzing [933, 1541, 784]. aneurysms [1142, 1146]. animal [256]. anisotropic [482, 797, 192, 1510]. annealing [549]. announcements [936]. annual [373]. annulue [874]. anomalous [1417]. anomaly [873]. anorexia [326]. antiretroviral [1080, 1209, 1219]. Antisoliton [213]. antisolitons [987]. Apollonius [451]. apparent [1137]. apparent-slip [1137]. appearance [911]. Application [154, 872, 1818, 87, 1541, 1532, 405, 549, 836, 363, 1138, 480, 1318, 1547, 1032, 1057, 1495, 1542, 1401, 51, 1431, 605, 412, 1467, 1216, 10, 251, 1224, 553, 383, 1543, 944, 56, 424, 1348, 741, 1405, 1817, 1140, 1846, 1640, 1059, 1448, 1831, 1004]. Applications [827, 1198, 750, 1894, 22, 825, 286, 953, 1303, 1696, 135, 118, 1277, 1507, 760, 35, 738, 1337, 86, 1629, 399, 1105, 1715, 834, 1430, 1365]. Applied [1504, 477, 1791, 1879, 1398, 32, 1625, 600, 1560, 595, 560, 1720]. Applying [1695]. approach [97, 1753, 1518, 1497, 1613, 1665, 1432, 62, 1791, 1778, 364, 1304, 1399, 1114, 1610, 1644, 1678, 1241, 1079, 727, 978, 1266, 1764, 495, 301, 685, 159, 1053, 414, 1059, 1597, 1061, 1208, 1694, 386, 1334, 1417, 1485, 273, 1182, 86, 173, 1515, 295, 1024, 176, 278, 794, 276, 219, 327, 1102, 468, 965, 767, 1771, 1050, 1583, 1490, 1002, 1249, 893, 1166, 115, 673, 1742]. approaches [1892, 275, 958, 663, 556, 1482, 1782]. approximants [441]. Approximate [1296, 285, 1201, 1032, 709, 1623, 1700, 294, 1602]. Approximating [1400, 1412]. Approximation [1625, 1720, 78, 484, 1397, 1423, 1531, 43, 432, 1827, 126, 986, 164, 553, 632, 1629, 1832, 1600, 1650]. approximations [73, 1321, 1226, 304]. April [572, 737, 910, 1085, 1283, 1428, 49]. aquifer [954]. aquifers [356]. Arbitrary [1388, 1602]. Arbitrary-level [1388]. arch [1820]. Architectural [1054]. architecture [1057, 1619, 1618, 133, 411, 278]. architectures [426, 71]. area [1368]. ARFIMA [368]. arising [1068, 138, 221]. arithmetic [218]. arithmetical [835]. ARMA [698]. ARMAX [1766]. arms [934]. Arnold [1236]. Arnoldi [1877]. array [1697, 871, 1154]. arresting [1005]. ARTEMIS2122 [584]. arteriole [1179]. Artificial [1484, 869, 309, 1487]. ASEAN [1743, 1774, 941]. ASEAN-5 [1743, 1774]. Asia [375, 1662, 940, 1771]. Asian [673]. Aspects [1905, 1622, 474, 1577, 530, 1620]. assembly [1879, 1782]. Assessing [958, 1703, 957, 965]. Assessment [1017, 1488, 711, 631, 20]. Asset [1761, 1787]. assignment [280, 12]. assignments [359]. assimilates [715]. assimilation [1565]. associated [1135, 1755, 593, 178]. association [62]. Associative [17, 1887]. asymmetric [1262, 642, 1764]. asymmetry [936, 664]. Asymptotic [1778, 1272, 1919, 1707, 1887]. asymptotically [1768]. asynchronous [1398, 1120, 606]. atmosphere [1200, 864, 865, 857, 64]. atmospheric [1194, 956, 868]. atoms [1910]. attractive [505]. attractor [895]. attribute [891, 1203]. augmentation [118]. August [76, 281, 824, 1503]. Australia [627, 659, 929, 951, 1753, 375, 367, 672, 1748, 661, 959, 373, 676, 1473]. Australian [1735, 933, 1757, 1471]. Author [216, 270, 307, 351, 436, 501, 563, 624, 701, 764, 822, 887, 949, 1011, 1064, 1123, 1173, 1189]. Authors [1045]. autocorrelation [1768]. Automata [297, 1619, 1618, 1060, 784, 15, 1149]. automate [70]. automated [1058]. Automatic [826, 545, 1388, 340, 722]. Automatically [1040]. automation [1614]. automatized [1593]. automotive [412]. autonomous [1103, 1032]. autoregressions [369]. autoregressive [671, 1757, 1077, 1875, 1872]. autotuning [13]. avalanche [890]. average [1117]. Averaged [1264, 35, 632]. averages [961]. Averaging [1273]. avoidance [381]. avoiding [1706]. axi [198]. axi-symmetric [198]. axis [1309, 1308, 1725]. axisymmetric [746]. axonemal [1863].

B [1742, 1460]. bacillus [920]. background [509]. Bäcklund [1915, 1270]. backprojection [795]. backscattering [533, 540]. Backward [979, 1531, 396]. Baikonur [870]. balance [1574]. balanced [103]. balancing [340]. ballistic [540]. balls [751]. BAM [1658]. band [1596, 1607]. bandwidth [632]. bank [942]. bar [611]. Barbanis [33]. barriers [531]. based [837, 1395, 733, 565, 759, 1613, 1040, 1665, 954, 534, 779, 280, 316, 1791, 475, 157, 1055, 1610, 1057, 1286, 770, 1325, 630, 442, 919, 732, 1439, 9, 1619, 1618, 787, 1053, 1546, 1431, 913, 761, 1741, 583, 593, 1680, 7, 881, 1510, 133, 1181, 1485, 895, 597, 639, 417, 1768, 1310, 1042, 326, 1728, 219, 246, 556, 528, 557, 846, 1561, 1780, 1690, 852, 449, 1869, 1058, 252, 1663, 1617, 1496, 1545, 1550, 1323, 586]. bases [1408, 26, 836]. basin [355]. basis [1573, 859, 1412]. batch [693, 882, 303]. battle [652, 652]. Bayes [1206]. Bayesian [1738, 668, 1478, 1892, 1665, 369, 1707, 237, 1515, 965, 633]. BBGKY [1167]. beam [153, 395, 47]. beams [832]. Bearing [1545, 1322, 279]. bearings [280]. bearings-only [280]. bed [1922, 913]. beds [1825]. beef [1750]. behavior [1816, 1674, 1147, 9, 103, 1054, 1066, 45, 1919]. behavior-based [9]. behaviors [9]. behaviour [1145, 1077, 516]. behavioural [1795]. belt [967]. Beltrami [28]. benchmark [1693]. bends [972]. Bernoulli [47]. Bernstein [1397, 1412]. Bernstein-type [1397]. best [650, 1395, 1727]. between [1262, 395, 672, 1889, 318, 1559, 1464, 1774, 707, 104, 172, 296, 100, 557, 874, 698, 1460]. beyond [1831]. bi [503, 1887]. bi-directional [503, 1887]. bias [369, 371, 541]. bibliometric [966]. Bickley [513]. bidimensional [78]. Bienayme [1340]. Biennial [930, 952, 627, 659, 929, 951]. Bifurcation [1795, 1204, 1632, 974, 329, 1841, 124, 1529]. Bifurcations [1258, 911, 277, 613, 1147, 1704, 1259]. bilateral [1394]. bilinear [40]. binary [1136, 988, 1652]. Bingham [470]. binomial [1654]. bio [431]. bio- [431]. biochemical [1598, 1599]. biochemistry [451]. biodegradable [719]. biofilms [1676]. bioheat [924]. biological [920]. biomathematical [653]. biomechanical [471, 468]. biomechanics [495]. biomimetics [732]. Biophysical [628]. bioprocesses [253, 18]. Biot [240]. biotechnology [960, 1773]. birth [920, 45]. birthday [165]. bistable [347]. Bivariate [1403, 1451, 1469, 933, 1216]. BKP [1915]. black [132, 252]. black-box [132, 252]. blackboard [70]. blades [175]. blending [1835, 1409]. Blind [237, 382]. Block [1506, 1551, 1700, 107, 1496]. Blockset [590]. blood [1147]. blow [1919]. board [429, 879, 437, 447, 456, 523, 564, 571, 579, 610, 625, 726, 648, 657, 680, 689, 702, 775, 790, 807, 823, 841, 849, 1037, 1046, 888, 898, 909, 917, 927, 950, 970, 991, 1012, 1020, 1029, 1065, 1074, 1084, 1093, 1124, 1132, 1174, 1202, 1212, 1222, 1232, 1255, 1276, 1282, 1290, 1299, 1313, 1352, 1360, 1376, 1391, 1415, 1427, 1443, 1455, 1493, 1502, 1514, 1536, 1636, 1525, 1580, 1669, 1688, 1713, 1746, 1777, 1798, 1812, 1838, 1866, 1884, 1899, 1928]. bodies [467, 344, 1170]. body [1508]. Boltzmann [1136, 1138, 1141, 1142, 1143, 1144, 1644, 1145, 1147, 1148, 1150, 1151, 1152, 1153, 1154, 1284, 1155, 1158, 1159, 1163, 1165, 1166, 1169, 1647, 1170]. Bond [95, 665, 101, 1762, 84, 1385, 942]. Bondgraphs [322]. bone [1180]. Book [322]. Boole [23]. Boolean [811]. boost [596]. bootstrap [1496]. bootstrapped [961]. Bootstrapping [1815, 1474]. borne [1032]. Borutzky [322]. Bose [1854, 1267, 1241, 505, 1852, 982, 988, 1271, 1000, 1004]. both [318, 1382]. bottom [503, 1925]. boundaries [1039]. boundary [1588, 1549, 1553, 804, 1572, 240, 138, 1601, 1705, 799, 442, 749, 59, 1404, 1071, 1179, 1180, 1829, 1649, 1527, 1856, 465, 1701]. boundary-fitted [1553]. boundary-layer [749]. bounded [319, 1531, 1679, 163, 987, 1639].

Boundedness [1651]. bounds [1830, 219]. Boussinesq [1250, 1237, 193, 1925, 1125]. Boussinesq-double [1125]. box [723, 1755, 132, 252]. BP [16]. bracket [508]. Bragg [1271]. brain [1692]. branching [829, 1339]. break [635]. breakage [1574]. breakdown [1215, 986]. breaking [974, 190]. breaks [1481, 1743, 1418]. breast [752]. Breathers [509, 197]. brewing [252]. bright [263]. Broad [1462, 1719]. Broad-market [1462]. broadband [582]. Brownian [1383, 1742]. brushless [588]. Bubble [1908]. buck [1890, 591]. building [1040, 1434]. buildings [103]. bundles [1246]. buoyancy [1916]. Burdekin [959]. Burgers [1638, 514, 1924, 1023]. Burr [1703]. bus [312]. business [1076, 1354, 1764]. bypass [1338].

c [153]. cable [1793]. CAE [94]. CAE-tool [94]. cage [601]. calculated [90]. Calculation [1903, 715, 1272, 1862]. calculations [41]. Calendar [215, 223, 234, 243, 259, 269, 299, 314, 428, 435, 446, 455, 570, 578, 609, 623, 848, 647, 656, 679, 688, 700, 763, 774, 806, 821, 840, 878, 725, 755, 990, 1019, 886, 897, 908, 916, 926, 948, 969, 1010, 1028, 1036, 1044, 1063, 1073, 1083, 1092, 1122, 1131, 1172, 1188, 1312, 1351, 1211, 1221, 1231, 1254, 1275, 1281, 1289, 1298, 1375, 1390, 1414, 1426, 1442, 1454, 1492, 1501, 1513, 1524, 1535, 1579, 1635, 1668, 1687, 1712, 1745, 1776, 1797, 1811, 1837, 1865, 1883]. Calendar [1898, 1927, 350, 401, 390, 500, 522, 562, 183, 1359]. Calender [324, 735, 789, 306, 380, 333]. calibration [1891, 867]. callus [256]. CALS [87]. Camassa [1853, 200]. camphoraceous [1431]. Canada [1076]. Canberra [627, 659]. cancellation [711]. cancelling [759, 109]. cancer [1692]. canyon [377]. canyons [377]. capacitance [158]. capacity [46, 1541]. capelin [1817]. capillaries [818]. capillary [1245, 731, 1239]. capillary-gravity [1245]. capital [640, 674, 938]. capture [669]. car [423]. carbon [956, 1605]. CARI [752]. Carlo [542, 532, 1518, 154, 533, 1469, 1101, 534, 1381, 536, 1819, 155, 617, 156, 396, 526, 786, 159, 540, 1596, 160, 880, 161, 541, 543, 1517, 166, 546, 169, 547, 1417, 171, 550, 551, 172, 552, 904, 176, 555, 1698, 729, 178, 557, 742, 179, 121, 180, 529, 559, 892, 921]. carriers [336]. carrying [1541]. Cartesian [493]. cartilaginous [492]. cartographic [1831]. CAS [27]. cascade [1551, 1890]. case [677, 671, 1758, 843, 1033, 1790, 367, 247, 717, 1908, 293, 831, 13, 119, 1719, 1488, 645, 1167]. Caspian [863]. catalyzed [1833]. cataract [1767]. catastrophic [1590]. catching [366]. catchment [628, 629, 959, 1368]. catchment-scale [628]. catchments [661, 1483, 355]. catenary [1570]. Cauchy [1874]. causal [672]. causality [1569, 698]. cavitation [747]. Cavity [983, 799]. CE [1328, 1511]. CE/SE [1511]. celestial [344]. cell [11, 1147, 493, 1575, 1421, 1680, 1102]. cell-based [1680]. cells [1166, 256]. cellular [1499, 1516, 1031, 1619, 1618, 12, 1794, 1651]. cement [419]. censored [1703, 1206, 1789]. centered [621]. centers [1022]. central [997]. centres [255]. cerate [1751]. Cerenkov [1267]. certain [1397, 1068, 1778, 165]. CESA [1049]. CEV [1519]. CFD [1507]. Chaem [1483]. Chain [808, 1819, 1852, 1592]. chains [199, 1646]. Challenges [931, 556]. chance [272]. change [662, 1892, 628, 1450, 147, 1702, 1419, 1382]. changed [1468]. changes [1079, 635, 868]. changing [1473]. channel [477, 1218, 1611, 12]. channels [619, 72]. Chaos [1089, 188, 1805, 337, 1867, 1538]. Chaotic [383, 336, 348, 1089, 894, 1554, 1420, 1653, 895, 45, 338, 697, 148, 384, 1550]. character [1015]. characteristic [1873, 1480]. characteristics [342, 715, 114, 487, 1612, 1278]. characterization [1816, 1405]. charge [161]. charged [1140]. Chebyshev [1917, 1824, 919, 1402, 1856]. Chebyshev-type [1824, 1402]. check [275]. cheese [721]. chemical [1193, 75, 230, 303, 431, 1160]. chemistry [1144]. chemostat [1126]. chemotherapy [124]. child [371]. China [1541, 1803, 644, 939, 1489, 1773, 1477]. Chinese [937, 1477, 1742]. Chiral [517]. chlorophyll [254]. Choice [1370, 1051]. choices [677]. Choquet [1373]. Christoffel [1116]. chromatographic [111]. Chua [1287, 1420, 1538, 812]. cigarette [1433]. circle [1178]. circuit [1040, 1610, 1616, 1695, 205, 1116, 1115, 1099]. circuit-based [1040]. circuits [1613, 90, 549]. circulation [64]. cities [921]. City [1541, 377]. CLAM [1488]. Clarifying [1559]. class [566, 759, 1089, 1052, 1031, 1265, 1041, 69, 1407, 1528, 1024, 567, 519]. classes [651]. classical [18, 161, 1831]. Classification [1423, 199]. classifier [20]. clearance [1349]. climate [377, 147, 714]. closed [911, 1785, 906, 77]. closed-form [1785]. closed-loop [906, 77]. clotting [1138, 1142]. cluster [583]. clustering [406, 148]. clusters [204]. clutch [1879]. clutter [154]. CNG [1722]. Cnoidal [1041, 974]. co [1730]. co-ordinates [1730]. coagulation [1139, 530]. coalescence [1574]. coarsening [1680]. coastal [1485, 356, 1488]. coated [1568]. coaxial [874]. code [832, 545, 115]. coded [1156]. coefficient [369, 1424, 1886, 81, 442, 1649, 803, 1862]. coefficients [1109, 1678, 432, 163, 1296, 1186, 1353, 140]. coercive [914, 393, 398]. cognitive [1892]. Cohen [1645, 1659, 1229]. coherent [994]. cointegrating [1562, 1357]. Cointegration [374, 1753, 1217, 1286, 1521, 373, 1464, 1771, 1496]. collaborative [1056]. collagen [256]. collapse [265, 1005]. collateral [1784]. collective [119, 1039]. Colligation [39]. collision [221, 1152, 1845, 381]. collocation [138, 1822, 149, 1183, 1701, 1087]. colloidal [1140, 1156]. colloids [558]. collusion [372]. color [1440, 1721, 1438]. colorable [117]. colored [730]. column [422]. combination [413]. combine [710]. Combined [1199, 544, 97, 1655, 358]. Combining [1718, 1101]. combustion [1193, 995]. Commemoration [498]. Comment [1932, 1929, 1933]. Comments [1930]. commerce [1436]. commodity [946]. common [958, 600]. communication [312, 1641]. Communications [296]. commutative [1367]. Commutativity [1915, 1270, 844]. Compact [519, 1572, 1612, 621, 239, 117]. compacton [1342]. Compactons [510, 567, 1007]. Comparative [935, 81, 599, 1105, 79]. Comparing [1482]. Comparison [1735, 75, 228, 354, 197, 35, 1906, 1126, 1469, 1142, 395, 618, 1245, 433, 1611, 707, 513, 172, 557, 1856]. compartment [1080, 1209, 1219]. compatibility [327]. compensation [584, 607, 600]. Competition [1126, 110]. compilation [135]. complementarity [221]. Complex [319, 302, 1371, 1852, 1554, 1258, 1259, 207, 41, 962, 1600, 1003, 1349, 1008]. complexes [855]. complexity [137, 654]. component [875, 782, 982, 133, 2]. component-based [133]. components [101, 843, 544]. composite [1323, 1279, 855, 1278]. Compound [1556, 1825]. comprehensive [1507]. compressible [462]. compression [1621]. compressor [462]. Comput [1932, 1931, 1929, 1209, 1219, 1933]. Computation [973, 993, 286, 1626, 285, 29, 467, 1881, 526, 1509, 1070, 828, 728, 1071, 203, 1910, 1682, 716]. Computational [1691, 1303, 1177, 135, 134, 1401, 130, 1616, 1561, 1622]. computationally [111]. Computations [1069, 473, 1679, 1153, 399]. Computer [287, 25, 833, 119, 1416, 343, 22, 241, 825, 834, 1568, 87, 430, 587, 808, 134, 629, 1654, 47, 23, 845, 1725, 55, 1577, 326, 278, 327, 846, 1446, 1715, 283, 1717]. Computer-aided [25]. computer-algebra [87]. computers [757]. Computing [478, 1321, 1214, 1720, 425, 1040, 310, 1691, 137, 1504, 767, 574, 1696]. concentration [146, 866, 716]. concentrations [956]. concept [1850, 199, 91]. concepts [430, 1054, 1430]. Concerning [122]. concrete [1266]. condensates [1854, 1267, 1241, 505, 982, 988, 1271, 1000, 1004]. condition [141, 1032, 1886, 1052, 59, 1071, 1445, 1545]. Conditional [1737, 1465, 1757, 1762, 1480, 1736, 1769, 637, 946, 1872, 1460]. conditioned [442, 539]. conditioner [1181]. Conditioning [696, 57]. conditions [1588, 90, 714, 465]. conduction [1295, 240, 1531, 1705]. conductivity [797, 1886, 1532]. conductors [797]. cones [1633]. Conference [930, 952, 973, 993, 1096, 36, 627, 659, 929, 951]. Confidence [1868]. confinement [744]. conflict [934]. conflicts [100]. Conformal [1842, 792]. conforming [1335]. congestion [1681]. Conic [451]. Conjugate [1326, 481]. conjunctive [1370]. connected [804, 606, 1112, 1930, 603, 1666]. connection [187]. connectivity [255]. cons [89]. conservation [483, 1193, 474, 201, 1849, 1345]. conservative [694, 1261, 750, 1002]. Considering [895]. consistent [1243, 1164]. Constant [1519, 1886, 520]. constituent [1760]. constitutive [1870]. Constrained [675, 1591, 576, 321, 1894, 1639]. constraint [388, 272, 1621]. constraints [1497, 1236, 1403]. Constructing [1201, 1522, 1769, 1164, 978]. Construction [810, 1835, 293, 1409, 813, 1022, 1384, 162, 27, 1834, 297, 866]. consumer [1437, 1734]. consumers [247]. consumption [711, 672]. contact [476, 914, 393, 883, 1150, 398, 317, 1278]. contagion [941]. containing [832]. contaminant [169, 547, 356]. contaminated [168]. contamination [1406]. content [932]. continental [662]. continuation [1631, 489]. Continuous [1266, 1783, 1779, 363, 1817, 638, 612, 1575, 1480, 69, 1339, 813]. continuous-time [638, 612, 1480, 69]. continuously [1207]. continuum [495]. Contopoulos [33]. contractive [817]. Control [592, 593, 591, 1081, 425, 1192, 1497, 1078, 1287, 1589, 587, 612, 706, 450, 272, 88, 1591, 1601, 61, 145, 32, 1364, 17, 495, 1420, 9, 321, 1112, 5, 419, 1366, 1110, 1704, 1781, 740, 420, 599, 10, 303, 597, 238, 417, 1060, 1599, 1520, 1642, 1684, 1639, 1175, 343, 1530, 1050, 1583, 232, 846, 431, 582, 697, 812, 1107, 15, 403, 381, 1732, 1663, 596, 1896, 1550, 339, 594, 716]. control-state [1591]. controlled [341, 1725, 742, 1051, 1446]. controller [1104, 1369, 422, 418, 1617]. controllers [566, 1108]. convection [1540, 1572, 240, 1873, 432, 728, 164, 1726, 1158, 1343, 1690, 621, 1249, 117]. convection-reaction [1726]. convective [1886, 1841, 1532]. Convergence [1774, 649, 1917, 1398, 486, 1408, 1445, 1730, 3]. convergent [320, 1416]. converter [591, 1097, 592, 587, 1369, 1102, 598, 1215]. Converters [1096, 581, 1095, 1890]. convolution [7]. cooperation [1372, 1593, 100]. cooperative [1066]. coopetition [1333]. COOPT [232]. coordinates [1725, 741]. copula [1740]. copula-threshold-GARCH [1740]. core [1262]. corner [974]. corners [477]. corporate [665, 942, 1734]. Corporation [944]. corrected [866]. correction [607, 1559, 1309, 1308]. correlated [50, 1517]. correlation [759, 1768, 1737]. correlations [1762, 946, 1460]. corresponding [860, 300]. Corrigendum [1931]. corrosion [1144, 1701]. corrugations [1802]. cosine [923]. Cost [338, 61, 1471]. cost-optimization [61]. costs [1795]. cough [257]. Could [1786]. count [1384, 943]. counterparts [69]. countries [934, 640, 1357, 1521]. country [945]. coupled [75, 348, 504, 1200, 506, 264, 1261, 1498, 1204, 1826, 512, 55, 487, 251, 1920, 816, 1566, 148, 1529]. coupling [592, 103, 1264, 1002]. Courant [1511]. covariance [1655, 1585]. Cover [736, 429, 437, 447, 456, 523, 564, 571, 579, 610, 625, 648, 657, 680, 689, 702, 775, 790, 807, 823, 841, 849, 879, 888, 898, 909, 917, 927, 950, 970, 991, 1012, 1020, 1029, 662, 1483]. cover-Editorial [429]. coverage [1707]. Cox [1708]. crack [798, 279]. crafts [1593]. creation [1907, 294, 1434]. credit [665, 1734]. creep [491]. crevice [1144]. crime [1472]. Criteria [1177, 1432, 1645, 370, 1447, 1366, 166, 1527, 1373, 2, 1658]. criterion [890, 1079]. critical [191, 1919]. Croatia [36]. Cronbach [1208]. cross [1845]. cross-modulation [1845]. crosses [988]. crowd [1227]. crowds [99]. crust [1540]. crystals [543]. cubature [557]. Cubic [1175, 440, 1262, 975, 138, 277, 1258, 1259, 1848, 1003, 621]. cubic-quintic [975]. cultural [1791, 416]. cultures [1575]. curd [721]. currency [940, 634, 1555, 1477]. current [1452, 1931, 1843, 1616, 361, 1916]. currents [1111, 857]. curve [1619, 1216, 1772]. curves [911, 1835, 826, 1725]. curvilinear [309, 1725, 1730]. customer [1436]. customized [1040]. cut [493]. CW [1258]. cycle [1076, 1354, 613, 1764, 986, 922, 1804]. cycles [771, 613, 1868]. cyclic [1808, 783]. Cyclides [1178]. cylinders [874].

daily [712, 1741]. dam [1820]. damage [1180]. damped [1842]. damping [262, 1918]. Danish [844]. Dark [1004]. data [50, 733, 662, 669, 29, 354, 116, 62, 280, 1201, 1424, 1384, 1893, 883, 937, 1621, 1522, 1827, 712, 1703, 1789, 1565, 1474, 639, 1224, 670, 943, 1832, 758, 729, 1639, 255, 867, 866, 1286]. Data-based [1286]. data-driven [1522]. dates [249]. day [936]. dc [591, 1097, 606, 1102, 1215, 588, 1369]. dc/dc [591, 1102, 1215]. dead [1420, 339]. dead-zone [1420]. death [920, 346]. death-diffusion [920]. debonding [1323]. decanano [540]. decay [1721]. December [143, 308, 315, 1047, 1361, 1581, 627, 659]. Decentralized [906]. decimation [733]. decision [1432, 63, 237, 965, 1888, 1488, 891, 1203]. decision-making [1888, 891, 1203]. decisions [861, 1371]. decline [1892]. decomposition [827, 810, 480, 476, 914, 1677, 1506, 34, 1327, 433, 573, 41, 392]. deconvolution [460]. decoupling [138]. DEDS [650]. deep [1244, 354, 1242, 1245, 210]. deep-water [210]. default [1784]. defect [412]. defective [141]. Defining [24, 835]. definition [341, 122]. deformation [860, 469]. deformations [1674]. deforming [1320]. degenerate [24, 1225]. degenerated [1650]. degradation [719]. degree [1197, 277, 568]. Delaunay [136]. Delay [1285, 814, 1881, 126, 1880, 691, 1792, 1564, 1658, 1779, 1896, 1558]. Delay-dependent [1285, 1880, 1658]. delay-integro-differential [1792]. delay-integro-differential-algebraic [1564]. delayed [1645, 1031, 1889, 1292, 809]. delays [1499, 312, 1516, 53, 1207, 1285, 420, 1887, 1228, 1641, 1659, 921, 1658, 1794, 1229]. delivery [1748, 1560, 131]. demand [1753, 643, 370, 1357, 1521, 373, 1766, 1490]. Dempster [1888]. denominator [1247]. denominators [1236]. dense [1060]. densities [249]. density [1735, 186, 1756, 620, 426, 551, 961, 45, 449]. density-dependent [45]. departing [838]. dependent [1090, 441, 472, 1532, 1285, 1880, 1726, 1699, 1649, 45, 803, 431, 1690, 1585, 1039, 1658]. deposit [547]. deposition [1750]. depressant [876]. depth [198, 846]. deregularization [92]. Deregulation [942, 372]. Derivation [1245, 1519, 1071, 999]. Derivative [1780, 602]. derivatives [535, 1400, 632, 1691]. derived [1752, 826, 256]. deriving [782, 832, 1763]. Descent [1627]. Describing [101, 169]. Description [249, 913, 719]. Descriptor [89, 1778]. Design [1104, 1616, 710, 1024, 423, 1683, 1347, 1571, 1497, 1613, 341, 1068, 1722, 1433, 732, 787, 422, 1127, 1614, 1056, 238, 1099, 1706, 327, 70, 836, 132, 967, 1054]. designing [1108]. designs [1120]. destabilisation [964]. detail [573]. detailed [1519]. detect [279, 636]. Detecting [1481]. Detection [1869, 316, 88, 1286, 1881, 1438, 1053, 412, 439, 409, 600, 408, 1405]. Detection-recognition [1869]. detectors [552]. determinant [1859]. determinants [676]. Determination [450, 1247, 1649, 801, 826, 1034, 1435]. determining [860, 1076, 123, 330]. deterministic [611, 173, 602]. developed [634]. developing [934, 640, 921]. Development [1724, 582, 1432, 249, 675, 859, 199, 94, 1330, 1418, 663, 852]. Developments [493, 939]. deviate [666]. device [541, 550, 1284]. devices [1068, 1609, 161, 546, 556]. DGFEM [1343]. diabetes [328]. diagnosability [1058]. Diagnosing [375]. diagnosis [316, 1053, 605, 144, 881, 411, 19, 326, 407, 403]. Diagnostics [637]. diagonalization [528]. diagram [141]. Diaphony [1033]. diesel [593]. difference [484, 186, 1143, 1068, 1444, 1243, 830, 53, 206, 514, 995, 813, 1856, 621, 1566, 117, 304]. differencing [368]. different [1676, 1022, 619, 924, 544, 318, 818, 1716, 172, 45, 674, 338, 1482, 255]. differential [285, 1632, 288, 1675, 653, 116, 1791, 1423, 1778, 1801, 1398, 441, 1243, 1326, 394, 829, 433, 199, 126, 903, 1296, 1781, 119, 1182, 89, 236, 1831, 1560, 232, 179, 574, 681, 1792, 1564, 649]. differential-algebraic [89]. differential-difference [1243]. differentiation [722]. diffusion [1320, 1201, 1321, 1572, 1678, 1886, 81, 1889, 920, 432, 1826, 1533, 164, 1786, 66, 1467, 995, 1924, 553, 554, 803, 1343, 1690, 621, 681, 892, 1563, 117, 1651]. diffusionless [514]. diffusions [1226]. diffusive [206]. diffusively [148]. Digital [598, 1118, 537, 538, 84, 583, 177, 178, 582]. dimension [974, 394]. dimensional [261, 1673, 1194, 535, 1355, 1197, 228, 1295, 388, 1242, 450, 1573, 1176, 1689, 786, 472, 197, 35, 543, 399, 554, 174, 1239, 1530, 557, 923, 1606, 529, 681, 893, 1167, 311]. dimensions [684, 1458, 651, 1925, 1205, 892]. DIMEX [1302]. Diophantine [1800]. dioxide [956]. dip [1922]. dip-slip [1922]. Dirac [508]. Direct [32, 1154, 1452, 1931, 1801, 1304, 1723, 686, 513, 1684, 80, 1574]. directional [1395, 503, 1887]. directly [1906]. Dirichlet [792]. disabled [115]. discharge [1678, 355]. discontinuities [192, 1405]. discontinuity [1067, 636]. discontinuous [474, 1689, 432, 235, 1680, 220, 1353]. Discounting [1590]. Discrepancy [157, 1033, 538, 548]. Discrete [1817, 811, 1827, 1102, 1345, 902, 682, 1672, 638, 1318, 1196, 768, 337, 1595, 1266, 1052, 695, 1402, 1059, 1329, 229, 981, 197, 509, 510, 1272, 126, 491, 620, 607, 1880, 443, 783, 1924, 69, 1475, 1767, 1060, 1409, 1050, 1887, 1731, 1920, 891, 1270]. discrete-time [638, 768, 69]. discrete-type [1475, 1767]. discreteness [1005]. Discretization [47, 320, 1596, 980, 1731]. discretizations [482, 979, 1265, 1182, 1911, 1730, 54]. discretized [475, 1905]. disease [1892, 1567]. disinfestation [723]. disordered [528]. dispatch [1791, 16]. dispatching [1447]. Dispersion [1911, 1853, 1858, 1148, 1251, 1356, 1903]. dispersion-managed [1356, 1903]. dispersive [684, 1539, 998, 1906, 211, 239, 567, 519]. displacement [1191, 696, 1922, 105]. dissipation [1244, 202, 1008]. Dissipative [1263, 202, 1157]. dissipator [309]. dissolution [904]. distance [1293, 698]. distances [985]. distillation [422]. distinguishing [797]. distortion [1101]. distributed [97, 1381, 312, 1592, 247, 1053, 1533, 1207, 962, 1887, 1896]. Distribution [1660, 639, 715, 114, 1106, 1380, 1067, 1495, 1542, 1594, 118, 538, 1785, 1152, 1385, 1716, 1224, 1661, 1042, 1641, 636, 866]. Distribution-free [639]. Distributions [1697, 1738, 1654, 1517, 1556, 729, 1386]. disturbances [870, 61, 145, 596]. divergence [1815]. diversity [1791]. dividend [531, 936]. divisible [312]. division [1518]. divorce [1770]. Dixon [1422]. Do [666, 669, 1419]. does [1771]. dollar [940]. Domain [476, 914, 982, 988, 1316, 1370, 98, 752, 804, 34, 1327, 1153, 20, 998, 1906, 460]. Domain-wall [988]. domains [1304, 41, 1451]. domestic [1790]. dominance [1517, 639]. Dominant [717]. doors [421]. doped [1751]. dose [1693]. double [759, 804, 1611, 1612, 179, 1125, 109, 1555]. double-connected [804]. double-gate [1612]. double-talk [759, 109]. double-threshold [1555]. doubly [293, 1106]. drag [1153]. drastic [559]. Dressed [205]. drift [638]. drifting [751]. drive [1104, 597]. driven [693, 1522, 1576]. drives [1078, 1118, 585, 1112, 604, 1100, 586, 424]. driving [1825]. drop [1151]. droplet [614, 815, 905]. dropping [1181]. drought [1224]. drug [1560]. Drying [1149, 714]. DSI [994]. DSII [994]. DSS [629]. DTC [597]. Dual [1218, 1262, 1097, 1801]. dual-core [1262]. duality [1367]. dually [1464]. duct [1644, 1170]. ducts [1193, 1328, 41]. due [1200, 1922, 604, 265]. dump [153]. duopoly [319, 1795]. duration [1089]. durations [944]. during [870, 1056, 904]. Dust [1139]. Dyadic [1198]. Dym [200]. Dymola [708]. Dynamic [802, 1599, 1520, 559, 322, 1896, 363, 1570, 1795, 1552, 1922, 1562, 919, 1825, 1098, 1740, 882, 420, 487, 597, 77, 1284, 1185, 100, 241, 2, 1278, 1203, 834, 1737]. dynamic-game [100]. Dynamical [1889, 1629, 811, 1674, 1622, 1623, 589, 1909, 920, 1633, 1694, 294, 170, 295, 297, 54, 817, 85]. dynamically [1584]. Dynamics [69, 1271, 319, 1752, 871, 1817, 1708, 954, 120, 192, 1238, 709, 1177, 1541, 362, 1764, 508, 405, 1508, 1151, 1559, 743, 962, 1157, 868, 1487, 1004, 1845, 1920, 1446, 1663].

e-commerce [1436]. earnings [1755, 936, 636]. East [328, 1662, 935, 940, 1771]. eastern [863, 856]. EBDF [769]. echo [759, 109]. ECLMS [109]. ecological [1829, 965, 855, 1896]. ecologically [861]. ecology [100]. econometrics [931]. Economic [367, 366, 1791, 362, 859, 1418, 16, 365]. economics [363]. economies [1784, 945]. economy [364, 781, 100]. ecosystem [858, 1548]. ecosystems [1485]. eddy [1821, 1644, 1154]. edge [1310, 1728]. Editorial [524, 1582, 1, 1429, 564, 429, 437, 447, 456, 523, 571, 579, 610, 625, 726, 648, 657, 680, 689, 702, 775, 790, 807, 823, 841, 849, 879, 1037, 1046, 888, 898, 909, 917, 927, 950, 970, 991, 1012, 1020, 1029, 1065, 1074, 1084, 1093, 1124, 1132, 1174, 1202, 1212, 1222, 1232, 1255, 1276, 1282, 1290, 1299, 1313, 1352, 1360, 1376, 1391, 1415, 1427, 1443, 1455, 1493, 1502, 1514, 1536, 1636, 1525, 1580, 1669, 1688, 1713, 1746, 1777, 1798, 1812, 1838, 1866, 1884, 1899, 1928]. education [1603]. Effect [1596, 748, 876, 371, 1575, 1419, 1606, 78, 1607, 1008, 1278]. effective [275]. effectiveness [1786]. Effects [168, 347, 1478, 768, 1200, 377, 956, 1119, 743, 361, 1699, 1768, 670, 517, 1005, 1916]. efficiencies [937]. efficiency [1146, 1471, 1532, 273]. Efficient [1805, 48, 1801, 11, 1723, 300, 396, 1870, 1621, 1724, 111, 318, 901, 528, 722, 1620]. EGARCH [1759]. eggs [711]. eigenproblems [1877]. eigenstructure [954]. eigenvalue [1214]. eigenvalues [141, 984]. Einstein [1854, 1267, 1241, 505, 1852, 982, 988, 1271, 1000, 1004]. elaboration [27]. Elastic [1252, 1863, 533, 398, 1353]. elasticity [696, 804, 914, 1399, 393, 883, 1519, 488]. Elasto [1279]. Elasto-plasticity [1279]. elastoplastic [1325]. Election [1804]. Electric [1096, 581, 1697, 1452, 1931, 1104, 585, 1095, 614, 815, 905, 1185, 1684, 424]. electrical [416, 1114, 606, 158, 802, 55]. electrified [1539]. electrode [1697]. electrodynamics [1238]. electroencephalography [1494]. electromagnetic [1315, 1118, 1120, 518]. electromechanical [604]. electromigration [555]. electron [533, 534, 156, 1607, 866]. Electronic [1614, 589, 1215]. electronics [642, 594]. electrons [528]. Electrophoretic [1140]. electrotherapy [1697]. element [106, 1193, 1319, 1197, 228, 1176, 474, 1873, 279, 492, 739, 235, 220, 1334, 1828, 399, 317, 464, 311]. elementary [1718]. Elements [1198, 1180, 1341, 1344, 1346, 140]. Elimination [835, 549, 836]. Ellipsoidal [276, 219]. Elliptic [1619, 1697, 1318, 617, 43, 1329, 1335, 1183, 271, 1345, 463, 140]. elliptical [1449]. embed [1458]. Embedded [267, 787, 1869]. emergency [63]. emerging [634]. emission [156]. Emissions [781, 872, 662, 694]. Empirical [668, 1206, 640, 1431, 167, 1480, 1386]. EMTP [1101]. emulator [598]. enclosures [728]. end [708]. endogeneity [371]. endogenous [943, 365]. endplate [361]. energetic [255]. Energy [1356, 1396, 782, 672, 593, 1919, 1585, 1916]. enforce [1051]. enforced [874]. enforcing [1163]. engine [1193]. Engineering [825, 1509, 738, 836, 132, 461]. enhance [1786]. enhancing [256]. ensemble [171]. ensembles [516]. Ensuring [1698]. Entropic [1136, 1155, 1163]. entropy [1756, 57, 370, 310]. enumerable [835]. envelope [986]. Envelopes [1685, 826]. environment [826, 1593, 861, 853, 1406, 574]. environmental [1484, 630, 1463]. environments [130, 133]. enzyme [1600]. epidemic [692, 1783, 397, 1520, 1558]. equal [1654]. equalization [237, 382]. equalizer [237]. equation [1252, 1673, 187, 1355, 1197, 189, 975, 976, 1237, 193, 1573, 1424, 1305, 330, 496, 505, 1243, 1601, 1853, 507, 396, 1857, 1914, 539, 264, 1261, 1498, 920, 196, 981, 739, 1248, 149, 202, 546, 53, 1258, 1259, 1848, 514, 1924, 1025, 884, 1710, 999, 1874, 1023, 1342, 452, 190, 232, 1856, 1026, 1003, 621, 681, 1862, 1164, 1165, 1006, 518, 463, 741, 1915, 1007, 1563, 1008, 1657, 1647, 1871, 1701, 1859, 117, 304, 1270, 1574]. Equations [973, 993, 503, 1752, 810, 285, 1608, 1315, 682, 1632, 288, 1250, 1240, 1195, 1675, 653, 116, 684, 1387, 1423, 450, 1778, 1801, 138, 1398, 123, 832, 309, 1873, 441, 65, 1823, 1870, 753, 1800, 1326, 919, 1840, 394, 829, 43, 432, 1245, 1905, 433, 1265, 651, 1041, 199, 126, 1086, 1087, 164, 903, 451, 235, 1296, 203, 511, 385, 274, 119, 512, 1726, 994, 1842, 89, 997, 554, 785, 409, 236, 1919, 1560, 813, 1690, 1002, 681, 1920, 816, 1348, 893, 892, 239, 567, 519, 1125, 1792, 1918, 1013, 649]. equidistributed [1643]. equilibria [1631]. Equilibrium [45, 1752, 75, 120, 1139, 1090, 934, 1761, 196, 77]. equispaced [1682]. equity [1459, 1461]. equivalent [1695]. Erlang [1654]. Ermakov [1248]. erosion [1748]. erratic [1236]. Erratum [1209, 1219]. Error [1549, 491, 1830, 477, 1319, 228, 1033, 371, 979, 1559, 1510, 575, 632, 1772]. Errors [1792, 1378]. Establishing [1478]. Estimates [692, 477, 1319, 157, 491, 180]. Estimating [635, 921, 81, 529]. Estimation [1479, 1380, 1693, 1471, 937, 667, 255, 1772, 253, 354, 1387, 250, 1294, 1594, 1401, 18, 1562, 118, 1707, 1522, 761, 1655, 1480, 1736, 1750, 1515, 1042, 803, 121, 368, 633, 1396]. estimator [396, 667]. estimators [1469, 1206, 1768, 943]. Estuary [35]. Eucommia [256]. Euler [309, 47, 649]. Eulerian [844]. Euro [942]. Europa [1540]. European [935]. evaluating [1434]. Evaluation [1432, 1679, 146, 218, 1737, 640, 1439, 1437, 1417, 1485, 600, 1386, 357, 2, 1429, 1430]. event [735, 584, 1595, 1052, 541, 443, 783, 1060, 1050]. Events [215, 223, 234, 243, 259, 269, 428, 455, 570, 578, 609, 623, 848, 647, 656, 700, 774, 789, 806, 821, 840, 878, 1019, 886, 897, 908, 916, 948, 969, 1010, 1028, 1036, 1044, 1063, 1073, 1083, 1092, 1122, 1131, 1172, 1188, 1211, 1221, 1231, 1254, 1275, 1281, 1289, 1298, 1375, 1390, 1414, 1426, 1442, 1454, 1492, 1501, 1513, 1524, 1535, 1579, 1635, 1668, 1687, 1712, 1745, 1776, 1797, 1811, 1837, 1865, 1883, 1898, 1927, 306, 350, 380, 401, 435, 446, 500, 562, 679, 688]. events [763, 990, 926, 1609, 1059, 522, 1359]. evidence [1357, 1735, 1076, 1418, 1077, 1419]. Evolution [973, 993, 1719, 1773, 191, 192, 1791, 651, 1334, 512, 515, 816]. Evolutionary [611, 1439, 414, 1487]. evolutive [1379]. ex [1698]. Exact [1932, 1830, 753, 1923, 1933, 684, 1852, 978]. examination [1892]. example [341, 959, 962]. Examples [466, 1908]. exchange [876, 932, 843, 1762, 941]. exchanger [1225]. exchanges [1749, 1056]. excitable [212]. excitated [440]. excitation [1539]. excitations [772, 568]. excited [342, 277]. Existence [1516, 1699, 1800, 472]. exogenous [365]. exothermic [303]. expectations [615]. expected [1742]. experience [1489]. experiment [529]. experimental [1120, 1115, 1105, 255, 1113]. Experiments [218, 42, 43, 1629, 560]. expert [27, 413, 326, 70]. expertise [855]. explanation [1487]. explicit [1696, 1870, 913, 1726, 1025, 1563]. Exploding [1003]. exploiting [1437]. exploration [1854, 1268]. exponent [1034, 1618]. Exponential [809, 1659, 1665, 1645, 1516, 360, 1658, 1229]. exponentially [1416]. exports [934]. expression [1785, 4]. Extended [291, 1909, 1859, 871, 188, 118, 509, 1655, 758, 360]. extending [1218]. extension [595]. exterior [1153]. External [1180, 277, 568, 982, 378]. extra [1693]. extracting [158]. extraction [1721]. extreme [1042].

fabric [1439]. face [1438, 621]. facilities [619, 72]. Fact [1459]. facto [1698]. factor [1547, 1758, 1377, 256]. factorization [288, 793, 1700]. factors [645]. factory [272]. FADE [1417]. failure [1236, 316]. failures [1458]. families [33, 1844, 165, 544]. family [953]. farming [858]. farms [661]. Fast [694, 1682, 901, 60, 254, 3]. FastDer [722]. faster [280]. Fat [664, 1750, 1386]. fat-tailed [1386]. Fault [407, 408, 88, 1881, 1053, 439, 144, 881, 411, 19, 409, 403, 1405]. faulting [1922]. faults [601, 605, 1113]. February [334, 690, 899, 1223, 1233, 1392, 31, 38]. Fed [1106, 1114]. Feedback [698, 1497, 237, 420]. feedbacks [383]. FEM [1335, 1345, 1388, 1348]. female [667]. femtosecond [265]. fermentation [252]. Fermi [1260]. fermion [1902]. ferromagnetic [467]. ferromagnetism [1327]. fertilizer [1568, 1788]. Feshbach [1241]. FETI [475, 1323]. FETI-based [1323]. FETs [742]. few [986]. few-cycle [986]. Feynman [539]. FGMRES [1677]. fibrous [1323]. field [1697, 120, 1303, 614, 815, 905, 1910, 1504, 784, 1606, 574, 241, 1161, 1162]. fieldless [467]. fields [1553, 747]. fifth [1041, 1404, 999]. fifth-order [1041, 1404]. filament [189, 1246]. filaments [508]. filling [1725]. film [904]. filter [1440, 1118, 1024]. filtered [795]. filtering [1015, 417]. filters [1655, 1105, 1107]. fin [1532]. final [1424]. finance [1463, 1640, 553]. financial [1733, 1755, 1660, 1386, 664]. find [1624]. Finite [106, 186, 1198, 1143, 1068, 1334, 399, 317, 464, 140, 484, 1302, 1319, 1197, 797, 189, 228, 768, 1387, 1809, 1176, 1444, 474, 1873, 1601, 1364, 279, 492, 739, 542, 235, 830, 220, 1332, 119, 1828, 1726, 206, 514, 995, 1698, 784, 1341, 469, 1856, 1346, 79, 1566, 1646, 311, 304, 1405]. Finite-difference [186, 1143, 1444, 1856]. Finite-element [317]. finite-gap [189]. finite-sample [1387]. fins [1532]. fire [1449, 1565]. firmness [250]. First [627, 929, 1383, 1424, 1400, 618, 394, 1114]. first-order [1424, 394]. first-passage [1383, 618]. fish [1817]. Fisher [1247]. fishways [1347]. fission [213]. fissured [1816]. fitted [1553]. Fitting [1893, 255]. Fitzhugh [1710]. five [1309, 1308, 1725]. five-axis [1309, 1308, 1725]. fixed [1119, 1768]. fixed-frequency [1119]. flame [751]. flatness [798]. flexibility [1719]. flexible [1835, 832, 1637, 414, 343, 1373]. floating [158]. floods [862]. Floquet [189].

Flow [1533, 876, 1818, 875, 477, 1319, 191, 1320, 746, 473, 1644, 1679, 1294, 1146, 99, 1328, 1447, 220, 1332, 1154, 426, 854, 41, 399, 748, 1338, 212, 838, 1730, 466, 874, 1511, 462, 1349, 1166, 107, 1168]. flows [485, 1809, 745, 640, 472, 1148, 1509, 749, 146, 1891, 1239, 1155, 747, 1159, 1161, 1162, 1617, 1167, 1170]. flowshop [359]. fluctuation [860]. fluctuations [1549]. fluid [186, 1315, 98, 875, 240, 1177, 508, 1332, 854, 997, 874, 1249, 462, 1166]. fluids [472, 1331, 1156]. FLUKA [153]. fluorescence [254, 255]. flutter [342]. flux [1751, 1100, 1916]. fluxes [146]. flying [343]. flywheel [593]. FMS [1061]. Fock [528]. focused [86]. focusing [202]. Fokker [893]. fold [733]. following [1452, 1931]. food [110, 415]. Forage [714]. force [1825, 1153, 424]. forced [772, 191, 1842]. forced-damped [1842]. forces [801]. forcing [1200]. Forecast [369, 1665]. Forecasting [1762, 1487, 954, 1662, 405, 859, 8, 1561, 403, 376, 1460]. forecasts [1735, 671, 56]. foreign [1762]. forest [1449, 1716]. Foreword [581, 282, 184, 225, 260, 502, 992, 971, 1234, 1256, 1839, 1901, 1715]. form [1674, 1267, 1242, 1785, 913, 294]. formalised [1056]. formalism [980]. formation [1851, 11, 1910]. formed [346]. forming [1673]. forms [116, 1007]. formula [1433]. formulae [142]. formulas [835]. formulation [508, 159]. formulations [399]. four [1395, 118, 1119, 611, 1310, 1007]. four-bar [611]. four-directional [1395]. four-parameter [118]. four-triangle [1310]. four-wire [1119]. Fourier [50, 1236, 1237, 884]. fourth [1317, 515, 621, 117]. fourth-order [1317, 515, 621, 117]. FPGA [586]. FPGA-based [586]. fractal [1141, 1829, 297]. Fractional [1630, 1025, 1691, 1159, 1648, 368]. fractions [864]. fracture [1475]. fracturing [1176]. frame [1831]. framework [631, 1724, 1508, 1614, 1056, 663, 1446, 1657]. frameworks [1486]. Fredholm [1013]. Free [1916, 1674, 1200, 1809, 1679, 639, 602, 343, 1051]. Free-Choice [1051]. free-flying [343]. free-form [1674]. freedom [277, 568]. Frenet [1831]. Frequency [54, 440, 772, 98, 1119, 1269]. friction [1109, 914, 883, 424]. frictionless [317]. Front [736, 429, 437, 447, 456, 523, 564, 571, 579, 610, 625, 648, 657, 680, 689, 702, 775, 790, 807, 823, 841, 849, 879, 888, 898, 909, 917, 927, 950, 970, 991, 1012, 1020, 1029, 1841, 1307, 1003]. fronts [1682]. frozen [799]. fuel [781, 1166]. full [253, 1596, 1585, 64]. full-band [1596]. full-Galerkin [64]. fully [583, 750, 582]. fully-digital [582]. function [718, 73, 450, 617, 782, 618, 539, 159, 1785, 1357, 1521, 636]. Functional [430, 218, 1444, 913, 302, 549, 551, 1058]. functionals [813]. functions [759, 811, 1363, 1197, 1573, 1626, 157, 142, 620, 1152, 1480, 691, 1247, 41, 1691, 236]. fundamental [1006]. fundamentals [666]. funding [1473]. fungal [719]. Furnace [743]. Further [1569, 630, 91]. Future [130, 129]. futures [1552, 1786, 634, 374]. Fuzzy [62, 9, 7, 616, 1586, 1372, 1497, 1370, 1432, 814, 1878, 364, 1803, 1294, 1369, 919, 695, 414, 405, 781, 593, 1436, 310, 385, 386, 8, 10, 410, 1515, 409, 1888, 846, 1561, 14, 15, 1550]. fuzzy-model-based [919].

G [125, 1357]. G-7 [1357]. G7 [1419]. GA [16]. Gabor [1440, 1869]. Galerkin [976, 1237, 1801, 474, 1689, 1498, 1086, 235, 1828, 819, 64, 1241]. Galton [1340, 1716]. Gambling [181]. game [319, 273, 100]. games [77]. gamma [1594, 118, 168, 1224]. gamma-contaminated [168]. Gap [1902, 189, 462]. GAR [1875]. GARCH [843, 1387, 1764, 1740, 1741, 1377, 633, 1555, 1742]. Gas [1149, 1168, 662, 120, 1139, 1156, 466, 1170]. Gas-kinetic [1168]. gasoline [372]. gate [1611, 1612]. gate-all-around [1611]. gauge [1719]. Gauss [142]. Gaussian [872, 831, 1661]. GCC [1521]. GDP [1076, 672]. General [1675, 934, 1518, 341, 1444, 753, 203, 310, 1183, 296, 1730, 64]. generalization [1831]. Generalized [1135, 1852, 1007, 953, 1737, 1250, 914, 1677, 1594, 118, 393, 829, 1507, 795, 397, 639, 884, 1224, 1042, 1445, 1875, 1026, 1918, 1872]. generate [712]. generated [467, 760, 1246, 170]. Generation [1720, 1732, 1672, 98, 136, 1922, 1304, 1506, 1307, 1509, 1654, 545, 1925, 265, 1504, 748, 2, 1915]. Generator [606, 1381, 416, 1551, 545, 1808, 1099, 392]. Generators [181, 1643, 165, 544, 593, 167, 548]. generic [1595, 431]. Genetic [77, 12, 1088, 11, 279, 1826, 422, 13, 382, 384]. Geo [963, 861]. geo-information [861]. Geo-temporal [963]. geodesic [1505]. geographic [1758]. geographical [1368]. Geoinformation [863, 856]. geologic [1816]. geological [57]. geomechanics [480]. Geometric [1816, 1742]. geometrical [767]. geometries [1608]. geometry [377, 490, 743, 1831, 558]. GeV [153]. GeV/c [153]. Gibbs [633]. Gibraltar [1821]. Gierer [195]. Ginzburg [975, 1258, 1259, 1848, 1003]. given [914, 1716]. glass [412]. Global [1031, 397, 1887, 180, 1228, 911, 853, 1615, 337, 1516, 589, 472, 247, 1117, 1780, 1918, 1229]. globally [1645]. Gompertz [1785]. Gordon [1355, 1573, 510, 203, 923, 1125]. governance [1773]. governed [920, 1845]. governments [645]. gradient [1752, 481, 1510, 716]. Gram [43, 432, 1859]. Gram-type [1859]. granular [1001]. granules [1568]. graph [101, 1385]. graphic [826]. graphical [1108]. graphics [1895]. graphs [811, 84, 1769, 95, 1058]. gratings [796]. gravity [1245, 208, 1239]. gravity-capillary [1239]. Greater [1477]. green [220, 617, 618, 539]. Greenhouse [707, 662, 248, 716]. greenhouses [706]. GreenLab [1448]. Grey [723]. Grey-box [723]. Grid [473, 1060, 1720, 757, 1672, 1305, 135, 1551, 1509, 830, 1504]. Grids [1198, 1553, 1505, 1177, 1304, 1506, 309, 1129, 318, 1182, 1730, 621, 1732]. grippers [1589]. Gröbner [26, 836]. Gross [505]. Grossberg [1645, 1659, 1229]. ground [860, 1546]. ground-based [1546]. ground-water [860]. groundwater [75, 954, 354, 628, 1749, 169]. group [978, 108, 1911]. grouped [639]. Grouping [11]. groups [674]. growth [718, 1452, 1931, 1076, 367, 249, 644, 1448, 1418, 366, 1556, 729, 256, 365]. guaranteed [51]. GUI [144]. guidance [1111]. Guide [1045]. guiding [1178]. gyroid [1145].

Haar [157, 40, 301, 685, 686, 903, 236]. Hadamard [828]. hadron [153]. Halton [1014]. Hamiltonian [284, 33, 1626, 979, 542, 1583]. Hammerstein [1806]. hand [1199]. handling [1637, 794]. hanging [1388]. haplotype [1782]. haptic [1098]. hard [798, 1236, 793]. hardware [1381, 1619, 1620]. hardware/software [1620]. harmful [864]. Harmless [1229]. Harmonic [1101, 1672, 1505, 1260, 864, 1107]. harmonics [600, 1113]. Harmonization [444]. Harry [200]. Hartree [528]. harvester [710]. harvesting [1779, 922]. Haselgrove [560]. having [206, 239]. hazard [1475, 1767]. heart [1284]. heat [876, 1295, 102, 467, 1305, 1821, 330, 1531, 496, 1705, 1225, 149, 104, 1649, 304]. heater [677]. heating [924]. heavy [1702]. heavy-tailed [1702]. Heckman [667]. Heckscher [641]. hedging [1740, 1786].

Heisenberg [980]. held [36]. Helmholtz [1197, 192, 739, 1874, 741]. Hemodynamic [1603]. HEMTs [540]. HERMES [1348]. Hermite [138, 1344]. heterogeneous [1191, 1795, 1149, 1906, 469, 96]. heteroscedastic [1757]. heteroscedasticity [637, 1872]. heteroskedastic [1557, 1517]. heteroskedasticity [1736]. Heuristics [1422]. hexagonal [208]. Hicks [613]. Hierarchic [1344, 1346]. Hierarchical [160, 1807, 846]. hierarchies [1263, 200]. hierarchy [200, 985]. High [425, 875, 1572, 1421, 174, 1570, 535, 1197, 1679, 149, 783, 137, 77, 70, 557, 1159, 529, 1205, 893, 892]. high-dimensional [535, 557, 529]. High-order [1572, 149, 77]. high-performance [137]. High-resolution [875, 1421]. Higher [884, 1756, 684, 90, 1858, 507, 1914, 651, 199, 162, 1344, 1346, 1647]. Higher-order [884, 199, 1344, 1346, 1647]. higher-rank [162]. highly [1353]. Hilbert [1015]. hillslopes [967]. hip [1475]. Hirota [1914]. HiSIM [1610]. HIV [1209, 1219, 124, 1080]. hodograph [978, 200]. Hölder [141]. hole [172]. Holm [1853, 200]. holonomic [767]. homogeneity [1815, 1709, 1474]. homogeneous [1627, 490, 1906]. Homogenisation [1543]. Homogenization [496, 732]. homotopy [1921, 1266, 1532]. Hong [1760]. Hopf [1632, 1259]. Hopfield [1207, 654]. horizon [253, 1784, 1639]. hospital [1767]. house [770]. hp} [1345, 1388, 1348]. hp}-FEM [1345, 1388, 1348]. Huang [1015]. human [674, 938]. humans [122]. hump [186]. Hurst [1034]. Hybrid [1793, 1554, 382, 1140, 894, 337, 18, 1098, 781, 1789, 707, 691, 8, 1185, 1684, 278, 1105, 1344, 423, 1487, 16, 384]. hybridization [414]. hydraulic [876, 1816, 101, 1176, 1695, 819]. hydro [384]. hydrodynamic [1141, 35, 55]. hydrodynamics [1332]. hydrogen [1751]. hydrology [967]. hydropower [1103]. hydrostatic [1194, 105]. hydrothermal [16]. Hyman [1342]. Hyperbolic [53, 227, 1302, 1225, 1650, 91]. hyperparameter [1042]. hypersonic [749]. hypotheses [1515]. Hypothesis [1765, 674]. Hysteresis [590, 1088, 496, 491, 339].

ICMS [663]. ideal [1164]. identical [72]. identifiability [275, 1705]. identifiable [717]. Identification [1109, 1846, 1088, 1829, 798, 565, 1806, 612, 46, 1678, 1826, 144, 1694, 906, 1115, 1435, 424, 1560, 1411, 3]. identify [799]. Identifying [1424]. IFC [756, 765, 1525, 1580]. II [1751, 1252, 1263, 1703, 1206, 22, 255]. ill [1196, 442]. ill-conditioned [442]. ill-posed [1196]. illumination [1440]. illusion [1459]. IM [597]. IMAC [734]. IMACS [353, 1359, 215, 223, 234, 243, 259, 269, 299, 314, 182, 214, 222, 233, 242, 258, 268, 298, 305, 313, 350, 380, 401, 428, 435, 446, 324, 333, 390, 323, 332, 349, 379, 389, 400, 427, 434, 445, 455, 500, 522, 562, 570, 578, 609, 623, 454, 499, 521, 561, 569, 577, 608, 622, 848, 647, 656, 679, 688, 700, 763, 774, 789, 806, 821, 840, 878, 725, 755, 646, 655, 678, 687, 699, 724, 754, 762, 773, 788, 805, 820, 839, 847, 877, 990, 1019]. IMACS [886, 897, 908, 916, 926, 948, 969, 1010, 1028, 1036, 1044, 930, 952, 885, 896, 907, 915, 925, 947, 968, 989, 1009, 1018, 1027, 1035, 1043, 973, 993, 1063, 1073, 1083, 1092, 1122, 1131, 1172, 1188, 1062, 1072, 1082, 1091, 1121, 1130, 1171, 1187, 1211, 1221, 1231, 1254, 1275, 1281, 1289, 1298, 1210, 1220, 1230, 1253, 1274, 1280, 1288, 1297, 1375, 1390, 1414, 1426, 1442, 1454, 1492, 1501, 1513, 1524, 1535, 1579, 1635, 1374, 1389, 1413, 1425, 1441, 1453, 1491, 1500, 1512, 1523, 1534, 1578, 1634]. IMACS [1668, 1687, 1712, 1745, 1776, 1797, 1811, 1837, 1865, 1883, 1898, 1927, 1667, 1686, 1711, 1744, 1775, 1796, 1810, 1836, 1864, 1882, 1897, 1926, 627, 659, 929, 951, 1358]. IMACS-update [182]. image [1438, 160, 7, 758, 1729]. images [1721]. imaging [752]. imbalance [1552]. immersed [1549]. immigration [1340]. immiscible [1191]. Immune [1439, 881]. Immune-based [1439]. Impact [1845, 665, 1743, 1483, 1080, 1209, 1219, 1099, 1770]. impacts [942, 965]. impedance [802]. imperfect [1707]. imperfections [604, 543]. implantation [786]. Implementation [1325, 29, 1055, 507, 539, 149, 1577]. implementations [1430]. implicit [1193, 1194, 1261, 1328, 149, 1726, 834]. implicitization [1216]. imploding [1576]. importance [110, 1763]. Imposing [1346]. Improved [1791, 221, 535, 1196, 1483, 1545]. improvements [1496]. Improving [1397, 57]. impulse [711]. impulses [1516, 1416, 1794]. Impulsive [1704, 697, 812, 1558, 1550, 1783, 1889, 1699, 1520, 1659, 1779, 1896]. IMSC [735]. Inbound [939]. incidence [668, 1783, 1080, 1209, 1219, 397, 1558]. incident [156]. incidents [1472]. Including [1749, 1527, 1585, 1113]. inclusion [1781, 1445]. incomplete [670]. incompressible [1679, 65, 1332, 1730]. incorporating [1549, 48, 378]. increased [1217]. Incremental [65, 1690]. indentation [1802]. independence [1769]. independent [912, 1112]. Index [37, 67, 112, 216, 270, 307, 150, 351, 436, 501, 563, 822, 949, 1123, 1189, 624, 701, 764, 887, 1011, 1064, 1173, 1465, 1760, 90, 139, 1703, 1436, 1875, 91]. India [672]. Indication [489]. indicators [228]. indices [180, 1780]. Indirect [419, 424]. individuals [1340]. Indonesia [672]. induced [186, 504, 156, 1324, 863, 345]. Induction [1106, 1078, 601, 467, 606, 604, 605, 439, 1100, 254, 4]. induction-generated [467]. inductor [1097, 1119]. Industrial [1758, 1437, 1107]. industry [87, 1547, 1640, 1773]. inelastic [1252]. inequalities [475]. infection [1032]. inference [1294, 39, 378, 1789, 639, 1772]. Inferring [1674]. infinite [974, 1784]. infinite-dimension [974]. infinite-horizon [1784]. infinitely [53]. inflation [1790, 1470]. inflection [1175]. influence [1749, 880, 747]. influenza [1567]. Information [936, 852, 532, 932, 861, 135, 1079, 378, 1741, 1510, 421, 1585, 1368, 645]. inhibitory [1126, 1499]. inhomogeneous [452, 1004]. Initial [1527]. Initial-boundary [1527]. injection [1722]. Injectivity [831]. inlet [1533]. inner [830]. innovation [1547, 1758]. innovations [1702, 1377]. innovative [960]. Input [644, 1705, 1420]. insensitive [1511]. insider [932]. insightful [836]. inspired [12]. instabilities [958]. Instability [1840, 749, 263, 192, 746, 1841, 345, 210, 747, 1576]. Instability-wave [749]. instantaneous [872, 1772]. Integrability [1243]. integrable [199]. integral [1824, 539, 1070, 1373, 1013]. integrals [529, 1656]. integrated [1613, 1458, 957, 631, 147, 1616, 549, 1486]. Integration [1368, 1488, 1518, 288, 1743, 1921, 1870, 1840, 1521, 1842, 997, 1054, 178, 557, 1157, 1343, 559, 1205, 673, 134, 1519]. integrators [229]. integro [119, 1792, 1564]. integro-differential [119]. intelligence [869, 1484]. Intelligent [1433, 406, 1437, 403, 1430, 17, 238, 1429]. Interacting [976, 120, 1819, 1843, 1052, 528]. Interaction [266, 1570, 1240, 858, 727, 1472, 344, 100, 742]. interactions [1908, 1272, 1904, 513]. interactive [1057, 1108]. interception [249]. interconnected [814]. interconnection [39]. interday [1709]. interdependence [1774]. interdependencies [1598]. Interdisciplinary [322]. interest [933, 1757]. interface [875, 1162, 1278]. interface-tracking [1162]. interfaces [1098]. interfacial [1245, 1239]. interferometry [860]. interior [894]. interlocking [26, 327]. intermittency [1015]. intermodal [358]. Internal [1860, 1193, 568, 417, 1113, 1916]. Internally [1246]. International [973, 993, 1096, 1754, 640, 1766, 1490, 1357]. Internetics [134]. interpolants [1917, 1400, 1822, 1402, 1727]. interpolating [1830]. interpolation [1820, 1824, 318, 1216, 1409, 1834]. interpolatory [1306]. interpretation [1850, 144, 1378]. intersections [1624, 921]. interstate [1753]. Interval [787, 785, 1735, 218, 1893, 276, 1228]. Interval-based [787]. intervals [310]. intervention [668]. Intra [1741, 936]. Intra-daily [1741]. intra-day [936]. Intraday [1708, 1709]. Intrinsic [1260]. Introduction [1048, 245, 704, 1484]. intrusion [1411]. intuition [84]. Invariant [1633, 911, 1440, 833, 760, 489, 13, 1226, 1729]. invariants [753]. Inverse [1494, 5, 792, 800, 1295, 1846, 1860, 1399, 1689, 330, 442, 802, 795, 86, 1661, 1871]. inversion [51, 491]. inverter [584]. investigate [962]. Investigation [876, 874, 453, 1593, 309, 833]. investigations [290]. inviscid [462]. involved [1070]. ion [786]. Ionosphere [870]. ionospheric [866]. ions [156]. IP [1597]. IRI [449, 866]. irreducible [760]. irreducible-invariant [760]. island [945]. isolation [1793, 88, 408]. isometric [1505]. isotherms [720]. isotropic [490, 998]. Issue [704, 627, 659, 929, 951, 1715, 1429, 245]. issues [1387]. Itô [274]. iterated [1698]. Iteration [481, 1026]. iterations [1445]. Iterative [304, 1399, 842, 620, 303, 117]. ITI’99 [36]. IVPs [1530].

J [255]. Jacobi [142]. January [151, 325, 457, 658, 889, 1213, 21, 1670]. Japan [1354, 1472, 1475, 1767]. Japanese [372, 1467, 942, 1770, 645]. JAVA [246]. jet [513]. Jindrich [498]. jitter [1903]. job [414, 1373]. job-shop [414, 1373]. Johns [1218]. Joint [1387, 1286, 1716, 1322, 620]. July [68, 402, 1133, 1300, 929, 951, 269, 1813]. Jump [1467, 1786, 819, 1405]. jumping [1326]. jumps [843, 1739, 1759]. junction [361]. June [58, 244, 391, 580, 776, 36, 1456, 1799].

Kac [539]. Kaczmarz [758]. Kadomtsev [1252, 518, 1859]. KAM [1236]. Kármán [479]. Kazakhstan [870, 862, 866]. KdV [1250, 187, 1858, 507, 1857, 1807, 1041, 985, 207, 1025, 1856, 1008]. KdV-type [1807]. KdV/NLS [187]. keep [1293]. Kelvin [192]. Kennet [660]. Kernel [961, 396, 1694]. kinematic [1071]. kinetic [156, 396, 256, 1168]. kinetics [1676, 1417]. kitchen [238]. Klein [510]. knee [1322]. Knot [189]. knowledge [837, 48, 770, 1431, 252]. knowledge-based [837, 770, 252]. Kong [1760]. Korean [676, 1464]. Korteweg [187, 506, 200, 1356, 1903, 1862, 1647]. KP [1263, 1243, 999]. KP-II [1263]. Kriging [1722]. Krylov [1615]. Kung [1929]. Kuramoto [1673]. Kusuoka [553]. Kutta [1302, 1423, 309, 274, 515, 1564]. kVA [1099].

labor [667]. laboratory [1268]. Lactobacillus [718]. Ladik [1907]. Lagrange [290]. Lagrangian [1194, 832, 508, 980, 146, 452]. Laguerre [142]. Lake [1488, 1829]. lakes [1487]. Lambert [73]. Laminar [41, 745]. laminated [317]. LAN [730]. Land [860, 662, 628, 957, 355]. Land-surface [860]. land-use [628]. Landau [975, 1258, 1259, 1848, 1003]. landfall [1803]. landuse [1483]. Laplace [28, 1214, 1701].

Laplacian [497]. Large [312, 1821, 1644, 757, 650, 1877, 475, 1631, 99, 1548, 1154, 1416, 1474, 232, 654, 582, 1845, 817]. large-scale [232, 654, 817]. laser [1324, 727, 743, 175]. laser-induced [1324]. laser-material [727]. latent [85]. laterally [35]. Lattice [1144, 1147, 1148, 1150, 1151, 1152, 1284, 1158, 1167, 1169, 1170, 1136, 1138, 1141, 1142, 1143, 1907, 1260, 1644, 1145, 162, 1153, 166, 1154, 417, 1155, 1156, 1159, 1165, 1166, 1647, 1149, 1163]. lattice-Boltzmann [1145, 1166]. lattices [348, 1909, 1910, 1000]. launches [870]. law [1106, 597]. laws [483, 474, 201, 831]. layer [191, 749, 1249]. layered [1621, 1543]. layout [770]. LB [1140, 1160]. LC [205]. lead [483]. leads [1562]. learn [122]. Learning [421, 9, 303, 410, 14, 15, 4]. Least [1183, 1197, 1380, 1873, 1408, 1574]. least-squares [1873, 1408, 1574]. Lee [1929]. left [1674]. Legendre [80, 1013]. lemniscates [1192]. length [1767]. Leslie [45]. Lessons [1458]. lethal [1126]. lettuce [716]. level [1452, 1931, 954, 61, 1119, 1664, 1919, 1388]. levels [1681]. Levitation [1111]. life [1930, 1789, 1666]. life-span [1930, 1666]. lifetime [1703]. lifting [1545]. light [249, 220]. like [92, 470, 1833]. likelihood [280, 1387, 18, 118, 1655, 1480, 1736, 667, 368]. likelihood-based [280]. limit [1853]. limited [72, 758]. limited-data [758]. limits [461]. Lindley [1495, 1542]. Lindstedt [33, 293]. line [565, 1252, 253, 430, 29, 612, 711, 1150, 1528, 904, 1732]. line-soliton [1252]. line-width [904]. Linear [1825, 1694, 962, 531, 1236, 1675, 1843, 894, 90, 1696, 1268, 709, 155, 1801, 300, 221, 1399, 34, 1643, 1763, 1800, 685, 321, 43, 433, 1272, 418, 1296, 144, 420, 1247, 397, 1842, 1115, 170, 398, 77, 211, 1602, 276, 219, 423, 515, 1641, 465, 1249, 408, 722, 461, 376, 91, 1792, 311]. linear-quadratic [1696]. linearities [277]. linearity [440, 1248, 1357, 204]. linearly [1261, 1920, 1566]. Lines [226, 106, 98, 1695, 555, 231]. LINEX [1206]. link [1589, 1780]. linking [662]. links [1486]. Liouville [442]. liquid [1143, 747, 1170]. liquids [893]. Lissajous [291]. listed [1464]. liveness [1051]. livestock [858]. LMI [1497, 1024]. LMS [683]. load [312, 1791, 8, 1641]. loading [241]. loadings [871]. loads [168, 71]. loan [1734]. loans [1734]. Lobachevsky [490]. Local [394, 980, 1728, 1137, 485, 193, 1305, 505, 1891, 1181, 549, 407, 645]. Localised [994]. localization [51]. Localized [261, 972, 976, 1260, 266]. locally [348, 1194]. location [1081]. loci [453]. locking [1264]. logic [62, 364, 1369, 414, 405, 781, 23, 593, 326, 1888, 846]. logical [39, 1718]. logistic [53, 514, 616]. Long [1459, 1912, 479, 1240, 976, 753, 1787, 985, 516, 1026]. Long-run [1459]. long-term [1787]. long-time [516]. longest [1310, 1728]. longest-edge [1310, 1728]. longitudinal [1154]. Looking [169]. loop [832, 906, 77]. loops [1642]. Lorentz [1238]. Lorenz [639]. loss [1016, 1206, 489]. losses [1100]. low [1078, 1452, 1931, 485, 1114, 1629]. low-level [1452, 1931]. low-rank [1629]. lowland [660]. LQG [768]. LRD [1479]. LSM [1111]. LSQR [1196]. lumped [97, 1749]. Lurie [1364]. Lyapunov [1363, 813]. lysimeter [354].

M [245, 125, 704]. M/G/1 [125]. M2SABI’01 [705]. Mach [485, 749]. machine [592, 74, 1120, 1309, 1308, 1117, 1106]. Machines [1096, 581, 1095, 601, 1114, 606, 1113, 1349]. machining [1879, 1725]. macro [362]. macro-economic [362]. macroeconomic [675]. macromolecules [1421]. made [1266]. Mae [1483]. maglev [1111]. magnet [980]. magnetic [1116, 1115, 747, 1161]. magnetostrictive [1379]. magnets [602]. Magnitude [1893]. maintenance [619, 72, 1575]. major [676, 660]. Makeham [1380, 1785]. Making [1784, 1888, 966, 891, 1203]. Malaysia [373]. Maldives [1490]. male [494]. man [74]. manage [1056]. managed [1356, 1903]. Management [1488, 861, 1722, 1458, 1590, 661, 630, 1059, 1485, 1185, 965, 660]. Managing [654, 1733, 27, 444]. Manakov [1852, 984]. maneuvering [154]. manifest [85]. manifestation [1238]. manifolds [1627, 833, 1682]. manipulators [1589]. manufacturing [1547, 415, 1637, 616]. many [53]. map [1754, 1782, 713]. Mapping [1804, 792, 1116, 41, 817]. mappings [831]. maps [284, 895, 170]. March [224, 703, 1075, 44, 1714]. marginally [717]. marine [1593, 1485]. market [1752, 1735, 665, 1462, 615, 1761, 372, 1467, 1555, 1742]. markets [375, 1552, 1743, 1762, 1292, 1559, 1774, 942, 634, 946, 1468, 1804, 1742]. Markov [669, 1819, 1354, 1664, 1646]. Markov-switching [669]. Markovian [649]. Maruyama [649]. mass [1339, 1000, 256]. massless [293]. matching [410]. material [727]. materials [261, 1379, 1323, 1279, 1334, 750, 1543, 1001]. Math [1932, 1931, 1929, 1209, 1219, 1933]. Mathematica [292, 1656]. Mathematical [871, 1568, 1452, 1931, 129, 652, 494, 1583, 721, 462, 837, 483, 587, 1696, 1567, 719, 180, 856]. MATHMOD [778]. MATLAB [590]. MATLAB/Power [590]. matrices [1626, 1624, 828, 1700, 1216]. Matrix [1182, 312, 1625, 43, 432, 45]. matroids [95]. matter [1854]. Maximally [1643]. Maximum [18, 118, 1655, 1480, 368, 1756, 280, 1387, 1329, 828, 1736, 667]. maximums [761]. Maxwell [1608, 1315, 1238, 1348]. May [766, 918, 928, 1094, 1291, 92]. MCM [525]. MD [1140]. MD/LB [1140]. mean [1481, 1557, 1321, 1702, 146, 1910, 1382]. means [799]. measles [1515]. measurability [54]. measure [1791, 370]. measured [761]. measurement [1424, 371, 241]. measurements [153, 280, 250, 711, 799, 442, 1649, 1585, 1586]. measures [1780]. Measuring [1354, 938]. mechanical [782, 833, 604, 1330, 1284]. mechanics [1135, 1850, 292, 1180, 461, 741]. mechanism [611, 1127]. mechanisms [131]. mechatronic [779]. media [1816, 797, 1676, 1320, 46, 487, 251, 998, 1906, 212, 469, 1158, 1913, 1004]. mediaeval [652]. medical [1138]. Mediterranean [714]. medium [1723, 266, 317]. Meinhardt [195]. Mel’nikov [190]. MEM [1741]. MEM-GARCH [1741]. membranes [1421]. memory [479, 396, 17, 1334, 750, 278, 1887, 1165]. Merton [1929, 1876]. mesenchyme [1330]. mesh [1395, 1818, 136, 485, 1307, 1328, 1507, 1510]. mesh-adaptation [1818]. meshes [482, 1396, 136, 1400, 1344, 79]. Meshkov [746, 747, 1576]. Meshless [201, 1337, 1577]. mesophase [1145]. Mesoscopic [1137]. metacontinuum [1238]. metal [467, 747]. metallic [555]. metals [374, 1468]. metamodel [1722]. Method [864, 810, 154, 484, 1136, 1193, 1549, 481, 1101, 33, 1236, 535, 1319, 1197, 800, 1140, 1295, 1805, 976, 1237, 894, 485, 476, 11, 1572, 1573, 1809, 474, 467, 1877, 1689, 1057, 1380, 1873, 1591, 793, 799, 769, 686, 1532, 1498, 1307, 321, 1328, 433, 739, 293, 94, 161, 149, 139, 1152, 1153, 1548, 1214, 235, 546, 220, 1880, 795, 310, 1828, 1726, 8, 1154, 1337, 554, 1710, 1408, 1832, 602, 1155, 80, 1342, 231, 1159, 1561, 1026, 1690, 621, 1162, 560, 636, 392, 1411, 1511, 817, 1125, 107, 16, 1647, 921, 891, 1871, 1701, 1013, 64, 649, 1545, 1574]. Method [226, 1706]. Methodological [859]. Methodology [322]. methods [531, 902, 336, 75, 677, 1302, 1318, 875, 669, 1675, 228, 1109, 1627, 1199, 1615, 1423, 1921, 1444, 1398, 588, 1822, 1630, 441, 32, 507, 279, 493, 1567, 1129, 1447, 201, 1086, 1087, 163, 1700, 830, 274, 294, 1842, 86, 1577, 1894, 794, 1874, 600, 56, 557, 1856, 529, 515, 865, 79, 1563, 1792, 1353, 1657, 1564, 1203, 834, 594]. metric [1505]. MeV [156]. MIC [1335]. mice [1452, 1931]. micro [1103, 378]. micro-simulation [378]. microbial [110, 729]. microchannel [745]. Microcontinuum [468]. microengine [175]. microflows [1137]. microresonators [983]. microscale [304]. microstructural [495]. microstructured [1001]. microstructures [1279]. microvascular [1147]. migration [1879, 179, 1365]. mill [708]. milling [1309, 1308]. mimetic [830, 1182]. MIMO [1615, 1364]. Minimal [1396, 1834]. Minimisation [1227, 1527]. minimization [864, 1574]. Minimum [632, 370, 961, 1731]. minors [828]. mirror [1616]. mitigation [1107]. Mixed [492, 1316, 1177, 1873, 1277, 1335, 1659]. mixing [745]. mixture [876, 1738, 1757, 831]. mixture-of-normal [1738]. mixtures [1136, 42, 1852]. MKP [518, 1263]. MKP-II [1263]. mobile [9, 381, 1365]. mobility [983, 1607]. Modal [513]. mode [1881, 1637, 1059, 1781, 1264]. mode-coupling [1264]. mode-locking [1264]. Model [1615, 1059, 880, 967, 1932, 591, 757, 692, 1851, 733, 1795, 1478, 1568, 1469, 1737, 693, 1194, 1674, 1109, 1379, 1540, 933, 1354, 1088, 1200, 1143, 1929, 715, 709, 858, 601, 1176, 1336, 74, 272, 782, 1114, 1610, 912, 370, 1470, 1592, 957, 1449, 61, 576, 124, 1401, 362, 675, 1763, 1749, 118, 1277, 919, 1279, 494, 1519, 247, 195, 320, 1889, 1567, 1245, 5, 1448, 980, 1826, 94, 490, 1330, 1876, 1787, 35, 815, 1357, 1703, 1559, 1789, 1436, 1080, 1209, 1219, 1863, 881, 1467, 1565, 168, 444, 1829, 397, 1612, 1077, 1767, 1699, 417, 603, 943]. model [1520, 1117, 1100, 1339, 1102, 100, 1560, 1584, 360, 1875, 1544, 254, 213, 558, 1411, 1923, 1933, 368, 1772, 856, 633, 107, 256, 1647, 1555, 1872, 1779, 64, 922, 1896, 866, 1558, 1550, 1529, 1742, 844]. model-based [881, 1550]. Model-order [1615]. Modelica [101]. Modeling [1096, 708, 1095, 1098, 111, 1611, 105, 1110, 1127, 1333, 108, 784, 586, 1585, 322, 1113, 1166, 1890, 186, 1892, 1452, 1931, 779, 1589, 590, 1076, 1793, 1068, 110, 1458, 1590, 1449, 863, 1119, 1597, 1695, 707, 8, 995, 997, 171, 998, 1906, 1054, 468, 1156, 1583, 96, 616, 742, 1386, 855, 311, 1692]. modelled [1483]. Modelling [97, 1753, 1733, 932, 930, 952, 1103, 662, 1755, 642, 1757, 1756, 1662, 643, 1820, 581, 102, 1057, 248, 672, 88, 1748, 1324, 727, 956, 1640, 959, 780, 913, 1331, 1330, 905, 882, 960, 627, 659, 951, 104, 1115, 1490, 720, 1734, 1349, 85, 1382, 1646, 1460, 387, 106, 718, 677, 1137, 1318, 653, 1721, 1484, 1320, 652, 1322, 628, 1323, 364, 1114, 63, 631, 25, 589, 661, 1225, 1328, 492, 712, 543, 1766, 1332, 1334, 1385, 750, 929, 1485, 55, 103, 854, 89, 211, 278, 1486, 469, 966, 1488, 721, 252, 893, 376, 462, 356, 459]. models [531, 871, 1735, 483, 872, 1465, 1738, 953, 1817, 75, 662, 1040, 638, 723, 1806, 98, 101, 328, 669, 671, 1461, 354, 814, 1557, 1878, 1321, 1696, 1444, 808, 1609, 1273, 371, 1693, 1595, 1384, 630, 1562, 958, 1603, 1783, 695, 1149, 147, 1825, 510, 1522, 378, 1739, 230, 146, 125, 1740, 1548, 1421, 1786, 707, 220, 1480, 1736, 443, 1891, 93, 783, 1750, 738, 103, 426, 667, 1475, 172, 1598, 962, 1768, 1769, 19, 944, 714, 1264, 45, 276, 219, 1833, 452, 634, 180, 1482, 255, 1734, 252, 698, 637, 1164, 819, 664, 1586]. modes [972, 1260, 213]. modification [370, 997]. Modified [1195, 200, 392, 518, 684, 1295, 799, 207, 1874, 1023]. Modular [1348]. Modulated [975, 709, 266]. modulation [1845]. module [9, 343]. moist [107]. MOL [227]. molecular [1139, 175]. moment [1647]. moments [1756, 1638, 1698, 1729]. momentum [1462]. monetary [1521, 1771, 673]. Money [1470, 1521, 643, 1357]. Monitoring [1079, 853, 413, 846, 862, 852, 1368, 660, 1545]. monolitic [27]. monotonicity [1403]. Monte [532, 1518, 154, 533, 1469, 1101, 534, 1381, 536, 1819, 155, 617, 156, 396, 526, 786, 159, 540, 1596, 160, 880, 161, 541, 542, 543, 1517, 166, 546, 169, 547, 1417, 171, 550, 551, 172, 552, 904, 176, 555, 1698, 729, 178, 557, 742, 179, 121, 180, 529, 559, 892, 921]. Montgomery [1618]. monthly [355]. monuments [1721]. mooring [106]. Moreau [1325]. Morocco [328]. morphological [715, 1330]. Morse [115]. MOS [1613]. MOSFET [1610, 1612]. MOSFETs [1611]. motion [287, 1383, 280, 832, 753, 1150, 1421, 175, 1039, 381, 1742]. motions [344]. motor [1078, 1104, 1112, 604, 605, 1100]. motors [588, 439, 595]. Moulton [1630]. movement [1336, 963, 1646]. movie [1755]. Moving [913, 1508, 871, 875, 139, 905, 381, 1496]. Moving-body [1508]. MPDAEs [1527]. MPI [488]. MSE [961]. MSSANZ [353, 930, 952, 627, 659, 929, 951]. MSSANZ/IMACS [353, 930, 952, 627, 659, 929, 951]. MULINO [629]. MULINO-DSS [629]. Mullins [1908]. Multi [1551, 1059, 1366, 202, 1692, 261, 186, 693, 592, 1381, 1432, 1589, 1676, 1303, 1398, 359, 1447, 411, 1842, 1102, 922]. multi-agent-architecture [411]. multi-batch [693]. Multi-block [1551]. multi-cell [1102]. multi-converter [592]. Multi-criteria [1366, 1432, 1447]. multi-cycle [922]. multi-dimensional [261]. multi-fluid [186]. Multi-focusing [202]. multi-link [1589]. multi-machine [592]. Multi-Model [1059]. multi-point [1381]. multi-pond [922]. multi-resolution [1692]. Multi-scale [1692, 1303]. multi-species [1676]. multi-splitting [1398]. multi-stage [359]. multi-symplectic [1842]. multibody [832]. multibus [1107]. multicell [598]. multidimensional [1302, 1909]. Multidisciplinary [1303]. Multifrequency [795]. multigate [1606]. multigrid [621, 1353]. multilevel [587, 485]. multinomial [1815]. multiobjective [1682]. Multiple [591, 671, 1656, 1203, 440, 566, 1481, 814, 1241, 1079, 1625, 1070, 1447, 420, 1664, 1485, 1642, 1228, 392, 1732, 891]. multiple-attribute [891]. multiple-control [1642]. multiple-use [1485]. multipopulation [416]. multiprocessor [71]. multiresolution [1729]. multiresponse [891]. Multiscale [1680, 1341]. multisensor [62]. multiserver [619, 72]. multispectral [1186]. multistep [1792]. multisymplectic [979, 1911]. Multisymplecticity [1849]. multitarget [280]. multivariable [906]. Multivariate [1665, 1463, 1465, 1764, 1736, 56, 1544, 1872, 1742]. multivector [116]. multiwavelets [1022]. municipal [405]. mutation [611]. mysterious [210].

Nagumo [1710]. Nahm [1905]. nano [1609]. nano-scale [1609]. nanotube [743, 1605]. nanowire [1606, 744]. Nash [1752]. natural [861, 1458, 1364, 728, 1158, 1249, 1718]. nature [1841, 873]. naval [19]. Navier [1316, 65]. Near [1395, 1727]. Near-best [1395]. nearly [1242]. Necas [498]. negative [1654, 1855, 1000]. negative-mass [1000]. neighbor [56]. Nessyahu [997]. net [443]. nets [1057, 537, 538, 1061, 882, 177, 1051, 730, 1053, 15]. Network [607, 1613, 1645, 1793, 48, 576, 1522, 1330, 1597, 20, 413, 1598, 1184, 1100, 1719, 965, 654, 297, 1641, 1487, 1663, 922, 1365]. network-based [1613, 1663]. networked [1642]. networks [347, 1499, 1847, 312, 316, 1516, 1031, 606, 5, 605, 302, 382, 707, 1207, 426, 69, 1416, 383, 1599, 19, 346, 1887, 407, 1228, 423, 809, 1368, 408, 1446, 2, 1659, 3, 1658, 1794, 1229, 1651]. networks-based [316]. Neumann [59, 463]. Neural [1613, 316, 605, 20, 257, 19, 1100, 14, 1663, 1499, 1645, 1793, 48, 1516, 18, 1031, 1522, 419, 418, 707, 1207, 426, 69, 413, 383, 1184, 654, 407, 1228, 423, 809, 1487, 1368, 408, 2, 1659, 3, 1658, 1794, 1229, 1651]. neural-classical [18]. Neuro [13, 1294]. Neuro-genetic [13]. neurocontrol [6]. neurologic [360]. neuromuscular [361]. neutral [1778, 1119, 1285, 1880, 1024, 1564]. neutron [1070]. News [182, 214, 222, 233, 242, 258, 268, 298, 305, 313, 323, 332, 349, 379, 389, 400, 427, 434, 445, 454, 499, 521, 561, 569, 577, 608, 622, 646, 655, 678, 687, 699, 724, 734, 754, 762, 773, 788, 805, 820, 839, 847, 877, 885, 896, 907, 925, 947, 968, 989, 1009, 1018, 1027, 1035, 1043, 1062, 1072, 1082, 1091, 1121, 1130, 1171, 1187, 1210, 1220, 1230, 1253, 1274, 1280, 1288, 1297, 1358, 1311, 1350, 1374, 1389, 1413, 1425, 1441, 1453, 1491, 1500, 1512, 1523, 1534, 1578, 1634, 1667]. News [1686, 1711, 1744, 1775, 1796, 1810, 1836, 1864, 1882, 1897, 1926, 944]. Newton [286, 1628, 1236, 1591, 799]. Newtonian [1145, 1331]. next [56]. NICE [1870]. Nipp [331]. nitrate [716]. nitriding [1678]. NLPDES [1844]. NLS [187, 263, 190]. NLSE [1845]. NLSOFT [1108]. no [794]. no-response [794]. node [2]. nodes [1388]. Noise [711, 761, 345, 50, 1118, 768, 1120, 1546, 1533, 1869]. Noise-induced [345]. noisy [1827, 1538]. Non [1610, 1145, 144, 204, 515, 1249, 376, 311, 440, 261, 531, 483, 75, 1236, 1850, 1032, 193, 709, 467, 277, 1886, 1763, 1800, 43, 1272, 72, 1331, 1633, 418, 119, 1335, 397, 1182, 1115, 398, 998, 1649, 211, 748, 1066, 212, 1017, 466, 423, 595, 465, 408, 722, 1164, 461, 71, 520, 22]. non-autonomous [1032]. non-conforming [1335]. non-constant [1886, 520]. non-cooperative [1066]. non-equilibrium [75]. non-existence [1800]. non-ferromagnetic [467]. non-ideal [1164]. non-identical [72]. non-isotropic [998].

Non-linear [144, 515, 1249, 376, 311, 531, 709, 1763, 43, 1272, 418, 397, 1115, 398, 211, 423, 465, 408, 722, 461]. non-linearities [277]. Non-linearity [204, 440]. non-local [193]. Non-Newtonian [1145, 1331]. non-probabilistic [1850]. Non-quasi-static [1610]. non-resonant [261]. non-ruin [119]. non-sinusoidal [595]. non-smooth [1633]. non-solenoidal [212]. non-standard [483, 1649, 22]. non-traditional [1017]. non-uniform [1182, 748, 71]. non-uniqueness [1236, 466]. nonalgebraic [289]. nonautonomous [1651]. noncompact [519]. Nonequilibrium [540]. nonexistence [490]. nonextensive [1135]. nonhomogeneous [1531]. Nonlinear [973, 993, 1843, 342, 1539, 94, 1575, 511, 1251, 1102, 1913, 772, 972, 565, 1497, 696, 262, 1363, 187, 191, 684, 814, 192, 275, 568, 1398, 1273, 289, 1470, 32, 1369, 1870, 1909, 301, 1554, 1840, 264, 1261, 1498, 196, 1861, 1108, 1204, 1329, 229, 981, 433, 509, 510, 1265, 651, 198, 1653, 1366, 202, 546, 1904, 1694, 1246, 205, 512, 169, 750, 206, 995, 1924, 294, 1528, 884, 1264, 803, 1583, 180, 1343, 1912, 1920, 816, 1663, 239, 567, 596, 1013, 1558, 1529]. nonlinearities [669, 1326, 1482]. Nonlinearity [1248, 1262, 638, 263, 1557]. nonlinearly [1566, 519]. nonlocal [1410, 484]. nonnegativity [1345]. Nonoverlapping [34]. nonparametric [669]. nonsmooth [1316, 1591]. Nonstandard [484, 1444, 1567, 1715, 1199, 206, 514, 995]. nonuniform [1328, 1129, 545]. Nordsieck [1129]. Normal [294, 1738, 632]. North [959]. northern [863, 355]. note [66, 1648]. Novel [192, 546, 1617, 115]. November [127, 448, 626, 850, 1038, 1537, 1900]. nozzle [709]. NPC [596]. NSFD [1247]. Nuclear [413, 1318, 1585]. Number [181, 481, 485, 1643, 165, 167, 1808, 548, 1716, 1340, 1159, 392, 1511]. numbers [1381, 141, 749]. numeric [290]. Numerical [1410, 1854, 1138, 1250, 1319, 653, 1553, 1379, 1676, 42, 1322, 57, 1689, 123, 46, 330, 1531, 1601, 124, 977, 1857, 1146, 745, 472, 264, 1498, 43, 903, 1751, 487, 854, 996, 997, 470, 471, 175, 211, 1720, 1605, 1338, 1023, 1833, 556, 132, 923, 1343, 1026, 865, 465, 1161, 1163, 131, 356, 1566, 1576, 1697, 483, 1315, 1240, 752, 1355, 191, 1721, 804, 1199, 1268, 1573, 1016, 395, 1609, 1853, 1870, 769, 1567, 924, 433, 163, 513, 1154, 994, 265, 1025, 1629, 173, 1602, 1342, 1730, 1912, 1650, 560, 1249, 530, 1563, 1617]. numerically [1725]. Nutrient [660, 1016]. nutrients [1568].

O [255]. oats [714]. object [97]. object-oriented [97]. objectives [122]. objects [297, 381]. observability [1052]. observables [542]. observation [913]. observations [1051]. observer [895, 598]. observer-based [895]. observers [253]. obsolescence [955]. obstacle [1394, 802, 795, 1706]. obstacles [793, 59, 1732]. obtain [1407]. ocean [1860, 1200]. oceanic [1878]. oceanographic [504]. October [113, 438, 1526, 1885]. ODE [1125]. ODEs [1199, 289, 769]. odours [1431]. off [781]. offensive [1126]. offerings [1459]. office [1755]. official [1481]. Ohlin [641]. oil [876, 1336]. oligopoly [1752]. OLS [1562]. On-line [565, 253, 612, 711]. on-site [1260]. One [529, 1194, 974, 1197, 388, 277, 568, 1689, 796, 197, 554, 1894, 1530, 681, 1167]. one-degree-of-freedom [277]. one-dimensional [1194, 1197, 388, 1689, 197, 554, 1530, 681, 1167]. one-shot [1894]. one-sided [974]. only [280]. open [1644, 832]. open-loop [832]. operating [1078, 1059, 596]. Operation [869, 1571, 359, 922]. operations [606]. Operator [844, 75, 1370, 442, 28, 1214, 1017, 1378, 1203]. operators [1395, 1268, 1824, 1306, 1266, 1625, 611, 491, 1182, 1435]. Optical [262, 986, 261, 1910, 1000, 865, 517]. Optimal [1081, 1347, 1571, 955, 1371, 732, 1740, 426, 1684, 431, 1107, 922, 565, 1014, 612, 1591, 1601, 32, 321, 783, 1515, 232, 515, 716]. optimality [414]. optimisation [894, 1592, 254, 716]. optimisation-oriented [1592]. Optimization [619, 1788, 740, 1309, 1308, 1185, 602, 1871, 217, 1683, 902, 1104, 1109, 1627, 869, 894, 1722, 102, 1305, 300, 61, 32, 1279, 247, 147, 611, 1724, 1366, 1616, 1614, 830, 1024, 1682, 1894, 1100, 654, 431, 817, 891]. optimizing [1482]. Option [1929, 1876, 1519, 1559]. options [957, 1292, 139]. optoelectronic [1068]. orbit [742]. Orbital [1855, 344]. order [1317, 1355, 1615, 1572, 1424, 1801, 912, 1679, 1400, 1326, 1914, 394, 1041, 199, 149, 1129, 1404, 274, 1726, 103, 884, 997, 961, 77, 999, 1875, 1344, 1346, 515, 621, 1648, 1647, 117]. ordered [1475, 645]. ordinary [285, 288, 1675, 653, 1398, 1326, 394, 1296]. ordinates [1730]. organic [1825]. organizing [1879, 1782, 713]. oriented [97, 1592, 96]. Orthogonal [109, 1407, 1451]. orthogonality [1346]. orthogonalization [1602]. oscillation [1794]. Oscillations [339]. oscillator [346]. oscillators [347, 348, 291, 346, 812, 148]. oscillatory [1846]. Ostrovsky [1006]. other [826, 166]. outbound [676, 1489]. outdegrees [1716]. outer [1179]. output [733, 1364, 1643, 644, 420]. overset [1508, 1730]. oversized [408]. overview [1733]. OWA [1435]. Oxygen [1179].

P [255]. package [232]. packages [1371]. packet [191]. Padé [1342, 819]. Pages [49, 76, 143, 31, 38, 21, 58, 68, 44, 127, 113, 82, 281, 308, 315, 151, 244, 224, 334, 325, 391, 402, 448, 438, 572, 457, 580, 626, 737, 824, 690, 658, 776, 703, 766, 850, 910, 1047, 899, 889, 918, 928, 1038, 1021, 1030, 1085, 1133, 1075, 1094, 1283, 1361, 1223, 1233, 1213, 1300, 1291, 1428, 1503, 1581, 1392, 1456, 1537, 1526, 1670, 1799, 1813, 1714, 1900, 1885]. Painlevé [199]. panel [937, 1521, 1419, 670, 1354]. pantograph [1570]. paper [708]. Papers [627, 659, 929, 951]. parabolic [1410, 1318, 450, 1424, 123, 1919, 741, 1657, 1871]. paradigm [312, 135]. paradox [1476]. Parallel [532, 757, 1819, 155, 507, 1307, 149, 728, 1320, 546, 1700, 167, 237, 273, 278, 178, 117]. parallelization [1809]. Parameter [912, 1594, 1705, 1750, 1560, 254, 817, 3, 768, 1624, 123, 537, 1631, 18, 118, 717, 1654, 880, 900, 906, 1598, 729, 1845, 368]. parameter-influence [880]. Parameters [1411, 1396, 668, 1088, 1380, 880, 1538, 865]. Parametric [1789, 772, 1835, 277, 568, 549, 1409, 1435]. parametrization [1505]. Pareto [414, 1682]. Pareto-optimality [414]. parity [409]. Park [595]. part [708, 1309, 1308]. Partial [148, 1400, 441, 433, 1560, 681]. partially [1051]. Particle [164, 1850, 894, 1260, 1332, 556, 1157]. particle-based [556]. particles [1148, 104, 1845]. partition [1728]. partitioning [136, 1620]. partitions [1310]. parts [793]. passage [1383, 618]. passenger [838]. Passive [409, 606, 1264]. past [129]. Pasta [1260]. patch [662]. patches [1889, 1341]. patents [642]. path [958, 539, 1309, 1308, 273, 1528]. path-integral [539]. paths [534]. patients [1767]. Pattern [1431, 1673, 1146, 410]. pattern-forming [1673]. patterned [1904]. Patterns [346, 939, 204, 212, 1054]. pavement [241, 449]. payload [343]. PC [583]. PC-cluster [583]. PDAEs [91]. PDE [230, 206, 995, 1247, 1409, 132, 271, 722].

PDEs [227, 1631, 979, 1894, 560]. PDM [1055]. pedestrian [921]. pedestrians [99]. pelagic [1817]. percolation [251]. percolative [176]. perfect [797]. Performance [600, 2, 1570, 425, 312, 1119, 1596, 165, 1703, 137]. performances [1206]. period [1642]. Periodic [682, 33, 1499, 1860, 1273, 1516, 796, 1279, 509, 293, 1207, 1699, 816, 1007]. periodicity [809]. permanent [843, 1242, 1419]. permeation [1751]. permuted [1652]. persistence [1462]. persons [115]. perspective [1372]. perturbation [262, 1914, 331, 190]. perturbations [768, 842]. perturbed [1287, 193, 1852, 1731, 1792]. pest [1520]. pest-epidemic [1520]. Petri [1053, 1057, 1061, 443, 882, 1051, 15, 730]. petroleum [1191]. Petrov [819, 1801]. Petviashvili [1252, 518, 1859]. Pfaffianization [1270]. pharmaceutical [1773]. pharmacology [451]. Phase [558, 1191, 1137, 42, 1324, 1112, 439, 599, 1808, 854, 603, 1105, 1161, 1162, 1167, 596]. phase-field [1162]. phase/switch/level [1119]. Phenomena [973, 993, 483, 1137, 1369, 1450, 1861, 738, 251, 555, 744]. phenomenon [1129]. phi [1007]. phi-four [1007]. Philippines [672]. photon [552]. photonic [543]. physical [1378]. Physically [96]. physics [1138, 738]. physiology [1603]. phytoplankton [1016]. pi [1695]. PID [13]. piecewise [170]. piezoelectric [487]. PIM [357]. PIMTRACS [357]. Ping [355]. Pinney [1248]. pinning [195]. Pipeline [52, 693, 93]. Pitaevskii [505]. pitch [1204, 1529]. pixel [552]. placement [1128]. planar [1304]. Planck [893]. plane [453, 1244, 804, 977, 1258, 340]. planning [381]. plant [1448, 10, 108]. plantations [967]. plants [566, 419, 255]. plasma [186, 1315, 265, 997]. plasma-Maxwell [1315]. plasticity [1279]. plastics [719]. plates [479]. plume [872, 743]. pneumatic [1109]. Poincare [293, 33]. point [801, 662, 1381, 1702, 166, 795, 170, 1175]. point-forces [801]. point/patch [662]. points [1081, 1481, 1076, 1354, 145, 1326, 635, 1293]. pointwise [1585]. Poisson [953, 668, 1542, 1551, 539, 546, 296]. Pole [1128, 1113]. policies [1051]. policy [1547, 957, 675, 966, 645, 716]. pollutant [1818, 1069]. pollutants [114, 63]. polluted [1406]. Pollution [1826, 757, 1081, 872, 864]. Poly [337]. Poly-quadratic [337]. polycrystalline [156]. polygonal [1528]. polyhedron [331]. polyhedrons [286]. polymer [1886, 464]. polymeric [893]. polymers [1833]. Polynomial [1808, 1867, 1363, 1199, 1805, 1820, 1407, 295, 1445, 1175, 1412]. polynomials [1451]. polyphase [588]. polytopic [1781]. pond [922]. population [1892, 124, 1339, 1574]. populations [1815, 1406, 1066]. poroacoustic [1861]. porous [1676, 1320, 46, 1825, 728, 251, 1158, 1913, 1169]. Port [779]. Port-based [779]. Portfolio [1465, 1737, 1787, 1490]. portfolios [1734]. posed [1196]. posedness [1858]. poseness [1918]. Positive [1499, 1000]. Positive- [1000]. positivity [1163]. positivity-enforcing [1163]. post [1678, 1489, 1698]. post-1978 [1489]. post-discharge [1678]. posteriori [477, 1319]. potential [1697, 1610, 1260, 1546]. potentials [982]. poultry [1640]. powder [720]. Powell [1396]. Power [1106, 894, 1217, 135, 589, 829, 583, 740, 108, 1185, 1684, 1105, 600, 582, 449, 1107, 463, 1620, 16, 1890, 1215, 594, 590]. power-efficient [1620]. powers [1517]. practical [922]. practice [1322]. pre [1152, 1181]. pre-collision [1152]. pre-conditioner [1181]. precalcination [419]. precedents [1770]. precision [1586]. preconditioned [1196]. preconditioners [696, 480]. Preconditioning [432, 739, 1677, 43, 1335]. predator [1444, 1889, 1699, 1779]. predator-prey [1889]. Prediction [715, 277, 1120, 713, 355, 406, 1930, 845, 56, 1666]. Predictive [252, 1478, 675, 1366, 418, 1639]. Preface [1393, 1814, 1747, 1457, 128, 1362, 1301, 1314, 152, 1134, 791, 1604, 352, 335, 458, 1190, 1671, 851, 83, 777]. preference [677, 1437]. preprocessing [230]. presence [1664, 600, 1161]. present [129]. Presentation [597]. preserving [460, 515]. Presidential [1804]. prespecified [1128]. pressure [472]. pressure- [472]. pressurized [1695]. prevalence [1080, 1209, 1219]. prey [1444, 1889, 1699, 1779]. Priaralye [856]. Price [372, 1090, 1464, 1742, 1446]. Price-controlled [1446]. prices [666, 247, 1077, 1419, 944]. pricing [535, 1929, 1519, 1292, 1876]. primary [122]. principal [1214, 2]. principle [344, 122, 1731]. principles [1329]. priori [48, 793, 125]. probabilistic [1850]. probabilities [1585]. probability [125, 620]. probed [1271]. probit [1469, 1475, 645]. problem [217, 1394, 1410, 482, 1316, 1236, 875, 800, 804, 1295, 1846, 388, 1638, 914, 11, 1068, 416, 1090, 329, 1399, 1689, 330, 1531, 359, 415, 32, 442, 1450, 393, 883, 130, 829, 924, 1908, 139, 792, 420, 133, 398, 470, 383, 1099, 1684, 1874, 864, 803, 317, 78, 1856, 1731, 463, 1782, 891, 1203, 1871]. Problems [867, 837, 650, 287, 1317, 1318, 497, 1196, 1319, 1791, 476, 1677, 1572, 155, 617, 138, 1723, 300, 221, 1591, 686, 292, 414, 1329, 331, 802, 1447, 1404, 451, 1680, 488, 740, 1681, 1335, 549, 854, 895, 471, 553, 1183, 1682, 1527, 80, 600, 1373, 178, 557, 1343, 1345, 852, 465, 461, 741, 79, 140]. procedure [1305, 46, 1407, 1530, 1915]. procedures [483, 1068, 123]. process [1678, 1788, 1385, 1832, 1340, 719, 1888, 1759]. processes [1830, 1138, 1397, 1757, 1387, 1201, 842, 111, 913, 230, 712, 620, 783, 906, 55, 713, 553, 246, 530, 464]. processing [1895, 901, 867]. processor [296]. processors [1447, 1700]. product [310]. production [74, 61, 145, 937, 406, 882, 1058, 256]. productive [1340]. productivity [1476, 1547, 1758, 367, 1473]. products [312, 110, 1437, 1703, 830]. Professor [498]. profile [607, 1377, 748, 1916]. profiles [818, 1186, 1916]. profits [1462]. program [29]. programmable [1618]. programming [430, 1694, 1185]. programs [430, 1717]. progressive [1789]. progressively [1703, 1206]. projection [1394, 485, 1241, 695]. Projections [177, 1184]. Projective [1420, 1554]. projects [1432, 576, 444]. promises [1784]. promoting [645]. Proof [536]. Propagation [1807, 198, 752, 536, 749, 1269, 985, 731, 1925, 998, 211]. propellers [988]. Properties [943, 1316, 1140, 1141, 1022, 842, 1664, 207, 1875, 431]. proportional [1475, 1767]. pros [89]. Proteomic [1847]. provide [1447]. province [1803]. proving [1718]. PS [255]. Pseudo [1385, 170, 167, 1736, 1808]. Pseudo-bond [1385]. pseudo-random [167, 1808]. Pseudorandom [181, 1643, 548, 1652]. Pseudospectral [1710, 507]. PSI [1465]. PSPC [296]. Public [1547]. Pula [36]. pulsatile [745]. pulsating [1179]. pulse [709, 584, 265, 998, 1545]. pulses [261]. pump [105]. purification [1571]. PWM [599].

QMC [893]. QoS [1446]. quadrangles [1505]. quadrangular [1304]. Quadratic [1822, 1086, 1087, 440, 261, 1696, 337, 277, 1400, 321, 77, 1731]. quadrilateral [1344]. Qualitative [842, 1316]. quality [902, 425, 869, 711, 537, 729, 600, 1652]. quanta [361]. Quantification [752, 1805]. quantifier [836]. quantifiers [835]. quantile [1785, 1042]. quantity [729]. Quantum [551, 744, 1588, 534, 782, 396, 526, 159, 1910, 1605, 767, 1606, 741]. quantum-kinetic [396]. quantum-waveguide [1588]. quartic [440, 1395, 277]. Quasi [527, 163, 55, 1874, 1293, 1205, 892, 1850, 1505, 1610, 1824, 1400, 1822, 1678, 505, 1306, 1402, 1112, 166, 554, 1600, 1727, 178, 557, 121, 529, 1845]. quasi- [557]. Quasi-hydrodynamic [55]. quasi-interpolants [1400, 1822, 1402, 1727]. quasi-interpolation [1824]. quasi-interpolatory [1306]. quasi-isometric [1505]. quasi-local [505]. Quasi-Monte [892, 166, 178, 121, 529]. quasi-particle [1850]. quasi-particles [1845]. Quasi-random [1293, 1205, 554]. Quasi-randomized [163]. Quasi-regression [527]. Quasi-reversibility [1874]. quasi-six-phase [1112]. quasi-steady [1678]. quasi-steady-state [1600]. quasilinear [1317]. quasiperiodic [509, 54]. Quasiperiodicity [1215]. Queen [383]. Queensland [959]. quintic [1262, 975, 1258, 1259, 1848, 1003]. quote [1708].

R [1235, 1257]. R&D [273]. Radau [142]. radial [1573, 311]. radiances [1186]. Radiation [868, 120, 1267, 1609]. radiative [248]. radioactive [547]. radionuclides [179]. rail [358]. rail/road [358]. Railway [26, 327, 1719]. rainfall [1749, 962]. rainfall-runoff [1749, 962]. Random [1696, 902, 218, 1478, 1381, 1321, 1638, 1273, 61, 145, 842, 158, 880, 900, 165, 545, 167, 1808, 170, 554, 670, 1867, 1293, 559, 1205, 392]. random-walk [158]. randomization [179]. randomized [163]. randomly [342]. Randomness [1652, 890, 559]. range [476, 1483, 1741, 1912]. range-based [1741]. ranges [218, 1631]. rank [1765, 162, 1629]. rank-size [1765]. rapidly [565]. rarefied [1168]. Rasetti [508]. rate [1709, 1757, 843, 1076, 1142, 1929, 1876, 1770, 1558]. rates [933, 706, 1783, 941, 45]. ratings [935, 945]. rational [831]. rationality [319]. rationalized [236]. ray [731, 818, 552]. Rayleigh [1401]. RBF [382]. RBFNN [565]. reachability [1364]. reaction [75, 432, 1726, 514, 554, 1599, 1585, 681, 1651]. reaction/groundwater [75]. reactions [1600]. Reactive [1637, 894, 1017]. reactor [303, 413, 1160]. reactors [70, 431]. Real [585, 587, 63, 72, 73, 1076, 584, 619, 1597, 583, 831, 410, 1060, 1156, 582, 1412]. real-coded [1156]. Real-time [585, 587, 72, 584, 619, 1597, 583, 1060, 582]. Realistic [954]. reality [1786]. realization [103]. realizations [1857]. Realized [1760]. reboot [1707]. receding [1639]. receding-horizon [1639]. recharge [354]. recirculation [70]. recognition [1440, 1438, 1431, 1681, 257, 1729, 1869, 115]. Reconfigurable [417]. reconfiguration [597]. reconstructing [1871]. Reconstruction [796, 1186, 1701, 1830, 758, 865]. recovery [800]. rectifier [1119, 596]. rectifiers [599]. Recurrent [265, 48, 1516, 5, 1487]. Recursive [1589, 410, 1407, 1717, 1834, 392]. recursively [835]. red [220]. red-and-green [220]. reduced [103, 558]. Reducing [1528, 760]. Reduction [595, 1118, 912, 1146, 1546, 94, 1129, 559]. reductions [1142, 1615]. redundancy [549]. redundant [1110, 1545]. Refinable [1306, 1727]. Refinement [1310, 485, 473, 1728]. refinement/alignment [473]. reform [367]. reforming [25]. refrigerator [238]. Regge [508]. regime [986, 1473]. region [1078, 1531, 935, 873]. regional [631, 645, 873]. regional-scale [631]. regions [1754, 1721, 1128, 1183, 1916]. regression [953, 1708, 1562, 527, 1765, 683, 1694, 635, 1768]. regularization [1295, 460, 271, 1657]. regularized [976, 753, 1152, 1026]. regulators [1847]. regulatory [1330, 1598]. reinforced [383]. Reinforcement [4]. Related [946, 911, 1022, 978, 794]. relation [1146, 385]. relational [14]. relations [1552, 1251]. relationship [672, 1559]. relationships [426]. relaxation [1398]. relay [339]. release [1568, 63, 547, 361]. reliability [1114, 1206, 176]. Reliable [883, 1070, 1653]. reluctance [1104, 1117]. Remarks [984, 1367, 1861]. remodeling [1180]. remote [865]. renewable [615]. repairable [1707]. replacement [1322]. replication [880]. replicator [363, 1752, 1833]. Reply [1933]. Report [36, 560]. reported [636]. repository [1318]. representation [1450]. representative [362]. reputation [1466]. rescaled [1034]. research [1473]. reservoir [869]. reservoirs [1191, 57]. residual [1496]. residual-based [1496]. resist [904]. resistance [876]. resolution [875, 1421, 1692]. resolvent [155]. resolver [1118]. Resolving [1476]. Resonance [440, 772, 497, 291, 1263, 1241]. resonances [504, 568]. resonant [261, 344]. Resource [74, 576, 1458, 615, 100]. resources [861, 1541, 629]. respect [729]. response [628, 711, 1444, 221, 1693, 1292, 552, 794]. responses [906]. rest [1601]. result [1408]. resultant [1422]. resultants [587]. resulting [1410]. results [786, 1840, 265, 1115, 637, 1918]. retail [372]. Retaining [84]. retrieval [532]. return [1461, 1552, 1462, 1742]. returns [1555]. reversibility [1874]. reversible [1844]. Review [599, 322, 86, 1577]. revisited [613, 1248, 23]. Revisiting [1521]. revolving [874]. Reynolds [1159]. RF [1610]. rheology [398, 470]. rich [953]. Richtmyer [746, 747, 1576]. Riemann [1638]. Riemmanian [1505]. right [1199, 1703]. right-hand [1199]. rigid [1508, 340]. rigorous [999]. ring [1367, 1099]. rise [255]. risk [1733, 1754, 1755, 1737, 1803, 371, 1693, 935, 1640, 1227, 945, 1734, 1742, 1460]. risks [1590]. River [1294, 35, 1081, 660]. rivers [1347]. RKHPU [1337]. RKL [1932, 1933, 1923]. RLS [901]. RNA [1833]. RNA-like [1833]. road [1681, 358, 357]. roads [1748, 1831]. Robin [800]. robot [51, 122]. robotic [343]. robotics [1587]. robots [24, 9, 1110, 381]. Robust [566, 768, 303, 1639, 381, 1304, 1447, 386, 1024, 409, 1228, 1663]. Robustness [10, 219]. rocket [870]. rods [1154]. role [1562]. roll [1204, 1529]. rolling [845, 1385]. roof [1434]. root [1387, 1079, 1655, 1664, 1419, 1445]. roots [1286, 1412]. Rosenau [1342]. ROSGRO [715]. Rossby [191]. rotating [47, 1271]. rotational [741]. rotationally [1729]. Rotations [1309]. rotor [279, 340]. rotor-bearing [279]. rough [412]. roughness [904, 449]. routes [327]. ruin [531, 119]. rule [1765, 7, 557]. rule-based [7]. rules [1795, 162, 166]. run [1459]. Runge [1302, 1423, 309, 274, 515, 1564]. runoff [1749, 962]. runs [757].

S [814, 1878, 695, 386, 1550]. S&P [376]. S.I. [283, 226, 185, 404]. SABI’01 [704]. SABI’99 [245]. Sabin [1396]. saddle [958]. saddle-path [958]. safeguard [168]. safety [853]. salient [1113]. salinity [661, 965]. sample [1387, 1496]. sampled [761]. samples [1206]. Sampling [1642, 1081, 1698, 559, 633]. Sandalwood [1431]. SAR [860]. SARS [1277]. satellite [867]. satisfaction [1436]. saturable [262]. saturated [1115, 595]. saturating [263]. saturation [1783]. Scalability [475]. Scale [495, 650, 1315, 1303, 628, 1877, 1609, 629, 631, 589, 959, 486, 232, 654, 461, 817, 1692]. scales [1241]. scattered [796]. scattering [1860, 1218, 59, 1904, 795, 1271, 1906, 1186, 1732]. scenario [974]. Schauder [1408]. Scheduling [1061, 416, 359, 415, 576, 414, 783, 1642, 1373, 403, 384]. schematic [1080, 1209, 1219]. scheme [1078, 1250, 1195, 1194, 1355, 536, 316, 1679, 1870, 695, 1261, 1621, 206, 514, 995, 1247, 997, 553, 1157, 923, 1690, 1002, 621, 1566, 1168, 1167, 117, 1545]. schemes [395, 309, 1218, 750, 884, 12, 681, 1163]. Schrödinger [1273, 264, 1261, 1498, 196, 229, 981, 1265, 202, 884, 1002, 1920, 816, 1566]. Schwarz [1318, 1116]. Science [825, 134, 130, 780]. Scientific [1720, 1504, 296, 574]. scrambled [538]. screen [846]. screen-based [846]. screening [783]. screens [59]. screw [462, 1349]. screw-type [1349]. SDE [173]. SE [1328, 1511]. sea [1922, 504, 863]. sealing [462]. search [1752, 902, 504, 900, 549]. seasonal [860]. seasoned [1459]. seawater [1411]. Second [274, 659, 951, 1400, 1326, 1726, 997, 1875]. second-order [1726, 997]. secondary [122, 857]. section [108]. sector [960]. sectoral [674]. securities [1292]. sediment [1748, 1483, 959, 1522]. SEIR [1783, 1558, 320]. Sekerka [1908]. Selected [627, 659, 951, 929]. Selecting [650]. Selection [1729, 1469, 370, 770, 717, 900, 1787, 166, 632, 961, 1042, 1832]. Self [341, 1879, 1782, 772, 1243, 751, 1269, 204, 383, 1310, 6, 1617, 713]. self-consistent [1243]. self-drifting [751]. self-excitations [772]. self-feedbacks [383]. self-frequency [1269]. Self-organizing [1879, 1782, 713]. self-similar [1310]. self-similarity [204]. Self-synchronization [341]. self-tuning [6, 1617]. Semantic [1486]. Semi [1194, 1193, 914, 727, 393, 1328, 161, 303, 1842, 398, 1270]. semi-analytical [727]. semi-batch [303]. semi-classical [161]. semi-coercive [914, 393, 398]. semi-discrete [1270]. semi-implicit [1193, 1194, 1328]. Semi-Lagrangian [1194]. semi-linear [1842]. semiconductor [217, 1609, 983, 161, 546, 171, 176, 556]. semiconductors [1324, 159, 55]. Semidiscretization [92]. semilinear [1919]. Semipalatinsk [873]. sensibility [1434]. sensing [865]. sensitive [1411]. sensitivities [1598]. Sensitivity [1090, 729, 469, 179, 771, 752, 1750, 180, 1780]. sensorless [1078]. sensory [1437, 1429, 1430]. separate [696, 675]. separated [793, 1125]. separation [42, 111]. September [82, 1021, 1030]. sequence [1014]. sequences [170, 177, 178, 1652]. Sequential [147]. serial [1768]. Series [1112, 1481, 671, 933, 1461, 1660, 919, 1438, 1930, 1736, 635, 931, 864, 56, 296, 1561, 1386, 1666, 1382]. services [1597, 1595]. Set [51, 733, 300, 412, 1474, 794, 276, 219]. set-handling [794]. set-theoretic [276, 219]. sets [166, 835, 54]. setting [60]. setup [1309]. several [395, 59, 444]. severe [596]. Sextic [1404]. Seychelles [1490]. Shafer [1888]. shaft [279]. shallow [810, 1195, 1069, 1809, 1336, 1823, 198, 1895, 1086, 1087, 235, 1891, 1487, 79]. shallow-water [1895]. Shape [217, 1192, 1347, 1197, 732, 1654, 1724, 1334, 750, 1894, 1434]. shapes [1846, 1863, 41]. share [1742]. shared [278]. shares [1460]. shear [191, 1843, 1142, 472, 1148]. shear-dependent [472]. sheet [1851]. sheets [464]. Shell [575]. Shift [760, 1269, 573, 60]. Shifted [919]. Shifted-Chebyshev [919]. Shifting [1369]. shifts [1664, 674, 1862]. ship [1204, 20, 1529]. shock [205, 206, 1545, 1315]. shocks [375, 1419, 1555, 1576]. shop [414, 1447, 1373]. short [1240, 1929, 1611, 1876, 8, 1099, 1912, 16, 384]. short-channel [1611]. short-circuit [1099]. short-term [8, 16, 384]. shortage [619]. shot [1894]. Showing [1800]. shrimp [922]. shrinkage [527]. shunt [607, 1105, 1107]. shunting [1499]. SiC [172]. sided [974]. sides [1199]. sideways [1657]. sigmoid [707]. signal [336, 460]. signalized [921]. signals [1869]. signature [797]. significance [683, 1208]. significant [1482]. silicon [1607]. sill [1916]. similar [1310]. similarity [204]. Simple [693, 536, 782, 495, 1330, 318, 1789, 310, 19]. simplex [1752, 382]. Simplification [28, 394]. Simplified [588, 1872]. simplifying [295]. Simul [1929, 1209, 1219]. simulate [70]. simulated [549, 913]. Simulating [1153, 1406, 1450, 1867, 1159]. Simulation [1932, 531, 1096, 1931, 1383, 263, 1593, 706, 581, 1336, 442, 377, 920, 1895, 439, 731, 818, 627, 659, 951, 361, 1584, 340, 449, 1933, 825, 1587, 217, 97, 1570, 1103, 1136, 1193, 872, 1315, 1040, 1568, 1665, 1138, 723, 534, 535, 87, 1381, 536, 1104, 191, 653, 585, 312, 1676, 1142, 1809, 1095, 74, 272, 1016, 102, 156, 584, 1144, 808, 1821, 1609, 1644, 588, 641, 1761, 1286, 1449, 145, 576, 25, 81, 977, 1595, 1384, 18, 630, 786, 495, 1147, 1148, 264, 1149, 540, 1596, 1098, 111, 780, 378, 230, 880, 541]. simulation [105, 47, 1447, 1208, 845, 443, 513, 169, 547, 52, 1154, 929, 55, 444, 1829, 86, 471, 172, 552, 553, 174, 175, 1042, 1284, 555, 278, 1504, 1706, 1155, 586, 358, 838, 1833, 343, 556, 1158, 1888, 923, 1606, 1730, 1017, 1026, 574, 1912, 241, 1161, 1162, 637, 1446, 71, 819, 660, 79, 1566, 464, 1168, 1169, 891, 1170, 730, 1576, 387, 930, 952, 1401, 1105]. simulation-based [1042]. simulation-optimization [891]. simulations [153, 1588, 1673, 1553, 1303, 504, 57, 395, 1610, 1551, 745, 472, 1150, 1508, 1151, 1416, 550, 1698, 654, 121, 1160, 1511]. simulator [583, 246, 846, 582, 357]. Simulink [586]. Simultaneous [136, 803, 1891, 361]. Simultaneously [1076]. sine [1355, 1573, 923, 1125, 203]. sine-Gordon [1355, 1573, 923, 1125].

Singapore [932, 366]. single [733, 1465, 1097, 920, 982, 125, 552, 1105]. single-component [982]. single-phase [1105]. single-species [920]. Singular [1624, 827, 1608, 1632, 40, 331]. singular-value [827]. Singularity [1851, 460]. singularity-preserving [460]. singularly [193, 1731, 1792]. sinh [1125]. sinh-Gordon [1125]. sinusoidal [595]. site [1260, 873]. situations [442]. Sivashinsky [1673]. six [1112]. size [1765]. sizes [1517]. skew [1626]. skew-symmetric [1677, 1626]. skin [78]. skinny [478]. slick [1336]. Sliding [1881, 1781]. slip [1137, 1922]. slopes [977]. slow [986]. slow-envelope [986]. slowly [198]. sludge [246]. Small [945, 1496, 476, 1853, 1236]. SMES [602]. Smoluchowski [452]. smooth [1633, 1732, 54, 633]. Smoothed [1332]. Smoothing [50, 1396, 1665, 1305, 786]. Snake [1848]. snow [780]. Sobolev [1873, 463]. Social [30]. socioeconomic [1458]. soft [793]. Software [825, 1101, 129, 453, 1432, 1108, 296, 232, 1620, 1371]. soil [723]. solenoidal [212]. solid [405, 104, 1170]. solids [1001, 528]. Solitary [1844, 506, 1006, 1932, 1250, 1846, 1860, 193, 977, 1807, 203, 207, 1855, 1912, 1923, 1933, 239, 267, 1863]. solitary-wave [1846, 239]. Soliton [1267, 1914, 1920, 520, 1315, 1252, 262, 504, 1539, 1907, 983, 1269, 1910, 997, 516, 1003, 1004, 518]. soliton-induced [504]. Solitons [987, 1262, 186, 263, 1237, 1917, 194, 1263, 985, 986, 1902, 1848, 996, 210, 266, 1000, 1001, 923, 517, 1007, 267]. Solution [902, 691, 398, 236, 1752, 1193, 1315, 1250, 1319, 804, 1638, 1572, 1573, 1801, 1144, 1723, 1399, 1689, 330, 1531, 1763, 393, 883, 769, 1498, 739, 139, 1404, 903, 1296, 470, 1699, 173, 1710, 1183, 1602, 1338, 852, 465, 560, 1006, 117, 304, 1574]. solutions [1697, 1752, 1932, 810, 682, 1252, 478, 1499, 189, 684, 1846, 1696, 193, 1539, 1852, 1516, 753, 978, 1857, 1800, 919, 472, 829, 981, 433, 509, 651, 1041, 293, 1447, 1207, 512, 1258, 994, 206, 1924, 294, 1025, 785, 1919, 1003, 1923, 1933, 239, 1007, 1125, 1008, 1859]. solve [1791, 11, 1874, 107]. solved [1345]. solvent [1886]. solver [1823, 1325, 1327, 488]. solvers [227, 132, 722]. Solving [837, 1608, 362, 1763, 1828, 891, 75, 1068, 65, 507, 396, 958, 1450, 130, 686, 331, 385, 133, 1530, 557, 1482, 621, 1657]. Some [911, 1022, 275, 1367, 1861, 1664, 944, 1051, 1918, 696, 474, 114, 1273, 651, 344, 1517, 1700, 854, 471, 852, 637, 461, 1387, 1908]. soot [1169]. sorption [169, 720]. sound [798, 861, 793, 740, 748]. sound-hard [798, 793]. sound-soft [793]. source [227, 1321, 123, 330, 1826, 795, 803, 1915, 1918, 1871]. sources [1494, 1243, 247, 959]. South [1464]. southeast [375, 1748]. sowing [249]. Space [480, 853, 1725, 862, 110, 832, 1925, 276, 219, 343, 1834, 431, 852, 1113]. Space-filling [1725]. spaces [684, 463]. spacings [165]. SPAI [1181]. Spain [1754]. span [1930, 1666]. Spanish [1719]. Spatial [114, 1472, 1662, 1090, 924, 1265, 1829, 266, 54, 866, 1308]. spatially [188, 247, 266, 803, 452, 1004]. spatially-modulated [266]. Spatio [212, 654, 1646]. Spatio-temporal [212, 654, 1646]. Spatiotemporal [348, 188]. Special [245, 704, 627, 659, 825, 929, 951, 1715, 1429]. species [1193, 1676, 920]. specific [70]. specification [1763, 1717]. specifications [1872]. Spectra [1317, 189, 1268, 1623, 868]. Spectral [321, 1657, 976, 1033, 1801, 449]. SpectrUW [1268]. Speculation [964]. speech [901]. speed [1103, 1078, 1118, 586]. speeding [30]. sphere [1086, 1087, 209]. Spheres [618, 575]. Spherical [746, 1823, 864]. spherically [1295]. Spike [195]. spillover [956]. spin [742]. spline [1395, 1400, 1822, 1404, 1086, 1087, 1731]. splines [1403, 1827, 1685, 1409, 1175]. split [75, 1873, 507, 884, 1684, 923, 1017, 1378]. split-operator [75]. split-step [507, 884]. splitting [1195, 844, 1398]. spray [709, 720]. spread [1449]. spreading [692, 1788, 713]. spreads [665, 1757]. spurious [1569]. square [1321, 1655, 1116, 1170]. square-root [1655]. squares [1197, 1380, 1873, 1408, 1183, 1574]. squirrel [601]. Stability [482, 1097, 1244, 814, 300, 309, 1225, 1880, 207, 209, 219, 15, 1564, 1529, 1497, 287, 1370, 1363, 768, 1645, 1878, 337, 1778, 1539, 1516, 290, 695, 1031, 489, 1285, 386, 397, 1855, 1699, 45, 276, 100, 1887, 1228, 515, 809, 1648, 1163, 1659, 1890, 1658, 1229, 1651, 1129]. stabilization [1111]. stabilized [1319, 996]. stabilizer [566]. Stable [1242, 1269, 681, 800, 1563]. Stablewise [1364]. stage [359, 397, 1779]. stage-structured [1779]. stages [1890]. standard [483, 1649, 22]. star [1070]. State [1052, 40, 1497, 253, 832, 1678, 1591, 913, 1548, 1699, 1600, 276, 219, 784, 1164]. state-dependent [1699]. state-feedback [1497]. state-space [832, 276]. stated [677]. states [196]. static [1610, 654]. station [1103]. Stationary [1794, 614, 550]. statistic [1034]. Statistical [1638, 196, 378, 683, 1135, 1892, 615, 1401, 1208, 639, 944, 636]. statistics [1815, 1616]. stator [601]. stay [1767]. steady [1069, 1678, 1600, 1249]. steam [723]. steel [1385]. steers [1750]. Stefan [1540, 470]. Stefan-like [470]. stent [1146]. stented [1142, 1146]. step [536, 507, 1789, 906, 884, 667, 1159]. step-stress [1789]. steppe [858]. stepping [819]. steps [1412]. stepsize [1530]. stiff [1805, 769, 301, 685]. stiffness [495]. Stochastic [771, 452, 530, 650, 902, 347, 1465, 1738, 1549, 653, 1479, 1423, 576, 842, 1401, 1857, 1448, 1739, 126, 1787, 146, 1517, 1214, 1828, 273, 444, 1339, 1833, 1868, 1157, 813, 1887, 1544, 1656, 1659, 649]. stock [666, 1735, 665, 375, 1743, 1762, 1559, 1774, 1467, 1077, 1419, 944, 1888, 938, 1804, 1555]. stocks [1760, 1464]. Stokes [801, 482, 1316, 1677, 65]. stops [395]. storage [869, 593]. Stosszahlansatz [1135]. STR [1357]. straight [430, 29, 1532, 731, 1528]. straight-line [430, 29]. straining [212]. Strait [1821]. strategies [1126, 1485, 469, 1779, 922]. strategy [1641, 1896, 1365]. stratification [141]. stream [25, 399]. stream-tube [399]. streets [1267]. stress [1789, 1607]. strict [890]. string [1601]. strip [1528]. strip-wise [1528]. Strong [126, 1002, 344, 515]. strong-stability-preserving [515]. strontium [1751]. Structural [1622, 1418, 1743, 1079, 842, 1279, 635, 1419, 1769, 673]. structure [565, 29, 116, 1878, 18, 644, 1596, 1904, 895, 1584, 1607, 1169]. structure-adaptation [565]. Structured [1216, 328, 768, 110, 1080, 1209, 1219, 219, 1779]. structures [871, 592, 30, 1533, 176, 567, 519, 744, 1278, 1353]. studies [671, 880, 996, 551, 360]. Study [1540, 1058, 1434, 677, 1370, 1240, 1379, 1141, 359, 415, 1145, 1853, 81, 540, 1119, 1448, 924, 1517, 1015, 599, 1751, 994, 1115, 904, 1605, 1105, 327, 582, 1488, 79, 239, 567]. stunting [371]. Sturm [442]. style [1718]. sub [1483, 959]. sub-catchment [959]. sub-catchments [1483]. subdivision [694]. submicron [1613]. subregion [1128]. Subsampling [1702]. subsets [761]. subsonic [749]. subspace [1628, 1615]. subspaces [760, 489]. Successes [1458]. sugar [406]. Sugeno [919]. suitability [1771]. suitable [1763, 1700]. Sum [548]. Sum-discrepancy [548]. Sun [757]. sunflower [249]. superconducting [602, 1111]. supercritical [1841, 1840]. superfluids [1902]. superlattice [744]. Superlinear [1398]. Superposition [1891]. supersonic [192]. supervised [713]. supervision [1057]. supervisory [1060, 1051]. supply [1592, 667]. supply-chain [1592]. support [63, 1694, 1488, 239, 1617]. supportability [74]. supported [479]. Supporting [411, 27]. suppression [1546]. surface [1851, 860, 1843, 1200, 628, 1809, 1610, 1679, 904, 865, 1916, 1802, 387]. surface-potential-based [1610]. surfaces [533, 1396, 1835, 745, 1150, 987, 1409, 904]. surges [863]. surrounding [936]. survey [483]. survival [1478, 45]. suspension [423]. suspensions [1140]. sustainability [100]. sustainable [629, 859, 852]. sustained [202]. SVD [827, 1623]. SVL [1099]. SVM [597]. SVM-DTC [597]. SW [860]. Swan [35]. swap [1757]. swarm [894]. swine [257]. switch [1119]. Switched [730, 1104, 1880, 1117]. Switching [1354, 1097, 669, 1660, 1466, 943, 649]. Sydney [114]. Symbolic [285, 1921, 290, 230, 1040, 29, 832, 294, 4]. Symbolically [90, 24]. symmetric [1262, 1295, 1022, 1677, 1626, 198, 1602]. symmetric/skew [1677]. symmetric/skew-symmetric [1677]. symmetries [285]. symmetry [190]. symmetry-breaking [190]. Symplectic [229, 203, 1840, 1842]. Symplectification [284]. Synchronization [347, 319, 336, 348, 341, 1287, 1089, 337, 1554, 1420, 1653, 66, 895, 346, 1538, 812, 148]. synchronizing [338]. synchronous [595, 1113]. syneresis [721]. Synthesis [1888, 160, 743, 1717, 1050, 256]. synthetic [712]. System [590, 1446, 772, 650, 1571, 801, 733, 33, 1316, 1287, 1109, 316, 768, 1793, 1032, 1722, 1778, 74, 277, 568, 1398, 1055, 1057, 1541, 1294, 370, 1852, 770, 63, 506, 1225, 279, 1450, 1707, 1554, 619, 292, 787, 1438, 920, 604, 405, 859, 1098, 1546, 1431, 984, 1507, 72, 583, 593, 1930, 144, 386, 27, 207, 170, 257, 89, 413, 326, 343, 70, 423, 1650, 855, 697, 357, 1368, 1356, 1903, 15, 1348, 660, 1566, 1666, 1564, 148, 384, 339]. Systems [581, 289, 631, 440, 837, 97, 1262, 284, 565, 1497, 696, 811, 1370, 1302, 1250, 592, 1363, 1089, 779, 188, 814, 1484, 1805, 612, 1844, 1615, 1143, 337, 1622, 275, 1095, 1623, 1016, 782, 1539, 1371, 832, 415, 145, 1364, 589, 290, 661, 1637, 1800, 769, 1052, 40, 301, 685, 1420, 1807, 394, 833, 1059, 229, 509, 1653, 1633, 1895, 418, 163, 1285, 1880, 1781, 443, 882, 420, 1694, 691, 487, 294, 1925, 89, 1060, 295, 1024, 1642, 1602, 26, 1538, 12, 338, 1050, 1560, 96, 1584, 232, 1888, 616, 297, 582, 14, 852, 1648, 408]. systems [1107, 1058, 54, 817, 1663, 85, 16, 834, 1550, 594, 1096, 322].

T [814, 1878, 695, 386, 603, 1550].

T-connected [603]. Tadmor [997]. Taha [185, 1235, 1257]. tailed [1702, 1386]. tails [1846, 664]. Taiwan [860, 1547, 1758, 843, 1640, 1418]. Takagi [919]. taking [861, 110]. tale [1477]. talk [759, 109]. tall [478]. Taming [137]. tangency [451]. tangential [192]. tape [1127]. tapered [1193, 1328]. target [154, 797, 1790]. Tarsier [663]. task [122]. tau [441]. taxonomy [630]. TCP [1617]. technical [483]. technique [733, 138, 541, 1616, 1786, 549]. Techniques [178, 638, 98, 869, 1484, 1506, 1723, 1433, 406, 739, 880, 599, 52, 1717, 1718]. technogenic [855]. technological [366]. tell [1771]. Temperature [845, 718, 1532, 542, 707, 1385]. temperature-dependent [1532]. Temporal [1461, 1662, 963, 212, 654, 1378, 1646]. tension [1851]. term [1321, 330, 1873, 1787, 8, 803, 1772, 16, 1871, 384]. terminals [358, 838]. terms [227, 1918, 1651]. ternary [42]. terrains [381]. terrestrial [865]. Test [670, 873, 716, 638, 890, 1217, 1033, 1015, 1786, 167, 1789, 1664, 548, 794, 1496]. Testing [1557, 1067, 1739, 1357, 941, 1474, 1768, 1759, 1742, 1815, 1387, 1765, 1515].

Tests [181, 1702, 165, 1517, 549, 1419, 669]. textile [425]. texture [1440]. th [1801]. th-order [1801]. Thailand [1662, 1458, 672, 661, 1483, 355]. Theil [370]. their [480, 1484, 57, 1624, 69, 327, 180, 1293, 1780]. thematic [867]. theorem [1135, 1632, 641, 1325]. theoretic [276, 219]. Theoretical [1378, 1266, 1840, 1115]. theory [1236, 587, 1844, 1914, 159, 1070, 293, 412, 513, 251, 551, 1657, 973, 993, 1888]. therapy [1080, 1209]. there [1790]. Thermal [240, 1114, 1549, 377, 1532, 10, 103, 1155, 873]. thermistor [1410]. thermo [393, 883, 55, 398]. thermo-elastic [398]. thermo-elasticity [393, 883]. thermo-electrical [55]. Thermodynamical [542]. thermodynamically [1164]. thermodynamics [25]. thermoelastic [388]. thermoelasticity [750]. thermoforming [464]. Thiab [1235, 1257]. Third [973, 993, 1355]. those [483]. Three [1191, 1377, 1239, 1606, 1295, 1242, 450, 1176, 1260, 786, 1119, 439, 599, 603, 255, 596, 311]. Three-dimensional [1239, 1606, 1295, 1242, 450, 786, 311]. Three-factor [1377]. Three-phase [1191, 1119, 439, 599, 603, 596]. three-phase/switch/level [1119]. threshold [933, 1217, 1032, 1740, 1559, 1077, 210, 1544, 1555, 1496]. Thresholds [1907]. through-process [1385]. throughput [783]. thulium [1751]. Tikhonov [1295, 271]. Time [783, 998, 56, 819, 363, 566, 565, 1628, 638, 98, 752, 1195, 1089, 536, 1499, 1481, 585, 671, 933, 1461, 1756, 768, 814, 587, 612, 337, 1090, 584, 1660, 441, 1881, 1601, 63, 1364, 589, 40, 619, 72, 1597, 126, 620, 583, 573, 13, 1285, 1480, 1736, 1880, 420, 635, 119, 1726, 691, 69, 410, 962, 1906, 1060, 1649, 931, 838, 1102, 516, 1050, 813, 1561, 431, 582, 1386, 1228, 1690, 1039, 1382, 1658, 1779, 1896, 1794, 1558, 339, 1053]. time-delay [814]. time-delays [420]. time-dependent [441, 1726, 1649, 1690, 1039]. time-domain [752]. Time-optimal [783]. time-series [1461]. time-shift-detail [573]. time-step [536]. Time-stepping [819]. time-varying [566, 565, 1628, 1499, 1756, 768, 40, 1285, 1880, 691, 962, 1228]. timed [1061, 443]. times [1383]. Timing [145, 1903, 1862]. tires [845, 311]. tissue [495]. tissues [492]. Toda [1852, 1270]. tokens [882]. Tomato [250]. tomographic [758]. tomography [802]. tool [1570, 453, 629, 144, 1309, 1308, 358, 1488, 94]. tooth [1330]. tori [1868]. Tornado [857]. torque [1104, 595]. torus [1215]. total [1547, 1758, 1322, 1600]. tourism [1753, 1662, 1418, 373, 939, 1489, 1766, 945, 1490]. tourist [1754, 963]. tourists [1646]. Townes [996]. Townsville [929, 951]. traces [1825]. tracking [154, 1628, 280, 1364, 1528, 963, 1162, 1039]. tract [494]. traction [424]. trade [1708, 613, 781]. trade-off [781]. trades [932]. trading [1709, 1552, 1888]. traditional [1017]. traffic [220, 1681, 426, 357, 71]. trained [9]. training [48, 1184]. trains [1267]. trajectory [102, 611, 1706]. trampling [1227]. transcription [1684]. transcriptional [1847]. transdermal [1560]. transfer [102, 248, 924, 104, 1649, 1356]. transform [50, 1860, 1015, 60, 1869]. transformation [978, 200, 1528, 595, 1915, 1270]. transformations [288, 1324, 1643]. transformer [603]. transient [552, 1906, 1511]. transistors [1605, 1606]. transition [633]. Transitions [1262, 1137, 558]. transitory [843]. translation [1723]. Transmission [972, 336, 604, 105, 1464, 818, 1460]. transonic [342]. Transparent [1588]. transport [1193, 75, 1818, 534, 1069, 1821, 1057, 159, 540, 161, 1127, 1179, 169, 251, 171, 172, 1605, 358, 1017, 356, 1365, 304]. transportation [343]. transverse [210]. trapezoidal [728]. trapped [1267]. trapping [1802]. travel [676]. Traveling [1924, 1259, 209]. Traveling-wave [1924]. Travelling [651, 512, 1236, 1125]. Travelling-wave [512]. traverse [421]. treatment [119, 1023, 528]. treatments [1475]. tree [967]. Trend [120, 1401]. trending [931]. triangle [1310, 1728]. triangular [1304, 830, 1510, 1154, 1451, 1344]. triangular/quadrilateral [1344]. tridiagonal [1700]. trio [1477]. Trotter [1157]. truck [168]. trucks [30]. truncated [284, 1542]. truncation [1874, 1378]. trunsored [1277]. Trust [1372]. TSK [1370]. Tsunami [1922, 1925]. tube [399]. tubular [1685]. tumor [1697, 1452, 1931]. tuning [6, 1617]. tunnel [248]. turbochargers [19]. turbulent [1821, 1644, 146]. Turing [1841]. turning [1076, 1354]. TVD [1726]. twisters [857]. Two [1315, 568, 61, 589, 461, 772, 875, 1355, 228, 1573, 1624, 249, 1483, 958, 472, 619, 394, 1889, 1654, 1112, 197, 72, 1787, 35, 543, 831, 854, 1925, 397, 667, 486, 399, 211, 663, 923, 340, 1161, 1162, 1487, 1167, 117, 1412]. two-asset [1787]. two-component [875]. two-dimensional [1355, 228, 1573, 472, 197, 543, 399, 923]. two-fluid [1315]. two-frequency [772]. Two-level [61]. two-motor [1112]. two-parameter [1624]. two-phase [854, 1161, 1162, 1167]. two-plane [340]. Two-scale [1315, 461, 486]. two-stage [397]. two-step [667]. two-steps [1412]. Two-to-one [568]. two-way [211]. type [1396, 1397, 1824, 1067, 978, 1402, 1807, 1265, 1703, 274, 1475, 1767, 295, 1349, 91, 267, 1859, 1206, 1789]. Type-I [1789]. Type-II [1206]. types [189, 753]. typhoon [1803]. Tzafestas [404].

U.S. [1755, 1076]. UK [666, 1076, 660]. Ulam [1260]. ulmoides [256]. Ultracold [1910]. ultrasound [752]. ultraspherical [1801]. ultraspherical-dual-Petrov [1801]. UML [1055]. unbiased [1768]. uncertain [1806, 1089, 1201, 883, 1285, 1880, 1887]. uncertainties [768, 1639]. uncertainty [1805, 276, 219, 1490]. unconditionally [320]. uncontrolled [275]. uncoupled [1566]. undergoing [1841, 469]. underground [871, 364]. underperformance [1459]. Underwriter [1466]. underwriting [942]. undoped [1612]. undulation [595]. uneven [503]. unfolding [1061]. ungauged [355]. uniaxial [1607]. unidirectional [466]. unified [1614]. uniform [1395, 1396, 1917, 1400, 1150, 1182, 748, 71]. unilateral [388, 914, 883, 60]. unilaterally [479]. union [940, 1128, 1771, 1477]. uniqueness [1236, 466]. unit [1387, 1079, 1286, 1664, 1419]. units [1895]. universe [490]. universities [1471]. unknown [1481, 123, 330, 1869]. unknowns [65, 1690]. unmodeled [1663]. unsealed [1748]. unsteady [1069, 1572, 102, 1328]. Unstructured [718, 1165, 1307, 1509, 1510, 79]. update [182]. upon [1916]. Upper [355]. upscaling [1816]. upstream [592]. upwind [1679]. Upwinding [231]. urinary [494]. USA [642, 366]. usage [1322, 247]. Use [966, 1484, 628, 57, 629, 1485, 355]. used [1697, 611, 1098]. Using [1440, 638, 279, 292, 712, 1447, 901, 1717, 837, 566, 218, 1830, 733, 677, 1040, 1118, 98, 1363, 1674, 669, 1319, 1432, 228, 1109, 312, 354, 869, 587, 1860, 1615, 250, 1921, 1573, 1336, 1723, 1057, 1380, 1644, 1294, 727, 1369, 1120, 694, 539, 1619, 5, 781, 1826, 1522, 422, 1508, 1330, 1895, 1703, 903, 1206, 1559, 1616, 1614, 237, 382, 144, 882, 1828, 1116, 1181, 783, 1154, 1750, 421, 713, 410, 1767, 41, 1099, 1408, 1185, 1684, 729, 1264, 1100, 409, 586, 360, 423, 515, 255, 865, 1411, 408, 368, 15, 633, 1166, 85, 1646, 891, 1496, 1701]. using [64, 866].

vacation [125]. vaccination [1783, 1515, 1558]. vacuum [1589]. Validated [1631]. Validation [1511, 733, 1105, 1113]. valuation [139]. value [827, 677, 804, 1572, 138, 1404, 1042, 1527, 1856, 1460]. value-at-risk [1460]. values [73, 1624, 1117]. valve [1284]. vapor [1143]. VAR [1769, 673]. variability [168, 1382]. Variable [1184, 1103, 105, 1296, 1925, 586, 1862, 1125]. variable-coefficient [1862]. variable-displacement [105]. variable-speed [586]. variables [291, 1548, 85, 891]. Variance [171, 1034, 1519, 761]. variate [545]. variation [163]. Variational [981, 475, 1304, 686, 1403, 1827, 80, 1026]. various [228, 93]. varying [566, 565, 1628, 1499, 1756, 768, 40, 198, 1285, 1880, 691, 962, 1584, 1228]. Vector [1893, 369, 312, 1032, 300, 1827, 1559, 1700, 1694, 1845]. vector-borne [1032]. vectors [1625, 760]. vehicle [1185, 1684]. vehicles [1104, 781]. velocity [692, 1142, 192, 1146, 1911, 520]. ventilation [706]. ventilator [1099]. ventricle [1674]. Verification [93, 1319, 1611, 1717, 1560, 1511]. versus [88, 1521, 179, 1163]. vertical [1186, 1249, 1405]. vertices [1716]. very [985]. vetches [714]. VI [1720]. via [1735, 50, 1835, 1540, 1844, 976, 145, 1643, 40, 293, 1655, 1740, 420, 1808, 1528, 1698, 1639, 236, 346, 923, 407, 3]. vibration [1793, 250]. Vibrations [395, 604]. vicinity [871]. view [170]. viewpoint [34]. virtual [710]. virtually [278]. viscoelastic [479, 1238, 240]. viscosities [472]. viscous [1109, 874, 1511]. visual [421]. visualization [694, 1720]. Visualizing [1598]. Vlasov [1608]. volatile [1825]. volatilities [1867]. Volatility [1764, 634, 1465, 1738, 665, 1479, 642, 843, 1076, 1552, 1557, 1760, 956, 1463, 1739, 1741, 1544, 376, 664, 1468, 1460]. voltage [1114, 607, 1107]. Volterra [682, 1792, 1013]. Volume [270, 307, 351, 436, 563, 624, 701, 764, 822, 887, 949, 1011, 1064, 1123, 1173, 1189, 1302, 1552, 1809, 694, 1332, 1726, 79, 501]. volumetric [1494]. Vortex [1802, 1851, 1854, 189, 1267, 508, 748]. Vries [187, 506, 200, 1356, 1903, 1862, 1647]. vulnerability [1590].

Walk [618, 575, 158, 554]. wall [1179, 988]. walls [982]. wandering [760]. warped [1527]. waste [1318, 405, 547]. wastewater [1571]. water [810, 677, 1818, 1195, 1069, 1244, 860, 869, 1242, 628, 1809, 1336, 46, 1541, 1823, 957, 629, 247, 1245, 198, 1895, 1086, 1087, 235, 1828, 1891, 108, 208, 209, 210, 720, 79, 387]. Watson [1716, 1340]. Wave [973, 993, 1904, 738, 1932, 1240, 752, 974, 191, 1844, 976, 1850, 1846, 193, 1601, 753, 796, 1840, 749, 509, 1272, 651, 1041, 913, 1863, 512, 1258, 206, 1924, 1842, 1849, 1026, 1923, 1933, 741, 79, 239, 1125, 1916, 1918, 267]. wave-structure [1904]. wave/breaking [974]. wave/corner [974]. waveform [1398]. waveguide [972, 1588]. waveguides [1904]. Wavelet [1181, 460, 301, 685, 686, 573, 1694, 60]. Wavelet-based [1181]. wavelets [1830, 40, 903, 511, 1407, 80, 368, 1013]. wavepackets [187]. waves [1250, 1236, 1244, 1847, 503, 1843, 975, 976, 1860, 1242, 506, 977, 1909, 1807, 920, 1245, 198, 203, 1246, 205, 1259, 1828, 207, 1925, 1855, 208, 209, 211, 868, 1912, 1913, 1249, 816, 1006, 387]. way [211]. weak [1423, 274, 1226]. Weakly [187, 193]. weather [712]. Web [1595]. weighing [354]. weight [1322, 142]. weight-bearing [1322]. weighted [533, 961, 463]. Weights [1435]. Well [1858, 1918]. Well-posedness [1858]. well-poseness [1918]. wet [708]. Weyl [559]. WGENK [712]. while [1450]. whose [483]. Wick [1828]. Wick-stochastic [1828]. width [709, 904]. width-modulated [709]. Wiener [1806]. Wigner [534, 159]. Wigner-function [159]. wild [140]. wildland [1565]. wind [863, 593]. wind-induced [863]. window [542]. wire [1793, 1119]. wise [1528]. within [826, 874, 660, 1657, 857]. Wolfgang [322]. wordlength [768]. works [108]. world [84]. worst [1033]. worst-case [1033].

X [731, 818, 552]. X-ray [731, 818, 552]. XII [1703].

Yellow [504]. yen [940, 942]. yield [1483, 1772]. yielding [1259]. Yiwu [1541]. yoghurt [720]. Yosida [1325]. ytterbium [1751].

Zealand [375, 643, 364, 367, 672, 960, 938, 365]. Zero [1542]. Zero-truncated [1542]. zeros [1445]. Zhejiang [1803]. zonal [191]. zone [1420, 339]. zoom [1891]. zooplankton [56].

References

Tzafestas:2000:E

[1] Spyros G. Tzafestas, Pierre Borne, and Elpida S. Tzafestas. Editorial. Mathematics and Computers in Simulation, 51(3–4):143– 144, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001123.

Tzafestas:2000:PED

[2] E. S. Tzafestas, A. Nikolaidou, and S. G. Tzafestas. Performance evaluation and dynamic node generation criteria for ‘principal component analysis’ neural networks. Mathematics and Computers in Simulation, 51(3– 4):145–156, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001135.

Yadaiah:2000:PIN

[3] N. Yadaiah, L. Sivakumar, and B. L. Deekshatulu. Parameter identification via neural networks with fast convergence. Mathematics and Computers in Simulation, 51(3–4):157– 167, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001147.

Vogiatzis:2000:RLS

[4] Dimitrios Vogiatzis and Andreas Stafylopatis. Reinforcement learning for symbolic expression induction. Mathematics and Computers in Simulation, 51(3–4):169– 179, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001159.

Kambhampati:2000:IMC

[5] C. Kambhampati, R. J. Craddock, M. Tham, and K. Warwick. Inverse model control using recurrent networks. Mathematics and Computers in Simulation, 51(3–4):181– 199, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001160.

Potocnik:2000:AST

[6] Primož Potočnik and Igor Grabec. Adaptive self-tuning neurocontrol. Mathematics and Computers in Simulation, 51(3–4):201–207, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475499001172.

Looney:2000:FRB

[7] Carl G. Looney. Fuzzy and rulebased image convolution. Mathematics and Computers in Simulation, 51 (3–4):209–219, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001184.

Mastorocostas:2000:HFM

[8] P. A. Mastorocostas, J. B. Theocharis, S. J. Kiartzis, and A. G. Bakirtzis. A hybrid fuzzy modeling method for short-term load forecasting. Mathematics and Computers in Simulation, 51 (3–4):221–232, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001196.

Izumi:2000:FBB

[9] Kiyotaka Izumi and Keigo Watanabe. Fuzzy behavior-based control trained by module learning to acquire the adaptive behaviors of mobile robots. Mathematics and Computers in Simulation, 51(3–4):233– 243, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001202.

Matko:2000:RFC

[10] D. Matko, I. Škrjanc, and G. Mušič. Robustness of fuzzy control and its application to a thermal plant. Mathematics and Computers in Simulation, 51(3–4):245–255, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475499001214.

DeLit:2000:GGA

[11] P. De Lit, E. Falkenauer, and A. Delchambre. Grouping genetic algorithms: an efficient method to solve the cell formation problem. Mathematics and Computers in Simulation, 51 (3–4):257–271, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001226.

Sandalidis:2000:GIC

[12] Harilaos G. Sandalidis, Peter Stavroulakis, and Joe Rodriguez-Tellez. Genetic inspired channel assignment schemes for cellular systems. Mathematics and Computers in Simulation, 51(3– 4):273–286, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001238.

Lima:2000:NGP

[13] João M. G. Lima and António E. Ruano. Neuro-genetic PID autotuning: time invariant case. Mathematics and Computers in Simulation, 51 (3–4):287–300, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549900124X.

Stamou:2000:NFR

[14] G. B. Stamou and S. G. Tzafestas. Neural fuzzy relational systems with a new learning algorithm. Mathematics and Computers in Simulation, 51 (3–4):301–314, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001263.

Tzafestas:2000:SAA

[15] S. G. Tzafestas and G. G. Rigatos. Stability analysis of an adaptive fuzzy control system using Petri Nets and learning automata. Mathematics and Computers in Simulation, 51(3–4): 315–339, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001275.

Xiangping:2000:HMG

[16] Meng Xiangping, Zhang Huaguang, and Tan Wanyu. A hybrid method of GA and BP for short-term economic dispatch of hydrothermal power systems. Mathematics and Computers in Simulation, 51(3–4):341– 348, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001287.

Hattori:2000:AMI

[17] Motonobu Hattori and Masafumi Hagiwara. Associative memory for intelligent control. Mathematics and Computers in Simulation, 51(3–4): 349–374, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001299.

Hanomolo:2000:MLP

[18] A. Hanomolo, Ph. Bogaerts, J. Graefe, M. Cherlet, J. Wérenne, and R. Hanus. Maximum likelihood parameter estimation of a hybrid neural-classical structure for the simulation of bioprocesses. Mathematics and Computers in Simulation, 51(3–4):375– 385, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001305.

Pantelelis:2000:NNS

[19] Nikos G. Pantelelis, Andreas E. Kanarachos, and Nikos Gotzias. Neural networks and simple models for the fault diagnosis of naval turbochargers. Mathematics and Computers in Simulation, 51(3–4):387– 397, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001317.

Lisowski:2000:NNC

[20] Józef Lisowski, Andrzej Rak, and Wojciech Czechowicz. Neural network classifier for ship domain assessment. Mathematics and Computers in Simulation, 51(3–4):399– 406, January 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001329.

Anonymous:2000:PJa

[21] Anonymous. Pages 143–416 (January 2000). Mathematics and Computers in Simulation, 51(3–4):??, January 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Steinberg:2000:NSA

[22] Stanly Steinberg. Non-standard Applications of Computer Algebra II. Mathematics and Computers in Simulation, 51(5):417–418, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001330.

Laita:2000:BLR

[23] Luis M. Laita, L. de Ledesma, Eugenio Roanes-Lozano, A. Pérez, and A. Brunori. Boole’s logic revisited from computer algebra. Mathematics and Computers in Simulation, 51 (5):419–439, February 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001342.

Gutierrez:2000:DDR

[24] Jaime Gutierrez and Tomas Recio. Defining degenerate robots symbolically. Mathematics and Computers in Simulation, 51(5):441– 449, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001354.

Grotendorst:2000:CAM

[25] J. Grotendorst and J. Dornseiffer. Computer-aided modelling and simulation of the thermodynamics of stream reforming. Mathematics and Computers in Simulation, 51(5):451– 471, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001366.

Roanes-Lozano:2000:RIS

[26] Eugenio Roanes-Lozano, Eugenio Roanes-Macías, and Luis M. Laita. Railway interlocking systems and Gröbner bases. Mathematics and Computers in Simulation, 51(5):473– 481, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001378.

Marlewski:2000:CAE

[27] Adam Marlewski and Magdalena Hajdasz. A CAS aid to the elaboration of the expert system supporting the managing the monolitic construction. Mathematics and Computers in Simulation, 51(5):483– 488, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847549900138X.

Hashiguchi:2000:SLB

[28] Hiroki Hashiguchi, Shigekazu Nakagawa, and Naoto Niki. Simplification of the Laplace–Beltrami operator. Mathematics and Computers in Simulation, 51(5):489– 496, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001391.

Castano:2000:DSS

[29] Bonifacio Castaño, Joos Heintz, Juan Llovet, and Raquel Martínez. On the data structure straight-line program and its implementation in symbolic computation. Mathematics and Computers in Simulation, 51(5):497– 528, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001408.

Faia:2000:SSS

[30] Michael A. Faia. Social structures and speeding trucks. Mathematics and Computers in Simulation, 51(5): 529–545, February 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847549900141X.

Anonymous:2000:PFa

[31] Anonymous. Pages 417–556 (February 2000). Mathematics and Computers in Simulation, 51(5):??, February 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Goto:2000:DOM

[32] Norihiro Goto and Hiroyasu Kawable. Direct optimization methods applied to a nonlinear optimal control problem. Mathematics and Computers in Simulation, 51(6):557– 577, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499001457.

Benbachir:2000:LPM

[33] Saâd Benbachir. Lindstedt–Poincaré method and periodic families of the Barbanis–Contopoulos Hamiltonian system. Mathematics and Computers in Simulation, 51(6):579– 596, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499001470.

Hadjidimos:2000:NDD

[34] A. Hadjidimos, D. Noutsos, and M. Tzoumas. Nonoverlapping domain decomposition:: a linear algebra viewpoint. Mathematics and Computers in Simulation, 51(6):597– 625, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475499001482.

Kurup:2000:CTD

[35] Rajendra G. Kurup, David P. Hamilton, and Robert L. Phillips. Comparison of two 2-dimensional, laterally averaged hydrodynamic model applications to the Swan River Estuary. Mathematics and Computers in Simulation, 51(6):627–638, February 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001469.

Kalpic:2000:RIC

[36] D. Kalpic and V. Hljuz Dabric. Report of the ITI’99 Conference held in Pula, Croatia, 15–18 June 1999. Mathematics and Computers in Simulation, 51 (6):639–640, February 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001494.

Anonymous:2000:Ia

[37] Anonymous. Index. Mathematics and Computers in Simulation, 51(6): 651–654, February 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001452.

Anonymous:2000:PFb

[38] Anonymous. Pages 557–654 (February 2000). Mathematics and Computers in Simulation, 51(6):??, February 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Franksen:2000:CLI

[39] Ole Immanuel Franksen and Peter Falster. Colligation or the logical inference of interconnection. Mathematics and Computers in Simulation, 52(1):1–9, March 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475499001524.

Hsiao:2000:SAT

[40] Chun-Hui Hsiao and Wen-June Wang. State analysis of time-varying singular bilinear systems via Haar wavelets. Mathematics and Computers in Simulation, 52(1):11–20, March 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475499001500.

Normandin:2000:LFC

[41] Magdeleine Normandin, Jean-Robert Clermont, and Ahmad Mahmoud. Laminar flow calculations in ducts of complex shapes using mapping functions and decomposition of domains. Mathematics and Computers in Simulation, 52(1):21–39, March 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475499001512.

Copetti:2000:NEP

[42] M. I. M. Copetti. Numerical experiments of phase separation in ternary mixtures. Mathematics and Computers in Simulation, 52(1):41– 51, March 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475499001536.

Juncu:2000:NEP

[43] Gheorghe Juncu and Constantin Popa. Numerical experiments with preconditioning by Gram matrix approximation for non-linear elliptic equations. Mathematics and Computers in Simulation, 52(1):53–71, March 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001439.

Anonymous:2000:PM

[44] Anonymous. Pages 1–85 (March 2000). Mathematics and Computers in Simulation, 52(1):??, March 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pykh:2000:ESC

[45] Yu. A. Pykh and S. S. Efremova. Equilibrium, stability and chaotic behavior in Leslie matrix models with different density-dependent birth and survival rates. Mathematics and Computers in Simulation, 52(2):87– 112, April 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001440.

Fatullayev:2000:NPI

[46] Afet Fatullayev and Emine Can. Numerical procedure for identification of water capacity of porous media. Mathematics and Computers in Simulation, 52(2):113–120, April 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400001464.

Kuo:2000:DCS

[47] Chung-Feng Jeffrey Kuo and Shu-Chyuarn Lin. Discretization and computer simulation of a rotating Euler–Bernoulli beam. Mathematics and Computers in Simulation, 52 (2):121–135, April 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400001488.

Dimopoulos:2000:ERN

[48] K. P. Dimopoulos, C. Kambhampati, and R. Craddock. Efficient recurrent neural network training incorporating a priori knowledge. Mathematics and Computers in Simulation, 52(2):137–162, April 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000149X.

Anonymous:2000:PAa

[49] Anonymous. Pages 87–174 (April 2000). Mathematics and Computers in Simulation, 52(2):??, April 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Amato:2000:SDC

[50] Umberto Amato and Italia De Feis. Smoothing data with correlated noise via Fourier transform. Mathematics and Computers in Simulation, 52(3– 4):175–196, June 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001476.

Jaulin:2000:SIA

[51] L. Jaulin, E. Walter, O. Lévêque, and D. Meizel. Set inversion for χalgorithms, with application to guaranteed robot localization. Mathematics and Computers in Simulation, 52 (3–4):197–210, June 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001506.

Matko:2000:PST

[52] Drago Matko, Gerhard Geiger, and Withold Gregoritza. Pipeline simulation techniques. Mathematics and Computers in Simulation, 52(3–4): 211–230, June 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540000152X.

Liu:2000:HLD

[53] Pingzhou Liu and Xiaoying Cui. Hyperbolic logistic difference equation with infinitely many delays. Mathematics and Computers in Simulation, 52 (3–4):231–250, June 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001531.

Vladimirov:2000:FMA

[54] Igor Vladimirov, Nikolai Kuznetsov, and Phil Diamond. Frequency measurability, algebras of quasiperiodic sets and spatial discretizations of smooth dynamical systems. Mathematics and Computers in Simulation, 52(3–4): 251–272, June 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001543.

Melnik:2000:QHM

[55] R. V. N. Melnik and Hao He. Quasihydrodynamic modelling and computer simulation of coupled thermoelectrical processes in semiconductors. Mathematics and Computers in Simulation, 52(3–4):273–287, June 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001610.

Reick:2000:TSP

[56] Christian H. Reick and Bernd Page. Time series prediction by multivariate next neighbor methods with application to zooplankton forecasts. Mathematics and Computers in Simulation, 52(3–4):289–310, June 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001567.

Delay:2000:NSG

[57] Frédérick Delay and Jean-Luc Lamotte. Numerical simulations of geological reservoirs: improving their conditioning through the use of entropy. Mathematics and Computers in Simulation, 52(3–4):311–331, June 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001579.

Anonymous:2000:PJb

[58] Anonymous. Pages 175–343 (30 June 2000). Mathematics and Computers in Simulation, 52(3–4):??, June 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Krutitskii:2000:ASS

[59] P. A. Krutitskii. Acoustic scattering by several obstacles and screens with Neumann boundary condition. Mathematics and Computers in Simulation, 52 (5–6):345–360, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001518.

Pham:2000:USS

[60] Joseph N. Q. Pham. A unilateral shift setting for the fast wavelet transform. Mathematics and Computers in Simulation, 52(5–6):361– 379, July 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001592.

Golenko-Ginzburg:2000:TLC

[61] Dimitri Golenko-Ginzburg, Gonik Aharon, and Sitniakovski Shimon. Two-level cost-optimization production control model under random disturbances. Mathematics and Computers in Simulation, 52(5–6):381– 398, July 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001634.

Chen:2000:FLA

[62] Y. M. Chen and H. C. Huang. Fuzzy logic approach to multisensor data association. Mathematics and Computers in Simulation, 52(5–6):399– 412, July 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001622.

Graber:2000:RTM

[63] Werner K. Graber and Fritz Gassmann. Real time modelling as an emergency decision support system for accidental release of air pollutants. Mathematics and Computers in Simulation, 52 (5–6):413–426, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400001646.

Yu:2000:GCM

[64] Jin-Yi Yu. A general circulation model of the atmosphere using the full-Galerkin method. Mathematics and Computers in Simulation, 52(5– 6):427–443, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000166X.

Garcia:2000:IUS

[65] Salvador Garcia. Incremental unknowns for solving the incompressible Navier–Stokes equations. Mathematics and Computers in Simulation, 52 (5–6):445–489, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001580.

Mackevicius:2000:NSD

[66] Vigirdas Mackevičius. A note on synchronization of diffusion. Mathematics and Computers in Simulation, 52 (5–6):491–495, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001609.

Anonymous:2000:Ib

[67] Anonymous. Index. Mathematics and Computers in Simulation, 52(5– 6):507–508, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540000197X.

Anonymous:2000:PJc

[68] Anonymous. Pages 345–508 (15 July 2000). Mathematics and Computers in Simulation, 52(5–6):??, July 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mohamad:2000:DCD

[69] S. Mohamad and K. Gopalsamy. Dynamics of a class of discrete-time neural networks and their continuoustime counterparts. Mathematics and Computers in Simulation, 53(1–2): 1–39, August 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001683.

Sanders:2000:SBE

[70] D. A. Sanders and A. D. Hudson. A specific blackboard expert system to simulate and automate the design of high recirculation airlift reactors. Mathematics and Computers in Simulation, 53(1–2):41–65, August 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001695.

Veglis:2000:ASM

[71] A. A. Veglis and A. S. Pombortsis. Analytical simulation of multiprocessor architectures under nonuniform traffic loads. Mathematics and Computers in Simulation, 53(1– 2):67–83, August 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001749.

Kreimer:2000:RTM

[72] Joseph Kreimer. Real-time multiserver system with two non-identical channels and limited maintenance facilities. Mathematics and Computers in Simulation, 53(1–2):85–94, August 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001713.

Barry:2000:AAR

[73] D. A. Barry, J.-Y Parlange, L. Li, H. Prommer, C. J. Cunningham, and F. Stagnitti. Analytical approximations for real values of the Lambert W -function. Mathematics and Computers in Simulation, 53(1–2): 95–103, August 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001725.

Dimitri:2000:RSS

[74] Golenko-Ginzburg Dimitri, Sitniakovski Shimon, and Papic Ljubisa. Resource supportability simulation model for a man–machine production system. Mathematics and Computers in Simulation, 53(1–2):105– 112, August 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001750.

Barry:2000:CSO

[75] D. A. Barry, K. Bajracharya, M. Crapper, H. Prommer, and C. J. Cunningham. Comparison of split-operator methods for solving coupled chemical non-equilibrium reaction/groundwater transport models. Mathematics and Computers in Simulation, 53(1–2): 113–127, August 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001828.

Anonymous:2000:PAb

[76] Anonymous. Pages 1–138 (15 August 2000). Mathematics and Computers in Simulation, 53(1–2):??, August 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ozyildirim:2000:GAC

[77] Süheyla Özyıldırım. Genetic algorithm for closed-loop equilibrium of high-order linear–quadratic dynamic games. Mathematics and Computers in Simulation, 53(3):139–147, September 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001555.

Shu:2000:ABP

[78] H. Z. Shu. Approximation of a bidimensional problem with skin effect. Mathematics and Computers in Simulation, 53(3):149–170, September 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001816.

Wang:2000:CSF

[79] Ji-Wen Wang and Ru-Xun Liu. A comparative study of finite volume methods on unstructured meshes for simulation of 2D shallow water wave problems. Mathematics and Computers in Simulation, 53(3):171–184, September 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001737.

Razzaghi:2000:LWD

[80] M. Razzaghi and S. Yousefi. Legendre wavelets direct method for variational problems. Mathematics and Computers in Simulation, 53(3):185– 192, September 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001701.

Guerra:2000:CSS

[81] Maria Letizia Guerra and Luciano Stefanini. A comparative simulation study for estimating diffusion coefficient. Mathematics and Computers in Simulation, 53(3):193–203, September 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001968.

Anonymous:2000:PS

[82] Anonymous. Pages 139–213 (30 September 2000). Mathematics and Computers in Simulation, 53(3):??, September 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Troch:2000:P

[83] I. Troch and F. Breitenecker. Preface. Mathematics and Computers in Simulation, 53(4–6):215–217, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002068.

Karnopp:2000:RAI

[84] Dean Karnopp. Retaining analog intuition in a digital world with bond graphs. Mathematics and Computers in Simulation, 53(4–6):219–226, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002081.

Willems:2000:MDS

[85] Jan C. Willems. Modelling dynamical systems using manifest and latent variables. Mathematics and Computers in Simulation, 53(4–6):227– 237, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002093.

Murray-Smith:2000:ISA

[86] D. J. Murray-Smith. The inverse simulation approach: a focused review of methods and applications. Mathematics and Computers in Simulation, 53(4–6):239– 247, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000210X.

Braun:2000:ACA

[87] S. Braun. Application of computeralgebra simulation (CALS) in industry. Mathematics and Computers in Simulation, 53(4–6):249–257, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002111.

Frank:2000:MFD

[88] P. M. Frank, E. Alcorta Garcıá, and B. Köppen-Seliger. Modelling for fault detection and isolation versus modelling for control. Mathematics and Computers in Simulation, 53(4–6):259– 271, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002123.

Muller:2000:DSP

[89] P. C. Müller. Descriptor systems: pros and cons of system modelling by differential-algebraic equations. Mathematics and Computers in Simulation, 53(4–6):273–279, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002135.

Clauss:2000:SCH

[90] C. Clauß, P. Schwarz, B. Straube, and W. Vermeiren. Symbolically calculated higher index conditions for linear circuits. Mathematics and Computers in Simulation, 53(4–6):281– 286, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002147.

Wagner:2000:FIC

[91] Yvonne Wagner. A further index concept for linear PDAEs of hyperbolic type. Mathematics and Computers in Simulation, 53(4–6):287– 291, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002159.

Gunther:2000:SMA

[92] Michael Günther. Semidiscretization may act like a deregularization. Mathematics and Computers in Simulation, 53(4–6):293–301, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002160.

Matko:2000:VVP

[93] Drago Matko, Gerhard Geiger, and Withold Gregoritza. Verification of various pipeline models. Mathematics and Computers in Simulation, 53(4–6): 303–308, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002172.

Kordt:2000:NMR

[94] M. Kordt and J. Ackermann. Nonlinear model reduction — method and CAE-tool development. Mathematics and Computers in Simulation, 53(4–6): 309–321, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002184.

Reibiger:2000:BGM

[95] A. Reibiger and H. Loose. Bond graphs and matroids. Mathematics and Computers in Simulation, 53(4–6): 323–332, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002196.

Schwarz:2000:POM

[96] Peter Schwarz. Physically oriented modeling of heterogeneous systems. Mathematics and Computers in Simulation, 53(4–6):333–344, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002202.

Aime:2000:MSC

[97] M. L. Aime and C. Maffezzoni. Modelling and simulation of combined lumped and distributed systems by an object-oriented approach. Mathematics and Computers in Simulation, 53(4–6): 345–351, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002263.

Book:2000:AGT

[98] Wayne J. Book and Cody Watson. Alternatives in the generation of time domain models of fluid lines using frequency domain techniques. Mathematics and Computers in Simulation, 53(4–6):353–365, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002275.

Hughes:2000:FLC

[99] R. L. Hughes. The flow of large crowds of pedestrians. Mathematics and Computers in Simulation, 53(4–6): 367–370, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002287.

Scheffran:2000:DIB

[100] Jürgen Scheffran. The dynamic interaction between economy and ecology: Cooperation, stability and sustainability for a dynamic-game model of resource conflicts. Mathematics and Computers in Simulation, 53(4–6): 371–380, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002299.

Borutzky:2000:DBG

[101] W. Borutzky, B. Barnard, and J. U. Thoma. Describing bond graph models of hydraulic components in Modelica. Mathematics and Computers in Simulation, 53(4–6):381–387, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002305.

Dinkelmann:2000:MSU

[102] M. Dinkelmann, M. Wächter, and G. Sachs. Modelling and simulation of unsteady heat transfer for aerospacecraft trajectory optimization. Mathematics and Computers in Simulation, 53(4–6):389–394, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002317.

Menezo:2000:ABR

[103] C. Ménézo, H. Bouia, J. J. Roux, and J. Virgone. Adaptation of the balanced realization to the coupling of reduced order models for the modelling of the thermal behavior of buildings. Mathematics and Computers in Simulation, 53(4–6):395– 401, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002329.

Mihalyko:2000:MHT

[104] Cs. Mihálykó, B. G. Lakatos, and T. Blickle. Modelling heat transfer between solid particles. Mathematics and Computers in Simulation, 53(4–6): 403–408, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002330.

Kugi:2000:MSH

[105] Andreas Kugi, Kurt Schlacher, Heinz Aitzetmüller, and Gottfried Hirmann. Modeling and simulation of a hydrostatic transmission with variabledisplacement pump. Mathematics and Computers in Simulation, 53(4–6):409– 414, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002342.

Aamo:2000:FEM

[106] O. M. Aamo and T. I. Fossen. Finite element modelling of mooring lines. Mathematics and Computers in Simulation, 53(4–6):415–422, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002354.

Woloszyn:2000:ABM

[107] Monika Woloszyn, Gilles Rusaouën, Jean-Jacques Roux, and Thierry Dagusé. Adapting block method to solve moist air flow model. Mathematics and Computers in Simulation, 53(4–6): 423–428, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002366.

Meusburger:2000:MWW

[108] Mathias Meusburger, Kurt Schlacher, and Alfons Sillaber. Modeling of the works water section of a power plant group. Mathematics and Computers in Simulation, 53(4–6):429– 435, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002378.

Yamashita:2000:OEA

[109] K. Yamashita, A. Shimabukuro, M. R. Asharif, and H. Miyagi. Orthogonal ECLMS algorithm for doubletalk echo cancelling. Mathematics and Computers in Simulation, 53(4–6): 437–442, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000238X.

Dens:2000:ITS

[110] E. J. Dens and J. F. Van Impe. On the importance of taking space into account when modeling microbial competition in structured food products. Mathematics and Computers in Simulation, 53(4–6):443– 448, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002391.

Klatt:2000:MCE

[111] Karsten-Ulrich Klatt, Guido Dünnebier, and Sebastian Engell. Modeling and computationally efficient simulation of chromatographic separation processes. Mathematics and Computers in Simulation, 53(4–6):449–455, October 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002408.

Anonymous:2000:Ic

[112] Anonymous. Index. Mathematics and Computers in Simulation, 53(4–6): 473–475, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002755.

Anonymous:2000:PO

[113] Anonymous. Pages 215–475 (30 October 2000). Mathematics and Computers in Simulation, 53(4–6):??, October 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Duc:2000:SDC

[114] Hiep Duc, Ian Shannon, and Merched Azzi. Spatial distribution characteristics of some air pollutants in Sydney. Mathematics and Computers in Simulation, 54(1–3):1–21, November 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001658.

Yang:2000:NAA

[115] Cheng-Hong Yang, Ching-Hsing Luo, Yuan-Long Jeang, and Gwo-Jia Jon. A novel approach to adaptive Morse code recognition for disabled persons. Mathematics and Computers in Simulation, 54(1–3):23–32, November 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001804.

Chard:2000:MDS

[116] Jeffrey A. Chard and Vadim Shapiro. A multivector data structure for differential forms and equations. Mathematics and Computers in Simulation, 54(1–3): 33–64, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001981.

Zhang:2000:TCF

[117] Jun Zhang, Lixin Ge, and Jules Kouatchou. A two colorable fourthorder compact difference scheme and parallel iterative solution of the 3D convection diffusion equation. Mathematics and Computers in Simulation, 54(1–3):65–80, November 30, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002056.

Hirose:2000:MLP

[118] Hideo Hirose. Maximum likelihood parameter estimation by model augmentation with applications to the extended four-parameter generalized gamma distribution. Mathematics and Computers in Simulation, 54(1–3):81– 97, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002019.

Makroglou:2000:CTI

[119] Athena Makroglou. Computer treatment of the integro-differential equations of collective non-ruin; the finite time case. Mathematics and Computers in Simulation, 54(1–3):99– 112, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002020.

Bove:2000:TED

[120] A. Bove, S. DeMartino, and G. Lauro. Trend to equilibrium in the dynamics of a gas interacting with a radiation field. Mathematics and Computers in Simulation, 54(1–3):113– 129, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002032.

Snyder:2000:AEQ

[121] William C. Snyder. Accuracy estimation for quasi-Monte Carlo simulations. Mathematics and Computers in Simulation, 54(1–3):131– 143, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002044.

Potkonjak:2000:CPS

[122] Veljko Potkonjak, Spyros Tzafestas, and Dragan Kostic. Concerning the primary and secondary objectives in robot task definition — the “learn from humans” principle. Mathematics and Computers in Simulation, 54(1–3):145– 157, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002251.

Fatullayev:2000:NPD

[123] Afet Fatullayev and Emine Can. Numerical procedures for determining unknown source parameter in parabolic equations. Mathematics and Computers in Simulation, 54(1–3):159– 167, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002214.

Gumel:2000:NBA

[124] A. B. Gumel, E. H. Twizell, and P. Yu. Numerical and bifurcation analyses for a population model of HIV chemotherapy. Mathematics and Computers in Simulation, 54(1–3):169– 181, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002226.

Labzovski:2000:PVP

[125] Shmuel N. Labzovski, Abraham Mehrez, and Ilia B. Frenkel. The a priori vacation probability in the M/G/1 single vacation models. Mathematics and Computers in Simulation, 54(1–3): 183–188, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002238.

Kuchler:2000:SDT

[126] Uwe Küchler and Eckhard Platen. Strong discrete time approximation of stochastic differential equations with time delay. Mathematics and Computers in Simulation, 54(1–3):189– 205, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540000224X.

Anonymous:2000:PN

[127] Anonymous. Pages 1–221 (30 November 2000). Mathematics and Computers in Simulation, 54(1–3):??, November 30, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Boisvert:2000:P

[128] Ronald Boisvert, Wayne Dyksen, and Elias Houstis. Preface. Mathematics and Computers in Simulation, 54(4– 5):223, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000183X.

Boisvert:2000:MSP

[129] Ronald F. Boisvert. Mathematical software: past, present, and future. Mathematics and Computers in Simulation, 54(4–5):227–241, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001853.

Houstis:2000:FPS

[130] Elias N. Houstis and John R. Rice. Future problem solving environments for computational science. Mathematics and Computers in Simulation, 54(4–5): 243–257, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400001877.

Trefethen:2000:NAD

[131] Anne E. Trefethen and Brian Ford. Numerical algorithm delivery mechanisms. Mathematics and Computers in Simulation, 54(4–5):259–268, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002561.

Schonauer:2000:NED

[132] Willi Schönauer. Numerical engineering: design of PDE black-box solvers. Mathematics and Computers in Simulation, 54(4–5):269–277, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001889.

Marinescu:2000:CBA

[133] Dan C. Marinescu and Ladislau Bölöni. A component-based architecture for problem solving environments. Mathematics and Computers in Simulation, 54(4–5):279–293, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001890.

Fox:2000:CSI

[134] Geoffrey C. Fox. From computational science to Internetics: Integration of science with computer science. Mathematics and Computers in Simulation, 54(4–5):295–306, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001907.

Eigenmann:2000:TCP

[135] Rudolf Eigenmann and Michael J. Voss. Towards a compilation paradigm for computational applications on the information power grid. Mathematics and Computers in Simulation, 54(4–5): 307–320, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001919.

Chrisochoides:2000:SMG

[136] Nikos Chrisochoides and Démian Nave. Simultaneous mesh generation and partitioning for Delaunay meshes. Mathematics and Computers in Simulation, 54(4–5):321–339, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001920.

Oldehoeft:2000:TCH

[137] Rod Oldehoeft. Taming complexity in high-performance computing. Mathematics and Computers in Simulation, 54(4–5):341–357, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400001932.

Dyksen:2000:NDT

[138] Wayne R. Dyksen and Robert E. Lynch. A new decoupling technique for the Hermite cubic collocation equations arising from boundary value problems. Mathematics and Computers in Simulation, 54(4–5):359– 372, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400001944.

Koulisianis:2000:MIM

[139] M. D. Koulisianis and T. S. Papatheodorou. A ‘moving index’ method for the solution of the American options valuation problem. Mathematics and Computers in Simulation, 54(4–5):

373–381, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001762.

Yang:2000:FEE

[140] Daoqi Yang. Finite elements for elliptic problems with wild coefficients. Mathematics and Computers in Simulation, 54(4–5):383–395, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001774.

Chaitin-Chatelin:2000:AHC

[141] F. Chaitin-Chatelin, A. Harrabi, and A. Ilahi. About Hölder condition numbers and the stratification diagram for defective eigenvalues. Mathematics and Computers in Simulation, 54(4–5):397– 402, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001786.

Gautschi:2000:GRF

[142] Walter Gautschi. Gauss–Radau formulae for Jacobi and Laguerre weight functions. Mathematics and Computers in Simulation, 54(4–5):403– 412, December 15, 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001798.

Anonymous:2000:PD

[143] Anonymous. Pages 223–423 (15 December 2000). Mathematics and Computers in Simulation, 54(4–5):??, December 15, 2000. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liu:2001:NLS

[144] Jui-Jung Liu, Shan-Jen Cheng, I-Chung Kung, Hui-Chen Chang, and S. A. Billings. Non-linear system identification and fault diagnosis using a new GUI interpretation tool. Mathematics and Computers in Simulation, 54(6): 425–449, January 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002743.

Golenko-Ginzburg:2001:TCP

[145] Dimitri Golenko-Ginzburg and Zohar Laslo. Timing control points via simulation for production systems under random disturbances. Mathematics and Computers in Simulation, 54(6): 451–458, January 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400001993.

Kurbanmuradov:2001:EMC

[146] O. Kurbanmuradov, Ü Rannik, K. Sabelfeld, and T. Vesala. Evaluation of mean concentration and fluxes in turbulent flows by Lagrangian stochastic models. Mathematics and Computers in Simulation, 54(6):459– 476, January 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002731.

Janssen:2001:SOI

[147] Marco A. Janssen. Sequential optimization of integrated climate change models. Mathematics and Computers in Simulation, 54(6):477–489, January 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002780.

Yanchuk:2001:PSC

[148] Sergiy Yanchuk, Yuri Maistrenko, and Erik Mosekilde. Partial synchronization and clustering in a system of diffusively coupled chaotic oscillators. Mathematics and Computers in Simulation, 54(6):491–508, January 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002767.

Kouatchou:2001:PIH

[149] Jules Kouatchou. Parallel implementation of a high-order implicit collocation method for the heat equation. Mathematics and Computers in Simulation, 54(6):509–519, January 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002809.

Anonymous:2001:I

[150] Anonymous. Index. Mathematics and Computers in Simulation, 54(6): 533–535, January 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000313X.

Anonymous:2001:PJa

[151] Anonymous. Pages 425–535 (15 January 2001). Mathematics and Computers in Simulation, 54(6):??, January 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Dimov:2001:P

[152] I. Dimov. Preface. Mathematics and Computers in Simulation, 55(1– 3):1–2, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002573.

Agosteo:2001:FSM

[153] S. Agosteo, C. Birattari, A. Foglio Para, M. Silari, and L. Ulrici. FLUKA simulations and measurements for a dump for a 250 GeV/c hadron beam. Mathematics and Computers in Simulation, 55(1–3):3–14, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540000241X.

Angelova:2001:AMC

[154] D. S. Angelova, Tz. A. Semerdjiev, V. P. Jilkov, and E. A. Semerdjiev. Application of a Monte Carlo method for tracking maneuvering target in clutter. Mathematics and Computers in Simulation, 55(1–3): 15–23, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002421.

Dimov:2001:PRM

[155] I. Dimov, V. Alexandrov, and A. Karaivanova. Parallel resolvent Monte Carlo algorithms for linear algebra problems. Mathematics and Computers in Simulation, 55(1–3): 25–35, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002433.

Dubus:2001:MCS

[156] A. Dubus, M. Rösler, and O. Benka. Monte Carlo simulation of kinetic electron emission induced by MeV He^{+} and He^{++} ions incident on polycrystalline aluminium. Mathematics and Computers in Simulation, 55(1–3): 37–48, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002445.

Entacher:2001:DEB

[157] Karl Entacher. Discrepancy estimates based on Haar functions. Mathematics and Computers in Simulation, 55(1–3): 49–57, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002457.

Iverson:2001:FRW

[158] Ralph B. Iverson and Yannick L. Le Coz. A floating random-walk algorithm for extracting electrical capacitance. Mathematics and Computers in Simulation, 55(1–3):59–66, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002469.

Jacoboni:2001:WFF

[159] Carlo Jacoboni, Andrea Bertoni, Paolo Bordone, and Rossella Brunetti. Wigner-function formulation for quantum transport in semiconductors: theory and Monte Carlo approach. Mathematics and Computers in Simulation, 55(1–3):67–78, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475400002470.

Keller:2001:HMC

[160] Alexander Keller. Hierarchical Monte Carlo image synthesis. Mathematics and Computers in Simulation, 55(1– 3):79–92, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002482.

Kosina:2001:MCM

[161] H. Kosina and M. Nedjalkov. The Monte Carlo method for semi-classical charge transport in semiconductor devices. Mathematics and Computers in Simulation, 55(1–3):93–102, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002494.

Langtry:2001:CHR

[162] Timothy N. Langtry. A construction of higher-rank lattice rules. Mathematics and Computers in Simulation, 55(1–3): 103–111, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002500.

Lecot:2001:QRN

[163] Christian Lécot. Quasi-randomized numerical methods for systems with coefficients of bounded variation. Mathematics and Computers in Simulation, 55(1–3):113–121, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002512.

Lecot:2001:PAC

[164] Christian Lécot and Wolfgang Ch. Schmid. Particle approximation of convection–diffusion equations. Mathematics and Computers in Simulation, 55(1–3):123–130, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002524.

LEcuyer:2001:PBS

[165] Pierre L’Ecuyer and Richard Simard. On the performance of birthday spacings tests with certain families of random number generators. Mathematics and Computers in Simulation, 55(1–3): 131–137, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002536.

Lemieux:2001:SCL

[166] Christiane Lemieux and Pierre L’Ecuyer. On selection criteria for lattice rules and other quasi-Monte Carlo point sets. Mathematics and Computers in Simulation, 55(1–3):139–148, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002548. The

Second IMACS Seminar on Monte Carlo Methods (Varna, 1999).

Liang:2001:NET

[167] Yufeng Liang and P. A. Whitlock. A new empirical test for parallel pseudo-random number generators. Mathematics and Computers in Simulation, 55(1–3):149–158, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540000255X.

Marseguerra:2001:EMV

[168] M. Marseguerra, E. Padovani, and E. Zio. Effects of model variability on the safeguard analysis of gamma-contaminated truck loads. Mathematics and Computers in Simulation, 55(1–3):159–166, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002585.

Marseguerra:2001:LMC

[169] M. Marseguerra and E. Zio. Looking at Monte Carlo simulation for describing nonlinear sorption in groundwater contaminant transport. Mathematics and Computers in Simulation, 55(1–3): 167–176, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002597.

Mori:2001:PRS

[170] Makoto Mori. Pseudo random sequences generated by piecewise linear maps from the view point of dynamical system. Mathematics and Computers in Simulation, 55(1–3):177– 189, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002603.

Nedjalkov:2001:VEM

[171] M. Nedjalkov and H. Kosina. Variance of the ensemble Monte Carlo algorithm for semiconductor transport modeling. Mathematics and Computers in Simulation, 55(1–3):191– 198, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002615.

Nilsson:2001:CBD

[172] H.-E. Nilsson, E. Bellotti, M. Hjelm, and K. Brennan. A comparison between different Monte Carlo models in simulation of hole transport in 4H-SiC. Mathematics and Computers in Simulation, 55(1–3):199– 208, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002627.

Ogawa:2001:DAN

[173] Shigeyoshi Ogawa. On a deterministic approach to the numerical solution of the SDE. Mathematics and Computers in Simulation, 55(1–3):209– 214, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002639.

Okten:2001:HDS

[174] Giray Ökten. High dimensional simulation. Mathematics and Computers in Simulation, 55(1–3):215–222, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002640.

Ota:2001:NSM

[175] Masahiro Ota, Tomohiko Nakao, and Moriyoshi Sakamoto. Numerical simulation of molecular motion around laser microengine blades. Mathematics and Computers in Simulation, 55(1–3): 223–230, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002652.

Pennetta:2001:MCP

[176] C. Pennetta, L. Reggiani, and Gy. Trefán. A Monte Carlo percolative approach to reliability analysis of semiconductor structures. Mathematics and Computers in Simulation, 55(1–3): 231–238, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002664.

Schmid:2001:PDN

[177] Wolfgang Ch. Schmid. Projections of digital nets and sequences. Mathematics and Computers in Simulation, 55(1–3):239–247, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002676.

Schmid:2001:TPQ

[178] Wolfgang Ch. Schmid and Andreas Uhl. Techniques for parallel quasi-Monte Carlo integration with digital sequences and associated problems. Mathematics and Computers in Simulation, 55(1–3):249–257, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002688.

Smidts:2001:SAM

[179] O. F. Smidts and O. Roussille. Sensitivity analysis in the migration of radionuclides: differential Monte Carlo versus double randomization. Mathematics and Computers in Simulation, 55(1–3): 259–270, February 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540000269X.

Sobol:2001:GSI

[180] I. M. Sobol’. Global sensitivity indices for nonlinear mathematical models and their Monte Carlo estimates. Mathematics and Computers in Simulation, 55(1–3):271–280, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002706.

Wegenkittl:2001:GTP

[181] Stefan Wegenkittl. Gambling tests for pseudorandom number generators. Mathematics and Computers in Simulation, 55(1–3):281–288, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002718.

Anonymous:2001:NIU

[182] Anonymous. News of IMACSupdate. Mathematics and Computers in Simulation, 55(1–3):289–294, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401002798.

Anonymous:2001:C

[183] Anonymous. CALENDAR. Mathematics and Computers in Simulation, 55(1–3):295–305, February 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401002804.

Taha:2001:Fa

[184] Thiab Taha, Jerry Bona, and Anne de Bouard. Foreword. Mathematics and Computers in Simulation, 55 (4–6):307, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002737.

Anonymous:2001:T

[185] Anonymous. (Taha S.I.). Mathematics and Computers in Simulation, 55(4– 6):307–648, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Baboolal:2001:FDM

[186] S. Baboolal. Finite-difference modeling of solitons induced by a density hump in a plasma multi-fluid. Mathematics and Computers in Simulation, 55(4– 6):309–316, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003104.

Boyd:2001:WNW

[187] John P. Boyd and Guan-Yu Chen. Weakly nonlinear wavepackets in the Korteweg–de Vries equation: the KdV/ NLS connection. Mathematics and Computers in Simulation, 55(4–6): 317–328, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002913.

Cai:2001:SCS

[188] David Cai, David W. McLaughlin, and Jalal Shatah. Spatiotemporal chaos in spatially extended systems. Mathematics and Computers in Simulation, 55(4–6):329–340, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002998.

Calini:2001:KTF

[189] Annalisa M. Calini and Thomas A. Ivey. Knot types, Floquet spectra, and finite-gap solutions of the vortex filament equation. Mathematics and Computers in Simulation, 55(4– 6):341–350, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003013.

Schober:2001:MAS

[190] C. M. Schober and A. Calini. Mel’nikov analysis of a symmetrybreaking perturbation of the NLS equation. Mathematics and Computers in Simulation, 55(4–6):351– 364, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540000286X.

Campbell:2001:NSN

[191] L. J. Campbell and S. A. Maslowe. A numerical simulation of the nonlinear critical layer evolution of a forced Rossby wave packet in a zonal shear flow. Mathematics and Computers in Simulation, 55(4–6):365– 375, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002937.

Choudhury:2001:NDN

[192] S. Roy Choudhury and Kevin G. Brown. Novel dynamics in the nonlinear evolution of the Kelvin– Helmholtz instability of supersonic anisotropic tangential velocity discontinuities. Mathematics and Computers in Simulation, 55(4–6):377– 391, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400003049.

Daripa:2001:WNL

[193] Prabir Daripa and Ranjan K. Dash. Weakly non-local solitary wave solutions of a singularly perturbed Boussinesq equation. Mathematics and Computers in Simulation, 55(4–6): 393–405, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002883.

Donagi:2001:AS

[194] Ron Y. Donagi and Emma Previato. Abelian solitons. Mathematics and Computers in Simulation, 55(4– 6):407–418, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003062.

Iron:2001:SPG

[195] David Iron and Michael J. Ward. Spike pinning for the Gierer–Meinhardt model. Mathematics and Computers in Simulation, 55(4–6):419–431, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400003037.

Jordan:2001:SES

[196] Richard Jordan and Christophe Josserand. Statistical equilibrium states for the nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 55(4–6):433– 447, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002925.

Kevrekidis:2001:COD

[197] P. G. Kevrekidis, K. Ø. Rasmussen, and A. R. Bishop. Comparison of one-dimensional and two-dimensional discrete breathers. Mathematics and Computers in Simulation, 55(4–6): 449–462, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002986.

Killen:2001:PAS

[198] S. M. Killen and R. S. Johnson. Propagation of axi-symmetric nonlinear shallow water waves over slowly varying depth. Mathematics and Computers in Simulation, 55(4–6): 463–472, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002949.

King:2001:DCP

[199] R. Bruce King. The development of the concept of Painlevé chains for the classification of integrable higher-order differential equations. Mathematics and Computers in Simulation, 55(4– 6):473–482, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003128.

Kraenkel:2001:MKV

[200] R. A. Kraenkel and A. I. Zenchuk. Modified Korteweg–de Vries hierarchy with hodograph transformation: Camassa–Holm and Harry–Dym hierarchies. Mathematics and Computers in Simulation, 55(4–6):483– 491, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400003086.

Lanson:2001:MMC

[201] N. Lanson and J. P. Vila. Meshless methods for conservation laws. Mathematics and Computers in Simulation, 55(4–6):493–501, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475400002858.

LeMesurier:2001:MFS

[202] Brenton J. LeMesurier. Multi-focusing and sustained dissipation in the dissipative nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 55(4–6):503–517, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400003074.

Lu:2001:SCS

[203] Xiaowu Lu. Symplectic computation of solitary waves for general Sine– Gordon equations. Mathematics and Computers in Simulation, 55(4–6): 519–532, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003001.

Ludu:2001:NLS

[204] A. Ludu and J. P. Draayer. Nonlinearity and self-similarity: patterns and clusters. Mathematics and Computers in Simulation, 55(4–6):533– 540, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002950.

Malfliet:2001:DSW

[205] Willy Malfliet and Bjorn Rombouts. Dressed shock waves in a nonlinear LC circuit. Mathematics and Computers in Simulation, 55(4–6): 541–547, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003025.

Mickens:2001:NFD

[206] Ronald E. Mickens. A nonstandard finite difference scheme for a nonlinear PDE having diffusive shock wave solutions. Mathematics and Computers in Simulation, 55(4–6): 549–555, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003098.

Miller:2001:SPS

[207] Judith R. Miller. Stability properties of solitary waves in a complex modified KdV system. Mathematics and Computers in Simulation, 55(4– 6):557–565, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400003116.

Nicholls:2001:HGW

[208] David P. Nicholls. On hexagonal gravity water waves. Mathematics and Computers in Simulation, 55(4– 6):567–575, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002962.

Panayotaros:2001:STW

[209] Panayotis Panayotaros. Stability of traveling water waves on the sphere. Mathematics and Computers in Simulation, 55(4–6):577–584, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002895.

Pelinovsky:2001:MTT

[210] Dmitry E. Pelinovsky. A mysterious threshold for transverse instability of deep-water solitons. Mathematics and Computers in Simulation, 55(4– 6):585–594, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002871.

Pelloni:2001:NMT

[211] Beatrice Pelloni and Vassilios A. Dougalis. Numerical modelling of twoway propagation of non-linear dispersive waves. Mathematics and Computers in Simulation, 55(4–6): 595–606, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400003050.

Ramos:2001:STP

[212] J. I. Ramos. Spatio-temporal patterns in excitable media with nonsolenoidal flow straining. Mathematics and Computers in Simulation, 55(4– 6):607–619, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002974.

Stoitcheva:2001:AMF

[213] G. Stoitcheva, A. Ludu, and J. P. Draayer. Antisoliton model for fission modes. Mathematics and Computers in Simulation, 55(4–6):621– 625, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002901.

Anonymous:2001:NIa

[214] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 55(4–6):627–632, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401003007.

Anonymous:2001:ICEa

[215] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 55(4–6):633–644, March 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003019.

Anonymous:2001:AI

[216] Anonymous. Author index. Mathematics and Computers in Simulation, 55(4–6):645–647, March 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401003020.

Abouchabaka:2001:SOS

[217] J. Abouchabaka, R. Aboulaich, O. Guennoun, A. Nachaoui, and A. Souissi. Shape optimization for a simulation of a semiconductor problem. Mathematics and Computers in Simulation, 56(1):1– 16, March 29, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400001956.

Alt:2001:EEF

[218] R. Alt and J.-L. Lamotte. Experiments on the evaluation of functional ranges using a random interval arithmetic. Mathematics and Computers in Simulation, 56(1):17– 34, March 29, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002779.

Qiu:2001:SRB

[219] Zhiping Qiu, Peter C. Müller, and Andreas Frommer. Stability robustness bounds for linear state– space models with structured uncertainty based on ellipsoidal settheoretic approach. Mathematics and Computers in Simulation, 56(1):35– 53, March 29, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475400002007.

Liu:2001:DFE

[220] Ru-Xun Liu, Hong Li, and Zhi-Feng Wang. The discontinuous finite element method for red-and-green light models for the traffic flow. Mathematics and Computers in Simulation, 56 (1):55–67, March 29, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475400002834.

ElKahoui:2001:IAL

[221] M’hammed El Kahoui, Andreas Weber, and Bernd Eberhardt. Improved algorithms for linear complementarity problems arising from collision response. Mathematics and Computers in Simulation, 56(1):69– 93, March 29, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002810.

Anonymous:2001:NIb

[222] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 56(1):95–100, March 29, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002981.

Anonymous:2001:ICEb

[223] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 56(1):101–112, March 29, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401002993.

Anonymous:2001:PM

[224] Anonymous. Pages 1–112 (29 March 2001). Mathematics and Computers in Simulation, 56(1):??, March 29, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Taha:2001:Fb

[225] Thiab Taha, J. Mac Hyman, Linda Petzold, and William Schiesser. Foreword. Mathematics and Computers in Simulation, 56(2):113, May 13, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003032.

Anonymous:2001:ML

[226] Anonymous. (Method of Lines S.I.). Mathematics and Computers in Simulation, 56(2):113–222, May 13, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ahmad:2001:MSH

[227] I. Ahmad and M. Berzins. MOL solvers for hyperbolic PDEs with source terms. Mathematics and Computers in Simulation, 56(2):115– 125, May 13, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401002841.

Cao:2001:CTD

[228] Weiming Cao, Weizhang Huang, and Robert D. Russell. Comparison of two-dimensional r-adaptive finite element methods using various error indicators. Mathematics and Computers in Simulation, 56(2):127– 143, May 13, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401002853.

Karpeev:2001:SID

[229] D. A. Karpeev and C. M. Schober. Symplectic integrators for discrete nonlinear Schrödinger systems. Mathematics and Computers in Simulation, 56 (2):145–156, May 13, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002865.

Kohler:2001:SPS

[230] R. Köhler, A. Gerstlauer, and M. Zeitz. Symbolic preprocessing for simulation of PDE models of chemical processes. Mathematics and Computers in Simulation, 56(2):157–170, May 13, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401002877.

Saucez:2001:UML

[231] Philippe Saucez, W. E. Schiesser, and Alain Vande Wouwer. Upwinding in the method of lines. Mathematics and Computers in Simulation, 56(2): 171–185, May 13, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002889.

Serban:2001:CSP

[232] Radu Serban and Linda R. Petzold. COOPT — a software package for optimal control of large-scale differential– algebraic equation systems. Mathematics and Computers in Simulation, 56 (2):187–203, May 13, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002890.

Anonymous:2001:NIc

[233] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 56(2):205–210, May 13, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003494.

Anonymous:2001:ICEc

[234] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 56(2):211–222, May 13, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003500.

Li:2001:DGF

[235] Hong Li and RuXun Liu. The discontinuous Galerkin finite element method for the 2D shallow water equations. Mathematics and Computers in Simulation, 56(3):223– 233, June 4, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401002774.

Razzaghi:2001:SDE

[236] M. Razzaghi and Y. Ordokhani. Solution of differential equations via rationalized Haar functions. Mathematics and Computers in Simulation, 56 (3):235–246, June 4, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401002786.

Lin:2001:BEU

[237] H. Lin and K. Yamashita. Blind equalization using parallel Bayesian decision feedback equalizer. Mathematics and Computers in Simulation, 56 (3):247–257, June 4, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002762.

Mraz:2001:DIC

[238] Miha Mraz. The design of intelligent control of a kitchen refrigerator. Mathematics and Computers in Simulation, 56(3):259–267, June 4, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002816.

Wazwaz:2001:SND

[239] A. M. Wazwaz. A study of nonlinear dispersive equations with solitary-wave solutions having compact support. Mathematics and Computers in Simulation, 56(3):269– 276, June 4, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401002919.

Demir:2001:TCV

[240] H. Demir. Thermal convection of viscoelastic fluid with Biot boundary conduction. Mathematics and Computers in Simulation, 56(3):277– 296, June 4, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401002944.

Sun:2001:CSF

[241] Lu Sun. Computer simulation and field measurement of dynamic pavement loading. Mathematics and Computers in Simulation, 56(3):297– 313, June 4, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540100297X.

Anonymous:2001:NId

[242] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 56(3):315–319, June 4, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003524.

Anonymous:2001:ICEd

[243] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 56(3):321– 330, June 4, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003536.

Anonymous:2001:PJb

[244] Anonymous. Pages 223–330 (4 June 2001). Mathematics and Computers in Simulation, 56(3):??, June 4, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Farkas:2001:IMS

[245] I. Farkas and T. Nybrant. Introduction to M^{2} SABI’99 Special Issue. Mathematics and Computers in Simulation, 56(4–5):331–332, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003044.

Samuelsson:2001:JBS

[246] Pär Samuelsson, Mats Ekman, and Bengt Carlsson. A JAVA based simulator of activated sludge processes. Mathematics and Computers in Simulation, 56(4–5):333–346, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003056.

Ioslovich:2001:MGO

[247] Ilya Ioslovich and Per-Olof Gutman. A model for the global optimization of water prices and usage for the case of spatially distributed sources and consumers. Mathematics and Computers in Simulation, 56(4–5): 347–356, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003068.

Farkas:2001:MRP

[248] I. Farkas, P. Weihs, A. Biró, W. Laube, J. Eitzinger, and A. Wójcicki. Modelling of radiative PAR transfer in a tunnel greenhouse. Mathematics and Computers in Simulation, 56(4– 5):357–368, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100307X.

Ferreira:2001:DDL

[249] A. M. Ferreira and F. G. Abreu. Description of development, light interception and growth of sunflower at two sowing dates and two densities. Mathematics and Computers in Simulation, 56(4–5):369–384, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003081.

DeKetelaere:2001:TFE

[250] Bart De Ketelaere and Josse De Baerdemaeker. Tomato firmness estimation using vibration measurements. Mathematics and Computers in Simulation, 56(4–5):385– 394, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003093.

Meszaros:2001:NAP

[251] Cs. Mészáros, I. Farkas, and Á. Bálint. A new application of percolation theory for coupled transport phenomena through porous media. Mathematics and Computers in Simulation, 56(4– 5):395–404, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100310X.

Trelea:2001:PMB

[252] Ioan Cristian Trelea, Mariana Titica, Sophie Landaud, Eric Latrille, Georges Corrieu, and Arlette Cheruy. Predictive modelling of brewing fermentation: from knowledge-based to black-box models. Mathematics and Computers in Simulation, 56(4–5): 405–424, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003111.

Bogaerts:2001:LSE

[253] Ph. Bogaerts and R. Hanus. On-line state estimation of bioprocesses with full horizon observers. Mathematics and Computers in Simulation, 56(4– 5):425–441, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003123.

Stirbet:2001:POF

[254] Alexandrina D. Stirbet, Philipp Rosenau, Andreas C. Ströder, and Reto J. Strasser. Parameter optimisation of fast chlorophyll fluorescence induction model. Mathematics and Computers in Simulation, 56(4–5):443– 450, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003135.

Strasser:2001:EEC

[255] Reto J. Strasser and Alexandrina D. Stirbet. Estimation of the energetic connectivity of PS II centres in plants using the fluorescence rise O–J–I–P: Fitting of experimental data to three different PS II models. Mathematics and Computers in Simulation, 56(4– 5):451–462, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003147.

Xing:2001:KMG

[256] Xin-Hui Xing, Akinori Ono, Kazuhiko Miyanaga, Yasunori Tanji, and Hajime Unno. A kinetic model for growth of callus derived from Eucommia ulmoides aiming at mass production of a factor enhancing collagen synthesis of animal cells. Mathematics and Computers in Simulation, 56(4– 5):463–474, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003159.

Moshou:2001:NRS

[257] D. Moshou, A. Chedad, A. Van Hirtum, J. De Baerdemaeker, D. Berckmans, and H. Ramon. Neural recognition system for swine cough. Mathematics and Computers in Simulation, 56(4–5):475– 487, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003160.

Anonymous:2001:NIe

[258] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 56(4–5):489–494, June 11, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475401003573.

Anonymous:2001:ICEe

[259] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 56(4–5):495– 507, June 11, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003585.

Taha:2001:Fc

[260] Thiab R. Taha, Mark J. Ablowitz, Gino Biondini, Yuji Kodama, and Vladimir Zakharov. Foreword. Mathematics and Computers in Simulation, 56(6):509, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100355X.

Ablowitz:2001:LMD

[261] Mark J. Ablowitz, Gino Biondini, and Steve Blair. Localized multidimensional optical pulses in nonresonant quadratic materials. Mathematics and Computers in Simulation, 56(6):511–519, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003287.

Biswas:2001:OSP

[262] Anjan Biswas. Optical soliton perturbation with nonlinear damping and saturable amplifiers. Mathematics and Computers in Simulation, 56(6): 521–537, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003226.

Buslaev:2001:SIB

[263] V. S. Buslaev and V. E. Grikurov. Simulation of instability of bright solitons for NLS with saturating nonlinearity. Mathematics and Computers in Simulation, 56(6):539– 546, July 9, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003238.

Ismail:2001:NSC

[264] M. S. Ismail and Thiab R. Taha. Numerical simulation of coupled nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 56 (6):547–562, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100324X.

Mlejnek:2001:RFP

[265] M. Mlejnek, M. Kolesik, E. M. Wright, and J. V. Moloney. Recurrent femtosecond pulse collapse in air due to plasma generation: numerical results. Mathematics and Computers in Simulation, 56(6):563–570, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003251.

Ramos:2001:ISS

[266] J. I. Ramos. Interaction of spatial solitons with a localized spatiallymodulated medium. Mathematics and Computers in Simulation, 56(6): 571–583, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003263.

Yang:2001:ESN

[267] J. Yang, B. A. Malomed, D. J. Kaup, and A. R. Champneys. Embedded solitons: a new type of solitary wave. Mathematics and Computers in Simulation, 56(6):585– 600, July 9, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003275.

Anonymous:2001:NIf

[268] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 56(6):601–606, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003603.

Anonymous:2001:ICEf

[269] Anonymous. IMACS calendar of events — July 2001. Mathematics and Computers in Simulation, 56 (6):607–614, July 9, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003615.

Anonymous:2001:AIVa

[270] Anonymous. Author index of volume 56. Mathematics and Computers in Simulation, 56(6):615– 618, July 9, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003627.

Sheela:2001:TRE

[271] S. Sheela and Arindama Singh. Tikhonov regularization of an elliptic PDE. Mathematics and Computers in Simulation, 57(1–2):1–4, August 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540000272X.

Dimitri:2001:SFC

[272] Golenko-Ginzburg Dimitri, Sitniakovski Shimon, and Papic Ljubisa. A simulation factory control model under a chance constraint. Mathematics and Computers in Simulation, 57 (1–2):5–18, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002792.

Mehrez:2001:EPP

[273] Abraham Mehrez and Moshe Justman. On the efficiency of the parallel path R&D approach: a stochastic game analysis. Mathematics and Computers in Simulation, 57(1–2): 19–28, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475400002822.

Mackevicius:2001:SOW

[274] Vigirdas Mackevičius and Jurgis Navikas. Second order weak Runge– Kutta type methods for Itô equations. Mathematics and Computers in Simulation, 57(1–2):29–34, August 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475400002846.

Denis-Vidal:2001:SEA

[275] Lilianne Denis-Vidal, Ghislaine Joly-Blanchard, and Céline Noiret. Some effective approaches to check the identifiability of uncontrolled nonlinear systems. Mathematics and Computers in Simulation, 57(1–2):35–44, August 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401002749.

Qiu:2001:EST

[276] Zhiping Qiu, Peter C. Müller, and Andreas Frommer. Ellipsoidal settheoretic approach for stability of linear state-space models with interval uncertainty. Mathematics and Computers in Simulation, 57(1–2): 45–59, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100283X.

El-Bassiouny:2001:PBE

[277] A. F. El-Bassiouny and H. M. Abdelhafez. Prediction of bifurcations for external and parametric excited one-degree-of-freedom system with quadratic, cubic and quartic nonlinearities. Mathematics and Computers in Simulation, 57(1–2):61–80, August 15, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401002920.

Pipis:2001:HAM

[278] A. Pipis, G. Theodoropoulos, M. Stefanidakis, and D. Lioupis. A hybrid approach for the modelling and simulation of a virtually shared memory parallel computer architecture. Mathematics and Computers in Simulation, 57 (1–2):81–93, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401002907.

He:2001:UGA

[279] Yongyong He, Dan Guo, and Fulei Chu. Using genetic algorithms and finite element methods to detect shaft crack for rotor-bearing system. Mathematics and Computers in Simulation, 57(1– 2):95–108, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401002956.

Chen:2001:FDA

[280] Liang Chen and Naoyuki Tokuda. A faster data assignment algorithm for maximum likelihood-based multitarget motion tracking with bearingsonly measurements. Mathematics and Computers in Simulation, 57(1–2): 109–120, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401002932.

Anonymous:2001:PA

[281] Anonymous. Pages 1–134 (15 August 2001). Mathematics and Computers in Simulation, 57(1–2):??, August 15, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Edneral:2001:F

[282] Victor F. Edneral and Stanly Steinberg. Foreword. Mathematics and Computers in Simulation, 57(3–5):135– 137, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003329.

Anonymous:2001:CA

[283] Anonymous. (Computer Algebra S.I.). Mathematics and Computers in Simulation, 57(3–5):135–334, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Andrianov:2001:STM

[284] Serge Andrianov. Symplectification of truncated maps for Hamiltonian systems. Mathematics and Computers in Simulation, 57(3–5):139– 145, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003330.

Andrianov:2001:SCA

[285] Serge Andrianov. Symbolic computation of approximate symmetries for ordinary differential equations. Mathematics and Computers in Simulation, 57(3–5):147–153, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003342.

Aranson:2001:CAN

[286] Alexander B. Aranson. Computation and applications of the Newton polyhedrons. Mathematics and Computers in Simulation, 57(3–5):155– 160, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003354.

Banshchikov:2001:CAP

[287] Andrej Banshchikov and Larissa Bourlakova. Computer algebra and problems of motion stability. Mathematics and Computers in Simulation, 57(3–5):161–174, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003366.

Berkovich:2001:IOD

[288] Lev M. Berkovich. The integration of ordinary differential equations: factorization and transformations. Mathematics and Computers in Simulation, 57(3–5):175–195, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003378.

Ferro:2001:SNN

[289] Giuseppa Carra’ Ferro. Systems of nonalgebraic nonlinear ODEs. Mathematics and Computers in Simulation, 57(3–5):197–209, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540100338X.

Gutnik:2001:SNI

[290] Sergey A. Gutnik. Symbolic–numeric investigations for stability analysis of Lagrange systems. Mathematics and Computers in Simulation, 57(3–5):211– 215, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003391.

Elipe:2001:ELV

[291] Antonio Elipe. Extended Lissajous variables for oscillators in resonance. Mathematics and Computers in Simulation, 57(3–5):217–226, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003408.

Irtegov:2001:USM

[292] V. D. Irtegov and T. N. Titorenko. Using the system “Mathematica” in problems of mechanics. Mathematics and Computers in Simulation, 57(3–5): 227–237, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100341X.

Khrustalev:2001:CDP

[293] Oleg Khrustalev and Sergey Vernov. Construction of doubly periodic solutions via the Poincare–Lindstedt method in the case of massless φ^{4} theory. Mathematics and Computers in Simulation, 57(3–5):239–252, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003421.

Mikram:2001:NFM

[294] Jilali Mikram and Fouad Zinoun. Normal form methods for symbolic creation of approximate solutions of nonlinear dynamical systems. Mathematics and Computers in Simulation, 57(3–5): 253–289, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003433.

Palacian:2001:AAS

[295] Jesús Palacián and Patricia Yanguas. Analytical approach for simplifying dynamical systems of polynomial type. Mathematics and Computers in Simulation, 57(3–5):291– 305, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003445.

SanJuan:2001:CBP

[296] J. F. San Juan and Alberto Abad. Communications between the Poisson series processor PSPC and general scientific software. Mathematics and Computers in Simulation, 57(3–5):307– 315, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003457.

Severyanov:2001:AND

[297] Vasily M. Severyanov. Automata network dynamical systems for construction of fractal objects. Mathematics and Computers in Simulation, 57(3–5): 317–324, September 26, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003469.

Anonymous:2001:NIg

[298] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 57(3–5):325–326, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004153.

Anonymous:2001:ICa

[299] Anonymous. IMACS calendar. Mathematics and Computers in Simulation, 57(3–5):327–333, September 26, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004141.

El-Banna:2001:SLV

[300] Abou-Zaid H. El-Banna and Sana’a A. Zarea. Stability of linear vector optimization problems corresponding to an efficient set. Mathematics and Computers in Simulation, 57(6):335– 345, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002828.

Hsiao:2001:HWA

[301] Chun-Hui Hsiao and Wen-June Wang. Haar wavelet approach to nonlinear stiff systems. Mathematics and Computers in Simulation, 57(6):347– 353, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002750.

Li:2001:CFN

[302] Chunguang Li, Xiaofeng Liao, Zhongfu Wu, and Juebang Yu. Complex functional networks. Mathematics and Computers in Simulation, 57(6):355– 365, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401002968.

Mezghani:2001:RIL

[303] M. Mezghani, G. Roux, M. Cabassud, B. Dahhou, M. V. Le Lann, and G. Casamatta. Robust iterative learning control of an exothermic semibatch chemical reactor. Mathematics and Computers in Simulation, 57(6): 367–385, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003172.

Zhang:2001:ISF

[304] Jun Zhang and Jennifer J. Zhao. Iterative solution and finite difference approximations to 3D microscale heat transport equation. Mathematics and Computers in Simulation, 57(6):387– 404, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003196.

Anonymous:2001:NIh

[305] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 57(6):405–406, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401004220.

Anonymous:2001:CE

[306] Anonymous. calender of events. Mathematics and Computers in Simulation, 57(6):407–412, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401004232.

Anonymous:2001:AIVb

[307] Anonymous. Author index volume 57. Mathematics and Computers in Simulation, 57(6):413–415, December 3, 2001. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004244.

Anonymous:2001:PDa

[308] Anonymous. Pages 335–416 (3 December 2001). Mathematics and Computers in Simulation, 57(6):??, December 3, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ganzha:2001:SIR

[309] Victor G. Ganzha and Evgenii V. Vorozhtsov. Stability investigation of Runge–Kutta schemes with artificial dissipator on curvilinear grids for the Euler equations. Mathematics and Computers in Simulation, 58 (1):1–35, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003214.

Luoh:2001:SMC

[310] Leh Luoh and Wen-June Wang. A simple method for computing the entropy of the product of general fuzzy intervals. Mathematics and Computers in Simulation, 58(1):37– 49, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003184.

Yan:2001:NLT

[311] Xiangqiao Yan. Non-linear threedimensional finite element modeling of radial tires. Mathematics and Computers in Simulation, 58(1):51– 70, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003202.

Chan:2001:LMV

[312] S. K. Chan, V. Bharadwaj, and D. Ghose. Large matrix–vector products on distributed bus networks with communication delays using the divisible load paradigm: performance analysis and simulation. Mathematics and Computers in Simulation, 58(1): 71–92, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003299.

Anonymous:2001:NIi

[313] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 58(1):93–94, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401004256.

Anonymous:2001:ICb

[314] Anonymous. IMACS calendar. Mathematics and Computers in Simulation, 58(1):95–100, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401004268.

Anonymous:2001:PDb

[315] Anonymous. Pages 1–100 (19 December 2001). Mathematics and Computers in Simulation, 58(1):??, December 19, 2001. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Chen:2002:NNB

[316] Y. M. Chen and M. L. Lee. Neural networks-based scheme for system failure detection and diagnosis. Mathematics and Computers in Simulation, 58 (2):101–109, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003305.

Seyidmamedov:2002:FEA

[317] Zahir Seyidmamedov. Finite-element analysis of frictionless contact problem for a laminated medium. Mathematics and Computers in Simulation, 58 (2):111–123, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003317.

Li:2002:SEA

[318] Jichun Li and C. S. Chen. A simple efficient algorithm for interpolation between different grids in both 2D and 3D. Mathematics and Computers in Simulation, 58(2):125–132, January 7, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003482.

Agiza:2002:CDS

[319] H. N. Agiza, A. S. Hegazi, and A. A. Elsadany. Complex dynamics and synchronization of a duopoly game with bounded rationality. Mathematics and Computers in Simulation, 58 (2):133–146, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003470.

Jansen:2002:UCD

[320] H. Jansen and E. H. Twizell. An unconditionally convergent discretization of the SEIR model. Mathematics and Computers in Simulation, 58(2): 147–158, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003561.

Jaddu:2002:SMC

[321] Hussein Jaddu. Spectral method for constrained linear–quadratic optimal control. Mathematics and Computers in Simulation, 58(2):159–169, January 7, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003597.

Thoma:2002:BRW

[322] Jean U. Thoma. Book review: Wolfgang Borutzky, Bondgraphs, a Methodology for Modeling Interdisciplinary Dynamic Systems. Mathematics and Computers in Simulation, 58(2):171– 172, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100369X.

Anonymous:2002:NIa

[323] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 58(2):173–174, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401004359.

Anonymous:2002:ICa

[324] Anonymous. IMACS calender. Mathematics and Computers in Simulation, 58(2):175–182, January 7, 2002.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475401004360.

Anonymous:2002:PJa

[325] Anonymous. Pages 101–182 (7 January 2002). Mathematics and Computers in Simulation, 58(2):??, January 7, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Perez-Carretero:2002:LCA

[326] Cristina Pérez-Carretero, Luis M. Laita, Eugenio Roanes-Lozano, Luis Lázaro, Jesús González-Cajal, and Laura Laita. A logic and computer algebra-based expert system for diagnosis of anorexia. Mathematics and Computers in Simulation, 58(3):183– 202, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003706.

Roanes-Lozano:2002:CAA

[327] Eugenio Roanes-Lozano, Eugenio Roanes-Macıás, and Luis M. Laita. A computer algebra approach to the design of routes and the study of their compatibility in a railway interlocking. Mathematics and Computers in Simulation, 58(3):203–214, February 26, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003718.

Boutayeb:2002:ASM

[328] A. Boutayeb and M. Derouich. Age structured models for diabetes in East Morocco. Mathematics and Computers in Simulation, 58(3):215–229, February 26, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003688.

Elhajji:2002:ABP

[329] S. Elhajji and M. Errachid. Analysis of a bifurcation problem. Mathematics and Computers in Simulation, 58(3): 231–245, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003664.

Fatullayev:2002:NSI

[330] Afet Golayoglu Fatullayev. Numerical solution of the inverse problem of determining an unknown source term in a heat equation. Mathematics and Computers in Simulation, 58(3):247– 253, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003652.

Khanin:2002:NPA

[331] R. Khanin. On the Nipp polyhedron algorithm for solving singular perturbation problems. Mathematics and Computers in Simulation, 58(3):255– 272, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003548.

Anonymous:2002:NIb

[332] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 58(3):273–274, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475402000046.

Anonymous:2002:ICb

[333] Anonymous. IMACS calender. Mathematics and Computers in Simulation, 58(3):275–282, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475402000058.

Anonymous:2002:PF

[334] Anonymous. Pages 183–282 (26 February 2002). Mathematics and Computers in Simulation, 58(3):??, February 26, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mosekilde:2002:P

[335] Erik Mosekilde, Alexander L. Fradkov, and Ilya I. Blekhman. Preface. Mathematics and Computers in Simulation, 58(4–6):283–284, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540100372X.

Andrievsky:2002:ASM

[336] B. Andrievsky. Adaptive synchronization methods for signal transmission on chaotic carriers. Mathematics and Computers in Simulation, 58(4– 6):285–293, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003731.

Daafouz:2002:PQS

[337] Jamal Daafouz and Gilles Millerioux. Poly-quadratic stability and global chaos synchronization of discrete time hybrid systems. Mathematics and Computers in Simulation, 58(4– 6):295–307, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003743.

Sarasola:2002:CSD

[338] C. Sarasola, F. J. Torrealdea, A. d’Anjou, and M. Graña. Cost of synchronizing different chaotic systems. Mathematics and Computers in Simulation, 58(4– 6):309–327, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003755.

Zhusubaliyev:2002:ORC

[339] Zh. T. Zhusubaliyev and E. A. Soukhoterin. Oscillations in a relay control system with hysteresis and time dead zone. Mathematics and Computers in Simulation, 58(4–6):329– 350, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003767.

Sperling:2002:STP

[340] L. Sperling, B. Ryzhik, Ch. Linz, and H. Duckstein. Simulation of two-plane automatic balancing of a rigid rotor. Mathematics and Computers in Simulation, 58(4–6):351– 365, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003779.

Blekhman:2002:SSC

[341] I. I. Blekhman, A. L. Fradkov, O. P. Tomchina, and D. E. Bogdanov. Selfsynchronization and controlled synchronization: general definition and example design. Mathematics and Computers in Simulation, 58(4–6): 367–384, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003780.

Christiansen:2002:NCR

[342] L. E. Christiansen, T. Lehn-Schiøler, E. Mosekilde, P. Gránásy, and H. Matsushita. Nonlinear characteristics of randomly excited transonic flutter. Mathematics and Computers in Simulation, 58(4–6):385–405, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003792.

Rutkovsky:2002:CSA

[343] V. Yu. Rutkovsky, V. M. Sukhanov, V. M. Glumov, S. D. Zemlyakov, and S. D. Dodds. Computer simulation of an adaptive control system for a free-flying space robotic module with flexible payload’s transportation. Mathematics and Computers in Simulation, 58(4–6):407– 421, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003809.

Khentov:2002:PSI

[344] A. Khentov. On the principle of strong interaction for the resonant orbital motions of some celestial bodies. Mathematics and Computers in Simulation, 58(4–6):423– 434, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003810.

Kuznetsov:2002:NIA

[345] Sergey P. Kuznetsov. Noise-induced absolute instability. Mathematics and Computers in Simulation, 58(4–6): 435–442, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003822.

Rubchinsky:2002:PNO

[346] L. L. Rubchinsky, M. M. Sushchik, and G. V. Osipov. Patterns in networks of oscillators formed via synchronization and oscillator death. Mathematics and Computers in Simulation, 58(4– 6):443–467, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003834.

Anishchenko:2002:SEN

[347] V. S. Anishchenko, O. V. Sosnovtseva, A. S. Kopejkin, D. D. Matujshkin, and A. V. Klimshin. Synchronization effects in networks of stochastic bistable oscillators. Mathematics and Computers in Simulation, 58(4– 6):469–476, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003846.

Belykh:2002:SSL

[348] V. N. Belykh, I. V. Belykh, and K. V. Nelvidin. Spatiotemporal synchronization in lattices of locally coupled chaotic oscillators. Mathematics and Computers in Simulation, 58(4– 6):477–492, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003858.

Anonymous:2002:NIc

[349] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 58(4–6):493–494, March 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475402000071.

Anonymous:2002:ICEa

[350] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 58(4–6):495–502, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000083.

Anonymous:2002:AIVa

[351] Anonymous. Author index for volume 58. Mathematics and Computers in Simulation, 58(4–6):503– 505, March 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540200006X.

McAleer:2002:P

[352] Michael McAleer and Les Oxley. Preface. Mathematics and Computers in Simulation, 59(1–3):1–2, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003883.

Anonymous:2002:MI

[353] Anonymous. (MSSANZ/IMACS). Mathematics and Computers in Simulation, 59(1–3):1–278, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Chapman:2002:CME

[354] T. G. Chapman and R. W. Malone. Comparison of models for estimation of groundwater recharge, using data from a deep weighing lysimeter. Mathematics and Computers in Simulation, 59 (1–3):3–17, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003895.

Schreider:2002:PMD

[355] S. Yu Schreider, A. J. Jakeman, J. Gallant, and W. S. Merritt. Prediction of monthly discharge in ungauged catchments under agricultural land use in the Upper Ping basin, northern Thailand. Mathematics and Computers in Simulation, 59(1–3):19– 33, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003901.

Volker:2002:NMC

[356] Raymond E. Volker, Qi Zhang, and David A. Lockington. Numerical modelling of contaminant transport in coastal aquifers. Mathematics and Computers in Simulation, 59(1– 3):35–44, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003913.

Takahashi:2002:ERT

[357] Michiya Takahashi, Toshio Nakanishi, Isao Miyoshi, and Tomokazu Fujikura. An evaluation of the road traffic system simulator PIMTRACS by PIM. Mathematics and Computers in Simulation, 59(1–3):45–56, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003925.

Rizzoli:2002:STC

[358] Andrea E. Rizzoli, Nicoletta Fornara, and Luca Maria Gambardella. A simulation tool for combined rail/ road transport in intermodal terminals. Mathematics and Computers in Simulation, 59(1–3):57– 71, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003937.

Futatsuishi:2002:SMS

[359] Yoshihiro Futatsuishi, Ichie Watanabe, and Toshio Nakanishi. A study of the multi-stage flowshop scheduling problem with alternative operation assignments. Mathematics and Computers in Simulation, 59(1–3):73– 79, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401003949.

Shanmugam:2002:ANS

[360] Ramalingam Shanmugam, Alfred A. Bartolucci, and Karan P. Singh. The analysis of neurologic studies using an extended exponential model. Mathematics and Computers in Simulation, 59(1–3):81–85, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003950.

Naka:2002:SAE

[361] Takashi Naka and Naoto Sakamoto. Simulation analysis of the effects of the simultaneous release of quanta of acetylcholine on the endplate current at the neuromuscular junction. Mathematics and Computers in Simulation, 59(1–3):87–94, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003962.

Herbert:2002:SDR

[362] Ric D. Herbert and Peter J. Stemp. Solving the dynamics of a representative agent macro-economic model. Mathematics and Computers in Simulation, 59(1–3):95– 104, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003974.

Aiyoshi:2002:ACT

[363] Eitaro Aiyoshi, Ryota Horie, and Atsushi Maki. An application of the continuous time replicator dynamic to economics. Mathematics and Computers in Simulation, 59(1– 3):105–113, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003986.

Draeseke:2002:FLA

[364] Robert Draeseke and David E. A. Giles. A fuzzy logic approach to modelling the New Zealand underground economy. Mathematics and Computers in Simulation, 59(1–3): 115–123, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003998.

Zhu:2002:NZE

[365] Shengxiu Zhu and Les Oxley. New Zealand economic growth — endogenous or exogenous? Mathematics and Computers in Simulation, 59(1– 3):125–131, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004001.

Lim:2002:EGT

[366] Lee K. Lim and Michael McAleer. Economic growth and technological catching up by Singapore to the USA. Mathematics and Computers in Simulation, 59(1–3):133– 141, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004013.

Fare:2002:ERP

[367] Rolf Färe, Shawna Grosskopf, and Dimitri Margaritis. Economic reform and productivity growth: the case of Australia and New Zealand. Mathematics and Computers in Simulation, 59(1–3):143–152, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475401004025.

Tse:2002:MLE

[368] Y. K. Tse, V. V. Anh, and Q. Tieng. Maximum likelihood estimation of the fractional differencing parameter in an ARFIMA model using wavelets. Mathematics and Computers in Simulation, 59(1–3):153– 161, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004037.

Bewley:2002:FAC

[369] Ronald Bewley. Forecast accuracy, coefficient bias and Bayesian vector autoregressions. Mathematics and Computers in Simulation, 59(1–3): 163–169, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401004049.

Fukushige:2002:MSC

[370] Mototsugu Fukushige and Kosuke Oya. On the model selection criteria for demand system: Theil’s minimum entropy measure and its modification. Mathematics and Computers in Simulation, 59(1–3):171– 177, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401004050.

Gibson:2002:EEM

[371] John Gibson. The effect of endogeneity and measurement error bias on models of the risk of child stunting. Mathematics and Computers in Simulation, 59 (1–3):179–185, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401004062.

Goto:2002:PCD

[372] Ujo Goto and C. R. McKenzie. Price collusion and deregulation in the Japanese retail gasoline market. Mathematics and Computers in Simulation, 59(1–3):187–195, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004074.

Lim:2002:CAA

[373] Christine Lim and Michael McAleer. A cointegration analysis of annual tourism demand by Malaysia for Australia. Mathematics and Computers in Simulation, 59(1–3):197– 205, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004086.

Watkins:2002:CAM

[374] Clinton Watkins and Michael McAleer. Cointegration analysis of metals futures. Mathematics and Computers in Simulation, 59(1–3):207–221, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004098.

Chan:2002:DSS

[375] W. S. Chan and W. N. Liu. Diagnosing shocks in stock markets of southeast Asia, Australia, and New Zealand. Mathematics and Computers in Simulation, 59(1–3):223– 232, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004104.

Verhoeven:2002:NLM

[376] Peter Verhoeven, Berndt Pilgram, Michael McAleer, and Alistair Mees.

Non-linear modelling and forecasting of S&P 500 volatility. Mathematics and Computers in Simulation, 59(1– 3):233–241, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004116.

Herbert:2002:SEC

[377] Jan M. Herbert and Ric D. Herbert. Simulation of the effects of canyon geometry on thermal climate in city canyons. Mathematics and Computers in Simulation, 59(1–3): 243–253, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004128.

Klevmarken:2002:SIM

[378] N. Anders Klevmarken. Statistical inference in micro-simulation models: incorporating external information. Mathematics and Computers in Simulation, 59(1–3):255– 265, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540100413X.

Anonymous:2002:NId

[379] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 59(1–3):267–268, May 10, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475402000472.

Anonymous:2002:ICEb

[380] Anonymous. IMACS calender of events. Mathematics and Computers in Simulation, 59(1–3):269– 277, May 10, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000484.

Tzafestas:2002:RMP

[381] S. G. Tzafestas, M. P. Tzamtzi, and G. G. Rigatos. Robust motion planning and control of mobile robots for collision avoidance in terrains with moving objects. Mathematics and Computers in Simulation, 59 (4):279–292, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003512.

Lin:2002:HSG

[382] H. Lin and K. Yamashita. Hybrid simplex genetic algorithm for blind equalization using RBF networks. Mathematics and Computers in Simulation, 59(4):293–304, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401003640.

Ohta:2002:CNN

[383] Masaya Ohta. Chaotic neural networks with reinforced self-feedbacks and its application to N-Queen problem. Mathematics and Computers in Simulation, 59(4):305–317, June 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401003676.

Yuan:2002:HCG

[384] Xiaohui Yuan, Yanbin Yuan, and Yongchuan Zhang. A hybrid chaotic genetic algorithm for short-term hydro system scheduling. Mathematics and Computers in Simulation, 59 (4):319–327, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003639.

Luoh:2002:NAS

[385] Leh Luoh, Wen-June Wang, and Yi-Ke Liaw. New algorithms for solving fuzzy relation equations. Mathematics and Computers in Simulation, 59 (4):329–333, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475401003871.

Luoh:2002:NSA

[386] Leh Luoh. New stability analysis of T–S fuzzy system with robust approach. Mathematics and Computers in Simulation, 59(4):335– 340, June 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401004177.

vanGroesen:2002:MSS

[387] E. van Groesen and J. H. Westhuis. Modelling and simulation of surface water waves. Mathematics and Computers in Simulation, 59(4):

341–360, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004165.

Copetti:2002:ODT

[388] M. I. M. Copetti. A one-dimensional thermoelastic problem with unilateral constraint. Mathematics and Computers in Simulation, 59(4):361– 376, June 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004190.

Anonymous:2002:NIe

[389] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 59(4):377–378, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000587.

Anonymous:2002:ICc

[390] Anonymous. IMACS calendar. Mathematics and Computers in Simulation, 59(4):379–387, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000599.

Anonymous:2002:PJb

[391] Anonymous. Pages 279–388 (1 June 2002). Mathematics and Computers in Simulation, 59(4):??, June 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tang:2002:MDM

[392] Hui-Chin Tang. Modified decomposition method for multiple recursive random number generator. Mathematics and Computers in Simulation, 59 (5):453–458, June 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004281.

Hlavacek:2002:SGS

[393] Ivan Hlaváček and Jiří Nedoma. On a solution of a generalized semicoercive contact problem in thermoelasticity. Mathematics and Computers in Simulation, 60(1–2):1–17, July 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475401004335.

Ines:2002:LSF

[394] Bacha Inès. Local simplification of first-order systems of ordinary differential equations of dimension two. Mathematics and Computers in Simulation, 60(1–2):19–44, July 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540100444X.

Dumont:2002:VBB

[395] Yves Dumont. Vibrations of a beam between stops: numerical simulations and comparison of several numerical schemes. Mathematics and Computers in Simulation, 60(1–2):45– 83, July 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000034.

Gurov:2002:EBM

[396] T. V. Gurov and P. A. Whitlock. An efficient backward Monte Carlo estimator for solving of a quantum-kinetic equation with memory kernel. Mathematics and Computers in Simulation, 60(1–2):85–105, July 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004438.

Moghadas:2002:GST

[397] S. M. Moghadas and A. B. Gumel. Global stability of a two-stage epidemic model with generalized nonlinear incidence. Mathematics and Computers in Simulation, 60(1–2): 107–118, July 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000022.

Nedoma:2002:SSC

[398] Jiří Nedoma and Ivan Hlaváček. Solution of a semi-coercive contact problem in a non-linear thermo-elastic rheology. Mathematics and Computers in Simulation, 60(1–2):119– 127, July 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000332.

Normandin:2002:FES

[399] Magdeleine Normandin, Dana Grecov Radu, Ahmad Mahmoud, and Jean-Robert Clermont. Finite element and stream-tube formulations for flow computations — two-dimensional applications. Mathematics and Computers in Simulation, 60(1–2):129– 134, July 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475401004402.

Anonymous:2002:NIf

[400] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 60(1–2):135–136, July 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475402001313.

Anonymous:2002:ICEc

[401] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 60(1–2):137– 143, July 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001325.

Anonymous:2002:PJc

[402] Anonymous. Pages 1–144 (15 July 2002). Mathematics and Computers in Simulation, 60(1–2):??, July 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Tzafestas:2002:IFF

[403] Spyros G. Tzafestas and Elpida S. Tzafestas. Intelligent forecasting, fault diagnosis, scheduling and control. Mathematics and Computers in Simulation, 60(3–5):145–147, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000095.

Anonymous:2002:T

[404] Anonymous. (Tzafestas S.I.). Mathematics and Computers in Simulation, 60(3–5):145–422, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Karavezyris:2002:ASD

[405] Vassilios Karavezyris, Klaus-Peter Timpe, and Ruth Marzi. Application of system dynamics and fuzzy logic to forecasting of municipal solid waste. Mathematics and Computers in Simulation, 60(3–5):149–158, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000101.

Kaburlasos:2002:ICT

[406] V. G. Kaburlasos, V. Spais, V. Petridis, L. Petrou, S. Kazarlis, N. Maslaris, and A. Kallinakis. Intelligent clustering techniques for prediction of sugar production. Mathematics and Computers in Simulation, 60(3–5):159– 168, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000113.

Skoundrianos:2002:FDL

[407] E. N. Skoundrianos and S. G. Tzafestas. Fault diagnosis via local neural networks. Mathematics and Computers in Simulation, 60(3–5):169– 180, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000125.

Thomas:2002:FDI

[408] Philippe Thomas and Dimitri Lefebvre. Fault detection and isolation in non-linear systems by using oversized neural networks. Mathematics and Computers in Simulation, 60(3–5):181– 192, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000137.

Puig:2002:PRF

[409] Vicenç Puig and Joseba Quevedo. Passive robust fault detection using fuzzy parity equations. Mathematics and Computers in Simulation, 60(3–5):193– 207, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000149.

Mouchaweh:2002:RLR

[410] Moamar Sayed Mouchaweh, Arnaud Devillez, Gerard Villermain Lecolier, and Patrice Billaudel. Recursive learning in real time using fuzzy pattern matching. Mathematics and Computers in Simulation, 60(3–5):209– 216, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000150.

Marzi:2002:SFD

[411] R. Marzi and P. John. Supporting fault diagnosis through a multiagent-architecture. Mathematics and Computers in Simulation, 60(3–5):217– 224, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000162.

Lee:2002:ARS

[412] Seungkoo Lee and George Vachtsevanos. An application of rough set theory to defect detection of automotive glass. Mathematics and Computers in Simulation, 60(3–5):225– 231, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000174.

Nabeshima:2002:NRM

[413] K. Nabeshima, T. Suzudo, T. Ohno, and K. Kudo. Nuclear reactor monitoring with the combination of neural network and expert system. Mathematics and Computers in Simulation, 60(3–5): 233–244, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000186.

Kacem:2002:POA

[414] Imed Kacem, Slim Hammadi, and Pierre Borne. Pareto-optimality approach for flexible job-shop scheduling problems: hybridization of evolutionary algorithms and fuzzy logic. Mathematics and Computers in Simulation, 60(3–5):245–276, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000198.

Gargouri:2002:SSP

[415] E. Gargouri, S. Hammadi, and P. Borne. A study of scheduling problem in agro-food manufacturing systems. Mathematics and Computers in Simulation, 60(3–5):277– 291, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000204.

Digalakis:2002:MCA

[416] Jason G. Digalakis and Konstantinos G. Margaritis. A multipopulation cultural algorithm for the electrical generator scheduling problem. Mathematics and Computers in Simulation, 60(3–5):293–301, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000216.

Nikolakopoulos:2002:RIM

[417] G. Nikolakopoulos and A. Tzes. Reconfigurable internal model control based on adaptive lattice filtering. Mathematics and Computers in Simulation, 60(3–5):303–314, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000228.

Lazar:2002:NPC

[418] Mircea Lazar and Octavian Pastravanu. A neural predictive controller for non-linear systems. Mathematics and Computers in Simulation, 60(3–5): 315–324, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540200023X.

Koumboulis:2002:IAN

[419] F. N. Koumboulis and N. D. Kouvakas. Indirect adaptive neural control for precalcination in cement plants. Mathematics and Computers in Simulation, 60(3–5):325–334, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000241.

Lu:2002:CPL

[420] Guoping Lu and L. F. Yeung. H ∞-control problem for linear systems with multiple time-delays via dynamic output feedback. Mathematics and Computers in Simulation, 60(3–5):335– 345, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000253.

Monasterio:2002:LTD

[421] Iñaki Monasterio, Elena Lazkano, Iñaki Rañó, and Basilo Sierra. Learning to traverse doors using visual information. Mathematics and Computers in Simulation, 60(3–5):347– 356, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000277.

Kitsios:2002:CDD

[422] I. Kitsios and T. Pimenides. H _{∞} controller design for a distillation column using genetic algorithms. Mathematics and Computers in Simulation, 60(3–5): 357–367, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000289.

Spentzas:2002:DNL

[423] Konstantinos Spentzas and Stratis A. Kanarachos. Design of a non-linear hybrid car suspension system using neural networks. Mathematics and Computers in Simulation, 60(3–5):369– 378, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000290.

Rizzo:2002:IFF

[424] R. Rizzo and D. Iannuzzi. Indirect friction force identification for application in traction electric drives. Mathematics and Computers in Simulation, 60(3–5): 379–387, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000307.

Anagnostopoulos:2002:HPC

[425] C. Anagnostopoulos, I. Anagnostopoulos, D. Vergados, G. Kouzas, E. Kayafas, V. Loumos, and G. Stassinopoulos. High performance computing algorithms for textile quality control. Mathematics and Computers in Simulation, 60(3–5):389–400, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000319.

Messai:2002:ONN

[426] Nadhir Messai, Philippe Thomas, Dimitri Lefebvre, and Abdellah El Moudni. Optimal neural networks architectures for the flow–density relationships of traffic models. Mathematics and Computers in Simulation, 60(3–5):401– 409, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000320.

Anonymous:2002:NIg

[427] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 60(3–5):411–412, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540200157X.

Anonymous:2002:ICEd

[428] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 60(3–5):413–422, September 30, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001581.

Anonymous:2002:IFCa

[429] Anonymous. Inside front cover-Editorial board. Mathematics and Computers in Simulation, 60(3–5): ifc, September 30, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001556.

Bruno:2002:FPC

[430] N. Bruno, J. Heintz, G. Matera, and R. Wachenchauzer. Functional programming concepts and straightline programs in computer algebra. Mathematics and Computers in Simulation, 60(6):423–473, October 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402000356.

Smets:2002:OCB

[431] Ilse Y. Smets and Jan F. Van Impe. Optimal control of (bio-)chemical reactors: generic properties of time and space dependent optimization. Mathematics and Computers in Simulation, 60(6): 475–486, October 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000344.

Juncu:2002:PGM

[432] Gh. Juncu and C. Popa. Preconditioning by Gram matrix approximation for diffusion–convection– reaction equations with discontinuous coefficients. Mathematics and Computers in Simulation, 60(6):487– 506, October 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000630.

Kaya:2002:NCP

[433] Doğan Kaya and Asif Yokus. A numerical comparison of partial solutions in the decomposition method for linear and nonlinear partial differential equations. Mathematics and Computers in Simulation, 60(6):507– 512, October 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004384.

Anonymous:2002:NIh

[434] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 60(6):513–514, October 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475402001672.

Anonymous:2002:ICEe

[435] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 60(6):515–524, October 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001684.

Anonymous:2002:AIVb

[436] Anonymous. Author index for volume 60. Mathematics and Computers in Simulation, 60(6):525–528, October 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001696.

Anonymous:2002:IFCb

[437] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 60(6):ifc, October 15, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001611.

Anonymous:2002:PO

[438] Anonymous. Pages 423–528 (15 October 2002). Mathematics and Computers in Simulation, 60(6):??, October 15, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Liang:2002:SFD

[439] B. Liang, B. S. Payne, A. D. Ball, and S. D. Iwnicki. Simulation and fault detection of three-phase induction motors. Mathematics and Computers in Simulation, 61(1):1– 15, November 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000642.

Abdelhafez:2002:RMF

[440] H. M. Abdelhafez. Resonance of multiple frequency excitated systems with quadratic, cubic and quartic nonlinearity. Mathematics and Computers in Simulation, 61(1):17–34, November 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001428.

Garcia-Olivares:2002:AAT

[441] Antonio García-Olivares. Analytical approximants of time-dependent partial differential equations with tau methods. Mathematics and Computers in Simulation, 61(1):35–45, November 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001337.

Hasanov:2002:SIC

[442] Alemder Hasanov and Burhan Pektas. Simulation of ill-conditioned situations in inverse coefficient problem for the Sturm–Liouville operator based on boundary measurements. Mathematics and Computers in Simulation, 61 (1):47–52, November 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001349.

Lopez-Mellado:2002:ADE

[443] Ernesto López-Mellado. Analysis of discrete event systems by simulation of timed Petri net models. Mathematics and Computers in Simulation, 61 (1):53–59, November 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001374.

Menipaz:2002:HSM

[444] Ehud Menipaz and Avner Ben-Yair. Harmonization simulation model for managing several stochastic projects. Mathematics and Computers in Simulation, 61(1):61–66, November 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001350.

Anonymous:2002:NIi

[445] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 61(1):67–68, November 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475402002045.

Anonymous:2002:ICEf

[446] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 61(1):69–76, November 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002057.

Anonymous:2002:IFCc

[447] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 61(1):ifc, November 1, 2002. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001969.

Anonymous:2002:PN

[448] Anonymous. Pages 1–76 (1 November 2002). Mathematics and Computers in Simulation, 61(1):??, November 1, 2002. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sun:2003:SPR

[449] Lu Sun. Simulation of pavement roughness and IRI based on power spectral density. Mathematics and Computers in Simulation, 61(2):77– 88, January 1, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540100386X.

Dehghan:2003:DCF

[450] Mehdi Dehghan. Determination of a control function in three-dimensional parabolic equations. Mathematics and Computers in Simulation, 61(2): 89–100, January 1, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475401004347.

Lewis:2003:CTE

[451] Robert H. Lewis and Stephen Bridgett. Conic tangency equations and Apollonius problems in biochemistry and pharmacology. Mathematics and Computers in Simulation, 61(2):101– 114, January 1, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001222.

Sabelfeld:2003:SLM

[452] Karl Sabelfeld and Anastasia Kolodko. Stochastic Lagrangian models and algorithms for spatially inhomogeneous Smoluchowski equation. Mathematics and Computers in Simulation, 61(2):

115–137, January 1, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001416.

Botana:2003:STI

[453] F. Botana and J. L. Valcarce. A software tool for the investigation of plane loci. Mathematics and Computers in Simulation, 61(2):139–152, January 1, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001738.

Anonymous:2003:NIa

[454] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 61(2):153–154, January 1, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475402002136.

Anonymous:2003:ICEa

[455] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 61(2):155–160, January 1, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002148.

Anonymous:2003:IFCa

[456] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 61(2):ifc, January 1, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002094.

Anonymous:2003:PJa

[457] Anonymous. Pages 77–160 (1 January 2003). Mathematics and Computers in Simulation, 61(2):??, January 1, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Nedoma:2003:P

[458] Jiří Nedoma. Preface. Mathematics and Computers in Simulation, 61(3–6): 161–163, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001386.

Anonymous:2003:Ma

[459] Anonymous. (MODELLING 2001). Mathematics and Computers in Simulation, 61(3–6):161–624, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sanchez-Avila:2003:WDS

[460] C. Sánchez-Ávila. Wavelet domain signal deconvolution with singularitypreserving regularization. Mathematics and Computers in Simulation, 61(3–6): 165–176, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000733.

Vala:2003:TSL

[461] J. Vala. Two-scale limits in some non-linear problems of engineering mechanics. Mathematics and Computers in Simulation, 61(3–6):177– 185, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000745.

Vimmr:2003:MMC

[462] J. Vimmr. Mathematical modelling of compressible inviscid fluid flow through a sealing gap in the screw compressor. Mathematics and Computers in Simulation, 61(3–6):187– 197, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000757.

Voldrich:2003:NPE

[463] Josef Voldřich. Neumann problem for elliptic equation in Sobolev power weighted spaces. Mathematics and Computers in Simulation, 61(3–6):199– 207, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000769.

Warby:2003:FES

[464] M. K. Warby, J. R. Whiteman, W.- G. Jiang, P. Warwick, and T. Wright. Finite element simulation of thermoforming processes for polymer sheets. Mathematics and Computers in Simulation, 61(3–6):209–218, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000770.

Svacek:2003:NSP

[465] Petr Sváček and Karel Najzar. Numerical solution of problems with non-linear boundary conditions. Mathematics and Computers in Simulation, 61(3–6): 219–228, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000782.

Solin:2003:ENU

[466] Pavel Šolín and Karel Segeth. Examples of non-uniqueness of almostunidirectional gas flow. Mathematics and Computers in Simulation, 61(3–6): 229–237, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000794.

Dolezel:2003:FMC

[467] Ivo Doležel, Pavel Šolín, and Bohuš Ulrych. On a fieldless method for the computation of induction-generated heat in 3D non-ferromagnetic metal bodies. Mathematics and Computers in Simulation, 61(3–6):239–247, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000800.

Rosenberg:2003:MAB

[468] Josef Rosenberg and Robert Cimrman. Microcontinuum approach in biomechanical modeling. Mathematics and Computers in Simulation, 61(3–6): 249–260, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000812.

Rohan:2003:SSM

[469] Eduard Rohan. Sensitivity strategies in modelling heterogeneous media undergoing finite deformation. Mathematics and Computers in Simulation, 61(3–6): 261–270, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000824.

Nedoma:2003:NSSa

[470] Jiří Nedoma. Numerical solution of a Stefan-like problem in Bingham rheology. Mathematics and Computers in Simulation, 61(3–6):271– 281, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000836.

Nedoma:2003:NSSb

[471] J. Nedoma, Z. Klézl, J. Fousek, Z. Kestřánek, and J. Stehlík. Numerical simulation of some biomechanical problems. Mathematics and Computers in Simulation, 61(3–6):283– 295, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000848.

Hron:2003:NSG

[472] J. Hron, J. Málek, J. Nečas, and K. R. Rajagopal. Numerical simulations and global existence of solutions of two-dimensional flows of fluids with pressure- and shear-dependent viscosities. Mathematics and Computers in Simulation, 61(3–6):297– 315, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540200085X.

Felcman:2003:GRA

[473] Jiří Felcman. Grid refinement/ alignment in 3D flow computations. Mathematics and Computers in Simulation, 61(3–6):317–331, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000861.

Dolejsi:2003:SAD

[474] Vít Dolejší, Miloslav Feistauer, and Christoph Schwab. On some aspects of the discontinuous Galerkin finite element method for conservation laws. Mathematics and Computers in Simulation, 61(3–6):333– 346, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000873.

Dostal:2003:SFB

[475] Zdeněk Dostál and David Horák. Scalability and FETI based algorithm for large discretized variational inequalities. Mathematics and Computers in Simulation, 61(3–6):347– 357, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000885.

Danek:2003:DDM

[476] Josef Daněk. Domain decomposition method for contact problems with small range contact. Mathematics and Computers in Simulation, 61(3–6): 359–373, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000897.

Burda:2003:PEE

[477] Pavel Burda, Jaroslav Novotný, and Bedřich Sousedík. A posteriori error estimates applied to flow in a channel with corners. Mathematics and Computers in Simulation, 61(3–6): 375–383, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000903.

Brandts:2003:CTS

[478] Jan Brandts. Computing tall skinny solutions of AX −XB = C. Mathematics and Computers in Simulation, 61(3–6): 385–397, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402000915.

Bock:2003:USV

[479] Igor Bock and Ján Lovíšek. On unilaterally supported viscoelastic von Kármán plates with a long memory. Mathematics and Computers in Simulation, 61(3–6):399–407, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402000952.

Blaheta:2003:SDP

[480] R. Blaheta, P. Byczanski, O. Jakl, and J. Starý. Space decomposition preconditioners and their application in geomechanics. Mathematics and Computers in Simulation, 61(3–6):409– 420, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000964.

Axelsson:2003:INC

[481] Owe Axelsson. Iteration number for the conjugate gradient method. Mathematics and Computers in Simulation, 61(3–6):421–435, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475402000976.

Apel:2003:SDS

[482] Thomas Apel and H. Maharavo Randrianarivony. Stability of discretizations of the Stokes problem on anisotropic meshes. Mathematics and Computers in Simulation, 61(3–6):437– 447, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000988.

Ansorge:2003:TNP

[483] R. Ansorge. Towards numerical procedures for those technical phenomena whose mathematical models lead to non-standard conservation laws: a survey. Mathematics and Computers in Simulation, 61(3–6):449– 463, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540200099X.

Anguelov:2003:NFD

[484] Roumen Anguelov and Jean M.-S. Lubuma. Nonstandard finite difference method by nonlocal approximation. Mathematics and Computers in Simulation, 61(3–6):465– 475, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001064.

Core:2003:MLM

[485] Xavier Coré, Philippe Angot, and Jean-Claude Latché. A multilevel local mesh refinement projection method for low Mach number flows. Mathematics and Computers in Simulation, 61(3–6): 477–488, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001404.

Nechvatal:2003:TSC

[486] Luděk Nechvátal. On two-scale convergence. Mathematics and Computers in Simulation, 61(3–6):489– 495, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402000940.

Melnik:2003:NAD

[487] Roderick V. N. Melnik. Numerical analysis of dynamic characteristics of coupled piezoelectric systems in acoustic media. Mathematics and Computers in Simulation, 61(3–6):497– 507, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001003.

Lirkov:2003:MSE

[488] Ivan Lirkov. MPI solver for 3D elasticity problems. Mathematics and Computers in Simulation, 61(3–6):509– 516, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001040.

Liberda:2003:ISL

[489] Ondřej Liberda and Vladimír Janovský. Indication of a stability loss in the continuation of invariant subspaces. Mathematics and Computers in Simulation, 61(3–6):517–524, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001015.

Krizek:2003:NLG

[490] Michal Křížek and Jana Pradlová. On the nonexistence of a Lobachevsky geometry model of an isotropic and homogeneous universe. Mathematics and Computers in Simulation, 61(3–6): 525–535, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001027.

Kuhnen:2003:EED

[491] Pavel Krejčıánd Klaus Kuhnen. Error estimates for the discrete inversion of hysteresis and creep operators. Mathematics and Computers in Simulation, 61(3–6):537–548, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001039.

Kaasschieter:2003:MFE

[492] E. F. Kaasschieter, A. J. H. Frijns, and J. M. Huyghe. Mixed finite element modelling of cartilaginous tissues. Mathematics and Computers in Simulation, 61(3–6):549– 560, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001052.

Ingram:2003:DCC

[493] D. M. Ingram, D. M. Causon, and C. G. Mingham. Developments in Cartesian cut cell methods. Mathematics and Computers in Simulation, 61(3–6): 561–572, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001076.

Horak:2003:MMM

[494] M. Horák and J. Křen. Mathematical model of the male urinary tract. Mathematics and Computers in Simulation, 61(3–6):573–581, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001088.

Holecek:2003:SCA

[495] Miroslav Holeček, Olga Červená, and Fanny Poirier. Scale continuum approach in biomechanics: a simple simulation of a microstructural control of tissues’ stiffness. Mathematics and Computers in Simulation, 61(3–6): 583–590, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540200109X.

Francu:2003:HHE

[496] Jan Franců. Homogenization of heat equation with hysteresis. Mathematics and Computers in Simulation, 61(3–6): 591–597, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001106.

Bouchala:2003:RPL

[497] Jiří Bouchala. Resonance problems for p-Laplacian. Mathematics and Computers in Simulation, 61(3–6):599– 604, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001398.

Anonymous:2003:CPJ

[498] Anonymous. Commemoration for Professor Jindřich Nečas. Mathematics and Computers in Simulation, 61(3– 6):605, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000041.

Anonymous:2003:NIb

[499] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 61(3–6):607–608, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475402002653.

Anonymous:2003:ICEb

[500] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 61(3–6):609–619, January 30, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002665.

Anonymous:2003:AIVa

[501] Anonymous. Author index of Volume 61. Mathematics and Computers in Simulation, 61(3–6):621– 624, January 30, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402002677.

Taha:2003:F

[502] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 62 (1–2):1, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001945.

Chen:2003:EBD

[503] M. Chen. Equations for bidirectional waves over an uneven bottom. Mathematics and Computers in Simulation, 62(1–2):3–9, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001933.

Chin-Bing:2003:ACO

[504] S. A. Chin-Bing, A. Warn-Varnas, D. B. King, K. G. Lamb, M. Teixeira, and J. A. Hawkins. Analysis of coupled oceanographic and acoustic soliton simulations in the Yellow Sea: a search for soliton-induced resonances. Mathematics and Computers in Simulation, 62(1–2):11–20, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001921.

Garcia-Ripoll:2003:QLG

[505] Juan J. Garcıá-Ripoll, Vladimir V. Konotop, Boris Malomed, and Víctor M. Pérez-Garcıá. A quasi-local Gross– Pitaevskii equation for attractive Bose– Einstein condensates. Mathematics and Computers in Simulation, 62(1– 2):21–30, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001908.

Grimshaw:2003:SWC

[506] Roger Grimshaw and Gerard Iooss. Solitary waves of a coupled Korteweg– de Vries system. Mathematics and Computers in Simulation, 62(1–2): 31–40, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001891.

Guo:2003:PIS

[507] Jinhua Guo and Thiab R. Taha. Parallel implementation of the split-step and the pseudospectral methods for solving higher KdV equation. Mathematics and Computers in Simulation, 62(1–2): 41–51, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540200188X.

Holm:2003:RRD

[508] Darryl D. Holm. Rasetti–Regge Dirac bracket formulation of Lagrangian fluid dynamics of vortex filaments. Mathematics and Computers in Simulation, 62(1–2):53–63, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001878.

Kevrekidis:2003:BBP

[509] P. G. Kevrekidis and M. I. Weinstein. Breathers on a background: periodic and quasiperiodic solutions of extended discrete nonlinear wave systems. Mathematics and Computers in Simulation, 62(1–2):65–78, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001854.

Kevrekidis:2003:CDN

[510] P. G. Kevrekidis and V. V. Konotop. Compactons in discrete nonlinear Klein–Gordon models. Mathematics and Computers in Simulation, 62(1– 2):79–89, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001842.

Ludu:2003:NEW

[511] A. Ludu, R. F. O’Connell, and J. P. Draayer. Nonlinear equations and wavelets. Mathematics and Computers in Simulation, 62(1–2):91–99, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001830.

Malfliet:2003:TWS

[512] W. Malfliet. Travelling-wave solutions of coupled nonlinear evolution equations. Mathematics and Computers in Simulation, 62(1–2):101– 108, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001829.

Mallier:2003:MIB

[513] R. Mallier and M. Haslam. Modal interactions in a Bickley jet: comparison of theory with direct numerical simulation. Mathematics and Computers in Simulation, 62(1–2):109– 115, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001817.

Mickens:2003:NFD

[514] Ronald E. Mickens. A nonstandard finite difference scheme for the diffusionless Burgers equation with logistic reaction. Mathematics and Computers in Simulation, 62(1–2):117– 124, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402001805.

Spiteri:2003:NLE

[515] Raymond J. Spiteri and Steven J. Ruuth. Non-linear evolution using optimal fourth-order strong-stability-preserving Runge–Kutta methods. Mathematics and Computers in Simulation, 62(1–2):125–135, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402001799.

Salupere:2003:LTB

[516] A. Salupere, J. Engelbrecht, and P. Peterson. On the long-time behaviour of soliton ensembles. Mathematics and Computers in Simulation, 62(1–2):

137–147, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001787.

Torres-Silva:2003:CEO

[517] H. Torres-Silva and M. Zamorano. Chiral effects on optical solitons. Mathematics and Computers in Simulation, 62(1–2):149–161, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001775.

Veerakumar:2003:MKP

[518] V. Veerakumar and M. Daniel. Modified Kadomtsev–Petviashvili (MKP) equation and electromagnetic soliton. Mathematics and Computers in Simulation, 62(1–2):163–169, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001763.

Wazwaz:2003:CNS

[519] Abdul-Majid Wazwaz and Thiab Taha. Compact and noncompact structures in a class of nonlinearly dispersive equations. Mathematics and Computers in Simulation, 62(1–2):171– 189, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001751.

Zenchuk:2003:SNC

[520] Alexandre I. Zenchuk. Soliton with non-constant velocity. Mathematics and Computers in Simulation, 62(1–2): 191–201, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540200174X.

Anonymous:2003:NIc

[521] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 62(1–2):203–204, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540300017X.

Anonymous:2003:ICEc

[522] Anonymous. IMACS calendar of Events. Mathematics and Computers in Simulation, 62(1–2):205–215, February 13, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403000181.

Anonymous:2003:IFCb

[523] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 62(1–2): ifc, February 13, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000065.

Entacher:2003:E

[524] Karl Entacher, Wolfgang Ch. Schmid, and Andreas Uhl. Editorial. Mathematics and Computers in Simulation, 62 (3–6):217–218, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002197.

Anonymous:2003:Mb

[525] Anonymous. (MCM 2001). Mathematics and Computers in Simulation, 62 (3–6):217–572, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Heinrich:2003:MCQ

[526] Stefan Heinrich. From Monte Carlo to quantum computation. Mathematics and Computers in Simulation, 62(3– 6):219–230, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002392.

Jiang:2003:QRS

[527] Tao Jiang and Art B. Owen. Quasiregression with shrinkage. Mathematics and Computers in Simulation, 62 (3–6):231–241, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002537.

Schreiber:2003:HFB

[528] Michael Schreiber and Thomas Vojta. The Hartree–Fock based diagonalization — an efficient algorithm for the treatment of interacting electrons in disordered solids. Mathematics and Computers in Simulation, 62 (3–6):243–254, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402002331.

Sobol:2003:OME

[529] I. M. Sobol’ and D. I. Asotsky. One more experiment on estimating highdimensional integrals by quasi-Monte Carlo methods. Mathematics and Computers in Simulation, 62(3–6): 255–263, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402002288.

Wagner:2003:SAN

[530] Wolfgang Wagner. Stochastic, analytic and numerical aspects of coagulation processes. Mathematics and Computers in Simulation, 62(3–6): 265–275, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402002367.

Albrecher:2003:SMR

[531] Hansjörg Albrecher, Reinhold Kainhofer, and Robert F. Tichy. Simulation methods in ruin models with nonlinear dividend barriers. Mathematics and Computers in Simulation, 62(3– 6):277–287, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402002252.

Alexandrov:2003:PMC

[532] V. N. Alexandrov, I. T. Dimov, A. Karaivanova, and C. J. K. Tan. Parallel Monte Carlo algorithms for information retrieval. Mathematics and Computers in Simulation, 62(3– 6):289–295, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002525.

Atanassov:2003:NWM

[533] Emanouil Atanassov, Ivan Dimov, and Alain Dubus. A new weighted Monte Carlo algorithm for elastic electron backscattering from surfaces. Mathematics and Computers in Simulation, 62(3–6):297–305, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002549.

Bordone:2003:MCS

[534] P. Bordone, A. Bertoni, R. Brunetti, and C. Jacoboni. Monte Carlo simulation of quantum electron transport based on Wigner paths. Mathematics and Computers in Simulation, 62(3– 6):307–314, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002410.

Boyle:2003:ISM

[535] Phelim P. Boyle, Adam W. Kolkiewicz, and Ken Seng Tan. An improved simulation method for pricing high-dimensional American derivatives. Mathematics and Computers in Simulation, 62(3–6):315–322, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002483.

Bufler:2003:PST

[536] F. M. Bufler, A. Schenk, and W. Fichtner. Proof of a simple time-step propagation scheme for Monte Carlo simulation. Mathematics and Computers in Simulation, 62(3–6):323– 326, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002203.

Flahive:2003:QPD

[537] Mary E. Flahive. The quality parameter for digital ( t, m, s )- nets. Mathematics and Computers in Simulation, 62(3–6):327–333, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002409.

Hong:2003:DDS

[538] Hee Sun Hong, Fred J. Hickernell, and Gang Wei. The distribution of the discrepancy of scrambled digital ( t, m, s )-nets. Mathematics and Computers in Simulation, 62(3– 6):335–345, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475402002380.

Hwang:2003:FKP

[539] Chi-Ok Hwang, Michael Mascagni, and James A. Given. A Feynman–Kac path-integral implementation for Poisson’s equation using an h-conditioned Green’s function. Mathematics and Computers in Simulation, 62(3–6): 347–355, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002240.

Kalna:2003:NBT

[540] K. Kalna and A. Asenov. Nonequilibrium and ballistic transport, and backscattering in decanano HEMTs: a Monte Carlo simulation study. Mathematics and Computers in Simulation, 62(3–6):357–366, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475402002239.

Kosina:2003:EBT

[541] H. Kosina, M. Nedjalkov, and S. Selberherr. An event bias technique for Monte Carlo device simulation. Mathematics and Computers in Simulation, 62(3–6):367– 375, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002458.

Kroger:2003:TOF

[542] H. Kröger, X. Q. Luo, and K. J. M. Moriarty. Thermodynamical observables in a finite temperature window from the Monte Carlo Hamiltonian. Mathematics and Computers in Simulation, 62(3–6):377– 383, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002306.

Langtry:2003:MCM

[543] Tim Langtry, Lindsay Botten, Ara Asatryan, and Ross McPhedran. Monte Carlo modelling of imperfections in two-dimensional photonic crystals. Mathematics and Computers in Simulation, 62(3–6):385– 393, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002264.

LEcuyer:2003:CGC

[544] Pierre L’Ecuyer and Jacinthe Granger-Piché. Combined generators with components from different families. Mathematics and Computers in Simulation, 62(3–6):395–404, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002343. 3rd

IMACS Seminar on Monte Carlo Methods—MCM 2001 (Salzburg).

Leydold:2003:ACG

[545] Josef Leydold, Gerhard Derflinger, Günter Tirler, and Wolfgang Hörmann. An automatic code generator for nonuniform random variate generation. Mathematics and Computers in Simulation, 62(3–6):405–412, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002446.

Li:2003:NPA

[546] Yiming Li, Hsiao-Mei Lu, Ting-Wei Tang, and S. M. Sze. A novel parallel adaptive Monte Carlo method for nonlinear Poisson equation in semiconductor devices. Mathematics and Computers in Simulation, 62(3–6): 413–420, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002355.

Marseguerra:2003:MCS

[547] Marzio Marseguerra, Enrico Zio, Edoardo Patelli, Francesca Giacobbo, Giancarlo Ventura, and Giorgio Mingrone. Monte Carlo simulation of contaminant release from a radioactive waste deposit. Mathematics and Computers in Simulation, 62(3–6): 421–430, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002379.

Matsumoto:2003:SDT

[548] Makoto Matsumoto and Takuji Nishimura. Sum-discrepancy test on pseudorandom number generators. Mathematics and Computers in Simulation, 62(3– 6):431–442, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002276.

Mergenthaler:2003:ASA

[549] W. Mergenthaler, B. Mauersberg, J. Feller, L. J. Stuehler, W. T. O’Grady, and J. S. Ledford. Application of the simulated annealing local search technique to problems of redundancy elimination in functional and parametric tests of integrated circuits. Mathematics and Computers in Simulation, 62(3–6): 443–451, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002227.

Nedjalkov:2003:MCA

[550] M. Nedjalkov, H. Kosina, and S. Selberherr. Monte Carlo algorithms for stationary device simulations. Mathematics and Computers in Simulation, 62(3–6):453– 461, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540200246X.

Nekovee:2003:QMC

[551] Maziar Nekovee, W. Matthew C. Foulkes, and Richard J. Needs. Quantum Monte Carlo studies of density functional theory. Mathematics and Computers in Simulation, 62(3–6): 463–470, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002471.

Nilsson:2003:MCS

[552] Hans-Erik Nilsson, Ervin Dubaric, Mats Hjelm, and Urban Englund. Monte Carlo simulation of the transient response of single photon absorption in X-ray pixel detectors. Mathematics and Computers in Simulation, 62 (3–6):471–478, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002422.

Ninomiya:2003:NSS

[553] Syoiti Ninomiya. A new simulation scheme of diffusion processes: application of the Kusuoka approximation to finance problems. Mathematics and Computers in Simulation, 62(3– 6):479–486, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002513.

Ogawa:2003:QRW

[554] S. Ogawa and C. Lécot. A quasi-random walk method for onedimensional reaction–diffusion equations. Mathematics and Computers in Simulation, 62(3–6):487–494, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002434.

Pennetta:2003:MCS

[555] C. Pennetta, L. Reggiani, and E. Alfinito. Monte Carlo simulation of electromigration phenomena in metallic lines. Mathematics and Computers in Simulation, 62(3–6): 495–499, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002215.

Saraniti:2003:NCP

[556] M. Saraniti, J. Tang, S. M. Goodnick, and S. J. Wigger. Numerical challenges in particle-based approaches for the simulation of semiconductor devices. Mathematics and Computers in Simulation, 62(3–6): 501–508, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540200229X.

Schurer:2003:CBQ

[557] Rudolf Schürer. A comparison between (quasi-)Monte Carlo and cubature rule based methods for solving high-dimensional integration problems. Mathematics and Computers in Simulation, 62(3–6):509–517, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475402002501.

Strepp:2003:PTM

[558] W. Strepp, S. Sengupta, M. Lohrer, and P. Nielaba. Phase transitions in model colloids in reduced geometry. Mathematics and Computers in Simulation, 62(3–6):519– 527, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540200232X.

Sugita:2003:DRW

[559] Hiroshi Sugita. Dynamic random Weyl sampling for drastic reduction of randomness in Monte Carlo integration. Mathematics and Computers in Simulation, 62(3–6):529– 537, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002318.

Takemi:2003:RNE

[560] Tomoaki Takemi and Shigeyoshi Ogawa. Report on the numerical experiments of Haselgrove’s method applied to the numerical solution of PDEs. Mathematics and Computers in Simulation, 62(3–6):539– 552, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002495.

Anonymous:2003:NId

[561] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 62(3–6):553–554, March 3, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475403000259.

Anonymous:2003:ICEd

[562] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 62(3–6):555–565, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000260.

Anonymous:2003:AIVb

[563] Anonymous. Author index of volume 62. Mathematics and Computers in Simulation, 62(3–6):567– 571, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000272.

Anonymous:2003:IFCc

[564] Anonymous. Inside front cover — editorial Board. Mathematics and Computers in Simulation, 62(3–6): ifc, March 3, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540300020X.

Apostolikas:2003:LRB

[565] Giorgos Apostolikas and Spyros Tzafestas. On-line RBFNN based identification of rapidly time-varying nonlinear systems with optimal structureadaptation. Mathematics and Computers in Simulation, 63(1):1–13, April 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402001593.

Alonso-Quesada:2003:RAS

[566] S. Alonso-Quesada and M. de la Sen. Robust adaptive stabilizer of a class of time-varying plants using multiple controllers. Mathematics and Computers in Simulation, 63 (1):15–34, April 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402002070.

Wazwaz:2003:ASC

[567] Abdul-Majid Wazwaz. An analytic study of compactons structures in a class of nonlinear dispersive equations. Mathematics and Computers in Simulation, 63(1):35– 44, April 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475402002550.

El-Bassiouny:2003:TOI

[568] A. F. El-Bassiouny, M. M. Kamel, and A. Abdel-Khalik. Two-to-one internal resonances in nonlinear two degree of freedom system with parametric and external excitations. Mathematics and Computers in Simulation, 63(1):45–56, April 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475402001362.

Anonymous:2003:NIe

[569] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 63(1):57–58, April 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300048X.

Anonymous:2003:ICEe

[570] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 63(1):59–71, April 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000491.

Anonymous:2003:IFCd

[571] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 63(1):ifc, April 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403000405.

Anonymous:2003:PA

[572] Anonymous. Pages 1–72 (17 April 2003). Mathematics and Computers in Simulation, 63(1):??, April 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Levan:2003:WTS

[573] N. Levan and C. S. Kubrusly. A wavelet “time-shift-detail” decomposition. Mathematics and Computers in Simulation, 63(2):73–78, June 10, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403000375.

Spitaleri:2003:SCE

[574] Rosa Maria Spitaleri. A scientific computing environment for differential field simulation. Mathematics and Computers in Simulation, 63(2): 79–91, June 10, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000363.

Mascagni:2003:SEA

[575] Michael Mascagni and Chi-Ok Hwang. ϵ-Shell error analysis for “Walk On Spheres” algorithms. Mathematics and Computers in Simulation, 63(2): 93–104, June 10, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000387.

Golenko-Ginzburg:2003:RCS

[576] Dimitri Golenko-Ginzburg, Aharon Gonik, and Zohar Laslo. Resource constrained scheduling simulation model for alternative stochastic network projects. Mathematics and Computers in Simulation, 63(2): 105–117, June 10, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000508.

Anonymous:2003:NIf

[577] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 63(2):119–120, June 10, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000521.

Anonymous:2003:ICEf

[578] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 63(2):121–133, June 10, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000594.

Anonymous:2003:IFCe

[579] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 63(2): ifc, June 10, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403000545.

Anonymous:2003:PJb

[580] Anonymous. Pages 73–134 (10 June 2003). Mathematics and Computers in Simulation, 63(2):??, June 10, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Dessaint:2003:FMS

[581] Louis-A Dessaint and Kamal Al-Haddad. Foreword: Modelling and simulation of electric machines, converters and systems. Mathematics and Computers in Simulation, 63(3– 5):135, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001708.

Snider:2003:DBR

[582] L. A. Snider, H. T. Su, K. W. Chan, and Do Van Que. Development of a broadband real-time fully-digital simulator for the study and control of large power systems. Mathematics and Computers in Simulation, 63(3–5):137– 149, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000909.

Larose:2003:FDR

[583] C. Larose, S. Guerette, F. Guay, A. Nolet, T. Yamamoto, H. Enomoto, Y. Kono, Y. Hasegawa, and H. Taoka.

A fully digital real-time power system simulator based on PC-cluster. Mathematics and Computers in Simulation, 63(3–5):151–159, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000715.

Dufour:2003:ASP

[584] Christian Dufour, Jean Bélanger, and Simon Abourida. Accurate simulation of a 6-pulse inverter with real-time event compensation in ARTEMIS˘2122. Mathematics and Computers in Simulation, 63(3–5):161– 172, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000727.

Champagne:2003:RTS

[585] R. Champagne, L.-A. Dessaint, and H. Fortin-Blanchette. Real-time simulation of electric drives. Mathematics and Computers in Simulation, 63(3–5): 173–181, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300065X.

Ricci:2003:MSF

[586] Francesco Ricci and Hoang Le-Huy. Modeling and simulation of FPGAbased variable-speed drives using Simulink. Mathematics and Computers in Simulation, 63(3–5):183– 195, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000661.

Chiasson:2003:RTC

[587] John Chiasson, Leon Tolbert, Keith McKenzie, and Zhong Du. Real-time computer control of a multilevel converter using the mathematical theory of resultants. Mathematics and Computers in Simulation, 63(3–5):197– 208, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000673.

Figueroa:2003:SSM

[588] J. Figueroa, C. Brocart, J. Cros, and P. Viarouge. Simplified simulation methods for polyphase brushless DC motors. Mathematics and Computers in Simulation, 63(3–5):209– 224, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000685.

Gusia:2003:TTS

[589] S. Gusia, F. Labrique, D. Grenier, and H. Buyse. Two time scale global dynamical modelling of power electronic systems. Mathematics and Computers in Simulation, 63(3–5):225– 236, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000697.

Casoria:2003:HMM

[590] Silvano Casoria, Gilbert Sybille, and Patrice Brunelle. Hysteresis modeling in the MATLAB/Power System Blockset. Mathematics and Computers in Simulation, 63(3–5):237– 248, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000703.

Alejo:2003:MMC

[591] D. Alejo, P. Maussion, and J. Faucher. Multiple model control of a Buck dc/dc converter. Mathematics and Computers in Simulation, 63(3–5):249– 260, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000739.

Bouscayrol:2003:CSM

[592] A. Bouscayrol, B. Davat, B. de Fornel, B. François, J. P. Hautier, F. Meibody-Tabar, E. Monmasson, M. Pietrzak-David, H. Razik, E. Semail, and F. Benkhoris. Control structures for multi-machine multi-converter systems with upstream coupling. Mathematics and Computers in Simulation, 63(3–5): 261–270, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000740.

Leclercq:2003:CBF

[593] Ludovic Leclercq, Benoit Robyns, and Jean-Michel Grave. Control based on fuzzy logic of a flywheel energy storage system associated with wind and diesel generators. Mathematics and Computers in Simulation, 63(3–5):271– 280, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000752.

daSilva:2003:ACM

[594] J. Fernando A. da Silva, V. Fernão Pires, Sónia Ferreira Pinto, and J. Dionísio Barros. Advanced control methods for power electronics systems. Mathematics and Computers in Simulation, 63(3–5):281– 295, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000764.

Sturtzer:2003:RTU

[595] G. Sturtzer, D. Flieller, and J.-P. Louis. Reduction of torque undulation and extension of the Park’s transformation applied to non-sinusoidal saturated synchronous motors. Mathematics and Computers in Simulation, 63(3–5):297– 305, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000776.

Yacoubi:2003:NNC

[596] Loubna Yacoubi, Farhat Fnaiech, Louis-A. Dessaint, and Kamal Al-Haddad. New nonlinear control of three-phase NPC boost rectifier operating under severe disturbances. Mathematics and Computers in Simulation, 63(3–5):307–320, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000788.

Monmasson:2003:PCL

[597] E. Monmasson and J. P. Louis. Presentation of a control law for IM drive based on the dynamic reconfiguration of a DTC algorithm and a SVM-DTC algorithm. Mathematics and Computers in Simulation, 63(3–5):321– 333, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300079X.

Ruelland:2003:DEO

[598] R. Ruelland, G. Gateau, T. Meynard, and J. C. Hapiot. Digital emulator and observer of multicell converter. Mathematics and Computers in Simulation, 63(3–5):335– 347, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000806.

Malinowski:2003:RCS

[599] M. Malinowski, M. P. Kazmierkowski, and A. Trzynadlowski. Review and comparative study of control techniques for three-phase PWM rectifiers. Mathematics and Computers in Simulation, 63(3–5):349–361, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000818.

Rechka:2003:PEH

[600] Sanae Rechka, Éloi Ngandui, Jianhong Xu, and Pierre Sicard. Performance evaluation of harmonics detection methods applied to harmonics compensation in presence of common power quality problems. Mathematics and Computers in Simulation, 63(3–5): 363–375, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540300082X.

Devanneaux:2003:AMS

[601] V. Devanneaux, B. Dagues, J. Faucher, and G. Barakat. An accurate model of squirrel cage induction machines under stator faults. Mathematics and Computers in Simulation, 63(3–5):377– 391, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000831.

Picaud:2003:OSS

[602] V. Picaud, P. Hiebel, and J. M. Kauffmann. Optimization of SMES and superconducting magnets with a derivative free deterministic method. Mathematics and Computers in Simulation, 63(3–5):393–406, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403000843.

Olivier:2003:AMC

[603] Guy Olivier, Radu Cojocaru, and Anthony Lefèvre. Analytical model of a T-connected three-phase transformer. Mathematics and Computers in Simulation, 63(3–5):407–419, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000855.

Kanaan:2003:AEV

[604] Hadi Youssef Kanaan, Kamal Al-Haddad, and Gilles Roy. Analysis of the electromechanical vibrations in induction motor drives due to the imperfections of the mechanical transmission system. Mathematics and Computers in Simulation, 63(3–5):421– 433, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000867.

Kowalski:2003:NNA

[605] Czeslaw T. Kowalski and Teresa Orlowska-Kowalska. Neural networks application for induction motor faults diagnosis. Mathematics and Computers in Simulation, 63(3–5):435– 448, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000879.

Iannuzzi:2003:GOA

[606] D. Iannuzzi, E. Pagano, L. Piegari, and O. Veneri. Generator operations of asynchronous induction machines connected to ac or dc active/passive electrical networks. Mathematics and Computers in Simulation, 63(3–5):449– 459, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000880.

Lagace:2003:NVP

[607] P. J. Lagacé. Network voltage profile correction by discrete shunt compensation. Mathematics and Computers in Simulation, 63(3–5):461– 469, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000892.

Anonymous:2003:NIg

[608] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 63(3–5):471–472, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540300171X.

Anonymous:2003:ICEg

[609] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 63(3–5):473–481, November 17, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001721.

Anonymous:2003:IFCf

[610] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 63(3–5): ifc, November 17, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001617.

Kanarachos:2003:EAD

[611] A. Kanarachos, D. Koulocheris, and H. Vrazopoulos. Evolutionary algorithms with deterministic mutation operators used for the optimization of the trajectory of a fourbar mechanism. Mathematics and Computers in Simulation, 63(6):483– 492, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000612.

Chou:2003:LIO

[612] Jyh-Horng Chou, Jung-Hung Sun, and Jyh-Nan Shieh. On-line identification and optimal control of continuoustime systems. Mathematics and Computers in Simulation, 63(6):493–503, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540300051X.

Gallegati:2003:HTC

[613] M. Gallegati, L. Gardini, T. Puu, and I. Sushko. Hicks’ trade cycle revisited: cycles and bifurcations. Mathematics and Computers in Simulation, 63(6): 505–527, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000600.

Langemann:2003:DSE

[614] Dirk Langemann. A droplet in a stationary electric field. Mathematics and Computers in Simulation, 63(6):529– 539, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000624.

Foroni:2003:ASE

[615] Ilaria Foroni, Laura Gardini, and J. Barkley Rosser. Adaptive and statistical expectations in a renewable resource market. Mathematics and Computers in Simulation, 63(6):541– 567, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000636.

Sevastjanov:2003:FMM

[616] P. V. Sevastjanov and P. Róg. Fuzzy modeling of manufacturing and logistic systems. Mathematics and Computers in Simulation, 63(6):569– 585, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000648.

Dimov:2003:GFM

[617] I. T. Dimov and R. Y. Papancheva. Green’s function Monte Carlo algorithms for elliptic problems. Mathematics and Computers in Simulation, 63(6):587–604, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475403000946.

Hwang:2003:ACG

[618] Chi-Ok Hwang and Michael Mascagni. Analysis and comparison of Green’s function first-passage algorithms with “Walk on Spheres” algorithms. Mathematics and Computers in Simulation, 63(6):605–613, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475403000910.

Ianovsky:2003:ORT

[619] Edward Ianovsky and Joseph Kreimer. Optimization of real-time multiserver system with two different channels and shortage of maintenance facilities. Mathematics and Computers in Simulation, 63(6):615–627, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000922.

Ladde:2003:JPD

[620] G. S. Ladde and Bonita A. Lawrence. On joint probability density functions of discrete time iterative processes. Mathematics and Computers in Simulation, 63(6):629–650, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000934.

Sun:2003:FOC

[621] Haiwei Sun, Ning Kang, Jun Zhang, and Eric S. Carlson. A fourth-order compact difference scheme on face centered cubic grids with multigrid method for solving 2D convection diffusion equation. Mathematics and Computers in Simulation, 63(6):651– 661, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403000958.

Anonymous:2003:NIh

[622] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 63(6):663–664, November 24, 2003. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475403001782.

Anonymous:2003:ICEh

[623] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 63(6):665–673, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001794.

Anonymous:2003:AIVc

[624] Anonymous. Author index of volume 63. Mathematics and Computers in Simulation, 63(6):675–679, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001824.

Anonymous:2003:IFCg

[625] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 63(6):ifc, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001757.

Anonymous:2003:PN

[626] Anonymous. Pages 483–680 (24 November 2003). Mathematics and Computers in Simulation, 63(6):??, November 24, 2003. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

McAleer:2004:FSI

[627] Michael McAleer and Les Oxley. First special issue: Selected papers of the MSSANZ/IMACS 14th Biennial Conference on Modelling and Simulation, Canberra, Australia, December 2001. Mathematics and Computers in Simulation, 64(1):1– 2, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001150.

Dawes:2004:BMC

[628] W. R. Dawes, M. Gilfedder, G. R. Walker, and W. R. Evans. Biophysical modelling of catchment-scale surface water and groundwater response to land-use change. Mathematics and Computers in Simulation, 64 (1):3–12, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001162.

Giupponi:2004:MDC

[629] C. Giupponi, J. Mysiak, A. Fassio, and V. Cogan. MULINO-DSS: a computer tool for sustainable use of water resources at the catchment scale. Mathematics and Computers in Simulation, 64(1):13–24, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001174.

Hare:2004:FTT

[630] M. Hare and P. Deadman. Further towards a taxonomy of agentbased simulation models in environmental management. Mathematics and Computers in Simulation, 64(1): 25–40, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001186.

Greiner:2004:SFR

[631] R. Greiner. Systems framework for regional-scale integrated modelling and assessment. Mathematics and Computers in Simulation, 64(1):41–51, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001198.

Nishiyama:2004:MNA

[632] Y. Nishiyama. Minimum normal approximation error bandwidth selection for averaged derivatives. Mathematics and Computers in Simulation, 64 (1):53–61, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001204.

Wago:2004:BES

[633] Hajime Wago. Bayesian estimation of smooth transition GARCH model using Gibbs sampling. Mathematics and Computers in Simulation, 64(1): 63–78, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001216.

Sequeira:2004:VMC

[634] John M. Sequeira, Pang Chia Chiat, and Michael McAleer. Volatility models of currency futures in developed and emerging markets. Mathematics and Computers in Simulation, 64 (1):79–93, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001228.

Maekawa:2004:EBP

[635] Koichi Maekawa, Zonglu He, and Kianheng Tee. Estimating break points in a time series regression with structural changes. Mathematics and Computers in Simulation, 64 (1):95–101, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300123X.

Takeuchi:2004:SMD

[636] Yoshiyuki Takeuchi. On a statistical method to detect discontinuity in the distribution function of reported earnings. Mathematics and Computers in Simulation, 64(1):103– 111, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001241.

Tsui:2004:DCH

[637] Albert K. Tsui. Diagnostics for conditional heteroscedasticity models: some simulation results. Mathematics and Computers in Simulation, 64(1): 113–119, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001253.

Becker:2004:UDT

[638] R. Becker and A. S. Hurn. Using discrete-time techniques to test continuous-time models for nonlinearity in drift. Mathematics and Computers in Simulation, 64(1):121– 131, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001265.

Murasawa:2004:DFS

[639] Yasutomo Murasawa and Kimio Morimune. Distribution-free statistical inference for generalized Lorenz dominance based on grouped data. Mathematics and Computers in Simulation, 64(1):133–142, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001277.

Hoti:2004:EEI

[640] Suhejla Hoti. An empirical evaluation of international capital flows for developing countries. Mathematics and Computers in Simulation, 64(1): 143–160, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001289.

Fukiharu:2004:SHO

[641] Toshitaka Fukiharu. A simulation of the Heckscher–Ohlin theorem. Mathematics and Computers in Simulation, 64(1):161–168, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001290.

Chan:2004:MAV

[642] Felix Chan, Dora Marinova, and Michael McAleer. Modelling the asymmetric volatility of electronics patents in the USA. Mathematics and Computers in Simulation, 64(1): 169–184, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001307.

Choi:2004:MDM

[643] Daniel Choi and Les Oxley. Modelling the demand for money in New Zealand. Mathematics and Computers in Simulation, 64(1):185–191, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001319.

Hu:2004:IOS

[644] Baiding Hu and Michael McAleer. Input–output structure and growth in China. Mathematics and Computers in Simulation, 64(1):193–202, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001320.

Tsuji:2004:OPA

[645] Masatsugu Tsuji and Jong-Il Choe. An ordered probit analysis of factors promoting a regional information policy: the case of Japanese local governments. Mathematics and Computers in Simulation, 64(1):203– 212, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001332.

Anonymous:2004:NIa

[646] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 64(1):213–214, January 5, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475403001940.

Anonymous:2004:ICEa

[647] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 64(1):215–222, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001952.

Anonymous:2004:IFCa

[648] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 64(1):ifc, January 5, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540300185X.

Yuan:2004:CEM

[649] Chenggui Yuan and Xuerong Mao. Convergence of the Euler–Maruyama method for stochastic differential equations with Markovian switching. Mathematics and Computers in Simulation, 64(2):223–235, January 27, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001344.

Alrefaei:2004:SBS

[650] Mahmoud H. Alrefaei and Ameen J. Alawneh. Selecting the best stochastic system for large scale problems in DEDS. Mathematics and Computers in Simulation, 64(2):237–245, January 27, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540300154X.

Khater:2004:TWS

[651] A. H. Khater, W. Malfliet, and E. S. Kamel. Travelling wave solutions of some classes of nonlinear evolution equations in (1 + 1) and higher dimensions. Mathematics and Computers in Simulation, 64(2):247–258, January 27, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001563.

Clements:2004:MMM

[652] Richard R. Clements and Roger L. Hughes. Mathematical modelling of a mediaeval battle: the Battle of Agincourt, 1415. Mathematics and Computers in Simulation, 64(2):259– 269, January 27, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001551.

Carletti:2004:NSS

[653] M. Carletti, K. Burrage, and P. M. Burrage. Numerical simulation of stochastic ordinary differential equations in biomathematical modelling. Mathematics and Computers in Simulation, 64(2):271–277, January 27, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001575.

Serpen:2004:MST

[654] Gursel Serpen. Managing spatiotemporal complexity in Hopfield neural network simulations for large-scale static optimization. Mathematics and Computers in Simulation, 64(2):279– 293, January 27, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001587.

Anonymous:2004:NIb

[655] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 64(2):295–296, January 27, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475403002325.

Anonymous:2004:ICEb

[656] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 64(2):297–304, January 27, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002337.

Anonymous:2004:IFCb

[657] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 64(2):ifc, January 27, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002271.

Anonymous:2004:PJa

[658] Anonymous. Pages 223–304 (27 January 2004). Mathematics and Computers in Simulation, 64(2):??, January 27, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

McAleer:2004:SSI

[659] Michael McAleer and Les Oxley. Second special issue: Selected papers of the MSSANZ/IMACS 14th Biennial Conference on Modelling and Simulation, Canberra, Australia, December 2001. Mathematics and Computers in Simulation, 64(3–4):305– 306, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300096X.

Wade:2004:NMS

[660] A. J. Wade, P. G. Whitehead, H. P. Jarvie, C. Neal, H. Prior, and P. J. Johnes. Nutrient monitoring, simulation and management within a major lowland UK river system: the Kennet. Mathematics and Computers in Simulation, 64(3–4):307–317, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000971.

Hall:2004:AMS

[661] N. Hall, R. Greiner, and S. Yongvanit. Adapting modelling systems for salinity management of farms and catchments in Australia and Thailand. Mathematics and Computers in Simulation, 64(3–4):319–327, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403000983.

Barson:2004:MGG

[662] M. M. Barson, L. A. Randall, and S. C. Barry. Modelling greenhouse gas emissions from land cover change: linking continental data with point/ patch models. Mathematics and Computers in Simulation, 64(3–4):329– 337, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403000995.

Rahman:2004:TIT

[663] J. M. Rahman, S. M. Cuddy, and F. G. R. Watson. Tarsier and ICMS: two approaches to framework development. Mathematics and Computers in Simulation, 64(3–4):339– 350, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001009.

Verhoeven:2004:FTA

[664] Peter Verhoeven and Michael McAleer. Fat tails and asymmetry in financial volatility models. Mathematics and Computers in Simulation, 64(3–4):351– 361, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001010.

Bewley:2004:ISM

[665] Ronald Bewley, David Rees, and Paul Berg. The impact of stock market volatility on corporate bond credit spreads. Mathematics and Computers in Simulation, 64(3–4):363– 372, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001022.

Allen:2004:DUS

[666] D. E. Allen and W. Yang. Do UK stock prices deviate from fundamentals? Mathematics and Computers in Simulation, 64(3–4):373– 383, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001034.

Nawata:2004:EFL

[667] Kazumitsu Nawata. Estimation of the female labor supply models by Heckman’s two-step estimator and the maximum likelihood estimator. Mathematics and Computers in Simulation, 64(3–4):385– 392, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001046.

Bartolucci:2004:EBA

[668] Alfred A. Bartolucci, Karan P. Singh, and Ramalingham Shanmugam. Empirical Bayesian analysis of the Poisson intervention and incidence parameters. Mathematics and Computers in Simulation, 64(3–4):393– 399, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001058.

Breunig:2004:DMS

[669] Robert V. Breunig and Adrian R. Pagan. Do Markov-switching models capture nonlinearities in the data?: Tests using nonparametric methods. Mathematics and Computers in Simulation, 64(3–4):401– 407, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300106X.

Oya:2004:TRE

[670] Kosuke Oya. Test of random effects with incomplete panel data. Mathematics and Computers in Simulation, 64(3–4):409–419, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001071.

Chan:2004:MFA

[671] W. S. Chan, S. H. Cheung, and K. H. Wu. Multiple forecasts with autoregressive time series models: case studies. Mathematics and Computers in Simulation, 64(3–4):421–430, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001083.

Fatai:2004:MCR

[672] K. Fatai, Les Oxley, and F. G. Scrimgeour. Modelling the causal relationship between energy consumption and GDP in New Zealand, Australia, India, Indonesia, The Philippines and Thailand. Mathematics and Computers in Simulation, 64(3–4):431– 445, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001095.

Zhang:2004:AMI

[673] Zhaoyong Zhang, Kiyotaka Sato, and Michael McAleer. Asian monetary integration: a structural VAR approach. Mathematics and Computers in Simulation, 64(3–4):447–458, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001101.

Sakata:2004:AHC

[674] Kei Sakata and Colin McKenzie. The accumulation of human capital and the sectoral shifts hypothesis for different age groups. Mathematics and Computers in Simulation, 64(3–4):459– 465, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001113.

Herbert:2004:CMP

[675] Ric D. Herbert and Rod D. Bell. Constrained macroeconomic policy development with a separate predictive model. Mathematics and Computers in Simulation, 64(3–4):467– 476, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001125.

Lim:2004:MDK

[676] Christine Lim. The major determinants of Korean outbound travel to Australia. Mathematics and Computers in Simulation, 64(3–4):477– 485, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001137.

Bartels:2004:VUS

[677] R. Bartels, D. G. Fiebig, and A. McCabe. The value of using stated preference methods: a case study in modelling water heater choices. Mathematics and Computers in Simulation, 64(3–4):487–495, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001149.

Anonymous:2004:NIc

[678] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 64(3–4):497–498, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000138.

Anonymous:2004:ICEc

[679] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 64(3–4):499–506, February 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540400014X.

Anonymous:2004:IFCc

[680] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 64(3–4): ifc, February 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000059.

Teixeira:2004:SSP

[681] João Teixeira. Stable schemes for partial differential equations: the onedimensional reaction–diffusion equation. Mathematics and Computers in Simulation, 64(5):507–520, February 25, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001599.

Baker:2004:PSD

[682] Christopher T. H. Baker and Yihong Song. Periodic solutions of discrete Volterra equations. Mathematics and Computers in Simulation, 64(5):521– 542, February 25, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001733.

Leontitsis:2004:SSL

[683] Alexandros Leontitsis and Jenny Pange. Statistical significance of the LMS regression. Mathematics and Computers in Simulation, 64(5):543– 547, February 25, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001927.

Chen:2004:NES

[684] Yong Chen, Biao Li, and Hongqing Zhang. New exact solutions for modified nonlinear dispersive equations mK(m, n) in higher dimensions spaces. Mathematics and Computers in Simulation, 64(5):549–559, February 25, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001939.

Hsiao:2004:HWA

[685] C. H. Hsiao. Haar wavelet approach to linear stiff systems. Mathematics and Computers in Simulation, 64(5): 561–567, February 25, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002131.

Hsiao:2004:HWD

[686] Chun-Hui Hsiao. Haar wavelet direct method for solving variational problems. Mathematics and Computers in Simulation, 64(5):569–585, February 25, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002143.

Anonymous:2004:NId

[687] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 64(5):587–588, February 25, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404000436.

Anonymous:2004:ICEd

[688] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 64(5):589–596, February 25, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404000448.

Anonymous:2004:IFCd

[689] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 64(5):ifc, February 25, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404000382.

Anonymous:2004:PF

[690] Anonymous. Pages 507–596 (25 February 2004). Mathematics and Computers in Simulation, 64(5):??, February 25, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Marzban:2004:STV

[691] H. R. Marzban and M. Razzaghi. Solution of time-varying delay systems by hybrid functions. Mathematics and Computers in Simulation, 64(6): 597–607, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001836.

Alves:2004:ESV

[692] O. Alves, C. E. Ferreira, and F. P. Machado. Estimates for the spreading velocity of an epidemic model. Mathematics and Computers in Simulation, 64(6):609–616, March 10, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002209.

Blazic:2004:SMM

[693] Sašo Blažič, Drago Matko, and Gerhard Geiger. Simple model of a multibatch driven pipeline. Mathematics and Computers in Simulation, 64(6): 617–630, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002192.

Holmstrom:2004:FVV

[694] Mats Holmström. Fast visualization of volume emissions using conservative subdivision. Mathematics and Computers in Simulation, 64(6): 631–642, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002210.

Huang:2004:PSS

[695] Chih-Peng Huang and Yau-Tarng Juang. A projection scheme to stability analysis of discrete T–S fuzzy models. Mathematics and Computers in Simulation, 64(6):643–648, March 10, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002222.

Axelsson:2004:CAS

[696] O. Axelsson and J. Karátson. Conditioning analysis of separate displacement preconditioners for some nonlinear elasticity systems. Mathematics and Computers in Simulation, 64(6): 649–668, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403002234.

Sun:2004:ICN

[697] Jitao Sun. Impulsive control of a new chaotic system. Mathematics and Computers in Simulation, 64(6): 669–677, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403002246.

Triacca:2004:FCD

[698] Umberto Triacca. Feedback, causality and distance between ARMA models. Mathematics and Computers in Simulation, 64(6):679–685, March 10, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403002258.

Anonymous:2004:NIe

[699] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 64(6):687–688, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404000692.

Anonymous:2004:ICEe

[700] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 64(6):689–696, March 10, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000709.

Anonymous:2004:AIVa

[701] Anonymous. Author index of volume 64. Mathematics and Computers in Simulation, 64(6):697–701, March 10, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000710.

Anonymous:2004:IFCe

[702] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 64(6):ifc, March 10, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000643.

Anonymous:2004:PMa

[703] Anonymous. Pages 597–702 (10 March 2004). Mathematics and Computers in Simulation, 64(6):??, March 10, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Linker:2004:IMS

[704] Raphael Linker, Per Olof Gutman, and Joshua Dayan. Introduction to M^{2} SABI’01 special issue. Mathematics and Computers in Simulation, 65 (1–2):1–2, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001356.

Anonymous:2004:Ma

[705] Anonymous. (M2SABI’01). Mathematics and Computers in Simulation, 65 (1–2):1–210, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Dayan:2004:SCV

[706] J. Dayan, E. Dayan, Y. Strassberg, and E. Presnov. Simulation and control of ventilation rates in greenhouses. Mathematics and Computers in Simulation, 65(1–2):3–17, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001368.

Linker:2004:GTM

[707] Raphael Linker and Ido Seginer. Greenhouse temperature modeling: a comparison between sigmoid neural networks and hybrid models. Mathematics and Computers in Simulation, 65 (1–2):19–29, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300137X.

Bortolin:2004:MWE

[708] Gianantonio Bortolin, Stefan Borg, and Per Olof Gutman. Modeling of the wet end part of a paper mill with Dymola. Mathematics and Computers in Simulation, 65 (1–2):31–38, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001381.

Delen:2004:NLM

[709] R. Delen, L. Clijmans, J. Anthonis, and H. Ramon. A non-linear model to approximate the dynamics of a pulse width-modulated spray nozzle. Mathematics and Computers in Simulation, 65(1–2):39–48, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001393.

Maertens:2004:DVC

[710] K. Maertens and J. De Baerdemaeker. Design of a virtual combine harvester. Mathematics and Computers in Simulation, 65(1–2):49– 57, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540300140X.

DeKetelaere:2004:NCL

[711] Bart De Ketelaere, Koen Maertens, and Josse De Baerdemaeker. Noise cancellation in on-line acoustic impulse response measurements for the quality assessment of consumption eggs. Mathematics and Computers in Simulation, 65(1–2):59–67, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001411.

Kuchar:2004:UWG

[712] Leszek Kuchar. Using WGENK to generate synthetic daily weather data for modelling of agricultural processes. Mathematics and Computers in Simulation, 65(1–2):69– 75, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001423.

Moshou:2004:PSP

[713] Dimitrios Moshou, Koen Deprez, and Herman Ramon. Prediction of spreading processes using a supervised Self-Organizing Map. Mathematics and Computers in Simulation, 65(1–2):77– 85, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001435.

Pinheiro:2004:FDM

[714] A. C. Pinheiro and J. O. Peça. Forage drying models for oats and vetches under Mediterranean climate conditions. Mathematics and Computers in Simulation, 65(1–2):87– 100, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001447.

Dayan:2004:PCM

[715] E. Dayan, E. Presnov, and M. Fuchs. Prediction and calculation of morphological characteristics and distribution of assimilates in the ROS-GRO model. Mathematics and Computers in Simulation, 65(1–2):101– 116, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001459.

deGraaf:2004:TAG

[716] Stefan C. de Graaf, Johannes D. Stigter, and Gerrit van Straten. Test of ACW-gradient optimisation algorithm in computation of an optimal control policy for achieving acceptable nitrate concentration of greenhouse lettuce. Mathematics and Computers in Simulation, 65(1–2):117– 126, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001460.

Ioslovich:2004:DPS

[717] Ilya Ioslovich, Per-Olof Gutman, and Ido Seginer. Dominant parameter selection in the marginally identifiable case. Mathematics and Computers in Simulation, 65(1–2):127– 136, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001472.

Baati:2004:UMG

[718] L. Bâati, G. Roux, B. Dahhou, and J.-L. Uribelarrea. Unstructured modelling growth of Lactobacillus acidophilus as a function of the temperature. Mathematics and Computers in Simulation, 65(1–2):137– 145, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001484.

Sang:2004:MDF

[719] Byoung-In Sang, Katsutoshi Hori, and Hajime Unno. A mathematical description for the fungal degradation process of biodegradable plastics. Mathematics and Computers in Simulation, 65 (1–2):147–155, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001496.

Stencl:2004:MWS

[720] Jiri Stencl. Modelling the water sorption isotherms of yoghurt powder spray. Mathematics and Computers in Simulation, 65(1–2):157– 164, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475403001502.

Tijskens:2004:MMS

[721] E. Tijskens and J. De Baerdemaeker. Mathematical modelling of syneresis of cheese curd. Mathematics and Computers in Simulation, 65(1–2): 165–175, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475403001514.

Tijskens:2004:FEA

[722] E. Tijskens, D. Roose, H. Ramon, and J. De Baerdemaeker. FastDer++, efficient automatic differentiation for non-linear PDE solvers. Mathematics and Computers in Simulation, 65(1– 2):177–190, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403001526.

Berruto:2004:GBM

[723] R. Berruto, P. Gay, P. Piccarolo, and C. Tortia. Grey-box models for steam soil disinfestation simulation. Mathematics and Computers in Simulation, 65(1–2):191– 200, April 9, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001538.

Anonymous:2004:NIf

[724] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 65(1–2):201–202, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404000953.

Anonymous:2004:ICa

[725] Anonymous. IMACS calendar. Mathematics and Computers in Simulation, 65(1–2):203–209, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404000965.

Anonymous:2004:EB

[726] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 65(1–2):ifc, April 9, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000850.

Gospavic:2004:MLM

[727] Radovan Gospavic, Milesa Sreckovic, and Viktor Popov. Modelling of laser-material interaction using semianalytical approach. Mathematics and Computers in Simulation, 65(3): 211–219, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002362.

Kumar:2004:PCN

[728] B. V. Rathish Kumar and Bipin Kumar. Parallel computation of natural convection in trapezoidal porous enclosures. Mathematics and Computers in Simulation, 65(3):221–229, April 22, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002349.

Poschet:2004:SAM

[729] F. Poschet, K. Bernaerts, A. H. Geeraerd, N. Scheerlinck, B. M. Nicolaï, and J. F. Van Impe. Sensitivity analysis of microbial growth parameter distributions with respect to data quality and quantity by using Monte Carlo analysis. Mathematics and Computers in Simulation, 65(3): 231–243, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002350.

Zaitsev:2004:SLS

[730] D. A. Zaitsev. Switched LAN simulation by colored Petri nets. Mathematics and Computers in Simulation, 65 (3):245–249, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002374.

Liu:2004:SXRa

[731] Andong Liu. Simulation of X-ray propagation in a straight capillary. Mathematics and Computers in Simulation, 65(3):251–256, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000023.

Hoppe:2004:OSD

[732] Ronald H. W. Hoppe and Svetozara I. Petrova. Optimal shape design in biomimetics based on homogenization and adaptivity. Mathematics and Computers in Simulation, 65(3): 257–272, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000035.

Anagnostopoulos:2004:FDB

[733] Dimosthenis Anagnostopoulos, Vassilis Dalakas, and Mara Nikolaidou. A m-fold-decimation-based technique for model validation using a single system output data set. Mathematics and Computers in Simulation, 65(3): 273–288, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000175.

Anonymous:2004:NIg

[734] Anonymous. News of IMAC. Mathematics and Computers in Simulation, 65(3):289–290, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001041.

Anonymous:2004:ICEf

[735] Anonymous. IMSC calender of event. Mathematics and Computers in Simulation, 65(3):291–297, April 22, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001053.

Anonymous:2004:IFCf

[736] Anonymous. Inside front cover. Mathematics and Computers in Simulation, 65(3):ifc, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000989.

Anonymous:2004:PAa

[737] Anonymous. Pages 211–298 (22 April 2004). Mathematics and Computers in Simulation, 65(3):??, April 22, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Melnik:2004:WPP

[738] Roderick V. N. Melnik and Alex Povitsky. Wave phenomena in physics and engineering: new models, algorithms, and applications. Mathematics and Computers in Simulation, 65 (4–5):299–302, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000187.

Kechroud:2004:PTS

[739] Riyad Kechroud, Azzeddine Soulaimani, Yousef Saad, and Shivaraju Gowda. Preconditioning techniques for the solution of the Helmholtz equation by the finite element method. Mathematics and Computers in Simulation, 65(4– 5):303–321, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000199.

Loncaric:2004:OPP

[740] J. Lončarić and S. V. Tsynkov. Optimization of power in the problems of active control of sound. Mathematics and Computers in Simulation, 65(4– 5):323–335, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000205.

Voon:2004:HEP

[741] L. C. Lew Yan Voon and M. Willatzen. Helmholtz equation in parabolic rotational coordinates: application to wave problems in quantum mechanics and acoustics. Mathematics and Computers in Simulation, 65(4–5): 337–349, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000217.

Shen:2004:MCM

[742] Min Shen, Semion Saikin, Ming-C. Cheng, and Vladimir Privman. Monte Carlo modeling of spin FETs controlled by spin–orbit interaction. Mathematics and Computers in Simulation, 65 (4–5):351–363, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000229.

Lobao:2004:FGE

[743] Diomar Cesar Lobão and Alex Povitsky. Furnace geometry effects on plume dynamics in laser ablation for nanotube synthesis. Mathematics and Computers in Simulation, 65(4–5): 365–383, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000230.

Willatzen:2004:QCP

[744] M. Willatzen, R. V. N. Melnik, C. Galeriu, and L. C. Lew Yan Voon. Quantum confinement phenomena in nanowire superlattice structures. Mathematics and Computers in Simulation, 65(4–5):385–397, May 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404000242.

Hitt:2004:NSL

[745] Darren L. Hitt and Matthew Mc-Garry. Numerical simulations of laminar mixing surfaces in pulsatile microchannel flows. Mathematics and Computers in Simulation, 65(4–5): 399–416, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000254.

Dutta:2004:SRM

[746] Srabasti Dutta, James Glimm, John W. Grove, David H. Sharp, and Yongmin Zhang. Spherical Richtmyer– Meshkov instability for axisymmetric flow. Mathematics and Computers in Simulation, 65(4–5):417– 430, May 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000266.

Samulyak:2004:RMI

[747] Roman Samulyak and Yarema Prykarpatskyy. Richtmyer–Meshkov instability in liquid metal flows: influence of cavitation and magnetic fields. Mathematics and Computers in Simulation, 65 (4–5):431–446, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000278.

Povitsky:2004:EVP

[748] Alex Povitsky, Tinghui Zheng, and Georgios H. Vatistas. Effect of vortex profile on sound generation in a non-uniform flow. Mathematics and Computers in Simulation, 65(4– 5):447–468, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540400028X.

Jiang:2004:IWP

[749] Li Jiang, Chau-Lyan Chang, Meelan Choudhari, and Chaoqun Liu. Instability-wave propagation in boundarylayer flows at subsonic through hypersonic Mach numbers. Mathematics and Computers in Simulation, 65 (4–5):469–487, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000308.

Matus:2004:AFC

[750] P. Matus, R. V. N. Melnik, L. Wang, and I. Rybak. Applications of fully conservative schemes in nonlinear thermoelasticity: modelling shape memory materials. Mathematics and Computers in Simulation, 65(4–5):489– 509, May 11, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540400031X.

Kagan:2004:SDF

[751] L. Kagan, S. Minaev, and G. Sivashinsky. On self-drifting flame balls. Mathematics and Computers in Simulation, 65(4–5):511–520, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404000321.

Bounaim:2004:QCB

[752] A. Bounaïm, S. Holm, W. Chen, and Å. Ødegård. Quantification of the CARI breast imaging sensitivity by 2D/3D numerical time-domain ultrasound wave propagation. Mathematics and Computers in Simulation, 65 (4–5):521–534, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000333.

Hamdi:2004:ESI

[753] S. Hamdi, W. H. Enright, W. E. Schiesser, and J. J. Gottlieb. Exact solutions and invariants of motion for general types of regularized long wave equations. Mathematics and Computers in Simulation, 65(4– 5):535–545, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000345.

Anonymous:2004:NIh

[754] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 65(4–5):547–548, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404001132.

Anonymous:2004:ICb

[755] Anonymous. IMACS calendar. Mathematics and Computers in Simulation, 65(4–5):549–555, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404001144.

Anonymous:2004:Ia

[756] Anonymous. IFC. Mathematics and Computers in Simulation, 65 (4–5):ifc, May 11, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001089.

Alexandrov:2004:PRL

[757] V. N. Alexandrov, W. Owczarz, P. G. Thomson, and Z. Zlatev. Parallel runs of a large air pollution model on a grid of Sun computers. Mathematics and Computers in Simulation, 65 (6):557–577, May 17, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000473.

Popa:2004:KEA

[758] Constantin Popa and Rafal Zdunek. Kaczmarz extended algorithm for tomographic image reconstruction from limited-data. Mathematics and Computers in Simulation, 65(6):579–598, May 17, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540400045X.

Asharif:2004:NCA

[759] Mohammad Reza Asharif, Rui Chen, and Katsumi Yamashita. A new class of adaptive algorithm based on correlation functions for double-talk acoustic echo cancelling. Mathematics and Computers in Simulation, 65 (6):599–605, May 17, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000485.

Kubrusly:2004:SRS

[760] Carlos S. Kubrusly and Nhan Levan. Shift reducing subspaces and irreducibleinvariant subspaces generated by wandering vectors and applications. Mathematics and Computers in Simulation, 65(6):607–627, May 17, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000722.

Kosir:2004:NVE

[761] Andrej Košir, Aljo Mujčić, Nermin Suljanović, and Jurij F. Tasič. Noise variance estimation based on measured maximums of sampled subsets. Mathematics and Computers in Simulation, 65(6):629–639, May 17, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000734.

Anonymous:2004:NIi

[762] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 65(6):641–642, May 17, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001247.

Anonymous:2004:ICEg

[763] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 65(6):643–649, May 17, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001259.

Anonymous:2004:AIVb

[764] Anonymous. Author index of volume 65. Mathematics and Computers in Simulation, 65(6):651–655, May 17, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001302.

Anonymous:2004:Ib

[765] Anonymous. IFC. Mathematics and Computers in Simulation, 65(6): ifc, May 17, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001211.

Anonymous:2004:PMb

[766] Anonymous. Pages 557–656 (17 May 2004). Mathematics and Computers in Simulation, 65(6):??, May 17, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Samoilenko:2004:GAQ

[767] A. M. Samoilenko, Y. A. Prykarpatsky, Ufuk Taneri, A. K. Prykarpatsky, and D. L. Blackmore. A geometrical approach to quantum holonomic computing algorithms. Mathematics and Computers in Simulation, 66 (1):1–20, June 4, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000357.

Chen:2004:RSA

[768] Shinn-Horng Chen and Jyh-Horng Chou. Robust stability analysis for discrete-time LQG system under finite wordlength effects, noise uncertainties and time-varying structured parameter perturbations. Mathematics and Computers in Simulation, 66 (1):21–32, June 4, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000758.

Hojjati:2004:EAM

[769] G. Hojjati, M. Y. Rahimi Ardabili, and S. M. Hosseini. A-EBDF: an adaptive method for numerical solution of stiff systems of ODEs. Mathematics and Computers in Simulation, 66(1):33–41, June 4, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540400076X.

Gonzalez-Uriel:2004:KBS

[770] Ana González-Uriel and Eugenio Roanes-Lozano. A knowledgebased system for house layout selection. Mathematics and Computers in Simulation, 66(1):43–54, June 4, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000771.

Bashkirtseva:2004:SSC

[771] I. A. Bashkirtseva and L. B. Ryashko. Stochastic sensitivity of 3D-cycles. Mathematics and Computers in Simulation, 66(1):55–67, June 4, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404000783.

Abdelhafez:2004:RNF

[772] H. M. Abdelhafez. Resonance of a nonlinear forced system with two-frequency parametric and selfexcitations. Mathematics and Computers in Simulation, 66(1):69–83, June 4, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404000801.

Anonymous:2004:NIj

[773] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 66(1):85–86, June 4, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001429.

Anonymous:2004:ICEh

[774] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 66(1):87–93, June 4, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001430.

Anonymous:2004:IFCg

[775] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 66(1): ifc, June 4, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001375.

Anonymous:2004:PJb

[776] Anonymous. Pages 1–94 (4 June 2004). Mathematics and Computers in Simulation, 66(1):??, June 4, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Troch:2004:P

[777] I. Troch and F. Breitenecker. Preface. Mathematics and Computers in Simulation, 66(2–3):95–97, June 29, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403001988.

Anonymous:2004:Mb

[778] Anonymous. (4th MATHMOD). Mathematics and Computers in Simulation, 66(2–3):95–252, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Breedveld:2004:PBM

[779] Peter C. Breedveld. Port-based modeling of mechatronic systems. Mathematics and Computers in Simulation, 66 (2–3):99–128, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540300199X.

Kleemayr:2004:MSS

[780] K. Kleemayr. Modelling and simulation in snow science. Mathematics and Computers in Simulation, 66(2– 3):129–153, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002003.

Kheir:2004:EFE

[781] Naim A. Kheir, Mutasim A. Salman, and Niels J. Schouten. Emissions and fuel economy trade-off for hybrid vehicles using fuzzy logic. Mathematics and Computers in Simulation, 66(2– 3):155–172, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002015.

Dorninger:2004:SQM

[782] Dietmar Dorninger and Maciej Ma̧czyński. A simple quantum mechanical model for deriving the energy function of n-component systems. Mathematics and Computers in Simulation, 66(2– 3):173–179, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002027.

Mayer:2004:TOS

[783] E. Mayer and J. Raisch. Timeoptimal scheduling for high throughput screening processes using cyclic discrete event models. Mathematics and Computers in Simulation, 66(2– 3):181–191, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002039.

Reger:2004:MAF

[784] J. Reger and K. Schmidt. Modeling and analyzing finite state automata in the finite field F _{2}. Mathematics and Computers in Simulation, 66(2– 3):193–206, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002040.

Polyak:2004:ISI

[785] B. T. Polyak and S. A. Nazin. Interval solutions for interval algebraic equations. Mathematics and Computers in Simulation, 66(2–3):207– 217, June 29, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475403002052.

Heitzinger:2004:AST

[786] Clemens Heitzinger, Andreas Hössinger, and Siegfried Selberherr. An algorithm for smoothing three-dimensional Monte Carlo ion implantation simulation results. Mathematics and Computers in Simulation, 66(2–3): 219–230, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002064.

Jersak:2004:IBA

[787] M. Jersak, K. Richter, and R. Ernst. Interval-based analysis in embedded system design. Mathematics and Computers in Simulation, 66(2–3): 231–242, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475403002076.

Anonymous:2004:NIk

[788] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 66(2–3):243–244, June 29, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404001752.

Anonymous:2004:ICEi

[789] Anonymous. IMACS calender of events. Mathematics and Computers in Simulation, 66(2–3):245– 251, June 29, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001764.

Anonymous:2004:IFCh

[790] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 66(2–3): ifc, June 29, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001648.

Kirsch:2004:P

[791] Andreas Kirsch and Carlos Alves. Preface. Mathematics and Computers in Simulation, 66(4–5):253– 254, July 21, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404000497.

Kress:2004:IDP

[792] R. Kress. Inverse Dirichlet problem and conformal mapping. Mathematics and Computers in Simulation, 66 (4–5):255–265, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000503.

Grinberg:2004:FMO

[793] N. I. Grinberg and A. Kirsch. The factorization method for obstacles with a-priori separated sound-soft and sound-hard parts. Mathematics and Computers in Simulation, 66(4–5): 267–279, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000515.

Potthast:2004:SHA

[794] Roland Potthast. A set-handling approach for the no-response test and related methods. Mathematics and Computers in Simulation, 66(4–5): 281–295, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000527.

Luke:2004:MIO

[795] D. Russell Luke. Multifrequency inverse obstacle scattering: the point source method and generalized filtered backprojection. Mathematics and Computers in Simulation, 66(4– 5):297–314, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000539.

Hettlich:2004:RPG

[796] F. Hettlich. Reconstruction of periodic gratings from one scattered wave. Mathematics and Computers in Simulation, 66(4–5):315–324, July 21, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000540.

Cakoni:2004:TSD

[797] Fioralba Cakoni and David Colton. A target signature for distinguishing perfect conductors from anisotropic media of finite conductivity. Mathematics and Computers in Simulation, 66 (4–5):325–335, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000552.

Alves:2004:IFS

[798] C. J. S. Alves and P. Serranho. On the identification of the flatness of a sound-hard acoustic crack. Mathematics and Computers in Simulation, 66 (4–5):337–353, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000564.

Ha-Duong:2004:MFN

[799] T. Ha-Duong, M. Jaoua, and F. Menif. A modified frozen Newton method to identify a cavity by means of boundary measurements. Mathematics and Computers in Simulation, 66(4–5): 355–366, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000576.

Chaabane:2004:SRM

[800] S. Chaabane, C. Elhechmi, and M. Jaoua. A stable recovery method for the Robin inverse problem. Mathematics and Computers in Simulation, 66 (4–5):367–383, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000588.

Alves:2004:DPF

[801] C. J. S. Alves and A. L. Silvestre. On the determination of point-forces on a Stokes system. Mathematics and Computers in Simulation, 66(4– 5):385–397, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540400059X.

Kim:2004:DIO

[802] K. Y. Kim, B. S. Kim, M. C. Kim, and S. Kim. Dynamic inverse obstacle problems with electrical impedance tomography. Mathematics and Computers in Simulation, 66(4–5):399– 408, July 21, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404000606.

Rodrigues:2004:SES

[803] F. A. Rodrigues, H. R. B. Orlande, and G. S. Dulikravich. Simultaneous estimation of spatially dependent diffusion coefficient and source term in a nonlinear 1D diffusion problem. Mathematics and Computers in Simulation, 66 (4–5):409–424, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000618.

Chapko:2004:NSB

[804] Roman Chapko. On the numerical solution of a boundary value problem in the plane elasticity for a doubleconnected domain. Mathematics and Computers in Simulation, 66(4–5): 425–438, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540400062X.

Anonymous:2004:NIl

[805] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 66(4–5):439–440, July 21, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404002149.

Anonymous:2004:ICEj

[806] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 66(4–5):441– 447, July 21, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002150.

Anonymous:2004:IFCi

[807] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 66(4–5): ifc, July 21, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404002083.

Egli:2004:CMC

[808] R. Egli and N. F. Stewart. Chain models in computer simulation. Mathematics and Computers in Simulation, 66 (6):449–468, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000746.

Sun:2004:EPS

[809] Changyin Sun and Chun-Bo Feng. Exponential periodicity and stability of delayed neural networks. Mathematics and Computers in Simulation, 66(6): 469–478, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000795.

Al-Khaled:2004:CSS

[810] Kamel Al-Khaled and Fathi Allan. Construction of solutions for the shallow water equations by the decomposition method. Mathematics and Computers in Simulation, 66(6):479– 486, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404000813.

Barrett:2004:DDS

[811] Chris L. Barrett, William Y. C. Chen, and Michelle J. Zheng. Discrete dynamical systems on graphs and Boolean functions. Mathematics and Computers in Simulation, 66(6): 487–497, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000825.

Sun:2004:ICS

[812] Jitao Sun and Yinping Zhang. Impulsive control and synchronization of Chua’s oscillators. Mathematics and Computers in Simulation, 66(6): 499–508, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404000837.

Shaikhet:2004:CLF

[813] Leonid E. Shaikhet. Construction of Lyapunov functionals for stochastic difference equations with continuous time. Mathematics and Computers in Simulation, 66(6):509–521, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001065.

Chen:2004:SAF

[814] C. W. Chen, W. L. Chiang, and F. H. Hsiao. Stability analysis of T–S fuzzy models for nonlinear multiple time-delay interconnected systems. Mathematics and Computers in Simulation, 66(6):523–537, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001314.

Langemann:2004:MDE

[815] Dirk Langemann and Marcel Krüger. 3D model of a droplet in an electric field. Mathematics and Computers in Simulation, 66(6):539–549, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540400134X.

Tsang:2004:EPW

[816] S. C. Tsang and K. W. Chow. The evolution of periodic waves of the coupled nonlinear Schrödinger equations. Mathematics and Computers in Simulation, 66(6):551–564, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001326.

Wang:2004:POL

[817] Guanyu Wang. Parameter optimization in large-scale dynamical systems: a method of contractive mapping. Mathematics and Computers in Simulation, 66(6):565–576, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001351.

Liu:2004:SXRb

[818] Andong Liu and Yuzheng Lin. Simulation of X-ray transmission in capillaries with different profiles. Mathematics and Computers in Simulation, 66(6): 577–584, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001478.

Venutelli:2004:TSP

[819] Maurizio Venutelli. Time-stepping Padé–Petrov–Galerkin models for hydraulic jump simulation. Mathematics and Computers in Simulation, 66(6): 585–604, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540400148X.

Anonymous:2004:NIm

[820] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 66(6):605–606, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404002241.

Anonymous:2004:ICEk

[821] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 66(6):607–612, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002253.

Anonymous:2004:AIVc

[822] Anonymous. Author index of volume 66. Mathematics and Computers in Simulation, 66(6):613–616, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002265.

Anonymous:2004:IFCj

[823] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 66(6):ifc, August 12, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404002216.

Anonymous:2004:PAb

[824] Anonymous. Pages 449–616 (12 August 2004). Mathematics and Computers in Simulation, 66(6):??, August 12, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wester:2004:ACA

[825] Michael J. Wester. Applications of computer algebra in science, engineering, simulation and special software. Mathematics and Computers in Simulation, 67(1–2):1–2, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001491.

Botana:2004:ADE

[826] Francisco Botana and José L. Valcarce. Automatic determination of envelopes and other derived curves within a graphic environment. Mathematics and Computers in Simulation, 67(1– 2):3–13, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001508.

Akritas:2004:ASV

[827] Alkiviadis G. Akritas and Gennadi I. Malaschonok. Applications of singular-value decomposition (SVD). Mathematics and Computers in Simulation, 67(1–2):15–31, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540400151X.

Koukouvinos:2004:CMM

[828] C. Koukouvinos, E. Lappas, and M. Mitrouli. On the computation of maximum minors of Hadamard matrices. Mathematics and Computers in Simulation, 67(1–2):33–44, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001521.

Jakubi:2004:BPG

[829] Alejandro S. Jakubi. The branching problem in generalized power solutions to differential equations. Mathematics and Computers in Simulation, 67(1–2): 45–54, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001533.

Liska:2004:AOI

[830] Richard Liska, Mikhail Shashkov, and Victor Ganzha. Analysis and optimization of inner products for mimetic finite difference methods on a triangular grid. Mathematics and Computers in Simulation, 67(1–2):55–66, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001545.

Lazard:2004:IRR

[831] Daniel Lazard. Injectivity of real rational mappings: the case of a mixture of two Gaussian laws. Mathematics and Computers in Simulation, 67(1–2): 67–84, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001557.

Forehand:2004:LMC

[832] David I. M. Forehand, Raya Khanin, and Matthew P. Cartmell. A Lagrangian multibody code for deriving the symbolic state-space equations of motion for open-loop systems containing flexible beams. Mathematics and Computers in Simulation, 67(1–2): 85–98, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001569.

Irtegov:2004:CAI

[833] V. D. Irtegov and T. N. Titorenko. Computer algebra and investigation of invariant manifolds of mechanical systems. Mathematics and Computers in Simulation, 67(1–2):99–109, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001570.

Zehetleitner:2004:CAM

[834] Kurt Zehetleitner and Kurt Schlacher. Computer algebra methods for implicit dynamic systems and applications. Mathematics and Computers in Simulation, 67(1–2):111–123, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001582.

Matiyasevich:2004:EQA

[835] Yuri Matiyasevich. Elimination of quantifiers from arithmetical formulas defining recursively enumerable sets. Mathematics and Computers in Simulation, 67(1–2):125–133, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001594.

Sawada:2004:AGB

[836] Hiroyuki Sawada and Xiu-Tian Yan. Application of Gröbner bases and quantifier elimination for insightful engineering design. Mathematics and Computers in Simulation, 67(1–2):135– 148, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001600.

Ahmed:2004:SMP

[837] Maher Ahmed, Rabab Ward, and Nawwaf Kharma. Solving mathematical problems using knowledgebased systems. Mathematics and Computers in Simulation, 67(1–2):149– 161, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001612.

Roanes-Lozano:2004:ATS

[838] Eugenio Roanes-Lozano, Luis M. Laita, and Eugenio Roanes-Macıás. An accelerated-time simulation of departing passengers’ flow in airport terminals. Mathematics and Computers in Simulation, 67(1–2):163–172, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001624.

Anonymous:2004:NIn

[839] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 67(1–2):173–174, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404002460.

Anonymous:2004:ICEl

[840] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 67(1–2):175–180, September 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404002472.

Anonymous:2004:IFCk

[841] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 67(1–2): ifc, September 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002393.

Griffin:2004:QPS

[842] Byron L. Griffin and G. S. Ladde. Qualitative properties of stochastic iterative processes under random structural perturbations. Mathematics and Computers in Simulation, 67(3):181– 200, November 15, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001776.

Chen:2004:GJP

[843] Shyh-Wei Chen and Chung-Hua Shen. GARCH, jumps and permanent and transitory components of volatility: the case of the Taiwan exchange rate. Mathematics and Computers in Simulation, 67(3):201–216, November 15, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002046.

Dimov:2004:OSC

[844] I. Dimov, I. Faragó, Á. Havasi, and Z. Zlatev. Operator splitting and commutativity analysis in the Danish Eulerian Model. Mathematics and Computers in Simulation, 67(3):217– 233, November 15, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002198.

Lin:2004:TPR

[845] Yeong-Jyh Lin and Sheng-Jye Hwang. Temperature prediction of rolling tires by computer simulation. Mathematics and Computers in Simulation, 67(3): 235–249, November 15, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002319.

Shieh:2004:CSB

[846] J. S. Shieh, D. A. Linkens, and J. E. Peacock. A computer screenbased simulator for hierarchical fuzzy logic monitoring and control of depth of anaesthesia. Mathematics and Computers in Simulation, 67(3):251– 265, November 15, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002307.

Anonymous:2004:NIo

[847] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 67(3):267–268, November 15, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404002666.

Anonymous:2004:CIE

[848] Anonymous. Calendar of IMACS events. Mathematics and Computers in Simulation, 67(3):269–274, November 15, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404002678.

Anonymous:2004:IFCl

[849] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 67(3):ifc, November 15, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002630.

Anonymous:2004:PN

[850] Anonymous. Pages 181–274 (15 November 2004). Mathematics and Computers in Simulation, 67(3):??, November 15, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sultangazin:2004:P

[851] Umirzak M. Sultangazin. Preface. Mathematics and Computers in Simulation, 67(4–5):275–277, December 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404001788.

Sultangazin:2004:ISB

[852] Umirzak Sultangazin. Information systems based on space monitoring for solution of some problems of sustainable development. Mathematics and Computers in Simulation, 67(4–5):279– 290, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540400179X.

Chernyavsky:2004:SME

[853] G. M. Chernyavsky. Space monitoring of the environment and global safety. Mathematics and Computers in Simulation, 67(4–5):291–299, December 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001806.

Mika:2004:NMS

[854] Stanislav Mika and Marek Brandner. Numerical modelling of some two-phase fluid flow problems. Mathematics and Computers in Simulation, 67(4–5):301– 305, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001818.

Suimenbaev:2004:MSE

[855] B. T. Suimenbaev, I. I. Grachev, and V. V. Colovyev. The modeling system for ecological expertise of composite technogenic complexes. Mathematics and Computers in Simulation, 67(4–5): 307–315, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540400182X.

Veselov:2004:GMM

[856] V. V. Veselov, V. Yu. Panichkin, N. M. Zakharova, T. N. Vinnikova, and L. Yu. Trushel. Geoinformation and mathematical model of Eastern Priaralye. Mathematics and Computers in Simulation, 67(4–5):317– 325, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001831.

Yershin:2004:TTS

[857] Sh. A. Yershin and A. K. Yershina. Tornado, twisters — secondary currents within the atmosphere. Mathematics and Computers in Simulation, 67(4–5): 327–334, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001843.

Demberel:2004:IML

[858] S. Demberel, Nicholas N. Olenev, and Igor G. Pospelov. An interaction model for livestock farming and steppe ecosystem. Mathematics and Computers in Simulation, 67(4–5):335– 342, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001855.

Kenzheguzin:2004:MBF

[859] M. B. Kenzheguzin and B. K. Yessekina. Methodological basis of forecasting of sustainable development of economic system. Mathematics and Computers in Simulation, 67(4–5):343– 349, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001867.

Chang:2004:LSD

[860] C. P. Chang, T. Y. Chang, C. T. Wang, C. H. Kuo, and K. S. Chen. Land-surface deformation corresponding to seasonal ground-water fluctuation, determining by SAR interferometry in the SW Taiwan. Mathematics and Computers in Simulation, 67(4–5): 351–359, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001879.

Cheremissina:2004:GIE

[861] E. N. Cheremissina, L. E. Chesallov, and O. V. Mitrakova. The geoinformation environment for taking ecologically sound decisions on natural resources management. Mathematics and Computers in Simulation, 67(4–5): 361–364, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001880.

Spivak:2004:SMF

[862] L. Spivak, O. Arkhipkin, V. Pankratov, I. Vitkovskaya, and G. Sagatdinova. Space monitoring of floods in Kazakhstan. Mathematics and Computers in Simulation, 67(4–5):365– 370, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001892.

Ismailova:2004:GMW

[863] B. B. Ismailova. Geoinformation modeling of wind-induced surges on the northern–eastern Caspian Sea. Mathematics and Computers in Simulation, 67(4–5):371–377, December 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404001909.

Rafatov:2004:MSH

[864] Ramiz Rafatov. Method of spherical harmonic series in the problem of minimization of atmosphere pollution by fractions of harmful admixtures. Mathematics and Computers in Simulation, 67(4–5):379– 389, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001910.

Sultangazin:2004:NMR

[865] U. M. Sultangazin, A. H. Ahmedzhanov, and V. N. Glushko. Numerical methods of reconstruction of optical parameters of terrestrial surface and atmosphere using remote sensing. Mathematics and Computers in Simulation, 67(4–5): 391–397, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001922.

Zhantaev:2004:CSD

[866] Z. S. Zhantaev, J. K. Postoev, T. B. Stepanova, and V. S. Khachikyan. To the construction of spatial distribution of electron concentration over Kazakhstan, using the IRI model corrected by ionospheric data. Mathematics and Computers in Simulation, 67(4–5): 399–402, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001934.

Sultangazin:2004:PSD

[867] U. Sultangazin, A. Terekhov, and N. Muratova. Problems of satellite data calibration and thematic processing. Mathematics and Computers in Simulation, 67(4–5):403– 410, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001946.

Somsikov:2004:RCA

[868] V. M. Somsikov, K. E. Dungenbaeva, and B. Ganguly. Radiation changes in atmospheric waves dynamics and spectra. Mathematics and Computers in Simulation, 67(4–5):411– 417, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404001958.

Chaves:2004:OSR

[869] Paulo Chaves, Tsuneo Tsukatani, and Toshiharu Kojiri. Operation of storage reservoir for water quality by using optimization and artificial intelligence techniques. Mathematics and Computers in Simulation, 67(4–5):419– 432, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540400196X.

Drobzhev:2004:IDD

[870] V. I. Drobzhev, G. I. Gordienko, and S. N. Mukasheva. Ionosphere disturbances during rocket launches at Baikonur (Kazakhstan). Mathematics and Computers in Simulation, 67(4–5): 433–439, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001971.

Alexeyeva:2004:MMA

[871] L. A. Alexeyeva and G. K. Kaishybaeva. Mathematical models of array dynamics in the vicinity of extended underground structures at the action of moving loadings. Mathematics and Computers in Simulation, 67(4–5):441– 449, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001983.

Arystanbekova:2004:AGP

[872] N. Kh. Arystanbekova. Application of Gaussian plume models for air pollution simulation at instantaneous emissions. Mathematics and Computers in Simulation, 67(4–5):451– 458, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404001995.

Velikanov:2004:ANR

[873] A. E. Velikanov. About the nature of regional thermal anomaly in Semipalatinsk Test Site region. Mathematics and Computers in Simulation, 67(4–5): 459–465, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002009.

Sultangazin:2004:IVF

[874] A. U. Sultangazin and K. N. Umbetov. Investigation of viscous fluid enforced flow within the annulue between coaxial revolving cylinders. Mathematics and Computers in Simulation, 67(4–5): 467–472, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002010.

Brandner:2004:HRM

[875] Marek Brandner and Stanislav Mika. High-resolution methods for twocomponent fluid flow problem with moving interface. Mathematics and Computers in Simulation, 67(4–5):473– 483, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002022.

Aldyarov:2004:IED

[876] T. K. Aldyarov, U. K. Zhapbasbayev, and G. I. Ramazanova. Investigation of effect of depressant additives on hydraulic resistance and heat exchange in oil mixture flow. Mathematics and Computers in Simulation, 67(4–5):485– 492, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002034.

Anonymous:2004:NIp

[877] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 67(4–5):493–494, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404002836.

Anonymous:2004:ICEm

[878] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 67(4–5):495–500, December 3, 2004. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002848.

Anonymous:2004:IFCm

[879] Anonymous. Inside front cover — Editorial board. Mathematics and Computers in Simulation, 67(4–5): ifc, December 3, 2004. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002782.

Korn:2005:MRT

[880] Granino A. Korn. Model replication techniques for parameter-influence studies and Monte Carlo simulation with random parameters. Mathematics and Computers in Simulation, 67 (6):501–513, January 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002320.

Luh:2005:IMB

[881] Guan-Chun Luh and Wei-Chong Cheng. Immune model-based fault diagnosis. Mathematics and Computers in Simulation, 67(6):515–539, January 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002368.

Lopez-Mellado:2005:MBP

[882] Ernesto López-Mellado, Norma Villanueva-Paredes, and Hugo Almeyda-Canepa. Modelling of batch production systems using Petri nets with dynamic tokens. Mathematics and Computers in Simulation, 67(6):541–558, January 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540400237X.

Hlavacek:2005:RSU

[883] I. Hlaváček and J. Nedoma. Reliable solution of an unilateral contact problem with friction and uncertain data in thermo-elasticity. Mathematics and Computers in Simulation, 67 (6):559–580, January 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002538.

Muslu:2005:HOS

[884] G. M. Muslu and H. A. Erbay. Higherorder split-step Fourier schemes for the generalized nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 67(6):581–595, January 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540400254X.

Anonymous:2005:NIa

[885] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 67(6):597–598, January 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475404002939.

Anonymous:2005:ICEa

[886] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 67(6):599–604, January 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002940.

Anonymous:2005:AIVa

[887] Anonymous. Author index of volume 67. Mathematics and Computers in Simulation, 67(6):605–609, January 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002952.

Anonymous:2005:IFCa

[888] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 67(6):ifc, January 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002903.

Anonymous:2005:PJ

[889] Anonymous. Pages 501–610 (3 January 2005). Mathematics and Computers in Simulation, 67(6):??, January 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Castro:2005:SAC

[890] Julio Cesar Hernandez Castro, José María Sierra, Andre Seznec, Antonio Izquierdo, and Arturo Ribagorda. The strict avalanche criterion randomness test. Mathematics and Computers in Simulation, 68(1):1–7, February 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475404002551.

Yang:2005:SMS

[891] Taho Yang and Pohung Chou. Solving a multiresponse simulationoptimization problem with discrete variables using a multiple-attribute decision-making method. Mathematics and Computers in Simulation, 68 (1):9–21, February 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002563.

Venkiteswaran:2005:QMC

[892] G. Venkiteswaran and M. Junk. Quasi-Monte Carlo algorithms for diffusion equations in high dimensions. Mathematics and Computers in Simulation, 68(1):23–41, February 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404002587.

Venkiteswaran:2005:QAH

[893] G. Venkiteswaran and M. Junk. A QMC approach for high dimensional Fokker–Planck equations modelling polymeric liquids. Mathematics and Computers in Simulation, 68(1):43– 56, February 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002575.

Chuanwen:2005:HMC

[894] Jiang Chuanwen and Etorre Bompard. A hybrid method of chaotic particle swarm optimization and linear interior for reactive power optimisation. Mathematics and Computers in Simulation, 68(1):57–65, February 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002599.

Millerioux:2005:CAS

[895] G. Millerioux, F. Anstett, and G. Bloch. Considering the attractor structure of chaotic maps for observer-based synchronization problems. Mathematics and Computers in Simulation, 68(1):67–85, February 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002605.

Anonymous:2005:NIb

[896] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 68(1):87–88, February 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847540500008X.

Anonymous:2005:ICEb

[897] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 68(1):89–94, February 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000091.

Anonymous:2005:IFCb

[898] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 68(1):ifc, February 3, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000030.

Anonymous:2005:PF

[899] Anonymous. Pages 1–94 (3 February 2005). Mathematics and Computers in Simulation, 68(1):??, February 3, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kumar:2005:APS

[900] Rajeeva Kumar, Pierre T. Kabamba, and David C. Hyland. Analysis and parameter selection for an adaptive random search algorithm. Mathematics and Computers in Simulation, 68 (2):95–103, April 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404002617.

Papaodysseus:2005:UFR

[901] C. Papaodysseus, G. Roussopoulos, and A. Panagopoulos. Using a fast RLS adaptive algorithm for efficient speech processing. Mathematics and Computers in Simulation, 68(2):105– 113, April 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002885.

Alrefaei:2005:SQR

[902] Mahmoud H. Alrefaei and Ameen J. Alawneh. Solution quality of random search methods for discrete stochastic optimization. Mathematics and Computers in Simulation, 68(2):115– 125, April 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002745.

Lepik:2005:NSD

[903] Ü. Lepik. Numerical solution of differential equations using Haar wavelets. Mathematics and Computers in Simulation, 68(2):127–143, April 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475404002757.

Patsis:2005:MCS

[904] G. P. Patsis. Monte Carlo study of surface and line-width roughness of resist film surfaces during dissolution. Mathematics and Computers in Simulation, 68(2):145–156, April 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404002769.

Langemann:2005:MDM

[905] Dirk Langemann. Modelling a droplet moving in an electric field. Mathematics and Computers in Simulation, 68 (2):157–169, April 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404003015.

Mei:2005:DCL

[906] Hua Mei, Shaoyuan Li, Wen-Jian Cai, and Qiang Xiong. Decentralized closed-loop parameter identification for multivariable processes from step responses. Mathematics and Computers in Simulation, 68(2): 171–192, April 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404003027.

Anonymous:2005:NIc

[907] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 68(2):193–194, April 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000728.

Anonymous:2005:ICEc

[908] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 68(2):195–200, April 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540500073X.

Anonymous:2005:IFCc

[909] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 68(2):CO2, April 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000698.

Anonymous:2005:PA

[910] Anonymous. Pages 95–200 (20 April 2005). Mathematics and Computers in Simulation, 68(2):??, April 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Agliari:2005:SGB

[911] Anna Agliari, Laura Gardini, and Tönu Puu. Some global bifurcations related to the appearance of closed invariant curves. Mathematics and Computers in Simulation, 68 (3):201–219, May 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404003040.

Feng:2005:PIM

[912] Lihong Feng. Parameter independent model order reduction. Mathematics and Computers in Simulation, 68 (3):221–234, May 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475404003039.

Kleinert:2005:MSO

[913] T. Kleinert and J. Lunze. Modelling and state observation of Simulated Moving Bed processes based on an explicit functional wave form description. Mathematics and Computers in Simulation, 68(3):235– 270, May 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404003052.

Danek:2005:DDG

[914] J. Daněk, I. Hlaváček, and J. Nedoma. Domain decomposition for generalized unilateral semi-coercive contact problem with given friction in elasticity. Mathematics and Computers in Simulation, 68(3):271– 300, May 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404003076.

Anonymous:2005:NId

[915] Anonymous. New of IMACS. Mathematics and Computers in Simulation, 68(3):301–302, May 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000819.

Anonymous:2005:ICEd

[916] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 68(3):303– 308, May 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000820.

Anonymous:2005:IFCd

[917] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 68 (3):CO2, May 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000789.

Anonymous:2005:PMa

[918] Anonymous. Pages 201–308 (1 May 2005). Mathematics and Computers in Simulation, 68(3):??, May 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Ho:2005:SCS

[919] Wen-Hsien Ho and Jyh-Horng Chou. Shifted-Chebyshev series solutions of Takagi–Sugeno fuzzy-model-based dynamic equations. Mathematics and Computers in Simulation, 68(4):309– 316, May 16, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475404003088.

Jin:2005:SBW

[920] Yongdong Jin, Jiuping Xu, Wenhua Zhang, Jiuli Luo, and Qiwang Xu. Simulation of biological waves in single-species bacillus system governed by birth and death-diffusion dynamical equation. Mathematics and Computers in Simulation, 68(4): 317–327, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475404003064.

Yang:2005:EPD

[921] Jianguo Yang, Qingfeng Li, Zhaoan Wang, and Jinmei Wang. Estimating pedestrian delays at signalized intersections in developing cities by Monte Carlo method. Mathematics and Computers in Simulation, 68(4):

329–337, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000509.

Yu:2005:OHS

[922] Run Yu and PingSun Leung. Optimal harvesting strategies for a multi-cycle and multi-pond shrimp operation: a practical network model. Mathematics and Computers in Simulation, 68 (4):339–354, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000492.

Sheng:2005:NST

[923] Q. Sheng, A. Q. M. Khaliq, and D. A. Voss. Numerical simulation of twodimensional sine-Gordon solitons via a split cosine scheme. Mathematics and Computers in Simulation, 68 (4):355–373, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000480.

Karaa:2005:NSB

[924] Samir Karaa, Jun Zhang, and Fuqian Yang. A numerical study of a 3D bioheat transfer problem with different spatial heating. Mathematics and Computers in Simulation, 68(4): 375–388, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000674.

Anonymous:2005:NIe

[925] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 68(4):389–390, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001138.

Anonymous:2005:ICEe

[926] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 68(4):391–396, May 16, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540500114X.

Anonymous:2005:IFCe

[927] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 68 (4):CO2, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001102.

Anonymous:2005:PMb

[928] Anonymous. Pages 309–396 (16 May 2005). Mathematics and Computers in Simulation, 68(4):??, May 16, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

McAleer:2005:FSI

[929] Michael McAleer, Les Oxley, and David Post. First special issue: selected papers of the MSSANZ/IMACS 15th Biennial Conference on modelling and simulation, Townsville, Australia, July 2003. Mathematics and Computers in Simulation, 68(5–6): 397–399, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000315.

Anonymous:2005:MIBa

[930] Anonymous. MSSANZ/IMACS 15th Biennial Conference on Modelling and Simulation. Mathematics and Computers in Simulation, 68(5–6):397–592, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Phillips:2005:CTT

[931] Peter C. B. Phillips. Challenges of trending time series econometrics. Mathematics and Computers in Simulation, 68(5–6):401–416, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000418.

Ann:2005:MIC

[932] Wong Kie Ann, John M. Sequeira, and Michael McAleer. Modelling the information content in insider trades in the Singapore exchange. Mathematics and Computers in Simulation, 68 (5–6):417–428, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000443.

Chan:2005:BTT

[933] W. S. Chan and S. H. Cheung. A bivariate threshold time series model for analyzing Australian interest rates. Mathematics and Computers in Simulation, 68(5–6):429– 437, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000339.

Fukiharu:2005:GEA

[934] T. Fukiharu. General equilibrium analysis on arms exports to developing countries in conflict. Mathematics and Computers in Simulation, 68(5– 6):439–448, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000364.

Hoti:2005:CAR

[935] Suhejla Hoti. Comparative analysis of risk ratings for the East European region. Mathematics and Computers in Simulation, 68(5–6): 449–462, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000455.

How:2005:IAS

[936] Janice C. Y. How, Peter Verhoeven, and Caro X. Huang. Information asymmetry surrounding earnings and dividend announcements: an intra-day analysis. Mathematics and Computers in Simulation, 68(5–6): 463–473, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000467.

Hu:2005:ECA

[937] Baiding Hu and Michael McAleer. Estimation of Chinese agricultural production efficiencies with panel data. Mathematics and Computers in Simulation, 68(5–6):474–483, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000327.

Trinh:2005:MSH

[938] Le Thi Van Trinh, John Gibson, and Les Oxley. Measuring the stock of human capital in New Zealand. Mathematics and Computers in Simulation, 68(5–6):484– 497, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000376.

Lim:2005:ITD

[939] Christine Lim and Grace W. Pan. Inbound tourism developments and patterns in China. Mathematics and Computers in Simulation, 68(5–6): 498–506, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000340.

Lim:2005:DYC

[940] Lee K. Lim. A dollar or yen currency union in East Asia. Mathematics and Computers in Simulation, 68 (5–6):507–516, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405000388.

McAleer:2005:TCA

[941] Michael McAleer and Jason Chee Wei Nam. Testing for contagion in ASEAN exchange rates. Mathematics and Computers in Simulation, 68(5–6): 517–525, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540500039X.

McKenzie:2005:DBU

[942] C. R. McKenzie and Sumiko Takaoka. Deregulation of bank underwriting activities: impacts in the Euro– yen and Japanese corporate bond markets. Mathematics and Computers in Simulation, 68(5–6):526– 535, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000406.

Oya:2005:PEC

[943] Kosuke Oya. Properties of estimators of count data model with endogenous switching. Mathematics and Computers in Simulation, 68(5–6): 536–544, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540500042X.

Peiris:2005:SSM

[944] Shelton Peiris, David Allen, and Wenling Yang. Some statistical models for durations and an application to News Corporation stock prices. Mathematics and Computers in Simulation, 68 (5–6):545–552, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000352.

Shareef:2005:SIT

[945] Riaz Shareef and Suhejla Hoti. Small island tourism economies and country risk ratings. Mathematics and Computers in Simulation, 68(5–6): 553–566, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000431.

Watkins:2005:RCM

[946] Clinton Watkins and Michael McAleer. Related commodity markets and conditional correlations. Mathematics and Computers in Simulation, 68(5– 6):567–579, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000479.

Anonymous:2005:NIf

[947] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 68(5–6):581–582, May 26, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405001278.

Anonymous:2005:ICEf

[948] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 68(5–6):583– 588, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540500128X.

Anonymous:2005:AIVb

[949] Anonymous. Author index of volume 68. Mathematics and Computers in Simulation, 68(5–6):589– 591, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405001291.

Anonymous:2005:IFCf

[950] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 68(5–6): CO2, May 26, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405001229.

McAleer:2005:SSI

[951] Michael McAleer, Les Oxley, and David Post. Second special issue: Selected papers of the MSSANZ/ IMACS 15th Biennial Conference on Modelling and Simulation, Townsville, Australia, July 2003. Mathematics and Computers in Simulation, 69 (1–2):1–3, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405000625.

Anonymous:2005:MIBb

[952] Anonymous. MSSANZ/IMACS 15th Biennial Conference on Modelling and Simulation. Mathematics and Computers in Simulation, 69(1–2):1–222, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Bae:2005:RFG

[953] S. Bae, F. Famoye, J. T. Wulu, A. A. Bartolucci, and K. P. Singh. A rich family of generalized Poisson regression models with applications. Mathematics and Computers in Simulation, 69 (1–2):4–11, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000637.

Bidwell:2005:RFG

[954] V. J. Bidwell. Realistic forecasting of groundwater level, based on the eigenstructure of aquifer dynamics. Mathematics and Computers in Simulation, 69(1–2):12–20, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000583.

Carlaw:2005:OO

[955] Kenneth I. Carlaw. Optimal obsolescence. Mathematics and Computers in Simulation, 69(1–2):21– 45, June 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000522.

Hoti:2005:MSE

[956] Suhejla Hoti, Michael McAleer, and Felix Chan. Modelling the spillover effects in the volatility of atmospheric carbon dioxide concentrations. Mathematics and Computers in Simulation, 69(1–2):46–56, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000510.

Gilmour:2005:AIM

[957] J. K. Gilmour, R. A. Letcher, and A. J. Jakeman. Analysis of an integrated model for assessing land and water policy options. Mathematics and Computers in Simulation, 69(1– 2):57–77, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000534.

Herbert:2005:ATC

[958] Ric D. Herbert, Peter J. Stemp, and William E. Griffiths. Assessing two common approaches for solving models with saddle-path instabilities. Mathematics and Computers in Simulation, 69(1–2):78–89, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000546.

Kinsey-Henderson:2005:MSS

[959] Anne E. Kinsey-Henderson, David A. Post, and Ian P. Prosser. Modelling sources of sediment at subcatchment scale: an example from the Burdekin catchment, North Queensland, Australia. Mathematics and Computers in Simulation, 69(1–2): 90–102, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000571.

Marsh:2005:MIA

[960] Dan Marsh and Les Oxley. Modelling innovative activity in the New Zealand biotechnology sector. Mathematics and Computers in Simulation, 69(1– 2):103–112, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540500056X.

Nishiyama:2005:KOS

[961] Y. Nishiyama. Kernel order selection by minimum bootstrapped MSE for density weighted averages. Mathematics and Computers in Simulation, 69 (1–2):113–122, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000595.

Norton:2005:LTV

[962] J. P. Norton and J. G. Chanat. Linear time-varying models to investigate complex distributed dynamics: a rainfall-runoff example. Mathematics and Computers in Simulation, 69(1– 2):123–134, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000601.

OConnor:2005:GTT

[963] A. O’Connor, A. Zerger, and B. Itami. Geo-temporal tracking and analysis of tourist movement. Mathematics and Computers in Simulation, 69(1– 2):135–150, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405000613.

Radalj:2005:SD

[964] Kim F. Radalj and Michael McAleer. Speculation and destabilisation. Mathematics and Computers in Simulation, 69(1–2):151–161, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475405000650.

Sadoddin:2005:BDN

[965] A. Sadoddin, R. A. Letcher, A. J. Jakeman, and L. T. H. Newham. A Bayesian decision network approach for assessing the ecological impacts of salinity management. Mathematics and Computers in Simulation, 69 (1–2):162–176, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405000558.

Smith:2005:UBM

[966] K. Smith and D. Marinova. Use of bibliometric modelling for policy making. Mathematics and Computers in Simulation, 69(1–2):177– 187, June 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000649.

Ticehurst:2005:MDH

[967] J. L. Ticehurst, B. F. W. Croke, and A. J. Jakeman. Model design for the hydrology of tree belt plantations on hillslopes. Mathematics and Computers in Simulation, 69(1– 2):188–212, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000662.

Anonymous:2005:NIg

[968] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 69(1–2):213–214, June 20, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405001461.

Anonymous:2005:ICEg

[969] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 69(1–2):215– 221, June 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001473.

Anonymous:2005:IFCg

[970] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 69(1–2): CO2, June 20, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001369.

Taha:2005:Fb

[971] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 69(3–4):vii, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001540.

Agrotis:2005:TNL

[972] Maria Agrotis, Panayotis G. Kevrekidis, and Boris A. Malomed. Transmission of nonlinear localized modes through waveguide bends. Mathematics and Computers in Simulation, 69(3–4): 223–234, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405000157.

Anonymous:2005:TIIa

[973] Anonymous. Third IMACS International Conference on Nonlinear Evolution Equations and Wave Phenomena: Computation and Theory. Mathematics and Computers in Simulation, 69 (3–4):223–422, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Boyd:2005:CWC

[974] John P. Boyd. The cnoidal wave/ corner wave/breaking wave scenario: a one-sided infinite-dimension bifurcation. Mathematics and Computers in Simulation, 69(3–4):235– 242, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000169.

Choudhury:2005:MAW

[975] S. Roy Choudhury. Modulated amplitude waves in the cubic-quintic Ginzburg–Landau equation. Mathematics and Computers in Simulation, 69(3–4):243–256, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405000170.

Christou:2005:ILW

[976] M. A. Christou and C. I. Christov. Interacting localized waves for the regularized long wave equation via a Galerkin spectral method. Mathematics and Computers in Simulation, 69 (3–4):257–268, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000182.

Guyenne:2005:NSS

[977] Philippe Guyenne and David P. Nicholls. Numerical simulation of solitary waves on plane slopes. Mathematics and Computers in Simulation, 69 (3–4):269–281, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000194.

Hematulin:2005:NAR

[978] A. Hematulin and S. V. Meleshko. A new approach related with group analysis and hodograph type transformation for constructing exact solutions. Mathematics and Computers in Simulation, 69(3–4):282– 289, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000285.

Islas:2005:BEA

[979] A. L. Islas and C. M. Schober. Backward error analysis for multisymplectic discretizations of Hamiltonian PDEs. Mathematics and Computers in Simulation, 69(3–4):290– 303, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000200.

Karpeev:2005:LLF

[980] D. Karpeev and C. M. Schober. Local Lagrangian formalism and discretization of the Heisenberg magnet model. Mathematics and Computers in Simulation, 69(3–4):304– 321, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000212.

Kaup:2005:VSD

[981] D. J. Kaup. Variational solutions for the discrete nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 69(3–4):322– 333, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000297.

Kevrekidis:2005:DWS

[982] P. G. Kevrekidis, B. A. Malomed, D. J. Frantzeskakis, A. R. Bishop, H. E. Nistazakis, and R. Carretero-González. Domain walls of singlecomponent Bose–Einstein condensates in external potentials. Mathematics and Computers in Simulation, 69(3– 4):334–345, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000303.

Kheradmand:2005:CSM

[983] R. Kheradmand, L. A. Lugiato, G. Tissoni, M. Brambilla, and H. Tajalli. Cavity soliton mobility in semiconductor microresonators. Mathematics and Computers in Simulation, 69(3– 4):346–355, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000224.

Klaus:2005:REM

[984] Martin Klaus. Remarks on the eigenvalues of the Manakov system. Mathematics and Computers in Simulation, 69(3–4):356–367, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000236.

Leblond:2005:KHP

[985] H. Leblond. The KdV hierarchy and the propagation of solitons on very long distances. Mathematics and Computers in Simulation, 69(3– 4):368–377, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000248.

Leblond:2005:OSF

[986] Hervé Leblond, F. Sanchez, I. V. Mel’nikov, and D. Mihalache. Optical solitons in a few-cycle regime: Breakdown of slow-envelope approximation. Mathematics and Computers in Simulation, 69(3–4):378– 388, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540500025X.

Ludu:2005:SAB

[987] A. Ludu. Solitons and antisolitons on bounded surfaces. Mathematics and Computers in Simulation, 69(3– 4):389–399, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405000261.

Malomed:2005:DWC

[988] B. A. Malomed, H. E. Nistazakis, P. G. Kevrekidis, and D. J. Frantzeskakis. Domain-wall crosses and propellers in binary Bose–Einstein condensates. Mathematics and Computers in Simulation, 69(3–4):400– 412, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405000273.

Anonymous:2005:NIh

[989] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 69(3–4):413–414, June 24, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405001552.

Anonymous:2005:CIEa

[990] Anonymous. Calendar of IMACS events. Mathematics and Computers in Simulation, 69(3–4):415– 421, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001564.

Anonymous:2005:IFCh

[991] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 69(3–4): CO2, June 24, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001497.

Taha:2005:Fa

[992] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 69 (5–6):423, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000911.

Anonymous:2005:TIIb

[993] Anonymous. Third IMACS International Conference on Nonlinear Evolution Equations and Wave Phenomena: Computation and Theory. Mathematics and Computers in Simulation, 69(5–6):423–612, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

McConnell:2005:LCS

[994] M. McConnell, A. S. Fokas, and B. Pelloni. Localised coherent solutions of the DSI and DSII equations — a numerical study. Mathematics and Computers in Simulation, 69(5–6): 424–438, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000972.

Mickens:2005:NFD

[995] Ronald E. Mickens. A nonstandard finite difference scheme for a PDE modeling combustion with nonlinear advection and diffusion. Mathematics and Computers in Simulation, 69(5– 6):439–446, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000923.

Montesinos:2005:NSS

[996] Gaspar D. Montesinos and Víctor M. Pérez-García. Numerical studies of stabilized Townes solitons. Mathematics and Computers in Simulation, 69(5– 6):447–456, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000935.

Naidoo:2005:NIP

[997] R. Naidoo and S. Baboolal. Numerical integration of the plasma fluid equations with a modification of the second-order Nessyahu–Tadmor central scheme and soliton modeling. Mathematics and Computers in Simulation, 69(5–6):457–466, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000947.

Norton:2005:TDM

[998] G. Norton and J. Novarini. Time domain modeling of pulse propagation in non-isotropic dispersive media. Mathematics and Computers in Simulation, 69(5–6):467–476, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000959.

Paumond:2005:TRD

[999] Lionel Paumond. Towards a rigorous derivation of the fifth order KP equation. Mathematics and Computers in Simulation, 69(5–6):477– 491, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000960.

Sakaguchi:2005:PNM

[1000] H. Sakaguchi and B. A. Malomed.

Positive- and negative-mass solitons in Bose–Einstein condensates with optical lattices. Mathematics and Computers in Simulation, 69(5–6): 492–501, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000996.

Salupere:2005:SMS

[1001] A. Salupere, J. Engelbrecht, O. Ilison, and L. Ilison. On solitons in microstructured solids and granular materials. Mathematics and Computers in Simulation, 69(5–6): 502–513, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540500100X.

Sonnier:2005:SCS

[1002] W. J. Sonnier and C. I. Christov. Strong coupling of Schrödinger equations: Conservative scheme approach. Mathematics and Computers in Simulation, 69(5–6):514–525, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001011.

Soto-Crespo:2005:ESF

[1003] J. M. Soto-Crespo and Nail Akhmediev. Exploding soliton and front solutions of the complex cubic–quintic Ginzburg–Landau equation. Mathematics and Computers in Simulation, 69(5–6):526–536, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405001023.

Theocharis:2005:DSD

[1004] G. Theocharis, D. J. Frantzeskakis, P. G. Kevrekidis, R. Carretero-González, and B. A. Malomed. Dark soliton dynamics in spatially inhomogeneous media: Application to Bose– Einstein condensates. Mathematics and Computers in Simulation, 69(5– 6):537–552, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001035.

Tzirakis:2005:CAE

[1005] N. Tzirakis and P. G. Kevrekidis. On the collapse arresting effects of discreteness. Mathematics and Computers in Simulation, 69(5–6):553– 566, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405000984.

Varlamov:2005:SWF

[1006] Vladimir Varlamov and Yue Liu. Solitary waves and fundamental solution for Ostrovsky equation. Mathematics and Computers in Simulation, 69(5– 6):567–579, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001047.

Wazwaz:2005:GFP

[1007] Abdul-Majid Wazwaz. Generalized forms of the phi-four equation with compactons, solitons and periodic solutions. Mathematics and Computers in Simulation, 69(5–6):580–588, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001059.

Wu:2005:EDS

[1008] Jiahong Wu and Juan-Ming Yuan. The effect of dissipation on solutions of the complex KdV equation. Mathematics and Computers in Simulation, 69(5– 6):589–599, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001060.

Anonymous:2005:NIi

[1009] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 69(5–6):600–601, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475405001771.

Anonymous:2005:ICEh

[1010] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 69(5–6):602–607, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405001783.

Anonymous:2005:AIVc

[1011] Anonymous. Author index of volume 69. Mathematics and Computers in Simulation, 69(5–6):608– 611, August 5, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405001825.

Anonymous:2005:IFCi

[1012] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 69(5–6): CO2, August 5, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001722.

Yousefi:2005:LWM

[1013] S. Yousefi and M. Razzaghi. Legendre wavelets method for the nonlinear Volterra–Fredholm integral equations. Mathematics and Computers in Simulation, 70(1):1– 8, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540500090X.

Chi:2005:OHS

[1014] H. Chi, M. Mascagni, and T. Warnock.

On the optimal Halton sequence. Mathematics and Computers in Simulation, 70(1):9–21, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847540500087X.

Li:2005:SIT

[1015] Helong Li, Lihua Yang, and Daren Huang. The study of the intermittency test filtering character of Hilbert– Huang transform. Mathematics and Computers in Simulation, 70(1):22– 32, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000893.

Dimitrov:2005:ANS

[1016] Dobromir T. Dimitrov and Hristo V. Kojouharov. Analysis and numerical simulation of phytoplankton–nutrient systems with nutrient loss. Mathematics and Computers in Simulation, 70 (1):33–43, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405000881.

Simpson:2005:ANT

[1017] Matthew J. Simpson, Kerry A. Landman, and T. Prabhakar Clement. Assessment of a non-traditional operator split algorithm for simulation of reactive transport. Mathematics and Computers in Simulation, 70(1):44– 60, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001072.

Anonymous:2005:NIj

[1018] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 70(1):61–62, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405001928.

Anonymous:2005:CIEb

[1019] Anonymous. Calendar of IMACS events. Mathematics and Computers in Simulation, 70(1):63–67, September 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540500193X.

Anonymous:2005:IFCj

[1020] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 70(1):CO2, September 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001874.

Anonymous:2005:PSa

[1021] Anonymous. Pages 1–68 (1 September 2005). Mathematics and Computers in Simulation, 70(1):??, September 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Cui:2005:SPC

[1022] Li-Hong Cui. Some properties and construction of multiwavelets related to different symmetric centers. Mathematics and Computers in Simulation, 70 (2):69–89, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001084.

Ramadan:2005:NTM

[1023] Mohamed A. Ramadan and Talaat S.

El-Danaf. Numerical treatment for the modified Burgers equation. Mathematics and Computers in Simulation, 70 (2):90–98, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001205.

Park:2005:DRF

[1024] Ju H. Park. Design of robust H_{∞} filter for a class of neutral systems: LMI optimization approach. Mathematics and Computers in Simulation, 70(2): 99–109, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001655.

Momani:2005:ENS

[1025] Shaher Momani. An explicit and numerical solutions of the fractional KdV equation. Mathematics and Computers in Simulation, 70(2):110– 118, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001643.

Soliman:2005:NSG

[1026] A. A. Soliman. Numerical simulation of the generalized regularized long wave equation by He’s variational iteration method. Mathematics and Computers in Simulation, 70(2):119– 124, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001679.

Anonymous:2005:NIk

[1027] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 70(2):125–126, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405002028.

Anonymous:2005:ICEi

[1028] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 70(2):127–131, September 8, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540500203X.

Anonymous:2005:IFCk

[1029] Anonymous. Inside front cover — Editorial Board. Mathematics and Computers in Simulation, 70(2):CO2, September 8, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405001990.

Anonymous:2005:PSb

[1030] Anonymous. Pages 69–132 (8 September 2005). Mathematics and Computers in Simulation, 70(2):??, September 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Huang:2005:GSA

[1031] Chuangxia Huang, Lihong Huang, and Zhaohui Yuan. Global stability analysis of a class of delayed cellular neural networks. Mathematics and Computers in Simulation, 70(3):133– 148, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001667.

Coutinho:2005:ATC

[1032] F. A. B. Coutinho, M. N. Burattini, L. F. Lopez, and E. Massad. An approximate threshold condition for nonautonomous system: an application to a vector-borne infection. Mathematics and Computers in Simulation, 70(3):

149–158, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001680.

Dick:2005:DDS

[1033] Josef Dick and Friedrich Pillichshammer. Diaphony, discrepancy, spectral test and worst-case error. Mathematics and Computers in Simulation, 70(3): 159–171, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001692.

Cajueiro:2005:RVS

[1034] Daniel O. Cajueiro and Benjamin M. Tabak. The rescaled variance statistic and the determination of the Hurst exponent. Mathematics and Computers in Simulation, 70(3):172– 179, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405001709.

Anonymous:2005:NIl

[1035] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 70(3):180–181, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475405002181.

Anonymous:2005:ICEj

[1036] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 70(3):182–185, November 8, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002193.

Anonymous:2005:EBa

[1037] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 70(3):CO2, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405002156.

Anonymous:2005:PN

[1038] Anonymous. Pages 133–186 (8 November 2005). Mathematics and Computers in Simulation, 70(3):??, November 8, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Triandaf:2005:CMA

[1039] Ioana Triandaf and Ira B. Schwartz.

A collective motion algorithm for tracking time-dependent boundaries. Mathematics and Computers in Simulation, 70(4):187–202, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405001850.

Bastos:2005:ABC

[1040] J. Bastos and A. Monti. Automatically building customized circuitbased simulation models using symbolic computing. Mathematics and Computers in Simulation, 70(4):203– 220, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405001977.

Khater:2005:CWS

[1041] A. H. Khater, M. M. Hassan, and R. S. Temsah. Cnoidal wave solutions for a class of fifth-order KdV equations. Mathematics and Computers in Simulation, 70(4):221–226, December 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002107.

Park:2005:SBH

[1042] Jeong-Soo Park. A simulation-based hyperparameter selection for quantile estimation of the generalized extreme value distribution. Mathematics and Computers in Simulation, 70(4):227– 234, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002132.

Anonymous:2005:NIm

[1043] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 70(4):235–236, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405002302.

Anonymous:2005:ICEk

[1044] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 70(4):237–240, December 1, 2005. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002314.

Anonymous:2005:GA

[1045] Anonymous. Guide for authors.

Mathematics and Computers in Simulation, 70(4):241–246, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475405002119.

Anonymous:2005:EBb

[1046] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 70(4):CO2, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002272.

Anonymous:2005:PD

[1047] Anonymous. Pages 187–246 (1 December 2005). Mathematics and Computers in Simulation, 70(4):??, December 1, 2005. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Craye:2006:I

[1048] E. Craye, E. Niel, and O. Korbaa. Introduction. Mathematics and Computers in Simulation, 70(5–6):247– 249, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002405.

Anonymous:2006:C

[1049] Anonymous. (CESA 2003). Mathematics and Computers in Simulation, 70(5–6):247–442, February 24, 2006.

CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sava:2006:CSA

[1050] Alexandru Tiberiu Sava and Hassane Alla. A control synthesis approach for time discrete event systems. Mathematics and Computers in Simulation, 70(5–6):250–265, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002417.

Sreenivas:2006:SOS

[1051] Ramavarapu S. Sreenivas. Some observations on supervisory policies that enforce liveness in partially controlled Free-Choice Petri nets. Mathematics and Computers in Simulation, 70(5–6): 266–274, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002429.

Holloway:2006:SOC

[1052] Lawrence E. Holloway, Yu Gong, and Jeff Ashley. State observability and condition observability for a class of interacting discrete event systems. Mathematics and Computers in Simulation, 70(5–6):275– 286, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002430.

Jiroveanu:2006:DAF

[1053] G. Jiroveanu and R. K. Boel. A distributed approach for fault detection and diagnosis based on Time Petri Nets. Mathematics and Computers in Simulation, 70(5–6):287– 313, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002442.

Perronne:2006:ACD

[1054] Jean-Marc Perronne, Laurent Thiry, and Bernard Thirion. Architectural concepts and Design Patterns for behavior modeling and integration. Mathematics and Computers in Simulation, 70(5–6):314–329, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002454.

Eynard:2006:PSI

[1055] Benoît Eynard, Thomas Gallet, Lionel Roucoules, and Guillaume Ducellier. PDM system implementation based on UML. Mathematics and Computers in Simulation, 70(5–6):330– 342, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002466.

Lombard:2006:TFM

[1056] M. Lombard and L. Gzara Yesilbas.

Towards a framework to manage formalised exchanges during collaborative design. Mathematics and Computers in Simulation, 70(5–6):343– 357, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405002478.

Ezzedine:2006:MIS

[1057] Houcine Ezzedine, Abdelwaheb Trabelsi, and Christophe Kolski. Modelling of an interactive system with an agent-based architecture using Petri nets, application of the method to the supervision of a transport system. Mathematics and Computers in Simulation, 70(5–6):358–376, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540500248X.

Toguyeni:2006:SDA

[1058] Abdoul K. A. Toguyéni, Etienne Craye, and Larbi Sekhri. Study of the diagnosability of automated production systems based on functional graphs. Mathematics and Computers in Simulation, 70(5–6):377– 393, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405002491.

Kamach:2006:MMA

[1059] Oulaid Kamach, Laurent Piétrac, and Éric niel. Multi-model approach to discrete events systems: Application to operating mode management. Mathematics and Computers in Simulation, 70(5–6):394–407, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475405002508.

Nourelfath:2006:GAS

[1060] Mustapha Nourelfath and Ahmed Khoumsi. Grid automata and supervisory control of dense real-time discrete event systems. Mathematics and Computers in Simulation, 70(5–6): 408–418, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540500251X.

Lee:2006:SAF

[1061] Jong-Kun Lee and Ouajdi Korbaa. Scheduling analysis of FMS: an unfolding timed Petri nets approach. Mathematics and Computers in Simulation, 70(5–6):419–432, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002521.

Anonymous:2006:NIa

[1062] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 70(5–6):433–434, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000103.

Anonymous:2006:ICEa

[1063] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 70(5–6):435–439, February 24, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000115.

Anonymous:2006:AIVa

[1064] Anonymous. Author index of volume 70. Mathematics and Computers in Simulation, 70(5–6):440– 442, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000127.

Anonymous:2006:EBa

[1065] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 70(5–6):CO2, February 24, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475406000073.

Pudmenzky:2006:ANC

[1066] Alexander Pudmenzky. On the advantages of non-cooperative behavior in agent populations. Mathematics and Computers in Simulation, 71 (1):1–8, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405002090.

Garrod:2006:TDT

[1067] Neil Garrod, Sonja Ratej Pirkovic, and Aljosa Valentincic. Testing for discontinuity or type of distribution. Mathematics and Computers in Simulation, 71(1):9–15, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405002120.

Dehghan:2006:FDP

[1068] Mehdi Dehghan. Finite difference procedures for solving a problem arising in modeling and design of certain optoelectronic devices. Mathematics and Computers in Simulation, 71 (1):16–30, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002259.

Cai:2006:CSU

[1069] Li Cai, Jian-Hu Feng, Wen-Xian Xie, and Jun Zhou. Computations of steady and unsteady transport of pollutant in shallow water. Mathematics and Computers in Simulation, 71 (1):31–43, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540500234X.

Jezequel:2006:RCM

[1070] F. Jézéquel, F. Rico, J.-M. Chesneaux, and M. Charikhi. Reliable computation of a multiple integral involved in the neutron star theory. Mathematics and Computers in Simulation, 71 (1):44–61, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002375.

Lee:2006:DCK

[1071] S. H. Lee and B. K. Soni. The derivation and the computation of kinematic boundary condition. Mathematics and Computers in Simulation, 71 (1):62–72, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405002533.

Anonymous:2006:NIb

[1072] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 71(1):73–74, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000280.

Anonymous:2006:ICEb

[1073] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 71(1):75–79, March 7, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000292.

Anonymous:2006:EBb

[1074] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 71(1):CO2, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000231.

Anonymous:2006:PMa

[1075] Anonymous. Pages 1–80 (7 March 2006). Mathematics and Computers in Simulation, 71(1):??, March 7, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Chen:2006:SMV

[1076] Shyh-Wei Chen. Simultaneously modeling the volatility of the growth rate of real GDP and determining business cycle turning points: Evidence from the U.S., Canada and the UK. Mathematics and Computers in Simulation, 71 (2):87–102, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002557.

Narayan:2006:BUS

[1077] Paresh Kumar Narayan. The behaviour of US stock prices: Evidence from a threshold autoregressive model. Mathematics and Computers in Simulation, 71(2):103–108, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475405002545.

Beguenane:2006:NSS

[1078] Rachid Beguenane, Mohand A. Ouhrouche, and Andrzej M. Trzynadlowski. A new scheme for sensorless induction motor control drives operating in low speed region. Mathematics and Computers in Simulation, 71(2): 109–120, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000036.

Fukuda:2006:MUR

[1079] Kosei Fukuda. Monitoring unit root and multiple structural changes: an information criterion approach. Mathematics and Computers in Simulation, 71(2):121–130, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000048.

Lopez:2006:SAS

[1080] L. F. Lopez, F. A. B. Coutinho, M. N. Burattini, and E. Massad. A schematic age-structured compartment model of the impact of antiretroviral therapy on HIV incidence and prevalence. Mathematics and Computers in Simulation, 71(2):131–148, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540600005X. See

errata [1209, 1219].

Alvarez-Vazquez:2006:OLS

[1081] L. J. Alvarez-Vázquez, A. Martínez, M. E. Vázquez-Méndez, and M. A. Vilar. Optimal location of sampling points for river pollution control. Mathematics and Computers in Simulation, 71(2):149–160, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000176.

Anonymous:2006:NIc

[1082] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 71(2):161–162, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000656.

Anonymous:2006:ICEc

[1083] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 71(2):163–168, April 11, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000668.

Anonymous:2006:EBc

[1084] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 71(2):CO2, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000620.

Anonymous:2006:PA

[1085] Anonymous. Pages 87–168 (11 April 2006). Mathematics and Computers in Simulation, 71(2):??, April 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Layton:2006:QSMa

[1086] Anita T. Layton, Christina C.

Christara, and Kenneth R. Jackson. Quadratic spline methods for the shallow water equations on the sphere: Galerkin. Mathematics and Computers in Simulation, 71(3):175– 186, May 11, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475405002351.

Layton:2006:QSMb

[1087] Anita T. Layton, Christina C.

Christara, and Kenneth R. Jackson. Quadratic spline methods for the shallow water equations on the sphere: Collocation. Mathematics and Computers in Simulation, 71(3): 187–205, May 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475405002363.

Chwastek:2006:IHM

[1088] Krzysztof Chwastek and Jan Szczyglowski. Identification of a hysteresis model parameters with genetic algorithms. Mathematics and Computers in Simulation, 71(3):206– 211, May 11, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000024.

Bowong:2006:CSD

[1089] Samuel Bowong, Moukam Kakmeni, and Rodoumta Koina. Chaos synchronization and duration time of a class of uncertain chaotic systems. Mathematics and Computers in Simulation, 71 (3):212–228, May 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000218.

Donato:2006:SAT

[1090] M. B. Donato, M. Milasi, and C. Vitanza. Sensitivity analysis for time dependent spatial price equilibrium problem. Mathematics and Computers in Simulation, 71(3):229– 239, May 11, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000358.

Anonymous:2006:NId

[1091] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 71(3):240–241, May 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000814.

Anonymous:2006:ICEd

[1092] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 71(3):242–247, May 11, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000826.

Anonymous:2006:EBd

[1093] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 71(3):CO2, May 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000784.

Anonymous:2006:PMb

[1094] Anonymous. Pages 175–254 (11 May 2006). Mathematics and Computers in Simulation, 71(3):??, May 11, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Dessaint:2006:MSE

[1095] Louis-A. Dessaint and Kamal Al-Haddad. Modeling and simulation of electric machines, converters and systems. Mathematics and Computers in Simulation, 71(4–6):

255, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000607.

Anonymous:2006:ICM

[1096] Anonymous. 8th International Conference on Modeling and Simulation of Electric Machines, Converters and Systems. Mathematics and Computers in Simulation, 71(4–6):255–528, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Benadero:2006:SAS

[1097] L. Benadero, R. Giral, A. El Aroudi, and J. Calvente. Stability analysis of a single inductor dual switching dc–dc converter. Mathematics and Computers in Simulation, 71(4– 6):256–269, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000371.

Khatounian:2006:MSH

[1098] F. Khatounian, S. Moreau, J. P. Louis, E. Monmasson, F. Louveau, and J. M. Alexandre. Modeling and simulation of a hybrid dynamic system used in haptic interfaces. Mathematics and Computers in Simulation, 71(4–6): 270–281, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000401.

Oslinger:2006:ASC

[1099] José L. Oslinger, Jairo A. Palacios, and Guillermo A. Jaramillo. Analysis of a short-circuit ring impact problem and design of a new ventilator for a 625 kVA generator, using SVL. Mathematics and Computers in Simulation, 71 (4–6):282–289, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000462.

Pryymak:2006:NNF

[1100] Bogdan Pryymak, Juan M. Moreno-Eguilaz, and Juan Peracaula. Neural network flux optimization using a model of losses in induction motor drives. Mathematics and Computers in Simulation, 71(4–6):290– 298, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000486.

Bellomo:2006:HDA

[1101] Luis Daniel Bellomo and Guy Olivier.

Harmonic distortion analysis software combining EMTP and Monte Carlo method. Mathematics and Computers in Simulation, 71(4–6):299– 309, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000516.

Robert:2006:DTM

[1102] Bruno Robert and Abdelali El Aroudi.

Discrete time model of a multicell dc/dc converter: Nonlinear approach. Mathematics and Computers in Simulation, 71(4–6):310– 319, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000425.

Ansel:2006:MSA

[1103] A. Ansel and B. Robyns. Modelling and simulation of an autonomous variable speed micro hydropower station. Mathematics and Computers in Simulation, 71(4–6): 320–332, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000498.

Cajander:2006:DOT

[1104] David Cajander and Hoang Le-Huy. Design and optimization of a torque controller for a switched reluctance motor drive for electric vehicles by simulation. Mathematics and Computers in Simulation, 71(4–6):333– 344, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000541.

Rahmani:2006:CSS

[1105] Salem Rahmani, Kamal Al-Haddad, and Hadi Youssef Kanaan. A comparative study of shunt hybrid and shunt active power filters for singlephase applications: Simulation and experimental validation. Mathematics and Computers in Simulation, 71(4– 6):345–359, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000553.

ElKhil:2006:PDL

[1106] Séjir Khojet El Khil, Ilhem Slama-Belkhodja, Maria Pietrzak-David, and Bernard de Fornel. Power distribution law in a Doubly Fed Induction Machine. Mathematics and Computers in Simulation, 71(4–6): 360–368, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000589.

Tlusty:2006:OCS

[1107] Josef Tlustý, Pavel Santarius, Viktor Valouch, and Jiří Škramlík. Optimal control of shunt active power filters in multibus industrial power systems for harmonic voltage mitigation. Mathematics and Computers in Simulation, 71(4–6):369– 376, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540600036X.

Kaddouri:2006:NIG

[1108] A. Kaddouri, S. Blais, M. Ghribi, and O. Akhrif. NLSOFT: an interactive graphical software for designing nonlinear controllers. Mathematics and Computers in Simulation, 71 (4–6):377–384, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000383.

Carducci:2006:IVF

[1109] G. Carducci, N. I. Giannoccaro, A. Messina, and G. Rollo. Identification of viscous friction coefficients for a pneumatic system model using optimization methods. Mathematics and Computers in Simulation, 71(4– 6):385–394, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000413.

LeBoudec:2006:MAC

[1110] Brice Le Boudec, Maarouf Saad, and Vahé Nerguizian. Modeling and adaptive control of redundant robots. Mathematics and Computers in Simulation, 71(4–6):395– 403, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000449.

Sakamoto:2006:LGS

[1111] Tetsuzo Sakamoto and Masaki Nakayama. Levitation–guidance stabilization of superconducting maglev through LSM currents. Mathematics and Computers in Simulation, 71(4–6):404– 414, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000474.

Jones:2006:SCQ

[1112] Martin Jones and Emil Levi. Series connected quasi-six-phase twomotor drives with independent control. Mathematics and Computers in Simulation, 71(4–6):415–424, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000395.

Tu:2006:MEV

[1113] X. Tu, L.-A. Dessaint, M. El Kahel, and A. Barry. Modeling and experimental validation of internal faults in salient pole synchronous machines including space harmonics. Mathematics and Computers in Simulation, 71 (4–6):425–439, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000450.

Eme:2006:TME

[1114] A. Eme, R. Glises, D. Chamagne, J. M. Kauffmann, F. Chalon, and T. Péra. Thermal modelling for electrical machines fed with low voltage: First approach of a reliability model. Mathematics and Computers in Simulation, 71(4–6):440– 445, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000504.

Moreau:2006:MIN

[1115] Sandrine Moreau and Jean-Claude Trigeassou. Modelling and identification of a non-linear saturated magnetic circuit: Theoretical study and experimental results. Mathematics and Computers in Simulation, 71(4– 6):446–459, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000528.

Markovic:2006:SMC

[1116] Miroslav Marković, Marcel Jufer, and Yves Perriard. A square magnetic circuit analysis using Schwarz–Christoffel mapping. Mathematics and Computers in Simulation, 71(4–6):460– 465, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000565.

Pires:2006:AVG

[1117] A. J. Pires, J. F. Martins, P. J. Branco, and J. A. Dente. An average values global model for the switched reluctance machine. Mathematics and Computers in Simulation, 71(4–6): 466–475, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406000437.

Bellini:2006:DFS

[1118] Armando Bellini and Stefano Bifaretti. A digital filter for speed noise reduction in drives using an electromagnetic resolver. Mathematics and Computers in Simulation, 71(4–6): 476–486, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540600053X.

Kanaan:2006:SEN

[1119] Hadi Youssef Kanaan, Kamal Al-Haddad, and Farhat Fnaiech. A study on the effects of the neutral inductor on the modeling and performance of a four-wire three-phase/switch/level fixed-frequency rectifier. Mathematics and Computers in Simulation, 71(4– 6):487–498, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000577.

Hecquet:2006:PEN

[1120] M. Hecquet, A. Ait-Hammouda, M. Goueygou, P. Brochet, and A. Randria. Prediction of the electromagnetic noise of an asynchronous machine using experimental designs. Mathematics and Computers in Simulation, 71 (4–6):499–509, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000590.

Anonymous:2006:NIe

[1121] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 71(4–6):510–511, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475406000954.

Anonymous:2006:ICEe

[1122] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 71(4–6):512– 517, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000966.

Anonymous:2006:AIVb

[1123] Anonymous. Author index of volume 71. Mathematics and Computers in Simulation, 71(4–6):518– 521, June 19, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001042.

Anonymous:2006:EBe

[1124] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 71(4–6):CO2, June 19, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406000929.

Wazwaz:2006:VSO

[1125] Abdul-Majid Wazwaz. The variable separated ODE method for travelling wave solutions for the Boussinesq-double sine-Gordon and the Boussinesq-double sinh-Gordon equations. Mathematics and Computers in Simulation, 72(1):1–9, July 3, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000735.

Abell:2006:CCC

[1126] Martha L. Abell, James P. Braselton, and Lorraine M. Braselton. Competition in the chemostat: a comparison of inhibitory and lethal offensive strategies. Mathematics and Computers in Simulation, 72(1):10– 25, July 3, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000723.

Lee:2006:MDT

[1127] Wah Soon Lee and Machavaram Venkata Calapathy Rao. Modeling and design of tape transport mechanism. Mathematics and Computers in Simulation, 72(1):26– 37, July 3, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000747.

Maamri:2006:PPU

[1128] N. Maamri, O. Bachelier, and D. Mehdi. Pole placement in a union of regions with prespecified subregion allocation. Mathematics and Computers in Simulation, 72(1):38– 46, July 3, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000759.

Kulikov:2006:NMN

[1129] G. Yu. Kulikov and S. K. Shindin.

Nordsieck methods on nonuniform grids: Stability and order reduction phenomenon. Mathematics and Computers in Simulation, 72(1):47– 56, July 3, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406000760.

Anonymous:2006:NIf

[1130] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 72(1):57–58, July 3, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001157.

Anonymous:2006:ICEf

[1131] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 72(1):59–64, July 3, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001169.

Anonymous:2006:EBf

[1132] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 72(1):CO2, July 3, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001108.

Anonymous:2006:PJ

[1133] Anonymous. Pages 1–70 (3 July 2006). Mathematics and Computers in Simulation, 72(1):??, July 3, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Inamuro:2006:P

[1134] Takaji Inamuro, Bruce Boghosian, and Sauro Succi. Preface. Mathematics and Computers in Simulation, 72(2– 6):71, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001224.

Abe:2006:GSA

[1135] Sumiyoshi Abe. Generalized Stosszahlansatz and associated H-theorem in nonextensive statistical mechanics. Mathematics and Computers in Simulation, 72(2–6): 72–78, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001248.

Arcidiacono:2006:ELB

[1136] S. Arcidiacono, S. Ansumali, I. V.

Karlin, J. Mantzaras, and K. B. Boulouchos. Entropic lattice Boltzmann method for simulation of binary mixtures. Mathematics and Computers in Simulation, 72(2–6):79– 83, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001261.

Benzi:2006:MML

[1137] R. Benzi, L. Biferale, M. Sbragaglia, S. Succi, and F. Toschi. Mesoscopic modelling of local phase transitions and apparent-slip phenomena in microflows. Mathematics and Computers in Simulation, 72(2–6):84–88, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001297.

Bernsdorf:2006:NSC

[1138] J. Bernsdorf, S. E. Harrison, S. M.

Smith, P. V. Lawford, and D. R. Hose. Numerical simulation of clotting processes: a lattice Boltzmann application in medical physics. Mathematics and Computers in Simulation, 72(2–6): 89–92, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001327.

Brilliantov:2006:DCE

[1139] Nikolai V. Brilliantov and Frank Spahn. Dust coagulation in equilibrium molecular gas. Mathematics and Computers in Simulation, 72(2–6): 93–97, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001376.

Chatterji:2006:EPC

[1140] A. Chatterji and J. Horbach. Electrophoretic properties of charged colloidal suspensions: Application of a hybrid MD/LB method. Mathematics and Computers in Simulation, 72(2–6): 98–102, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540600139X.

Chopard:2006:LBSa

[1141] B. Chopard, H. Nguyen, and S. Stoll. A lattice Boltzmann study of the hydrodynamic properties of 3D fractal aggregates. Mathematics and Computers in Simulation, 72(2–6):103– 107, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540600142X.

Chopard:2006:LBSb

[1142] Bastien Chopard, Rafik Ouared, and Daniel A. Rüfenacht. A lattice Boltzmann simulation of clotting in stented aneurysms and comparison with velocity or shear rate reductions. Mathematics and Computers in Simulation, 72(2–6):108–112, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001467.

Cristea:2006:FDL

[1143] A. Cristea, G. Gonnella, A. Lamura, and V. Sofonea. Finite-difference lattice Boltzmann model for liquid– vapor systems. Mathematics and Computers in Simulation, 72(2–6):113– 116, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001522.

Ebihara:2006:LBS

[1144] Ken-Ichi Ebihara and Hideo Kaburaki. Lattice Boltzmann simulation of solution chemistry for crevice corrosion. Mathematics and Computers in Simulation, 72(2–6):117–123, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001236.

Giupponi:2006:NNB

[1145] G. Giupponi and P. V. Coveney.

Non-Newtonian behaviour of the gyroid mesophase: a lattice-Boltzmann study. Mathematics and Computers in Simulation, 72(2–6):124–127, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001285.

Hirabayashi:2006:NAF

[1146] Miki Hirabayashi, Makoto Ohta, Krisztina Baráth, Daniel A. Rüfenacht, and Bastien Chopard. Numerical analysis of the flow pattern in stented aneurysms and its relation to velocity reduction and stent efficiency. Mathematics and Computers in Simulation, 72(2–6):128–133, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001352.

Hyakutake:2006:LBS

[1147] Toru Hyakutake, Takeshi Matsumoto, and Shinichiro Yanase. Lattice Boltzmann simulation of blood cell behavior at microvascular bifurcations. Mathematics and Computers in Simulation, 72(2–6):134–140, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001388.

Inamuro:2006:LBS

[1148] T. Inamuro and T. Ii. Lattice Boltzmann simulation of the dispersion of aggregated particles under shear flows. Mathematics and Computers in Simulation, 72(2–6):141–146, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001431.

Jankovic:2006:LGA

[1149] D. Jankovic and D. A. Wolf-Gladrow.

Lattice Gas Automata: Drying simulation in heterogeneous models. Mathematics and Computers in Simulation, 72(2–6):147–155, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001492.

Jia:2006:LBS

[1150] Xinli Jia, J. B. McLaughlin, and K. Kontomaris. Lattice Boltzmann simulations of contact line motion on uniform surfaces. Mathematics and Computers in Simulation, 72(2–6):156– 159, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001558.

Kusumaatmaja:2006:LBS

[1151] H. Kusumaatmaja, A. Dupuis, and J. M. Yeomans. Lattice Boltzmann simulations of drop dynamics. Mathematics and Computers in Simulation, 72(2–6):160–164, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001571.

Latt:2006:LBM

[1152] Jonas Latt and Bastien Chopard. Lattice Boltzmann method with regularized pre-collision distribution functions. Mathematics and Computers in Simulation, 72(2–6):165–168, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001583.

Latt:2006:SED

[1153] Jonas Latt, Yannick Grillet, Bastien Chopard, and Peter Wittwer. Simulating an exterior domain for drag force computations in the lattice Boltzmann method. Mathematics and Computers in Simulation, 72(2–6):169– 172, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001273.

Mayer:2006:DNL

[1154] Gusztáv Mayer and Gábor Házi. Direct numerical and large eddy simulation of longitudinal flow along triangular array of rods using the lattice Boltzmann method. Mathematics and Computers in Simulation, 72(2–6): 173–178, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001303.

Prasianakis:2006:ELB

[1155] N. I. Prasianakis, S. S. Chikatamarla, I. V. Karlin, S. Ansumali, and K. Boulouchos. Entropic lattice Boltzmann method for simulation of thermal flows. Mathematics and Computers in Simulation, 72(2–6): 179–183, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001339.

Sakazaki:2006:MCF

[1156] Y. Sakazaki, S. Masuda, J. Onishi, Y. Chen, and H. Ohashi. The modeling of colloidal fluids by the realcoded lattice gas. Mathematics and Computers in Simulation, 72(2–6):184– 189, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001364.

Serrano:2006:STI

[1157] M. Serrano, G. De Fabritiis, P. Español, and P. V. Coveney. A stochastic Trotter integration scheme for dissipative particle dynamics. Mathematics and Computers in Simulation, 72(2–6):190– 194, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001443.

Seta:2006:LBS

[1158] Takeshi Seta, Eishun Takegoshi, and Kenichi Okui. Lattice Boltzmann simulation of natural convection in porous media. Mathematics and Computers in Simulation, 72(2–6):195– 200, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001455.

Shu:2006:FSL

[1159] C. Shu, X. D. Niu, Y. T. Chew, and Q. D. Cai. A fractional step lattice Boltzmann method for simulating high Reynolds number flows. Mathematics and Computers in Simulation, 72(2–6): 201–205, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001480.

Sullivan:2006:CRL

[1160] Simon P. Sullivan, Lynn F. Gladden, and Michael L. Johns. 3D chemical reactor LB simulations. Mathematics and Computers in Simulation, 72(2–6): 206–211, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001509.

Tagawa:2006:NST

[1161] T. Tagawa. Numerical simulation of two-phase flows in the presence of a magnetic field. Mathematics and Computers in Simulation, 72(2–6):212– 219, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001534.

Takada:2006:PFM

[1162] Naoki Takada, Masaki Misawa, and Akio Tomiyama. A phase-field method for interface-tracking simulation of twophase flows. Mathematics and Computers in Simulation, 72(2–6):220– 226, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540600125X.

Tosi:2006:NSE

[1163] F. Tosi, S. Ubertini, S. Succi, H. Chen, and I. V. Karlin. Numerical stability of Entropic versus positivity-enforcing Lattice Boltzmann schemes. Mathematics and Computers in Simulation, 72(2–6):227– 231, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001315.

Tuzel:2006:CTC

[1164] Erkan Tüzel, Thomas Ihle, and Daniel M. Kroll. Constructing thermodynamically consistent models with a non-ideal equation of state. Mathematics and Computers in Simulation, 72(2–6):232–236, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001340.

Ubertini:2006:ULB

[1165] S. Ubertini, G. Bella, and S. Succi.

Unstructured lattice Boltzmann equation with memory. Mathematics and Computers in Simulation, 72(2–6):237– 241, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001406.

Wang:2006:MFF

[1166] Lian-Ping Wang and Behnam Afsharpoya. Modeling fluid flow in fuel cells using the lattice-Boltzmann approach. Mathematics and Computers in Simulation, 72(2–6):242– 248, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001418.

Xu:2006:LBS

[1167] Aiguo Xu, Sauro Succi, and Bruce M. Boghosian. Lattice BBGKY scheme for two-phase flows: One-dimensional case. Mathematics and Computers in Simulation, 72(2–6):249–252, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001479.

Xu:2006:GKS

[1168] Kun Xu and Eswar Josyula. Gaskinetic scheme for rarefied flow simulation. Mathematics and Computers in Simulation, 72(2–6):253–256, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001510.

Yamamoto:2006:LBS

[1169] Kazuhiro Yamamoto, Shingo Satake, Hiroshi Yamashita, Naoki Takada, and Masaki Misawa. Lattice Boltzmann simulation on porous structure and soot accumulation. Mathematics and Computers in Simulation, 72(2–6):257– 263, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001546.

Yoshino:2006:LBS

[1170] Masato Yoshino and Yusuke Mizutani. Lattice Boltzmann simulation of liquid–gas flows through solid bodies in a square duct. Mathematics and Computers in Simulation, 72(2–6): 264–269, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540600156X.

Anonymous:2006:NIg

[1171] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 72(2–6):270–271, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475406001935.

Anonymous:2006:ICEg

[1172] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 72(2–6):272–276, September 9, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001947.

Anonymous:2006:AIVc

[1173] Anonymous. Author index of volume 72. Mathematics and Computers in Simulation, 72(2–6):277– 280, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001959.

Anonymous:2006:EBg

[1174] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 72(2–6):CO2, September 9, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847540600190X.

Rodriguez:2006:CPS

[1175] J. Daza Rodríguez and F. Tovar. Cubic polynomial A-splines: G^{1}, G^{2} and inflection point control. Mathematics and Computers in Simulation, 73(1): 133–141, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001832.

Devloo:2006:FEM

[1176] Philippe R. B. Devloo, Paulo Dore Fernandes, Sônia M. Gomes, Cedric Marcelo Augusto Ayala Bravo, and Renato Gomes Damas. A finite element model for three dimensional hydraulic fracturing. Mathematics and Computers in Simulation, 73(1):142– 155, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001790.

Doescher:2006:CMG

[1177] Erwin Doescher, Haroldo F. de Campos Velho, and Fernando M. Ramos. Criteria for mixed grids in computational fluid dynamics. Mathematics and Computers in Simulation, 73(1): 156–167, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001728.

Franquiz:2006:CGC

[1178] J. Franquiz, M. Paluszny, and F. Tovar. Cyclides and the guiding circle. Mathematics and Computers in Simulation, 73(1):168–174, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001649.

Limon:2006:OTO

[1179] Alfonso Limon, Silvia Bertuglia, Ali Nadim, and Peter Salamon. Oxygen transport from the outer boundary of a pulsating wall of an arteriole. Mathematics and Computers in Simulation, 73(1):175–182, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001650.

Martinez:2006:EBR

[1180] Gabriela Martínez, José Manuel García Aznar, Manuel Doblaré, and Miguel Cerrolaza. External bone remodeling through boundary elements and damage mechanics. Mathematics and Computers in Simulation, 73(1):183– 199, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001674.

Martinez:2006:WBS

[1181] J. Martínez and G. Larrazábal. Wavelet-based SPAI pre-conditioner using local dropping. Mathematics and Computers in Simulation, 73(1): 200–214, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001637.

Montilla:2006:MAM

[1182] Orestes Montilla, Carlos Cadenas, and José Castillo. Matrix approach to mimetic discretizations for differential operators on non-uniform grids. Mathematics and Computers in Simulation, 73(1):215–225, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001698.

Pereyra:2006:LSC

[1183] V. Pereyra and G. Scherer. Least squares collocation solution of elliptic problems in general regions. Mathematics and Computers in Simulation, 73(1):226–230, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S037847540600173X.

Pereyra:2006:VPN

[1184] V. Pereyra, G. Scherer, and F. Wong. Variable projections neural network training. Mathematics and Computers in Simulation, 73(1):231–243, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001753.

Perez:2006:OPM

[1185] Laura V. Pérez, Guillermo R. Bossio, Diego Moitre, and Guillermo O. García. Optimization of power management in an hybrid electric vehicle using dynamic programming. Mathematics and Computers in Simulation, 73(1): 244–254, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001807.

Souto:2006:RVP

[1186] R. P. Souto, H. F. Campos Velho, and S. Stephany. Reconstruction of vertical profiles of the absorption and scattering coefficients from multispectral radiances. Mathematics and Computers in Simulation, 73(1):255– 267, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001777.

Anonymous:2006:NIh

[1187] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 73(1):268–269, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475406002266.

Anonymous:2006:ICEh

[1188] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 73(1):270–274, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406002278.

Anonymous:2006:AIVd

[1189] Anonymous. Author index of volume 73. Mathematics and Computers in Simulation, 73(1):275–277, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540600228X.

Pereyra:2006:P

[1190] Victor Pereyra and Jose Castillo. Preface. Mathematics and Computers in Simulation, 73(1–4):1, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540600187X.

Abreu:2006:TPI

[1191] E. Abreu, J. Douglas, F. Furtado, D. Marchesin, and F. Pereira. Three-phase immiscible displacement in heterogeneous petroleum reservoirs. Mathematics and Computers in Simulation, 73(1–4):2–20, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001765.

Arcos:2006:SCL

[1192] Gabriel Arcos, Guillermo Montilla, José Ortega, and Marco Paluszny. Shape control of 3D lemniscates.

Mathematics and Computers in Simulation, 73(1–4):21–27, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475406001625.

Arnau:2006:SIC

[1193] J. M. Arnau, S. Jerez, L. Jódar, and J. V. Romero. A semi-implicit conservation element–solution element method for chemical species transport simulation to tapered ducts of internal combustion engine. Mathematics and Computers in Simulation, 73(1–4): 28–37, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001662.

Bourchtein:2006:SLS

[1194] Andrei Bourchtein. Semi-Lagrangian semi-implicit locally one-dimensional scheme for hydrostatic atmospheric model. Mathematics and Computers in Simulation, 73(1–4):38–51, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001741.

Bourchtein:2006:MTS

[1195] Andrei Bourchtein and Ludmila Bourchtein. Modified time splitting scheme for shallow water equations. Mathematics and Computers in Simulation, 73(1–4):52–64, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001704.

Bunse-Gerstner:2006:IPL

[1196] Angelika Bunse-Gerstner, Valia Guerra-Ones, and Humberto Madrid de La Vega. An improved preconditioned LSQR for discrete ill-posed problems. Mathematics and Computers in Simulation, 73(1–4):65–75, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001819.

Cadenas:2006:LSF

[1197] Carlos E. Cadenas, Javier J. Rojas, and Vianey Villamizar. A least squares finite element method with high degree element shape functions for one-dimensional Helmholtz equation. Mathematics and Computers in Simulation, 73(1–4):76–86, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001686.

Cardoso:2006:FED

[1198] Cláudio G. S. Cardoso, Maria Cristina Cunha, Anamaria Gomide, Denis J. Schiozer, and Jorge Stolfi. Finite elements on dyadic grids with applications. Mathematics and Computers in Simulation, 73(1–4):87–104, November 6, 2006. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406001844.

Chen-Charpentier:2006:CNN

[1199] Benito M. Chen-Charpentier, Dobromir T. Dimitrov, and Hristo V. Kojouharov. Combined nonstandard numerical methods for ODEs with polynomial right-hand sides. Mathematics and Computers in Simulation, 73(1–4): 105–113, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001789.

Claeyssen:2006:FSC

[1200] J. R. Claeyssen, C. Garibotti, and S. Vielmo. The free surface of a coupled ocean–atmosphere model due to forcing effects. Mathematics and Computers in Simulation, 73(1–4):114– 124, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001716.

Cortes:2006:CAD

[1201] J. C. Cortés, P. Sevilla-Peris, and L. Jódar. Constructing approximate diffusion processes with uncertain data. Mathematics and Computers in Simulation, 73(1–4):125– 132, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001820.

Anonymous:2006:EBh

[1202] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 73(1–4):ifc, November 6, 2006. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475406002217.

Yang:2007:MAD

[1203] Taho Yang, Mu-Chen Chen, and Chih-Ching Hung. Multiple attribute decision-making methods for the dynamic operator allocation problem. Mathematics and Computers in Simulation, 73(5):285–299, January 10, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000887.

Kamel:2007:BAN

[1204] M. M. Kamel. Bifurcation analysis of a nonlinear coupled pitch– roll ship. Mathematics and Computers in Simulation, 73(5):300–308, January 10, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000899.

Takhtamyshev:2007:QRI

[1205] George Takhtamyshev, Bart Vandewoestyne, and Ronald Cools. Quasirandom integration in high dimensions. Mathematics and Computers in Simulation, 73(5):309–319, January 10, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406000905.

Li:2007:EBE

[1206] Xiuchun Li, Yimin Shi, Jieqiong Wei, and Jian Chai. Empirical Bayes estimators of reliability performances using LINEX loss under progressively Type-II censored samples. Mathematics and Computers in Simulation, 73(5): 320–326, January 10, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540600108X.

Liu:2007:APS

[1207] Bingwen Liu. Almost periodic solutions for Hopfield neural networks with continuously distributed delays. Mathematics and Computers in Simulation, 73(5):327–335, January 10, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001613.

Leontitsis:2007:SAC

[1208] Alexandros Leontitsis and Jenny Pagge. A simulation approach on Cronbach’s alpha statistical significance. Mathematics and Computers in Simulation, 73(5):336–340, January 10, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002059.

Lopez:2007:ESAa

[1209] L. F. Lopez, F. A. B. Coutinho, M. N.

Burattini, and E. Massad. Erratum to “A schematic age-structured compartment model of the impact of antiretroviral therapy on HIV incidence and prevalence” [Math. Comput. Simul. 71 (2006) 131–148]. Mathematics and Computers in Simulation, 73 (5):341, January 10, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406001030. See

[1080].

Anonymous:2007:NIa

[1210] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 73(5):342–343, January 10, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475406002850.

Anonymous:2007:ICEa

[1211] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 73(5):344–349, January 10, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406002862.

Anonymous:2007:EBa

[1212] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 73(5):CO2, January 10, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002801.

Anonymous:2007:PJa

[1213] Anonymous. Pages 285–350 (10 January 2007). Mathematics and Computers in Simulation, 73(5):??, January 10, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Lejay:2007:CPE

[1214] Antoine Lejay and Sylvain Maire.

Computing the principal eigenvalue of the Laplace operator by a stochastic method. Mathematics and Computers in Simulation, 73(6):351–363, February 6, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001856.

Zhusubaliyev:2007:QTB

[1215] Zhanybai T. Zhusubaliyev, Evgeniy Soukhoterin, and Erik Mosekilde. Quasiperiodicity and torus breakdown in a power electronic dc/dc converter. Mathematics and Computers in Simulation, 73(6):364–377, February 6, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406001868.

Marco:2007:SMA

[1216] Ana Marco and José-Javier Martínez.

Structured matrices in the application of bivariate interpolation to curve implicitization. Mathematics and Computers in Simulation, 73(6): 378–385, February 6, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406001881.

Cook:2007:TCT

[1217] Steven Cook. A threshold cointegration test with increased power. Mathematics and Computers in Simulation, 73(6): 386–392, February 6, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002060.

Hein:2007:DSC

[1218] Steffen Hein. Dual scattering channel schemes extending the Johns algorithm. Mathematics and Computers in Simulation, 73(6):393–407, February 6, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406002084.

Lopez:2007:ESAb

[1219] L. F. Lopez, Francisco A. B. Coutinho, Marcelo N. Burattini, and Eduardo Massad. Erratum to “A schematic agestructured compartment model of the impact of antiretroviral on HIV incidence and prevalence” [Math. Comput. Simul. 71 (2006) 131–148]. Mathematics and Computers in Simulation, 73 (6):408, February 6, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002369. See

[1080].

Anonymous:2007:NIb

[1220] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 73(6):409–410, February 6, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475406002977.

Anonymous:2007:ICEb

[1221] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 73(6):411–415, February 6, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002989.

Anonymous:2007:EBb

[1222] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 73(6):CO2, February 6, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540600293X.

Anonymous:2007:PFa

[1223] Anonymous. Pages 351–422 (6 February 2007). Mathematics and Computers in Simulation, 73(6):??, February 6, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Nadarajah:2007:GGD

[1224] Saralees Nadarajah and Arjun K.

Gupta. A generalized gamma distribution with application to drought data. Mathematics and Computers in Simulation, 74(1):1–7, February 15, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002102.

Hanke:2007:SAD

[1225] Michael Hanke, K. Henrik, A. Olsson, and Magnus Strömgren. Stability analysis of a degenerate hyperbolic system modelling a heat exchanger. Mathematics and Computers in Simulation, 74 (1):8–19, February 15, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002114.

Mackevicius:2007:WAI

[1226] Vigirdas Mackevičius. On weak approximations of ( a, b )-invariant diffusions. Mathematics and Computers in Simulation, 74(1):20–28, February 15, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002126.

Lee:2007:MRT

[1227] Ris S. C. Lee and Roger L. Hughes. Minimisation of the risk of trampling in a crowd. Mathematics and Computers in Simulation, 74(1):29– 37, February 15, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002138.

Song:2007:GRS

[1228] Qiankun Song and Jinde Cao. Global robust stability of interval neural networks with multiple time-varying delays. Mathematics and Computers in Simulation, 74(1):38–46, February 15, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540600214X.

Zhao:2007:HDG

[1229] Weirui Zhao and Yong Tan. Harmless delays for global exponential stability of Cohen–Grossberg neural networks. Mathematics and Computers in Simulation, 74(1):47–57, February 15, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002096.

Anonymous:2007:NIc

[1230] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 74(1):58–59, February 15, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475406003077.

Anonymous:2007:ICEc

[1231] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 74(1):60–64, February 15, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406003089.

Anonymous:2007:EBc

[1232] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 74(1):CO2, February 15, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847540600303X.

Anonymous:2007:PFb

[1233] Anonymous. Pages 1–70 (15 February 2007). Mathematics and Computers in Simulation, 74(1):??, February 15, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Taha:2007:Fa

[1234] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 74(2–3):71, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540600276X.

Anonymous:2007:TRTa

[1235] Anonymous. Thiab R. Taha. Mathematics and Computers in Simulation, 74(2–3):71–264, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Boyd:2007:WNM

[1236] John P. Boyd. Why Newton’s method is hard for travelling waves: Small denominators, KAM theory, Arnold’s linear Fourier problem, non-uniqueness, constraints and erratic failure. Mathematics and Computers in Simulation, 74(2–3):72–81, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002746.

Christou:2007:FGM

[1237] M. A. Christou and C. I. Christov. Fourier–Galerkin method for 2D solitons of Boussinesq equation. Mathematics and Computers in Simulation, 74(2–3):82–92, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002722.

Christov:2007:MLE

[1238] C. I. Christov. Maxwell–Lorentz electrodynamics as a manifestation of the dynamics of a viscoelastic metacontinuum. Mathematics and Computers in Simulation, 74(2–3): 93–104, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002503.

Parau:2007:TDG

[1239] Emilian I. Părău, Jean-Marc Vanden-Broeck, and Mark J. Cooker. Threedimensional gravity and gravitycapillary interfacial flows. Mathematics and Computers in Simulation, 74 (2–3):105–112, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002527.

Borluk:2007:NSL

[1240] H. Borluk, G. M. Muslu, and H. A.

Erbay. A numerical study of the long wave–short wave interaction equations. Mathematics and Computers in Simulation, 74(2–3):113–125, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002540.

Frinkle:2007:GPM

[1241] Karl H. Frinkle. A Galërkin projection and multiple scales approach to Feshbach resonance in Bose–Einstein condensates. Mathematics and Computers in Simulation, 74(2–3):126– 134, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002564.

Craig:2007:STD

[1242] Walter Craig, Diane M. Henderson, Maribeth Oscamou, and Harvey Segur. Stable three-dimensional waves of nearly permanent form on deep water. Mathematics and Computers in Simulation, 74(2–3):135– 144, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002655.

Gegenhasi:2007:IDD

[1243] Gegenhasi and Xing-Biao Hu. Integrability of a differential-difference KP equation with self-consistent sources. Mathematics and Computers in Simulation, 74(2–3):145– 158, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002692.

Canney:2007:SPW

[1244] Nathan E. Canney and John D. Carter. Stability of plane waves on deep water with dissipation. Mathematics and Computers in Simulation, 74(2– 3):159–167, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002734.

Kalisch:2007:DCM

[1245] Henrik Kalisch. Derivation and comparison of model equations for interfacial capillary-gravity waves in deep water. Mathematics and Computers in Simulation, 74(2–3):168– 178, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002552.

Ludu:2007:IGN

[1246] A. Ludu and N. Hutchings. Internally generated nonlinear waves in filament bundles. Mathematics and Computers in Simulation, 74(2–3): 179–189, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540600259X.

Mickens:2007:DDF

[1247] Ronald E. Mickens. Determination of denominator functions for a NSFD scheme for the Fisher PDE with linear advection. Mathematics and Computers in Simulation, 74(2–3): 190–195, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002618.

Kevrekidis:2007:NLE

[1248] Panayotis G. Kevrekidis and Yannis Drossinos. Nonlinearity from linearity: the Ermakov–Pinney equation revisited. Mathematics and Computers in Simulation, 74(2–3):196– 202, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406002631.

Tang:2007:NLW

[1249] X.-H. Tang and C. I. Christov. Nonlinear waves of the steady natural convection in a vertical fluid layer: a numerical approach. Mathematics and Computers in Simulation, 74(2– 3):203–213, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002448.

Bona:2007:NSK

[1250] J. L. Bona, V. A. Dougalis, and D. E. Mitsotakis. Numerical solution of KdV–KdV systems of Boussinesq equations: I. The numerical scheme and generalized solitary waves. Mathematics and Computers in Simulation, 74 (2–3):214–228, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002643.

Ludu:2007:NDR

[1251] A. Ludu and P. G. Kevrekidis. Nonlinear dispersion relations. Mathematics and Computers in Simulation, 74(2– 3):229–236, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002680.

Biondini:2007:EIL

[1252] Gino Biondini and Sarbarish Chakravarty. Elastic and inelastic line-soliton solutions of the Kadomtsev–Petviashvili II equation. Mathematics and Computers in Simulation, 74(2–3):237– 250, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002473.

Anonymous:2007:NId

[1253] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 74(2–3):251–252, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407000778.

Anonymous:2007:ICEd

[1254] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 74(2–3):253–257, March 7, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540700078X.

Anonymous:2007:EBd

[1255] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 74(2–3):CO2, March 7, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000705.

Taha:2007:Fb

[1256] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 74 (4–5):265, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002771.

Anonymous:2007:TRTb

[1257] Anonymous. Thiab R. Taha. Mathematics and Computers in Simulation, 74(4–5):265–442, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Mancas:2007:BPW

[1258] Stefan C. Mancas and S. Roy Choudhury. Bifurcations of plane wave (CW) solutions in the complex cubic–quintic Ginzburg–Landau equation. Mathematics and Computers in Simulation, 74(4–5):266– 280, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002710.

Mancas:2007:CCQ

[1259] Stefan C. Mancas and S. Roy Choudhury. The complex cubic– quintic Ginzburg–Landau equation: Hopf bifurcations yielding traveling waves. Mathematics and Computers in Simulation, 74(4–5):281–291, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002424.

Feng:2007:ILM

[1260] Bao-Feng Feng and Youn-Sha Chan. Intrinsic localized modes in a three particle Fermi–Pasta–Ulam lattice with on-site harmonic potential. Mathematics and Computers in Simulation, 74(4–5):292–301, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475406002436.

Ismail:2007:LIC

[1261] M. S. Ismail and Thiab R. Taha. A linearly implicit conservative scheme for the coupled nonlinear Schrödinger equation. Mathematics and Computers in Simulation, 74(4–5):302– 311, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002461.

Albuch:2007:TBS

[1262] Lior Albuch and Boris A. Malomed. Transitions between symmetric and asymmetric solitons in dualcore systems with cubic–quintic nonlinearity. Mathematics and Computers in Simulation, 74(4–5):312– 322, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002497.

Francisco:2007:DHR

[1263] Meltem L. Y. Francisco, Jyh-Hao Lee, and Oktay K. Pashaev. Dissipative hierarchies and resonance solitons for KP-II and MKP-II. Mathematics and Computers in Simulation, 74(4– 5):323–332, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002539.

Proctor:2007:AMP

[1264] J. Proctor and J. Nathan Kutz. Averaged models for passive modelocking using nonlinear mode-coupling. Mathematics and Computers in Simulation, 74(4–5):333–342, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475406002576.

Kevrekidis:2007:CSD

[1265] P. G. Kevrekidis, S. V. Dmitriev, and A. A. Sukhorukov. On a class of spatial discretizations of equations of the nonlinear Schrödinger type. Mathematics and Computers in Simulation, 74(4– 5):343–351, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540600262X.

Hereman:2007:CDH

[1266] W. Hereman, B. Deconinck, and L. D. Poole. Continuous and discrete homotopy operators: a theoretical approach made concrete. Mathematics and Computers in Simulation, 74(4– 5):352–360, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002667.

Carretero-Gonzalez:2007:STV

[1267] R. Carretero-González, P. G. Kevrekidis, D. J. Frantzeskakis, B. A. Malomed, S. Nandi, and A. R. Bishop. Soliton trains and vortex streets as a form of Cerenkov radiation in trapped Bose– Einstein condensates. Mathematics and Computers in Simulation, 74(4– 5):361–369, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002679.

Deconinck:2007:SLN

[1268] Bernard Deconinck, Firat Kiyak, John D. Carter, and J. Nathan Kutz. SpectrUW: a laboratory for the numerical exploration of spectra of linear operators. Mathematics and Computers in Simulation, 74(4–5): 370–378, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002709.

Latas:2007:SSP

[1269] Sofia C. V. Latas and Mário F. S.

Ferreira. Stable soliton propagation with self-frequency shift. Mathematics and Computers in Simulation, 74(4– 5):379–387, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002606.

Zhao:2007:CPB

[1270] Jun-Xiao Zhao. Commutativity of Pfaffianization and Bäcklund transformation: the semi-discrete Toda equation. Mathematics and Computers in Simulation, 74(4–5):388– 396, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002588.

Muniz:2007:DRB

[1271] S. R. Muniz, D. S. Naik, M. Bhattacharya, and C. Raman. Dynamics of rotating Bose–Einstein condensates probed by Bragg scattering. Mathematics and Computers in Simulation, 74(4–5):397–404, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406002515.

Kevrekidis:2007:ACD

[1272] P. G. Kevrekidis, Avinash Khare, A. Saxena, I. Bena, and A. R. Bishop. Asymptotic calculation of discrete nonlinear wave interactions. Mathematics and Computers in Simulation, 74(4– 5):405–413, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002485.

Feng:2007:ASP

[1273] J. Feng and P. G. Kevrekidis. Averaging for some periodic and random nonlinear Schrödinger models. Mathematics and Computers in Simulation, 74(4–5):414–428, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540600245X.

Anonymous:2007:NIe

[1274] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 74(4–5):429–430, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407000912.

Anonymous:2007:ICEe

[1275] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 74(4–5):431–435, March 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000924.

Anonymous:2007:EBe

[1276] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 74(4–5):CO2, March 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000845.

Hirose:2007:MTM

[1277] Hideo Hirose. The mixed trunsored model with applications to SARS. Mathematics and Computers in Simulation, 74(6):443–453, April 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002151.

Wu:2007:ECI

[1278] James Shih-Shyn Wu, Jyh-Cheng Chang, and Jui-Pin Hung. The effect of contact interface on dynamic characteristics of composite structures. Mathematics and Computers in Simulation, 74(6):454–467, April 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002163.

Hoppe:2007:EPM

[1279] Ronald H. W. Hoppe and Svetozara I. Petrova. Elasto-plasticity model in structural optimization of composite materials with periodic microstructures. Mathematics and Computers in Simulation, 74(6):468– 480, April 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002175.

Anonymous:2007:NIf

[1280] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 74(6):481–482, April 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001048.

Anonymous:2007:ICEf

[1281] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 74(6):483–487, April 30, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540700105X.

Anonymous:2007:EBf

[1282] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 74(6):CO2, April 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000985.

Anonymous:2007:PA

[1283] Anonymous. Pages 443–494 (30 April 2007). Mathematics and Computers in Simulation, 74(6):??, April 30, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pelliccioni:2007:LBD

[1284] O. Pelliccioni, M. Cerrolaza, and M. Herrera. Lattice Boltzmann dynamic simulation of a mechanical heart valve device. Mathematics and Computers in Simulation, 75(1– 2):1–14, May 27, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002187.

Liu:2007:DDS

[1285] Xin-Ge Liu, Min Wu, Ralph Martin, and Mei-Lan Tang. Delaydependent stability analysis for uncertain neutral systems with timevarying delays. Mathematics and Computers in Simulation, 75(1–2):15– 27, May 27, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002199.

Fukuda:2007:JDU

[1286] Kosei Fukuda. Joint detection of unit roots and cointegration: Databased simulation. Mathematics and Computers in Simulation, 75(1–2):28– 36, May 27, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002345.

Botmart:2007:ACS

[1287] Thongchai Botmart and Piyapong Niamsup. Adaptive control and synchronization of the perturbed Chua’s system. Mathematics and Computers in Simulation, 75(1–2):37– 55, May 27, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406002357.

Anonymous:2007:NIg

[1288] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 75(1–2):56–57, May 27, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001188.

Anonymous:2007:ICEg

[1289] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 75(1–2):58– 62, May 27, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540700119X.

Anonymous:2007:EBg

[1290] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 75(1–2):CO2, May 27, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001127.

Anonymous:2007:PM

[1291] Anonymous. Pages 1–68 (27 May 2007). Mathematics and Computers in Simulation, 75(1–2):??, May 27, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kazmerchuk:2007:POS

[1292] Yuriy Kazmerchuk, Anatoliy Swishchuk, and Jianhong Wu. The pricing of options for securities markets with delayed response. Mathematics and Computers in Simulation, 75(3–4):69– 79, July 2, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002370.

Sobol:2007:QRP

[1293] I. M. Sobol and B. V. Shukhman.

Quasi-random points keep their distance. Mathematics and Computers in Simulation, 75(3–4):80– 86, July 2, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002382.

Firat:2007:RFE

[1294] Mahmut Firat and Mahmud Güngör.

River flow estimation using adaptive neuro fuzzy inference system. Mathematics and Computers in Simulation, 75(3–4):87–96, July 2, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002394.

Cheng:2007:MTR

[1295] Wei Cheng, Chu-Li Fu, and Zhi Qian. A modified Tikhonov regularization method for a spherically symmetric three-dimensional inverse heat conduction problem. Mathematics and Computers in Simulation, 75 (3–4):97–112, July 2, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002400.

Li:2007:ASL

[1296] Xian-Fang Li. Approximate solution of linear ordinary differential equations with variable coefficients. Mathematics and Computers in Simulation, 75 (3–4):113–125, July 2, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406002412.

Anonymous:2007:NIh

[1297] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 75(3–4):126–127, July 2, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001358.

Anonymous:2007:ICEh

[1298] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 75(3–4):128– 132, July 2, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540700136X.

Anonymous:2007:EBh

[1299] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 75(3–4):CO2, July 2, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001292.

Anonymous:2007:PJb

[1300] Anonymous. Pages 69–138 (2 July 2007). Mathematics and Computers in Simulation, 75(3–4):??, July 2, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Conti:2007:P

[1301] Costanza Conti, Rossana Morandi, Francesca Pistella, and Rosa Maria Spitaleri. Preface. Mathematics and Computers in Simulation, 75(5–6):139– 140, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406003326.

Bertolazzi:2007:DRK

[1302] Enrico Bertolazzi and Gianmarco Manzini. DIMEX Runge–Kutta finite volume methods for multidimensional hyperbolic systems. Mathematics and Computers in Simulation, 75(5–6):141– 160, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406003405.

Cheng:2007:MMS

[1303] Gary C. Cheng, Roy P. Koomullil, and Bharat K. Soni. Multidisciplinary and multi-scale computational field simulations — Algorithms and applications. Mathematics and Computers in Simulation, 75(5–6):161– 170, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406003363.

Egidi:2007:RDV

[1304] N. Egidi and P. Maponi. A robust direct variational approach for generation of quadrangular and triangular grids on planar domains. Mathematics and Computers in Simulation, 75(5–6): 171–181, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540600334X.

Egidi:2007:HEG

[1305] N. Egidi and P. Maponi. The heat equation as grid smoothing in a local optimization procedure. Mathematics and Computers in Simulation, 75(5–6): 182–190, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406003387.

Gori:2007:RIQ

[1306] Laura Gori, Francesca Pitolli, and Elisabetta Santi. Refinable interpolatory and quasi-interpolatory operators. Mathematics and Computers in Simulation, 75(5–6):191–199, September 5, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406003338.

Ito:2007:PUM

[1307] Yasushi Ito, Alan M. Shih, Anil K. Erukala, Bharat K. Soni, Andrey Chernikov, Nikos P. Chrisochoides, and Kazuhiro Nakahashi. Parallel unstructured mesh generation by an advancing front method. Mathematics and Computers in Simulation, 75(5–6):200– 209, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406003351.

Makhanov:2007:OCTb

[1308] Stanislav Makhanov. Optimization and correction of the tool path of the fiveaxis milling machine: Part 1. Spatial optimization. Mathematics and Computers in Simulation, 75(5–6):210– 230, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406003375.

Makhanov:2007:OCTa

[1309] Stanislav Makhanov. Optimization and correction of the tool path of the fiveaxis milling machine: Part 2: Rotations and setup. Mathematics and Computers in Simulation, 75(5–6):231– 250, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406003314.

Padron:2007:RBL

[1310] Miguel A. Padrón, José P. Suárez, and Ángel Plaza. Refinement based on longest-edge and self-similar fourtriangle partitions. Mathematics and Computers in Simulation, 75(5–6):251– 262, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406003399.

Anonymous:2007:Na

[1311] Anonymous. News. Mathematics and Computers in Simulation, 75(5–6):263– 264, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001814.

Anonymous:2007:Ca

[1312] Anonymous. Calendar. Mathematics and Computers in Simulation, 75(5–6): 265–270, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001826.

Anonymous:2007:EBi

[1313] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 75(5–6):CO2, September 5, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475407001747.

Danek:2007:P

[1314] Josef Danĕk, Marek Brandner, and Jiří Nedoma. Preface. Mathematics and Computers in Simulation, 76(1– 3):1–2, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407000134.

Baboolal:2007:TSN

[1315] S. Baboolal and R. Bharuthram. Twoscale numerical solution of the electromagnetic two-fluid plasma-Maxwell equations: Shock and soliton simulation. Mathematics and Computers in Simulation, 76(1–3):3–7, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000092.

Benes:2007:QPS

[1316] Michal Beneš. The qualitative properties of the Stokes and Navier–Stokes system for the mixed problem in a nonsmooth domain. Mathematics and Computers in Simulation, 76(1–3): 8–12, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001437.

Benedikt:2007:SFO

[1317] Jiří Benedikt. Spectra of fourth-order quasilinear problems. Mathematics and Computers in Simulation, 76(1– 3):13–17, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000390.

Blaheta:2007:SMD

[1318] R. Blaheta, R. Kohut, M. Neytcheva, and J. Starý. Schwarz methods for discrete elliptic and parabolic problems with an application to nuclear waste repository modelling. Mathematics and Computers in Simulation, 76(1– 3):18–27, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407000407.

Burda:2007:NSF

[1319] P. Burda, J. Novotný, and J. Šístek. Numerical solution of flow problems by stabilized finite element method and verification of its accuracy using a posteriori error estimates. Mathematics and Computers in Simulation, 76(1– 3):28–33, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407000171.

Cimrman:2007:MPD

[1320] Robert Cimrman and Eduard Rohan. On modelling the parallel diffusion flow in deforming porous media. Mathematics and Computers in Simulation, 76(1–3):34–43, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407000274.

Cortes:2007:CMS

[1321] J. C. Cortés, L. Jódar, L. Villafuerte, and R. J. Villanueva. Computing mean square approximations of random diffusion models with source term. Mathematics and Computers in Simulation, 76(1–3):44–48, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407000304.

Danek:2007:NMW

[1322] Josef Daněk, Jiří Nedoma, Ivan Hlaváček, Pavel Vavřík, and František Denk. Numerical modelling of the weight-bearing total knee joint replacement and usage in practice. Mathematics and Computers in Simulation, 76(1–3):49–56, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000183.

Dostal:2007:FBA

[1323] Zdeněk Dostál, David Horák, and Oldřich Vlach. FETI-based algorithms for modelling of fibrous composite materials with debonding. Mathematics and Computers in Simulation, 76(1– 3):57–64, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000389.

Gatskevich:2007:MLI

[1324] Elena Gatskevich, Petr Přikryl, and Gennadii Ivlev. Modelling laserinduced phase transformations in semiconductors. Mathematics and Computers in Simulation, 76(1–3):65–72, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000122.

Gruber:2007:IES

[1325] Peter Gruber and Jan Valdman. Implementation of an elastoplastic solver based on the Moreau–Yosida Theorem. Mathematics and Computers in Simulation, 76(1–3):73–81, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000225.

Herrmann:2007:CPS

[1326] Leopold Herrmann. Conjugate points of second order ordinary differential equations with jumping nonlinearities. Mathematics and Computers in Simulation, 76(1–3):82–85, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000353.

Janka:2007:DDS

[1327] Aleš Janka. A domain decomposition solver for ferromagnetism. Mathematics and Computers in Simulation, 76(1–3):86–93, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407000213.

Jerez:2007:NMS

[1328] S. Jerez, J. V. Romero, M. D. Roselló, and J. M. Arnau. A nonuniform mesh semi-implicit CE– SE method modelling unsteady flow in tapered ducts. Mathematics and Computers in Simulation, 76(1–3): 94–98, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000080.

Karatson:2007:DMP

[1329] János Karátson, Sergey Korotov, and Michal Křížek. On discrete maximum principles for nonlinear elliptic problems. Mathematics and Computers in Simulation, 76(1–3):99–108, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540700016X.

Kroc:2007:MMD

[1330] Jiří Kroc. Modelling of morphological development of tooth using simple regulatory network: mechanical model of mesenchyme. Mathematics and Computers in Simulation, 76(1–3): 109–115, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000109.

Kren:2007:MNN

[1331] Jiří Křen and Luděk Hynčík. Modelling of non-Newtonian fluids. Mathematics and Computers in Simulation, 76(1–3): 116–123, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000110.

Lobovsky:2007:SPH

[1332] Libor Lobovský and Jan Vimmr.

Smoothed particle hydrodynamics and finite volume modelling of incompressible fluid flow. Mathematics and Computers in Simulation, 76(1–3):124– 131, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001401.

Lopez-Gomez:2007:MC

[1333] J. López-Gómez and M. Molina-Meyer. Modeling coopetition. Mathematics and Computers in Simulation, 76(1–3): 132–140, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407000237.

Mahapatra:2007:FEA

[1334] D. Roy Mahapatra and R. V. N. Melnik. Finite element approach to modelling evolution of 3D shape memory materials. Mathematics and Computers in Simulation, 76(1–3):141– 148, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407000195.

Margenov:2007:MPN

[1335] S. Margenov and P. Minev. On a MIC(0) preconditioning of nonconforming mixed FEM elliptic problems. Mathematics and Computers in Simulation, 76(1–3):149–154, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407000328.

DiMartino:2007:SOS

[1336] B. Di Martino and M. Peybernes. Simulation of an oil slick movement using a shallow water model. Mathematics and Computers in Simulation, 76(1–3):

155–160, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700033X.

Mosova:2007:MRM

[1337] Vratislava Mošová. Meshless RKHPU method and its applications. Mathematics and Computers in Simulation, 76(1–3):161–165, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407000341.

Prokop:2007:NSB

[1338] V. Prokop and K. Kozel. Numerical solution of bypass flow. Mathematics and Computers in Simulation, 76(1–3): 166–170, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000201.

Rahimov:2007:SMC

[1339] Ibrahim Rahimov. On a stochastic model for continuous mass branching population. Mathematics and Computers in Simulation, 76(1–3):171– 176, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000316.

Rahimov:2007:NPI

[1340] I. Rahimov and S. Kurbanov. On the number of productive individuals in the Bienayme–Galton–Watson process with immigration. Mathematics and Computers in Simulation, 76(1–3):

177–180, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000365.

Rezzonico:2007:MAP

[1341] Vittoria Rezzonico, Alexei Lozinski, Marco Picasso, Jacques Rappaz, and Joël Wagner. Multiscale algorithm with patches of finite elements. Mathematics and Computers in Simulation, 76(1–3): 181–187, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001413.

Rus:2007:PNM

[1342] Francisco Rus and Francisco R. Villatoro. Padé numerical method for the Rosenau–Hyman compacton equation. Mathematics and Computers in Simulation, 76(1–3):188–192, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000250.

Sobotikova:2007:NID

[1343] Veronika Sobotíková and Miloslav Feistauer. Numerical integration in the DGFEM for nonlinear convection– diffusion problems. Mathematics and Computers in Simulation, 76(1–3):193– 197, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000286.

Solin:2007:HHO

[1344] Pavel Šolín and Karel Segeth. Hierarchic higher-order Hermite elements on hybrid triangular/quadrilateral meshes. Mathematics and Computers in Simulation, 76(1–3):198–204, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000262.

Solin:2007:DCN

[1345] Pavel Šolín, Tomáš Vejchodský, and Roberto Araiza. Discrete conservation of nonnegativity for elliptic problems solved by the hp-FEM. Mathematics and Computers in Simulation, 76(1–3): 205–210, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000249.

Solin:2007:IOH

[1346] Pavel Šolín, Tomáš Vejchodský, Martin Zítka, and Francisco Ávila. Imposing orthogonality to hierarchic higherorder finite elements. Mathematics and Computers in Simulation, 76(1–3): 211–217, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000377.

Alvarez-Vazquez:2007:OSD

[1347] L. J. Alvarez-Vázquez, A. Martínez, C. Rodríguez, M. E. Vázquez-Méndez, and M. A. Vilar. Optimal shape design for fishways in rivers. Mathematics and Computers in Simulation, 76(1–3):

218–222, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407000298.

Vejchodsky:2007:MHF

[1348] Tomáš Vejchodský, Pavel Šolín, and Martin Zítka. Modular hp-FEM system HERMES and its application to Maxwell’s equations. Mathematics and Computers in Simulation, 76(1–3): 223–228, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001425.

Vimmr:2007:MCC

[1349] Jan Vimmr. Modelling of complex clearance flow in screw-type machines. Mathematics and Computers in Simulation, 76(1–3):229–236, October 12, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407000079.

Anonymous:2007:Nb

[1350] Anonymous. News. Mathematics and Computers in Simulation, 76(1–3): 237–238, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407002236.

Anonymous:2007:Cb

[1351] Anonymous. Calendar. Mathematics and Computers in Simulation, 76(1–3): 239–242, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002248.

Anonymous:2007:EBj

[1352] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 76(1–3):CO2, October 12, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407002169.

Xiao:2007:AMM

[1353] Yingxiong Xiao, Ping Zhang, and Shi Shu. Algebraic multigrid methods for elastic structures with highly discontinuous coefficients. Mathematics and Computers in Simulation, 76(4): 249–262, December 1, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475406002758.

Chen:2007:MBC

[1354] Shyh-Wei Chen. Measuring business cycle turning points in Japan with the Markov Switching Panel model. Mathematics and Computers in Simulation, 76(4):263–270, December 1, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475406002783.

Bratsos:2007:TON

[1355] A. G. Bratsos. A third order numerical scheme for the two-dimensional sine-Gordon equation. Mathematics and Computers in Simulation, 76(4): 271–282, December 1, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475406003296.

Triki:2007:ETD

[1356] Houria Triki. Energy transfer in a dispersion-managed Korteweg–de Vries system. Mathematics and Computers in Simulation, 76(4):283–292, December 1, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475406003302.

Lee:2007:TLC

[1357] Chien-Chiang Lee, Pei-Fen Chen, and Chun-Ping Chang. Testing linearity in a cointegrating STR model for the money demand function: International evidence from G-7 countries. Mathematics and Computers in Simulation, 76(4):293–302, December 1, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540700002X.

Anonymous:2007:NIi

[1358] Anonymous. News of Imacs. Mathematics and Computers in Simulation, 76(4):303–304, December 1, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475407002765.

Anonymous:2007:ICEi

[1359] Anonymous. Imacs Calendar of Events. Mathematics and Computers in Simulation, 76(4):305–308, December 1, 2007. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002777.

Anonymous:2007:EBk

[1360] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 76(4):CO2, December 1, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002716.

Anonymous:2007:PD

[1361] Anonymous. Pages 249–314 (1 December 2007). Mathematics and Computers in Simulation, 76(4):??, December 1, 2007. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Borne:2008:P

[1362] Pierre Borne. Preface. Mathematics and Computers in Simulation, 76 (5–6):315, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001656.

Bouzaouache:2008:SAN

[1363] Hajer Bouzaouache and Naceur Benhadj Braiek. On the stability analysis of nonlinear systems using polynomial Lyapunov functions. Mathematics and Computers in Simulation, 76(5–6): 316–329, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001644.

Gruyitch:2008:SAO

[1364] Lyubomir T. Gruyitch and William Pratt Mounfield. Stablewise absolute output natural tracking control with finite reachability time: MIMO Lurie systems. Mathematics and Computers in Simulation, 76(5–6):330– 344, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001899.

Zgaya:2008:MSM

[1365] Hayfa Zgaya, Slim Hammadi, and Khaled Ghédira. A migration strategy of mobile agents for the transport network applications. Mathematics and Computers in Simulation, 76(5–6): 345–362, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001887.

Laabidi:2008:MCO

[1366] Kaouther Laabidi, Faouzi Bouani, and Mekki Ksouri. Multi-criteria optimization in nonlinear predictive control. Mathematics and Computers in Simulation, 76(5–6):363–374, January 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407001681.

DiLoreto:2008:SRD

[1367] M. Di Loreto, J. F. Lafay, and J. J. Loiseau. Some remarks on duality over a commutative ring. Mathematics and Computers in Simulation, 76(5–6):

375–387, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001693.

Thiery:2008:INN

[1368] Frédérik Thiery, Stéphane Grieu, Adama Traoré, Mathieu Barreau, and Monique Polit. Integration of neural networks in a geographical information system for the monitoring of a catchment area. Mathematics and Computers in Simulation, 76(5–6): 388–397, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001711.

Guesmi:2008:SNP

[1369] K. Guesmi, N. Essounbouli, A. Hamzaoui, J. Zaytoon, and N. Manamanni. Shifting nonlinear phenomena in a DC–DC converter using a fuzzy logic controller. Mathematics and Computers in Simulation, 76(5–6):398– 409, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700167X.

Benrejeb:2008:CCO

[1370] Mohamed Benrejeb, Anis Sakly, Kamel Ben Othman, and Pierre Borne. Choice of conjunctive operator of TSK fuzzy systems and stability domain study. Mathematics and Computers in Simulation, 76(5–6):410– 421, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001668.

Filip:2008:ODC

[1371] Florin Filip, Dumitru Popescu, and Mihaela Mateescu. Optimal decisions for complex systems — Software packages. Mathematics and Computers in Simulation, 76(5–6): 422–429, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001723.

Al-Mutairi:2008:TCF

[1372] M. S. Al-Mutairi, K. W. Hipel, and M. S. Kamel. Trust and cooperation from a fuzzy perspective. Mathematics and Computers in Simulation, 76(5–6): 430–446, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001632.

Saad:2008:CIC

[1373] Ihsen Saad, Slim Hammadi, Mohamed Benrejeb, and Pierre Borne. Choquet integral for criteria aggregation in the flexible job-shop scheduling problems. Mathematics and Computers in Simulation, 76(5–6):447– 462, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001905.

Anonymous:2008:NIa

[1374] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 76(5–6):463–464, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407003333.

Anonymous:2008:ICEa

[1375] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 76(5–6):465–469, January 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003345.

Anonymous:2008:EBa

[1376] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 76(5–6):CO2, January 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407003266.

Leung:2008:TFP

[1377] Pui-Lam Leung and Wing-Keung Wong. Three-factor profile analysis with GARCH innovations. Mathematics and Computers in Simulation, 77 (1):1–8, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000031.

Simpson:2008:TAP

[1378] Matthew J. Simpson and Kerry A.

Landman. Theoretical analysis and physical interpretation of temporal truncation errors in operator split algorithms. Mathematics and Computers in Simulation, 77(1):9–21, February 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407000055.

Cerimele:2008:NSE

[1379] M. M. Cerimele, F. Pistella, and V. Valente. Numerical study of an evolutive model for magnetostrictive materials. Mathematics and Computers in Simulation, 77(1):22– 33, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000067.

Feng:2008:EPM

[1380] Xinlong Feng, Guoliang He, and Abdurishit. Estimation of parameters of the Makeham distribution using the least squares method. Mathematics and Computers in Simulation, 77(1): 34–44, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000146.

Bruti-Liberati:2008:HGM

[1381] Nicola Bruti-Liberati, Filippo Martini, Massimo Piccardi, and Eckhard Platen. A hardware generator of multipoint distributed random numbers for Monte Carlo simulation. Mathematics and Computers in Simulation, 77 (1):45–56, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407000158.

Withers:2008:MTS

[1382] C. S. Withers, D. P. Krouse, C. P.

Pearson, and S. Nadarajah. Modelling time series when mean and variability both change. Mathematics and Computers in Simulation, 77(1): 57–63, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001255.

Burq:2008:SBM

[1383] Zaeem A. Burq and Owen D. Jones.

Simulation of Brownian motion at first-passage times. Mathematics and Computers in Simulation, 77(1):64– 71, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001267.

Hagmark:2008:CSC

[1384] Per-Erik Hagmark. On construction and simulation of count data models. Mathematics and Computers in Simulation, 77(1):72–80, February 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407001279.

Mandal:2008:PBG

[1385] Manisankar Mandal and Surjya K. Pal.

Pseudo-bond graph modelling of temperature distribution in a throughprocess steel rolling. Mathematics and Computers in Simulation, 77(1): 81–95, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001450.

So:2008:EEF

[1386] Mike K. P. So, Cathy W. S. Chen, Jen-Yu Lee, and Yi-Ping Chang. An empirical evaluation of fat-tailed distributions in modeling financial time series. Mathematics and Computers in Simulation, 77(1):96–108, February 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407001486.

Cook:2008:JML

[1387] Steven Cook. Joint maximum likelihood estimation of unit root testing equations and GARCH processes: Some finite-sample issues. Mathematics and Computers in Simulation, 77(1): 109–116, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001498.

Solin:2008:ALH

[1388] Pavel Šolín, Jakub Červený, and Ivo Doležel. Arbitrary-level hanging nodes and automatic adaptivity in the hp-FEM. Mathematics and Computers in Simulation, 77(1):117– 132, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001504.

Anonymous:2008:NIb

[1389] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 77(1):133–134, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407003448.

Anonymous:2008:ICEb

[1390] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 77(1):135–139, February 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540700345X.

Anonymous:2008:EBb

[1391] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 77(1):CO2, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407003382.

Anonymous:2008:PF

[1392] Anonymous. Pages 1–140 (15 February 2008). Mathematics and Computers in Simulation, 77(1):??, February 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Amaziane:2008:P

[1393] Brahim Amaziane, Domingo Barrera, and Driss Sbibih. Preface. Mathematics and Computers in Simulation, 77 (2–3):141, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002297.

Addou:2008:PAB

[1394] A. Addou, A. Lidouh, and B. Seddoug. A projection algorithm for bilateral obstacle problem. Mathematics and Computers in Simulation, 77(2– 3):142–150, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002303.

Ameur:2008:NBO

[1395] El Bachir Ameur, Domingo Barrera, María J. Ibáñez, and Driss Sbibih. Near-best operators based on a C^{2} quartic spline on the uniform fourdirectional mesh. Mathematics and Computers in Simulation, 77(2–3): 151–160, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002315.

Barrera:2008:MES

[1396] D. Barrera, M. A. Fortes, P. González, and M. Pasadas. Minimal energy C^{r}- surfaces on uniform Powell–Sabin type meshes: Estimation of the smoothing parameters. Mathematics and Computers in Simulation, 77(2–3): 161–169, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002327.

Cardenas-Morales:2008:ICB

[1397] D. Cárdenas-Morales and F. J.

Muñoz-Delgado. Improving certain Bernstein-type approximation processes. Mathematics and Computers in Simulation, 77(2–3):170–178, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002339.

El-Kyal:2008:SCA

[1398] M. El-Kyal and A. Machmoum. Superlinear convergence of asynchronous multi-splitting waveform relaxation methods applied to a system of nonlinear ordinary differential equations. Mathematics and Computers in Simulation, 77(2–3):179–188, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002340.

Ellabib:2008:IAS

[1399] A. Ellabib and A. Nachaoui. An iterative approach to the solution of an inverse problem in linear elasticity. Mathematics and Computers in Simulation, 77(2–3):189–201, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002352.

Foucher:2008:APD

[1400] Françoise Foucher and Paul Sablonnière.

Approximating partial derivatives of first and second order by quadratic spline quasi-interpolants on uniform meshes. Mathematics and Computers in Simulation, 77(2–3):202– 208, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002364.

Gutierrez:2008:TAC

[1401] R. Gutiérrez, R. Gutiérrez-Sánchez, and A. Nafidi. Trend analysis and computational statistical estimation in a stochastic Rayleigh model: Simulation and application. Mathematics and Computers in Simulation, 77(2– 3):209–217, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002376.

Ibanez-Perez:2008:CTD

[1402] María José Ibáñez-Pérez. On Chebyshev-type discrete quasi-interpolants. Mathematics and Computers in Simulation, 77(2–3):218–227, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475407002388.

Kouibia:2008:BVS

[1403] A. Kouibia and M. Pasadas. Bivariate variational splines with monotonicity constraints. Mathematics and Computers in Simulation, 77(2–3): 228–236, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700239X.

Lamnii:2008:SSS

[1404] A. Lamnii, H. Mraoui, D. Sbibih, and A. Tijini. Sextic spline solution of fifth-order boundary value problems. Mathematics and Computers in Simulation, 77(2–3):237– 246, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002613.

deSilanes:2008:NCF

[1405] María Cruz López de Silanes, María Cruz Parra, and Juan José Torrens. On a new characterization of finite jump discontinuities and its application to vertical fault detection. Mathematics and Computers in Simulation, 77(2– 3):247–256, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002406.

Marpeau:2008:SCP

[1406] F. Marpeau. Simulating the contamination of populations by a polluted environment. Mathematics and Computers in Simulation, 77(2–3): 257–265, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002418.

Moncayo:2008:RPO

[1407] M. Moncayo and R. J. Yáñez.

A recursive procedure to obtain a class of orthogonal polynomial wavelets. Mathematics and Computers in Simulation, 77(2–3):266– 273, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540700242X.

Palomares:2008:CRL

[1408] A. Palomares, M. Pasadas, V. Ramírez, and M. Ruiz Galán. A convergence result for a least-squares method using Schauder bases. Mathematics and Computers in Simulation, 77(2– 3):274–281, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002431.

Pasadas:2008:CBS

[1409] M. Pasadas and M. L. Rodríguez. Construction of blending surfaces by parametric discrete interpolation PDE splines. Mathematics and Computers in Simulation, 77(2–3):282– 290, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002443.

Ammi:2008:NAN

[1410] Moulay Rchid Sidi Ammi and Delfim F. M. Torres. Numerical analysis of a nonlocal parabolic problem resulting from thermistor problem. Mathematics and Computers in Simulation, 77 (2–3):291–300, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002455.

Tber:2008:PIS

[1411] Moulay Hicham Tber, Mohamed El Alaoui Talibi, and Driss Ouazar. Parameters identification in a seawater intrusion model using adjoint sensitive method. Mathematics and Computers in Simulation, 77(2–3): 301–312, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002467.

Zidna:2008:TSA

[1412] Ahmed Zidna and Dominique Michel.

A two-steps algorithm for approximating real roots of a polynomial in Bernstein basis. Mathematics and Computers in Simulation, 77(2–3): 313–323, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002479.

Anonymous:2008:NIc

[1413] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 77(2–3):324–325, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408000190.

Anonymous:2008:ICEc

[1414] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 77(2–3):326–330, March 7, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000207.

Anonymous:2008:EBc

[1415] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 77(2–3):CO2, March 7, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000128.

Mohamad:2008:CSE

[1416] Sannay Mohamad. Computer simulations of exponentially convergent networks with large impulses. Mathematics and Computers in Simulation, 77 (4):331–344, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001516.

Marseguerra:2008:MCE

[1417] M. Marseguerra and A. Zoia. Monte Carlo evaluation of FADE approach to anomalous kinetics. Mathematics and Computers in Simulation, 77 (4):345–357, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001541.

Lee:2008:SBT

[1418] Chien-Chiang Lee and Mei-Se Chien. Structural breaks, tourism development, and economic growth: Evidence from Taiwan. Mathematics and Computers in Simulation, 77(4): 358–368, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001553.

Narayan:2008:DSG

[1419] Paresh Kumar Narayan. Do shocks to G7 stock prices have a permanent effect?: Evidence from panel unit root tests with structural change. Mathematics and Computers in Simulation, 77(4):369–373, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001565.

Hung:2008:PSC

[1420] Yung-Ching Hung, Jun-Juh Yan, and Teh-Lu Liao. Projective synchronization of Chua’s chaotic systems with dead-zone in the control input. Mathematics and Computers in Simulation, 77(4):374–382, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001607.

Leiderman:2008:HRM

[1421] Karin Leiderman and Stanly Steinberg. High-resolution models of motion of macromolecules in cell membranes. Mathematics and Computers in Simulation, 77(4):383–399, April 4, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407001620.

Lewis:2008:HAD

[1422] Robert H. Lewis. Heuristics to accelerate the Dixon resultant. Mathematics and Computers in Simulation, 77 (4):400–407, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540700170X.

Debrabant:2008:CSR

[1423] Kristian Debrabant and Andreas Rößler. Classification of stochastic Runge–Kutta methods for the weak approximation of stochastic differential equations. Mathematics and Computers in Simulation, 77(4):408– 420, April 4, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407001929.

Deng:2008:ICF

[1424] Zui-Cha Deng, Jian-Ning Yu, and Liu Yang. Identifying the coefficient of first-order in parabolic equation from final measurement data. Mathematics and Computers in Simulation, 77 (4):421–435, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000086.

Anonymous:2008:NId

[1425] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 77(4):436–437, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000712.

Anonymous:2008:ICEd

[1426] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 77(4):438–442, April 4, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000724.

Anonymous:2008:EBd

[1427] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 77(4):CO2, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000657.

Anonymous:2008:PAa

[1428] Anonymous. Pages 331–442 (4 April 2008). Mathematics and Computers in Simulation, 77(4):??, April 4, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Zeng:2008:ESI

[1429] Xianyi Zeng and Yongsheng Ding. Editorial of the special issue on intelligent sensory evaluation. Mathematics and Computers in Simulation, 77 (5–6):vi–vii, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003503.

Zeng:2008:ISE

[1430] Xianyi Zeng, Da Ruan, and Ludovic Koehl. Intelligent sensory evaluation: Concepts, implementations, and applications. Mathematics and Computers in Simulation, 77(5–6): 443–452, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003515.

Kissi:2008:PRS

[1431] Mohamed Kissi, Mohammed Ramdani, Bernadette Bouchon-Meunier, and Driss Zakarya. Pattern recognition system based on empirical knowledge: Sandalwood and camphoraceous odours application. Mathematics and Computers in Simulation, 77 (5–6):453–463, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003527.

Buyukozkan:2008:ESD

[1432] Gülçin Büyüközkan and Da Ruan. Evaluation of software development projects using a fuzzy multi-criteria decision approach. Mathematics and Computers in Simulation, 77(5–6): 464–475, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003539.

Feng:2008:ITC

[1433] Tian-Jin Feng, Lin-Tao Ma, Xiang-Qian Ding, Ning Yang, and Xiezhong Xiao. Intelligent techniques for cigarette formula design. Mathematics and Computers in Simulation, 77 (5–6):476–486, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003540.

Tsutsumi:2008:SSC

[1434] Kazutoshi Tsutsumi and Keisuke Sasaki. Study on shape creation of building’s roof by evaluating aesthetic sensibility. Mathematics and Computers in Simulation, 77(5–6): 487–498, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003552.

Renaud:2008:WDO

[1435] Jean Renaud, Eric Levrat, and Christian Fonteix. Weights determination of OWA operators by parametric identification. Mathematics and Computers in Simulation, 77(5–6): 499–511, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003564.

Liu:2008:FMC

[1436] Xiaohong Liu, Xianyi Zeng, Yang Xu, and Ludovic Koehl. A fuzzy model of customer satisfaction index in e-commerce. Mathematics and Computers in Simulation, 77(5– 6):512–521, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003576.

Koehl:2008:ISE

[1437] Ludovic Koehl, Xianyi Zeng, Bin Zhou, and Yongsheng Ding. Intelligent sensory evaluation of industrial products for exploiting consumer’s preference. Mathematics and Computers in Simulation, 77(5–6):522– 530, May 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003588.

Jie:2008:FDR

[1438] Yang Jie, Ling Xufeng, Zhu Yitan, and Zheng Zhonglong. A face detection and recognition system in color image series. Mathematics and Computers in Simulation, 77(5–6): 531–539, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700359X.

Hou:2008:IBE

[1439] Caihong Hou, Yongsheng Ding, and Xianyi Zeng. Immune-based evolutionary algorithm for fabric evaluation. Mathematics and Computers in Simulation, 77(5–6):540– 549, May 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003606.

Al-Rawi:2008:UGF

[1440] Mohammed Al-Rawi and Jie Yang. Using Gabor filter for the illumination invariant recognition of color texture. Mathematics and Computers in Simulation, 77(5–6):550– 555, May 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003618.

Anonymous:2008:NIe

[1441] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 77(5–6):556–557, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847540800092X.

Anonymous:2008:ICEe

[1442] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 77(5–6):558– 562, May 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000931.

Anonymous:2008:EBe

[1443] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 77(5–6):CO2, May 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000852.

Dimitrov:2008:NFD

[1444] Dobromir T. Dimitrov and Hristo V.

Kojouharov. Nonstandard finitedifference methods for predator– prey models with general functional response. Mathematics and Computers in Simulation, 78(1):1– 11, June 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407001966.

Petkovic:2008:CCG

[1445] Miodrag S. Petković and Dušan M. Milošević. On the convergence condition of generalized root iterations for the inclusion of polynomial zeros. Mathematics and Computers in Simulation, 78(1):12– 26, June 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407001978.

Turk:2008:SDS

[1446] Tomaž Turk. System dynamics simulation of computer networks: Pricecontrolled QoS framework. Mathematics and Computers in Simulation, 78(1):27–39, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002005.

Kuo:2008:USM

[1447] Yiyo Kuo, Taho Yang, Chiwoon Cho, and Yao-Ching Tseng. Using simulation and multi-criteria methods to provide robust solutions to dispatching problems in a flow shop with multiple processors. Mathematics and Computers in Simulation, 78(1):40–56, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002017.

Kang:2008:ASS

[1448] M. Z. Kang, P. H. Cournède, P. de Reffye, D. Auclair, and B. G. Hu. Analytical study of a stochastic plant growth model: Application to the GreenLab model. Mathematics and Computers in Simulation, 78(1):57–75, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002029.

Glasa:2008:EMF

[1449] Jan Glasa and Ladislav Halada. On elliptical model for forest fire spread modeling and simulation. Mathematics and Computers in Simulation, 78(1):76–88, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002030.

Hernando:2008:SSR

[1450] Antonio Hernando, Luis De Ledesma, and Luis M. Laita. A system simulating representation change phenomena while problem solving. Mathematics and Computers in Simulation, 78(1):89–106, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002066.

Rababah:2008:BOP

[1451] Abedallah Rababah. Bivariate orthogonal polynomials on triangular domains. Mathematics and Computers in Simulation, 78(1):107– 111, June 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002078.

BerguesCabrales:2008:MMT

[1452] Luis Enrique Bergues Cabrales, Andrés Ramírez Aguilera, Rolando Placeres Jiménez, Manuel Verdecia Jarque, Héctor Manuel Camué Ciria, Juan Bory Reyes, Miguel Angel O’Farril Mateus, Fabiola Suárez Palencia, and Marisela González Ávila. Mathematical modeling of tumor growth in mice following low-level direct electric current. Mathematics and Computers in Simulation, 78(1):112– 120, June 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540700208X. See

corrigendum [1931].

Anonymous:2008:NIf

[1453] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 78(1):121–122, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001523.

Anonymous:2008:ICEf

[1454] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 78(1):123– 127, June 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001535.

Anonymous:2008:EBf

[1455] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 78(1):CO2, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001468.

Anonymous:2008:PJ

[1456] Anonymous. Pages 1–134 (June 2008).

Mathematics and Computers in Simulation, 78(1):??, June 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Argent:2008:P

[1457] Robert Argent, Michael McAleer, Les Oxley, and Andre Zerger. Preface. Mathematics and Computers in Simulation, 78(2–3):135–136, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408000232.

Ekasingh:2008:SFE

[1458] B. Ekasingh and R. A. Letcher. Successes and failures to embed socioeconomic dimensions in integrated natural resource management modeling: Lessons from Thailand. Mathematics and Computers in Simulation, 78 (2–3):137–145, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000244.

Allen:2008:LRU

[1459] D. E. Allen and V. Soucik. Long-run underperformance of seasoned equity offerings: Fact or an illusion? Mathematics and Computers in Simulation, 78(2–3):146–154, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000256.

daVeiga:2008:MVT

[1460] Bernardo da Veiga, Felix Chan, and Michael McAleer. Modelling the volatility transmission and conditional correlations between A and B shares in forecasting value-at-risk. Mathematics and Computers in Simulation, 78(2–3):155–171, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000268.

Chan:2008:TAE

[1461] W. S. Chan, S. H. Cheung, L. X. Zhang, and K. H. Wu. Temporal aggregation of equity return timeseries models. Mathematics and Computers in Simulation, 78(2–3): 172–180, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800027X.

Chow:2008:BMR

[1462] Ying-Foon Chow, Ming Liu, and Xinting Fan. Broad-market return persistence and momentum profits. Mathematics and Computers in Simulation, 78(2–3):181–188, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000281.

Hoti:2008:MVE

[1463] Suhejla Hoti, Michael McAleer, and Laurent L. Pauwels. Multivariate volatility in environmental finance. Mathematics and Computers in Simulation, 78(2–3):189– 199, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000293.

Lim:2008:CAP

[1464] Lee K. Lim. A cointegration analysis of price transmission between ADRs and dually listed South Korean stocks. Mathematics and Computers in Simulation, 78(2–3):200– 208, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800030X.

Asai:2008:PSI

[1465] Manabu Asai, Michael McAleer, and Bernardo de Veiga. Portfolio single index (PSI) multivariate conditional and stochastic volatility models. Mathematics and Computers in Simulation, 78(2–3):209–214, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000311.

McKenzie:2008:URS

[1466] C. R. McKenzie and Sumiko Takaoka.

Underwriter reputation and switching. Mathematics and Computers in Simulation, 78(2–3):215–222, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408000323.

Maekawa:2008:JDM

[1467] Koichi Maekawa, Sangyeol Lee, Takayuki Morimoto, and Ken ichi Kawai. Jump diffusion model with application to the Japanese stock market. Mathematics and Computers in Simulation, 78(2–3):223– 236, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000335.

Watkins:2008:HVM

[1468] Clinton Watkins and Michael McAleer. How has volatility in metals markets changed? Mathematics and Computers in Simulation, 78(2–3): 237–249, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000347.

Belkar:2008:MCC

[1469] R. Belkar and D. G. Fiebig. A Monte Carlo comparison of estimators for a bivariate probit model with selection. Mathematics and Computers in Simulation, 78(2–3): 250–256, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000359.

George:2008:MIN

[1470] Donald A. R. George and Les Oxley. Money and inflation in a nonlinear model. Mathematics and Computers in Simulation, 78(2–3): 257–265, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000360.

Horne:2008:ECE

[1471] Jocelyn Horne and Baiding Hu. Estimation of cost efficiency of Australian universities. Mathematics and Computers in Simulation, 78(2–3): 266–275, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000372.

Kakamu:2008:SIC

[1472] Kazuhiko Kakamu, Wolfgang Polasek, and Hajime Wago. Spatial interaction of crime incidents in Japan. Mathematics and Computers in Simulation, 78(2–3):276–282, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000384.

Marinova:2008:CRF

[1473] Dora Marinova and Peter Newman. The changing research funding regime in Australia and academic productivity. Mathematics and Computers in Simulation, 78(2–3): 283–291, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000396.

Morimune:2008:THL

[1474] K. Morimune and Y. Hoshino. Testing homogeneity of a large data set by bootstrapping. Mathematics and Computers in Simulation, 78(2–3): 292–302, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000402.

Nawata:2008:AHF

[1475] Kazumitsu Nawata, Ayako Nitta, Sonoko Watanabe, and Koichi Kawabuchi. An analysis of hip fracture treatments in Japan by the discretetype proportional hazard and ordered probit models. Mathematics and Computers in Simulation, 78 (2–3):303–312, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000414.

Carlaw:2008:RPP

[1476] Kenneth I. Carlaw and Les Oxley. Resolving the productivity paradox. Mathematics and Computers in Simulation, 78(2–3):313– 318, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000426.

Zhang:2008:GCC

[1477] Zhaoyong Zhang, Kiyotaka Sato, and Michael McAleer. Is Greater China a currency union?: a tale of the Chinese trio. Mathematics and Computers in Simulation, 78(2–3):319– 327, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000438.

Bartolucci:2008:EBP

[1478] Alfred A. Bartolucci, Sejong Bae, and Karan P. Singh. Establishing a Bayesian predictive survival model adjusting for random effects. Mathematics and Computers in Simulation, 78 (2–3):328–334, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800044X.

Casas:2008:ESV

[1479] Isabel Casas. Estimation of stochastic volatility with LRD. Mathematics and Computers in Simulation, 78 (2–3):335–340, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000451.

Liu:2008:MEL

[1480] Qingfeng Liu and Yoshihiko Nishiyama.

Maximum empirical likelihood estimation of continuous-time models with conditional characteristic functions. Mathematics and Computers in Simulation, 78(2–3):341– 350, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000463.

Cappelli:2008:DMM

[1481] Carmela Cappelli, Richard N. Penny, William S. Rea, and Marco Reale. Detecting multiple mean breaks at unknown points in official time series. Mathematics and Computers in Simulation, 78(2–3):351– 356, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000475.

Stemp:2008:CDA

[1482] Peter J. Stemp and Ric D. Herbert. Comparing different approaches for solving optimizing models with significant nonlinearities. Mathematics and Computers in Simulation, 78 (2–3):357–366, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000487.

Hartcher:2008:IIL

[1483] M. G. Hartcher and D. A. Post. The impact of improved landuse cover on the range of modelled sediment yield from two sub-catchments of the Mae Chaem, Thailand. Mathematics and Computers in Simulation, 78 (2–3):367–378, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000499.

Chen:2008:AIT

[1484] Serena H. Chen, Anthony J. Jakeman, and John P. Norton. Artificial Intelligence techniques: an introduction to their use for modelling environmental systems. Mathematics and Computers in Simulation, 78 (2–3):379–400, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000505.

McDonald:2008:ABM

[1485] A. D. McDonald, L. R. Little, R. Gray, E. Fulton, K. J. Sainsbury, and V. D. Lyne. An agent-based modelling approach to evaluation of multiple-use management strategies for coastal marine ecosystems. Mathematics and Computers in Simulation, 78(2–3): 401–411, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000517.

Rizzoli:2008:SLI

[1486] Andrea E. Rizzoli, Marcello Donatelli, Ioannis N. Athanasiadis, Ferdinando Villa, and David Huber. Semantic links in integrated modelling frameworks. Mathematics and Computers in Simulation, 78(2–3): 412–423, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000529.

Talib:2008:FEA

[1487] A. Talib, F. Recknagel, H. Cao, and D. T. van der Molen. Forecasting and explanation of algal dynamics in two shallow lakes by recurrent artificial neural network and hybrid evolutionary algorithm. Mathematics and Computers in Simulation, 78 (2–3):424–434, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000530.

Ticehurst:2008:IMD

[1488] Jenifer L. Ticehurst, Rebecca A. Letcher, and David Rissik. Integration modelling and decision support: a case study of the Coastal Lake Assessment and Management (CLAM) Tool. Mathematics and Computers in Simulation, 78(2–3):435–449, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000542.

Lim:2008:CPE

[1489] Christine Lim and Ying Wang. China’s post-1978 experience in outbound tourism. Mathematics and Computers in Simulation, 78(2–3):450– 458, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000554.

Shareef:2008:MIT

[1490] Riaz Shareef and Michael McAleer. Modelling international tourism demand and uncertainty in Maldives and Seychelles: a portfolio approach. Mathematics and Computers in Simulation, 78(2–3):459–468, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000566.

Anonymous:2008:NIg

[1491] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 78(2–3):469–470, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001882.

Anonymous:2008:ICEg

[1492] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 78(2–3):471– 474, July 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001894.

Anonymous:2008:EBg

[1493] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 78(2–3):CO2, July 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800181X.

Collar:2008:IEV

[1494] Andrés Fraguela Collar, Monserrat Morín Castillo, and Jacobo Oliveros Oliveros. Inverse electroencephalography for volumetric sources. Mathematics and Computers in Simulation, 78 (4):481–492, August 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002091.

Ghitany:2008:LDA

[1495] M. E. Ghitany, B. Atieh, and S. Nadarajah. Lindley distribution and its application. Mathematics and Computers in Simulation, 78 (4):493–506, August 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700211X.

Yang:2008:SSI

[1496] Zheng Yang, Zheng Tian, and Zixia Yuan. Small sample improvements in the threshold cointegration test using residual-based moving block bootstrap. Mathematics and Computers in Simulation, 78(4):507– 513, August 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002121.

Assawinchaichote:2008:FSF

[1497] Wudhichai Assawinchaichote, Sing Kiong Nguang, Peng Shi, and El-Kébir Boukas. H_{∞} fuzzy state-feedback control design for nonlinear systems with D-stability constraints: an LMI approach. Mathematics and Computers in Simulation, 78(4):514– 531, August 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002133.

Ismail:2008:NSC

[1498] M. S. Ismail. Numerical solution of coupled nonlinear Schrödinger equation by Galerkin method. Mathematics and Computers in Simulation, 78 (4):532–547, August 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002145.

Cai:2008:PAP

[1499] Mingshan Cai, Hong Zhang, and Zhaohui Yuan. Positive almost periodic solutions for shunting inhibitory cellular neural networks with timevarying delays. Mathematics and Computers in Simulation, 78(4):548– 558, August 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002285.

Anonymous:2008:NIh

[1500] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 78(4):559–560, August 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002024.

Anonymous:2008:ICEh

[1501] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 78(4):561– 565, August 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002036.

Anonymous:2008:EBh

[1502] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 78(4):CO2, August 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001961.

Anonymous:2008:PAb

[1503] Anonymous. Pages 481–572 (August 2008). Mathematics and Computers in Simulation, 78(4):??, August 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pistella:2008:ASC

[1504] Francesca Pistella, Rosa Maria Spitaleri, and Bharat K. Soni. Applied scientific computing: Grid generation and field simulation. Mathematics and Computers in Simulation, 78(5– 6):573–574, September 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001584.

Chumakov:2008:RMH

[1505] Gennadii A. Chumakov. Riemmanian metric of harmonic parametrization of geodesic quadrangles and quasiisometric grids. Mathematics and Computers in Simulation, 78(5–6): 575–592, September 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001596.

Egidi:2008:BDT

[1506] Nadaniela Egidi and Pierluigi Maponi. Block decomposition techniques in the generation of adaptive grids. Mathematics and Computers in Simulation, 78(5–6):593–604, September 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408001602.

Koomullil:2008:CGM

[1507] Roy Koomullil, Bharat Soni, and Rajkeshar Singh. A comprehensive generalized mesh system for CFD applications. Mathematics and Computers in Simulation, 78(5–6):605– 617, September 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001626.

Koomullil:2008:MBS

[1508] Roy Koomullil, Gary Cheng, Bharat Soni, Ralph Noack, and Nathan Prewitt. Moving-body simulations using overset framework with rigid body dynamics. Mathematics and Computers in Simulation, 78(5–6):618– 626, September 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001638.

Jahangirian:2008:AUG

[1509] A. Jahangirian and Y. Shoraka. Adaptive unstructured grid generation for engineering computation of aerodynamic flows. Mathematics and Computers in Simulation, 78(5–6):627– 644, September 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001614.

Marcuzzi:2008:AUT

[1510] F. Marcuzzi, M. Morandi Cecchi, and M. Venturin. An anisotropic unstructured triangular adaptive mesh algorithm based on error and error gradient information. Mathematics and Computers in Simulation, 78(5–6): 645–652, September 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540800164X.

Venkatachari:2008:VVC

[1511] Balaji Shankar Venkatachari, Gary C. Cheng, Bharat K. Soni, and S. C. Chang. Validation and verification of Courant number insensitive CE/ SE method for transient viscous flow simulations. Mathematics and Computers in Simulation, 78(5–6):653– 670, September 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001651.

Anonymous:2008:NIi

[1512] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 78(5–6):671–672, September 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475408002267.

Anonymous:2008:ICEi

[1513] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 78(5–6):673–677, September 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408002279.

Anonymous:2008:EBi

[1514] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 78(5–6):CO2, September 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S037847540800219X.

Ortega:2008:BAF

[1515] Neli R. S. Ortega, Eduardo Massad, and Cláudio José Struchiner. A Bayesian approach to fuzzy hypotheses testing for the estimation of optimal age for vaccination against measles. Mathematics and Computers in Simulation, 79(1):1– 13, October 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002480.

Gui:2008:EGE

[1516] Zhanji Gui, Xiao-Song Yang, and Weigao Ge. Existence and global exponential stability of periodic solutions of recurrent cellular neural networks with impulses and delays. Mathematics and Computers in Simulation, 79(1):14–29, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002492.

Lean:2008:SPS

[1517] Hooi-Hooi Lean, Wing-Keung Wong, and Xibin Zhang. The sizes and powers of some stochastic dominance tests: a Monte Carlo study for correlated and heteroskedastic distributions. Mathematics and Computers in Simulation, 79(1):30–48, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407002509.

Alrefaei:2008:AMC

[1518] Mahmoud H. Alrefaei and Houssam M. Abdul-Rahman. An adaptive Monte Carlo integration algorithm with general division approach. Mathematics and Computers in Simulation, 79 (1):49–59, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407002595.

Hsu:2008:CEV

[1519] Y. L. Hsu, T. I. Lin, and C. F. Lee. Constant elasticity of variance (CEV) option pricing model: Integration and detailed derivation. Mathematics and Computers in Simulation, 79(1):60–71, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407002601.

Pang:2008:DAP

[1520] Guoping Pang and Lansun Chen. Dynamic analysis of a pest-epidemic model with impulsive control. Mathematics and Computers in Simulation, 79(1):72–84, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407002674.

Lee:2008:MDF

[1521] Chien-Chiang Lee, Chun-Ping Chang, and Pei-Fen Chen. Money demand function versus monetary integration: Revisiting panel cointegration among GCC countries. Mathematics and Computers in Simulation, 79 (1):85–93, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002698.

Kisi:2008:CNN

[1522] Özgür Kisi. Constructing neural network sediment estimation models using a data-driven algorithm. Mathematics and Computers in Simulation, 79(1):94–103, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002819.

Anonymous:2008:NIj

[1523] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(1):104–105, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002681.

Anonymous:2008:ICEj

[1524] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(1):106– 110, October 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002693.

Anonymous:2008:IEBa

[1525] Anonymous. IFC — Editorial Board. Mathematics and Computers in Simulation, 79(1):CO2, October 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002632.

Anonymous:2008:PO

[1526] Anonymous. Pages 1–116 (October 2008). Mathematics and Computers in Simulation, 79(1):??, October 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pulch:2008:IBV

[1527] R. Pulch. Initial-boundary value problems of warped MPDAEs including minimisation criteria. Mathematics and Computers in Simulation, 79(2): 117–132, November 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002856.

Moustris:2008:RCP

[1528] George Moustris and Spyros G. Tzafestas. Reducing a class of polygonal path tracking to straight line tracking via nonlinear strip-wise affine transformation. Mathematics and Computers in Simulation, 79(2):133– 148, November 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003072.

Zhou:2008:SBA

[1529] Liangqiang Zhou and Fangqi Chen. Stability and bifurcation analysis for a model of a nonlinear coupled pitch–roll ship. Mathematics and Computers in Simulation, 79(2):

149–166, November 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700314X.

Salkuyeh:2008:PSC

[1530] Davod Khojasteh Salkuyeh, Faezeh Toutounian, and Hamed Shariat Yazdi. A procedure with stepsize control for solving n one-dimensional IVPs. Mathematics and Computers in Simulation, 79(2):167–176, November 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003187.

Feng:2008:NAS

[1531] Xiao-Li Feng, Zhi Qian, and Chu-Li Fu. Numerical approximation of solution of nonhomogeneous backward heat conduction problem in bounded region. Mathematics and Computers in Simulation, 79(2):177– 188, November 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003199.

Inc:2008:AHA

[1532] Mustafa Inc. Application of homotopy analysis method for fin efficiency of convective straight fins with temperature-dependent thermal conductivity. Mathematics and Computers in Simulation, 79(2):189–200, November 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407003205.

Kuptsov:2008:FDD

[1533] Pavel V. Kuptsov and Razvan A. Satnoianu. Flow and diffusion distributed structures with noise at the inlet. Mathematics and Computers in Simulation, 79(2):201–218, November 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407003217.

Anonymous:2008:NIk

[1534] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(2):219–220, November 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475408003005.

Anonymous:2008:ICEk

[1535] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(2):221–225, November 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003017.

Anonymous:2008:EBj

[1536] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(2):CO2, November 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002954.

Anonymous:2008:PN

[1537] Anonymous. Pages 117–232 (November 2008). Mathematics and Computers in Simulation, 79(2):??, November 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Salarieh:2008:ACS

[1538] Hassan Salarieh and Aria Alasty.

Adaptive chaos synchronization in Chua’s systems with noisy parameters. Mathematics and Computers in Simulation, 79(3):233–241, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003229.

El-Sayed:2008:NSE

[1539] M. F. El-Sayed. Nonlinear stability, excitation and soliton solutions in electrified dispersive systems. Mathematics and Computers in Simulation, 79(3): 242–257, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003230.

Cerimele:2008:SEC

[1540] Maria M. Cerimele, Daniela Mansutti, and Francesca Pistella. Study of Europa’s crust via a Stefan model with convection. Mathematics and Computers in Simulation, 79(3):258– 268, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003242.

Feng:2008:ASD

[1541] Li-Hua Feng, Xing-Cai Zhang, and Gao-Yuan Luo. Application of system dynamics in analyzing the carrying capacity of water resources in Yiwu City, China. Mathematics and Computers in Simulation, 79(3):269– 278, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700362X.

Ghitany:2008:ZTP

[1542] M. E. Ghitany, D. K. Al-Mutairi, and S. Nadarajah. Zero-truncated Poisson–Lindley distribution and its application. Mathematics and Computers in Simulation, 79(3):279–287, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003631.

Patricio:2008:HAL

[1543] M. Patrício, R. Mattheij, and G. de With. Homogenisation with application to layered materials. Mathematics and Computers in Simulation, 79(3):288–305, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003643.

So:2008:MTS

[1544] Mike K. P. So and C. Y. Choi. A multivariate threshold stochastic volatility model. Mathematics and Computers in Simulation, 79(3):306– 317, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000025.

Zhen:2008:BCM

[1545] Li Zhen, He Zhengjia, Zi Yanyang, and Chen Xuefeng. Bearing condition monitoring based on shock pulse method and improved redundant lifting scheme. Mathematics and Computers in Simulation, 79(3):318– 338, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000037.

Kinney:2008:PAN

[1546] W. A. Kinney and G. B. Smith. The potential for aircraft noise reduction through active suppression from a ground-based system. Mathematics and Computers in Simulation, 79(3): 339–351, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000049.

Chang:2008:PPI

[1547] Chia-Lin Chang and Stéphane Robin.

Public policy, innovation and total factor productivity: an application to Taiwan’s manufacturing industry. Mathematics and Computers in Simulation, 79(3):352–367, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000050.

Lawrie:2008:MAS

[1548] Jock Lawrie and John Hearne. A method for aggregating state variables in large ecosystem models. Mathematics and Computers in Simulation, 79(3):368–378, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000074.

Atzberger:2008:EAS

[1549] Paul J. Atzberger and Peter R. Kramer. Error analysis of a stochastic immersed boundary method incorporating thermal fluctuations. Mathematics and Computers in Simulation, 79(3):379–408, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475408000098.

Zhong:2008:ICF

[1550] Qishui Zhong, Jingfu Bao, Yongbin Yu, and Xiaofeng Liao. Impulsive control for T–S fuzzy model-based chaotic systems. Mathematics and Computers in Simulation, 79(3):409– 415, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000104.

Gross:2008:MBP

[1551] Andreas Gross and Hermann F. Fasel. Multi-block Poisson grid generator for cascade simulations. Mathematics and Computers in Simulation, 79(3):416– 428, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800058X.

Chen:2008:DRA

[1552] An-Sing Chen, Hung-Gay Fung, and Erin H. C. Kao. The dynamic relations among return volatility, trading imbalance, and trading volume in futures markets. Mathematics and Computers in Simulation, 79(3):429– 436, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000591.

Cerimele:2008:NSA

[1553] M. M. Cerimele, F. Pistella, and R. M. Spitaleri. Numerical simulations of acoustic fields on boundaryfitted grids. Mathematics and Computers in Simulation, 79(3):437–448, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000608.

Hu:2008:HPS

[1554] Manfeng Hu, Yongqing Yang, Zhenyuan Xu, and Liuxiao Guo. Hybrid projective synchronization in a chaotic complex nonlinear system. Mathematics and Computers in Simulation, 79(3): 449–457, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800061X.

Yang:2008:DTG

[1555] Yung-Lieh Yang and Chia-Lin Chang. A double-threshold GARCH model of stock market and currency shocks on stock returns. Mathematics and Computers in Simulation, 79(3):458– 474, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408000621.

Nadarajah:2008:CGD

[1556] Saralees Nadarajah. Compound growth distributions. Mathematics and Computers in Simulation, 79(3):475–488, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000633.

Chen:2008:TNM

[1557] Cathy W. S. Chen, Richard H. Gerlach, and Amanda P. J. Tai. Testing for nonlinearity in mean and volatility for heteroskedastic models. Mathematics and Computers in Simulation, 79(3): 489–499, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540800075X.

Zhao:2008:IVS

[1558] Zhong Zhao, Lansun Chen, and Xinyu Song. Impulsive vaccination of SEIR epidemic model with time delay and nonlinear incidence rate. Mathematics and Computers in Simulation, 79(3): 500–510, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000967.

Li:2008:CDR

[1559] Ming-Yuan Leon Li. Clarifying the dynamics of the relationship between option and stock markets using the threshold vector error correction model. Mathematics and Computers in Simulation, 79(3):511– 520, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001195.

Schittkowski:2008:PIM

[1560] Klaus Schittkowski. Parameter identification and model verification in systems of partial differential equations applied to transdermal drug delivery. Mathematics and Computers in Simulation, 79(3):521–538, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001201.

Singh:2008:CMF

[1561] S. R. Singh. A computational method of forecasting based on fuzzy time series. Mathematics and Computers in Simulation, 79(3):539–554, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001237.

Hayakawa:2008:RLD

[1562] Kazuhiko Hayakawa and Eiji Kurozumi. The role of “leads” in the dynamic OLS estimation of cointegrating regression models. Mathematics and Computers in Simulation, 79(3):555– 560, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001249.

Witek:2008:SEN

[1563] Marcin L. Witek, Joao Teixeira, and Piotr J. Flatau. On stable and explicit numerical methods for the advection– diffusion equation. Mathematics and Computers in Simulation, 79(3):561– 570, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001250.

Xu:2008:SRK

[1564] Y. Xu and J. J. Zhao. Stability of Runge–Kutta methods for neutral delay-integro-differential-algebraic system. Mathematics and Computers in Simulation, 79(3):571–583, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001262.

Mandel:2008:WFM

[1565] Jan Mandel, Lynn S. Bennethum, Jonathan D. Beezley, Janice L. Coen, Craig C. Douglas, Minjeong Kim, and Anthony Vodacek. A wildland fire model with data assimilation. Mathematics and Computers in Simulation, 79(3):584–606, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001572.

Wang:2008:NSN

[1566] Tingchun Wang, Tao Nie, Luming Zhang, and Fangqi Chen. Numerical simulation of a nonlinearly coupled Schrödinger system: a linearly uncoupled finite difference scheme. Mathematics and Computers in Simulation, 79(3):607–621, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001663.

Jodar:2008:NNM

[1567] Lucas Jódar, Rafael J. Villanueva, Abraham J. Arenas, and Gilberto C. González. Nonstandard numerical methods for a mathematical model for influenza disease. Mathematics and Computers in Simulation, 79(3): 622–633, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001675.

Basu:2008:MMC

[1568] S. K. Basu and Naveen Kumar. Mathematical model and computer simulation for release of nutrients from coated fertilizer granules. Mathematics and Computers in Simulation, 79(3): 634–646, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001687.

Cook:2008:FAS

[1569] Steven Cook. Further analysis of spurious causality. Mathematics and Computers in Simulation, 79(3):647– 651, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001699.

Alberto:2008:HPT

[1570] Angelines Alberto, Jesús Benet, Enrique Arias, David Cebrian, Tomás Rojo, and Fernando Cuartero. A high performance tool for the simulation of the dynamic pantograph– catenary interaction. Mathematics and Computers in Simulation, 79(3):652– 667, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001717.

Alvarez-Vazquez:2008:ODO

[1571] Lino J. Alvarez-Vázquez, Eva Balsa-Canto, and Aurea Martínez. Optimal design and operation of a wastewater purification system. Mathematics and Computers in Simulation, 79(3): 668–682, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001729.

Dehghan:2008:HOC

[1572] Mehdi Dehghan and Akbar Mohebbi. High-order compact boundary value method for the solution of unsteady convection–diffusion problems. Mathematics and Computers in Simulation, 79(3):683–699, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001730.

Dehghan:2008:NMS

[1573] Mehdi Dehghan and Ali Shokri. A numerical method for solution of the two-dimensional sine-Gordon equation using the radial basis functions. Mathematics and Computers in Simulation, 79(3):700–715, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001778.

Zhu:2008:LSM

[1574] Zhengjie Zhu, C. A. Dorao, and H. A.

Jakobsen. A least-squares method with direct minimization for the solution of the breakage–coalescence population balance equation. Mathematics and Computers in Simulation, 79(3): 716–727, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002097.

Kumar:2008:NAE

[1575] M. Vasudeva Kumar, A. Kienle, K. P.

Zeyer, and S. Pushpavanam. Nonlinear analysis of the effect of maintenance in continuous cell cultures. Mathematics and Computers in Simulation, 79(3): 728–748, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002103.

Zheng:2008:NSR

[1576] J. G. Zheng, T. S. Lee, and S. H. Winoto. Numerical simulation of Richtmyer–Meshkov instability driven by imploding shocks. Mathematics and Computers in Simulation, 79(3): 749–762, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002164.

Nguyen:2008:MMR

[1577] Vinh Phu Nguyen, Timon Rabczuk, Stéphane Bordas, and Marc Duflot. Meshless methods: a review and computer implementation aspects. Mathematics and Computers in Simulation, 79(3):763–813, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408000062.

Anonymous:2008:NIl

[1578] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(3):814–815, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475408003315.

Anonymous:2008:ICEl

[1579] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(3):816–819, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003327.

Anonymous:2008:IEBb

[1580] Anonymous. IFC — Editorial Board. Mathematics and Computers in Simulation, 79(3):CO2, December 1, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800325X.

Anonymous:2008:PD

[1581] Anonymous. Pages 233–826 (1 December 2008). Mathematics and Computers in Simulation, 79(3):??, December 1, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Troch:2008:E

[1582] Inge Troch and Felix Breitenecker.

Editorial. Mathematics and Computers in Simulation, 79(4):827–828, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000979.

Schlacher:2008:MMN

[1583] K. Schlacher. Mathematical modeling for nonlinear control: a Hamiltonian approach. Mathematics and Computers in Simulation, 79(4):829– 849, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001134.

Schwarz:2008:SSD

[1584] Peter Schwarz. Simulation of systems with dynamically varying model structure. Mathematics and Computers in Simulation, 79(4):850–863, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001146.

Tagesen:2008:MED

[1585] S. Tagesen. Modeling energy dependent nuclear reaction probabilities from pointwise measurements including full covariance information. Mathematics and Computers in Simulation, 79(4): 864–873, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001158.

Viertl:2008:FMP

[1586] Reinhard Viertl. Fuzzy models for precision measurements. Mathematics and Computers in Simulation, 79(4): 874–878, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540800116X.

Zlajpah:2008:SR

[1587] Leon Žlajpah. Simulation in robotics. Mathematics and Computers in Simulation, 79(4):879–897, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475408001183.

Arnold:2008:TBC

[1588] Anton Arnold and Maike Schulte.

Transparent boundary conditions for quantum-waveguide simulations. Mathematics and Computers in Simulation, 79(4):898–905, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408000980.

Callies:2008:RMC

[1589] R. Callies and S. Fronz. Recursive modeling and control of multilink manipulators with vacuum grippers. Mathematics and Computers in Simulation, 79(4):906–916, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408000992.

Ermoliev:2008:DCR

[1590] Y. Ermoliev, T. Ermolieva, G. Fischer, M. Makowski, S. Nilsson, and M. Obersteiner. Discounting, catastrophic risks management and vulnerability modeling. Mathematics and Computers in Simulation, 79(4):917– 924, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001006.

Gerdts:2008:NNM

[1591] Matthias Gerdts. A nonsmooth Newton’s method for control-state constrained optimal control problems. Mathematics and Computers in Simulation, 79(4):925–936, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001018.

Ghirardi:2008:OOM

[1592] Marco Ghirardi, Giuseppe Menga, and Nicola Sacco. An optimisationoriented model of distributed supplychain. Mathematics and Computers in Simulation, 79(4):937–946, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800102X.

Carrillo:2008:SEI

[1593] D. M. Carrillo, J. A. Lopez-Orozco, and J. M. Giron-Sierra. Simulation environment for the investigation of automatized cooperation of marine crafts. Mathematics and Computers in Simulation, 79(4):947–954, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001031.

Gomes:2008:PEG

[1594] O. Gomes, C. Combes, and A. Dussauchoy. Parameter estimation of the generalized gamma distribution. Mathematics and Computers in Simulation, 79(4):955–963, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001043.

Gyimesi:2008:WSG

[1595] M. Gyimesi. Web Services with generic simulation models for discrete event simulation. Mathematics and Computers in Simulation, 79(4):964– 971, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001055.

Karlowatz:2008:EBS

[1596] G. Karlowatz, W. Wessner, and H. Kosina. Effect of band structure discretization on the performance of full-band Monte Carlo simulation. Mathematics and Computers in Simulation, 79(4):972–979, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001067.

Kryvinska:2008:AAM

[1597] Natalia Kryvinska. An analytical approach for the modeling of real-time services over IP network. Mathematics and Computers in Simulation, 79(4): 980–990, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001079.

Noack:2008:VRI

[1598] S. Noack, A. Wahl, M. Haunschild, E. Qeli, B. Freisleben, and W. Wiechert. Visualizing regulatory interdependencies and parameter sensitivities in biochemical network models. Mathematics and Computers in Simulation, 79(4):

991–998, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001080.

Otero-Muras:2008:DAC

[1599] Irene Otero-Muras, Gábor Szederkényi, Katalin M. Hangos, and Antonio A. Alonso. Dynamic analysis and control of biochemical reaction networks. Mathematics and Computers in Simulation, 79(4):999–1009, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001092.

Pedersen:2008:TQS

[1600] Morten Gram Pedersen, Alberto M. Bersani, Enrico Bersani, and Giuliana Cortese. The total quasi-steady-state approximation for complex enzyme reactions. Mathematics and Computers in Simulation, 79(4):1010– 1019, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001109.

Gerdts:2008:NOC

[1601] Matthias Gerdts, Günter Greif, and Hans Josef Pesch. Numerical optimal control of the wave equation: optimal boundary control of a string to rest in finite time. Mathematics and Computers in Simulation, 79(4):1020– 1032, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001110.

Popa:2008:NSA

[1602] C. Popa. On numerical solution of arbitrary symmetric linear systems by approximate orthogonalization. Mathematics and Computers in Simulation, 79(4):1033–1038, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001122.

Hessinger:2008:HME

[1603] M. Hessinger, A. Holzinger, D. Leitner, and S. Wassertheurer. Hemodynamic models for education in physiology. Mathematics and Computers in Simulation, 79(4):1039–1047, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001171.

Li:2008:P

[1604] Yiming Li, Juncheng Cao, Pavel Dyshlovenko, Tatsuya Ezaki, Tor A. Fjeldly, Hans Kosina, Todd McKenzie, Hiroshi Mizuta, and Laurence T. Yang. Preface. Mathematics and Computers in Simulation, 79(4):1049– 1050, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408000578.

Pourfath:2008:NSQ

[1605] M. Pourfath, H. Kosina, and S. Selberherr. Numerical study of quantum transport in carbon nanotube transistors. Mathematics and Computers in Simulation, 79(4):1051–1059, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002583.

Shin:2008:TDQ

[1606] Mincheol Shin. Three-dimensional quantum simulation of multigate nanowire field effect transistors. Mathematics and Computers in Simulation, 79(4):1060–1070, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002868.

Ungersboeck:2008:EUS

[1607] E. Ungersboeck, W. Gös, S. Dhar, H. Kosina, and S. Selberherr. The effect of uniaxial stress on band structure and electron mobility of silicon. Mathematics and Computers in Simulation, 79(4):1071–1077, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002662.

Assous:2008:SVM

[1608] Franck Assous and Patrick Ciarlet. Solving Vlasov–Maxwell equations in singular geometries. Mathematics and Computers in Simulation, 79(4):1078– 1085, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002650.

Fedoseyev:2008:ANM

[1609] Alexander I. Fedoseyev, Marek Turowski, Michael L. Alles, and Robert A. Weller. Accurate numerical models for simulation of radiation events in nano-scale semiconductor devices. Mathematics and Computers in Simulation, 79(4):1086–1095, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002571.

Ezaki:2008:NQS

[1610] T. Ezaki, D. Navarro, Y. Takeda, N. Sadachika, G. Suzuki, M. Miura-Mattausch, H. J. Mattausch, T. Ohguro, T. Iizuka, M. Taguchi, S. Kumashiro, and S. Miyamoto. Non-quasi-static approach with surface-potential-based MOSFET model HiSIM for RF circuit simulations. Mathematics and Computers in Simulation, 79(4):1096– 1106, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002844.

Kolberg:2008:MVC

[1611] S. Kolberg, H. Børli, and T. A.

Fjeldly. Modeling, verification and comparison of short-channel double gate and gate-all-around MOS-FETs. Mathematics and Computers in Simulation, 79(4):1107–1115, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407002649.

Morris:2008:CMC

[1612] Hedley C. Morris and Alfonso Limon. A compact model for the I – V characteristics of an undoped doublegate MOSFET. Mathematics and Computers in Simulation, 79(4):1116– 1125, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407003096.

Avci:2008:NNB

[1613] M. Avci. Neural network-based design approach for submicron MOS integrated circuits. Mathematics and Computers in Simulation, 79(4):1126– 1136, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407003102.

Li:2008:EDA

[1614] Yiming Li, Shao-Ming Yu, and Yih-Lang Li. Electronic design automation using a unified optimization framework. Mathematics and Computers in Simulation, 79(4):1137–1152, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407003126.

Chu:2008:MOR

[1615] Chia-Chi Chu, Ming-Hong Lai, and Wu-Shiung Feng. Model-order reductions for MIMO systems using global Krylov subspace methods. Mathematics and Computers in Simulation, 79(4):1153–1164, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002546.

Li:2008:DOC

[1616] Yiming Li, Yih-Lang Li, and Shao-Ming Yu. Design optimization of a current mirror amplifier integrated circuit using a computational statistics technique. Mathematics and Computers in Simulation, 79(4):1165– 1177, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407003138.

Xiong:2008:NNA

[1617] N. Xiong, L. T. Yang, Y. Yang, X. Défago, and Y. He. A novel numerical algorithm based on selftuning controller to support TCP flows. Mathematics and Computers in Simulation, 79(4):1178–1188, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407002686.

Jeon:2008:MEA

[1618] Jun-Cheol Jeon and Kee-Young Yoo.

Montgomery exponent architecture based on programmable cellular automata. Mathematics and Computers in Simulation, 79(4):1189–1196, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407002558.

Jeon:2008:ECB

[1619] Jun-Cheol Jeon and Kee-Young Yoo. Elliptic curve based hardware architecture using cellular automata. Mathematics and Computers in Simulation, 79(4):1197–1203, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540700256X.

Wu:2008:AAP

[1620] Jigang Wu, Thambipillai Srikanthan, and Chengbin Yan. Algorithmic aspects for power-efficient hardware/ software partitioning. Mathematics and Computers in Simulation, 79(4):1204– 1215, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407002534.

Kamran:2008:ELD

[1621] Muhammad Kamran, Feng Shi, Yumin Xie, and Yizhuo Wang. An efficient layered data compression scheme with constraint analysis. Mathematics and Computers in Simulation, 79(4):1216– 1232, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407003084.

DelBuono:2008:SDS

[1622] N. Del Buono. Structural dynamical systems: Computational aspects. Mathematics and Computers in Simulation, 79(4):1233–1234, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001328.

Dieci:2008:SAA

[1623] L. Dieci and C. Elia. SVD algorithms to approximate spectra of dynamical systems. Mathematics and Computers in Simulation, 79(4):1235– 1254, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800133X.

Dieci:2008:SVT

[1624] Luca Dieci and Alessandro Pugliese.

Singular values of two-parameter matrices: an algorithm to accurately find their intersections. Mathematics and Computers in Simulation, 79(4):1255– 1269, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001341.

Hochbruck:2008:AMO

[1625] Marlis Hochbruck and Jörg Niehoff.

Approximation of matrix operators applied to multiple vectors. Mathematics and Computers in Simulation, 79(4):1270–1283, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001353.

DelBuono:2008:CFH

[1626] N. Del Buono, L. Lopez, and T. Politi. Computation of functions of Hamiltonian and skew-symmetric matrices. Mathematics and Computers in Simulation, 79(4):1284–1297, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001365.

Celledoni:2008:DMO

[1627] Elena Celledoni and Simone Fiori. Descent methods for optimization on homogeneous manifolds. Mathematics and Computers in Simulation, 79(4):1298–1323, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001377.

Baumann:2008:TVN

[1628] M. Baumann and U. Helmke. A timevarying Newton algorithm for adaptive subspace tracking. Mathematics and Computers in Simulation, 79(4):1324– 1345, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408001389.

Nonnenmacher:2008:DLR

[1629] Achim Nonnenmacher and Christian Lubich. Dynamical low-rank approximation: applications and numerical experiments. Mathematics and Computers in Simulation, 79(4):1346– 1357, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001390.

Galeone:2008:FAM

[1630] Luciano Galeone and Roberto Garrappa. Fractional Adams–Moulton methods. Mathematics and Computers in Simulation, 79(4):1358–1367, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001407.

Gameiro:2008:VCL

[1631] Marcio Gameiro, Jean-Philippe Lessard, and Konstantin Mischaikow. Validated continuation over large parameter ranges for equilibria of PDEs. Mathematics and Computers in Simulation, 79(4):1368–1382, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001419.

Beardmore:2008:HBT

[1632] R. Beardmore and K. Webster.

A Hopf bifurcation theorem for singular differential–algebraic equations. Mathematics and Computers in Simulation, 79(4):1383–1395, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408001420.

Kupper:2008:ICN

[1633] Tassilo Küpper. Invariant cones for non-smooth dynamical systems. Mathematics and Computers in Simulation, 79(4):1396–1408, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408001432.

Anonymous:2008:NIm

[1634] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(4):1409–1410, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003558.

Anonymous:2008:ICEm

[1635] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(4):1411–1414, December 15, 2008. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540800356X.

Anonymous:2008:EBk

[1636] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(4):CO2, December 15, 2008. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475408003480.

Hamani:2009:RMH

[1637] Nadia Hamani, Nathalie Dangoumau, and Etienne Craye. Reactive mode handling of flexible manufacturing systems. Mathematics and Computers in Simulation, 79(5):1421– 1439, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407003655.

Cunha:2009:SMS

[1638] M. Cristina C. Cunha and Fábio A.

Dorini. Statistical moments of the solution of the random Burgers– Riemann problem. Mathematics and Computers in Simulation, 79(5):1440– 1451, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002176.

Ramos:2009:RCR

[1639] C. Ramos, M. Martínez, J. Sanchis, and J. V. Salcedo. Robust constrained receding-horizon predictive control via bounded data uncertainties. Mathematics and Computers in Simulation, 79(5):1452– 1471, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002310.

Huang:2009:MRA

[1640] Biing-Wen Huang, Meng-Gu Chen, Chia-Lin Chang, and Michael McAleer. Modelling risk in agricultural finance: Application to the poultry industry in Taiwan. Mathematics and Computers in Simulation, 79(5):1472– 1487, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408002322.

Suresh:2009:NLD

[1641] S. Suresh, V. Mani, S. N. Omkar, H. J. Kim, and N. Sundararajan. A new load distribution strategy for linear network with communication delays. Mathematics and Computers in Simulation, 79(5):1488– 1501, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408002334.

Peng:2009:SPS

[1642] Chen Peng, Dong Yue, Zhou Gu, and Feng Xia. Sampling period scheduling of networked control systems with multiple-control loops. Mathematics and Computers in Simulation, 79(5): 1502–1511, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002346.

Harase:2009:MEP

[1643] Shin Harase. Maximally equidistributed pseudorandom number generators via linear output transformations. Mathematics and Computers in Simulation, 79(5):1512– 1519, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408002358.

Fernandino:2009:LES

[1644] M. Fernandino, K. Beronov, and T. Ytrehus. Large eddy simulation of turbulent open duct flow using a lattice Boltzmann approach. Mathematics and Computers in Simulation, 79 (5):1520–1526, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002383.

Chen:2009:NCG

[1645] Shengshuang Chen, Weirui Zhao, and Yong Xu. New criteria for globally exponential stability of delayed Cohen– Grossberg neural network. Mathematics and Computers in Simulation, 79 (5):1527–1543, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002395.

Xia:2009:MST

[1646] Jianhong (Cecilia) Xia, Panlop Zeephongsekul, and Colin Arrowsmith. Modelling spatio-temporal movement of tourists using finite Markov chains. Mathematics and Computers in Simulation, 79(5):1544– 1553, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002401.

Yan:2009:HOM

[1647] Guangwu Yan and Jianying Zhang.

A higher-order moment method of the lattice Boltzmann model for the Korteweg–de Vries equation. Mathematics and Computers in Simulation, 79(5):1554–1565, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002425.

Tavazoei:2009:NSF

[1648] Mohammad Saleh Tavazoei and Mohammad Haeri. A note on the stability of fractional order systems. Mathematics and Computers in Simulation, 79 (5):1566–1576, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002437.

Onyango:2009:DTD

[1649] T. T. M. Onyango, D. B. Ingham, D. Lesnic, and M. Slodička. Determination of a time-dependent heat transfer coefficient from non-standard boundary measurements. Mathematics and Computers in Simulation, 79 (5):1577–1584, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002449.

Stromgren:2009:NAD

[1650] Magnus Strömgren and Michael Hanke. On the numerical approximation of a degenerated hyperbolic system. Mathematics and Computers in Simulation, 79(5):1585– 1602, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002450.

Zhao:2009:BSN

[1651] Hongyong Zhao and Zisen Mao.

Boundedness and stability of nonautonomous cellular neural networks with reaction–diffusion terms. Mathematics and Computers in Simulation, 79 (5):1603–1617, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002462.

Tan:2009:RQP

[1652] Syn Kiat Tan and Sheng-Uei Guan.

Randomness quality of permuted pseudorandom binary sequences. Mathematics and Computers in Simulation, 79(5):1618–1626, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408002486.

Kuo:2009:RSN

[1653] Hang-Hong Kuo, Yi-You Hou, Jun-Juh Yan, and Teh-Lu Liao. Reliable synchronization of nonlinear chaotic systems. Mathematics and Computers in Simulation, 79(5):1627– 1635, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002760.

Jimenez:2009:CGE

[1654] F. Jiménez and P. Jodrá. On the computer generation of the Erlang and negative binomial distributions with shape parameter equal to two. Mathematics and Computers in Simulation, 79 (5):1636–1640, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002772.

Kulikova:2009:MLE

[1655] M. V. Kulikova. Maximum likelihood estimation via the extended covariance and combined square-root filters. Mathematics and Computers in Simulation, 79(5):1641– 1657, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540800284X.

Tocino:2009:MSI

[1656] A. Tocino. Multiple stochastic integrals with Mathematica. Mathematics and Computers in Simulation, 79 (5):1658–1667, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002851.

Xiong:2009:SRM

[1657] Xiang-Tuan Xiong, Chu-Li Fu, and Jin Cheng. Spectral regularization methods for solving a sideways parabolic equation within the framework of regularization theory. Mathematics and Computers in Simulation, 79(5): 1668–1678, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002863.

Yang:2009:NDD

[1658] Degang Yang, Xiaofeng Liao, Chunyan Hu, and Yong Wang. New delay-dependent exponential stability criteria of BAM neural networks with time delays. Mathematics and Computers in Simulation, 79(5):1679– 1697, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002905.

Wang:2009:ESI

[1659] Xiaohu Wang, Qingyi Guo, and Daoyi Xu. Exponential p-stability of impulsive stochastic Cohen–Grossberg neural networks with mixed delays. Mathematics and Computers in Simulation, 79(5):1698–1710, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002917.

Fukuda:2009:DSF

[1660] Kosei Fukuda. Distribution switching in financial time series. Mathematics and Computers in Simulation, 79(5): 1711–1720, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003054.

Nadarajah:2009:AIG

[1661] Saralees Nadarajah. An alternative inverse Gaussian distribution. Mathematics and Computers in Simulation, 79(5):1721–1729, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003066.

Chang:2009:MFT

[1662] Chia-Lin Chang, Songsak Sriboonchitta, and Aree Wiboonpongse. Modelling and forecasting tourism from East Asia to Thailand under temporal and spatial aggregation. Mathematics and Computers in Simulation, 79 (5):1730–1744, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003091.

Wang:2009:NNB

[1663] Dan Wang, Jialiang Huang, Weiyao Lan, and Xiaoqiang Li. Neural network-based robust adaptive control of nonlinear systems with unmodeled dynamics. Mathematics and Computers in Simulation, 79(5):1745– 1753, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540800311X.

Maki:2009:SPU

[1664] Daiki Maki. Some properties of a unit root test with multiple level shifts in the presence of Markov level shifts. Mathematics and Computers in Simulation, 79(5):1754– 1760, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003133.

Bermudez:2009:MES

[1665] José D. Bermúdez, Ana Corberán-Vallet, and Enriqueta Vercher. Multivariate exponential smoothing: a Bayesian forecast approach based on simulation. Mathematics and Computers in Simulation, 79(5):1761– 1769, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003145.

Wang:2009:LSP

[1666] Lichun Wang and Xuan Wang. The life-span prediction of a system connected in series. Mathematics and Computers in Simulation, 79(5):1770– 1777, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003157. See

comments [1930].

Anonymous:2009:NIa

[1667] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(5):1778–1779, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475408003996.

Anonymous:2009:ICEa

[1668] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(5):1780– 1783, January 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800400X.

Anonymous:2009:EBa

[1669] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(5):CO2, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003935.

Anonymous:2009:PJa

[1670] Anonymous. Pages 1421–1790 (January 2009). Mathematics and Computers in Simulation, 79(5):??, January 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Scherer:2009:P

[1671] Godela Scherer, Victor Pereyra, and Jose Castillo. Preface. Mathematics and Computers in Simulation, 79 (6):1791, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407003175.

Barrera-Sanchez:2009:ADH

[1672] Pablo Barrera-Sánchez, Longina Castellanos Noda, Francisco J. Domínguez-Mota, Guilmer F. González Flores, and Angel Pérez Domínguez. Adaptive discrete harmonic grid generation. Mathematics and Computers in Simulation, 79(6):1792– 1809, February 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407001930.

Blomgren:2009:RAD

[1673] Peter Blomgren, Antonio Palacios, and Scott Gasner. Recent advances in 2+1-dimensional simulations of the pattern-forming Kuramoto– Sivashinsky equation. Mathematics and Computers in Simulation, 79(6): 1810–1823, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003200.

Bravo:2009:ILV

[1674] Antonio Bravo, Rubén Medina, Gianfranco Passariello, and Mireille Garreau. Inferring the left ventricle dynamical behavior using a free-form deformations model. Mathematics and Computers in Simulation, 79(6): 1824–1833, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001449.

Butcher:2009:GLM

[1675] John Butcher. General linear methods for ordinary differential equations. Mathematics and Computers in Simulation, 79(6):1834–1845, February 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475407001462.

Chen-Charpentier:2009:NSM

[1676] Benito M. Chen-Charpentier, Dobromir T. Dimitrov, and Hristo V. Kojouharov. Numerical simulation of multi-species biofilms in porous media for different kinetics. Mathematics and Computers in Simulation, 79(6): 1846–1861, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540700153X.

DeCecchis:2009:FPS

[1677] Dany De Cecchis, Hilda López, and Brígida Molina. FGMRES preconditioning by symmetric/skew-symmetric decomposition of generalized Stokes problems. Mathematics and Computers in Simulation, 79(6):1862– 1877, February 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407001942.

Fraguela:2009:AID

[1678] A. Fraguela, J. A. Gómez, F. Castillo, and J. Oseguera. An approach for the identification of diffusion coefficients in the quasi-steady state of a post-discharge nitriding process. Mathematics and Computers in Simulation, 79(6):1878–1894, February 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407001528.

Ferreira:2009:EBH

[1679] V. G. Ferreira, F. A. Kurokawa, C. M. Oishi, M. K. Kaibara, A. Castelo, and J. A. Cuminato. Evaluation of a bounded high order upwind scheme for 3D incompressible free surface flow computations. Mathematics and Computers in Simulation, 79(6): 1895–1914, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001954.

Limon:2009:MCB

[1680] Alfonso Limon and Hedley Morris.

Multiscale cell-based coarsening for discontinuous problems. Mathematics and Computers in Simulation, 79(6): 1915–1925, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002042.

Lozano:2009:ARL

[1681] Angélica Lozano, Giuseppe Manfredi, and Luciano Nieddu. An algorithm for the recognition of levels of congestion in road traffic problems. Mathematics and Computers in Simulation, 79(6): 1926–1934, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407002054.

Pereyra:2009:FCE

[1682] Victor Pereyra. Fast computation of equispaced Pareto manifolds and Pareto fronts for multiobjective optimization problems. Mathematics and Computers in Simulation, 79(6): 1935–1947, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475407001474.

Alonso:2009:ADO

[1683] J. J. Alonso, P. LeGresley, and V. Pereyra. Aircraft design optimization. Mathematics and Computers in Simulation, 79(6):1948– 1958, February 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475407002108.

Perez:2009:OPS

[1684] Laura V. Pérez and Elvio A. Pilotta. Optimal power split in a hybrid electric vehicle using direct transcription of an optimal control problem. Mathematics and Computers in Simulation, 79(6): 1959–1970, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475407001590.

Paluszny:2009:ETS

[1685] Marco Paluszny and Francisco Tovar. Envelopes and tubular splines. Mathematics and Computers in Simulation, 79(6):1971–1976, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475407001917.

Anonymous:2009:NIb

[1686] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(6):1977–1978, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475409000159.

Anonymous:2009:ICEb

[1687] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(6):1979– 1982, February 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000160.

Anonymous:2009:EBb

[1688] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(6):CO2, February 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000093.

Epshteyn:2009:NSO

[1689] Y. Epshteyn, T. Khan, and B. Rivière.

Numerical solution of a one-dimensional inverse problem by the discontinuous Galerkin method. Mathematics and Computers in Simulation, 79(7): 1989–2000, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002814.

Song:2009:IUM

[1690] Lunji Song and Yujiang Wu. Incremental unknowns method based on the θ-scheme for time-dependent convection–diffusion equations. Mathematics and Computers in Simulation, 79(7):2001–2012, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002826.

Odibat:2009:CAC

[1691] Zaid M. Odibat. Computational algorithms for computing the fractional derivatives of functions. Mathematics and Computers in Simulation, 79 (7):2013–2020, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002838.

Zhang:2009:MSM

[1692] Le Zhang, L. Leon Chen, and Thomas S. Deisboeck. Multi-scale, multi-resolution brain cancer modeling. Mathematics and Computers in Simulation, 79(7):2021– 2035, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003169.

Gupta:2009:EER

[1693] Ramesh C. Gupta and Na Wang. Estimation of extra risk and benchmark dose in dose–response models. Mathematics and Computers in Simulation, 79(7):2036–2050, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003194.

Lu:2009:LPS

[1694] Zhao Lu, Jing Sun, and Kenneth R. Butts. Linear programming support vector regression with wavelet kernel: a new approach to nonlinear dynamical systems identification. Mathematics and Computers in Simulation, 79 (7):2051–2063, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003388.

Lin:2009:AEP

[1695] Kang Lin and Keith E. Holbert. Applying the equivalent pi circuit to the modeling of hydraulic pressurized lines. Mathematics and Computers in Simulation, 79(7):2064– 2075, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800339X.

Cortes:2009:RLQ

[1696] J. C. Cortés, L. Jódar, and L. Villafuerte. Random linear-quadratic mathematical models: Computing explicit solutions and applications. Mathematics and Computers in Simulation, 79 (7):2076–2090, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003765.

Aguilera:2009:DPE

[1697] Andrés Ramírez Aguilera, Luis Enrique Bergues Cabrales, Héctor Manuel Camué Ciria, Yudelmis Soler Pérez, Eduardo Roca Oria, Soraida Acosta Brooks, and Tamara Rubio González. Distributions of the potential and electric field of an electrode elliptic array used in tumor electrotherapy: Analytical and numerical solutions. Mathematics and Computers in Simulation, 79 (7):2091–2105, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003777.

Picard:2009:EFM

[1698] Richard R. Picard and Thomas E.

Booth. Ensuring finite moments in Monte Carlo simulations via iterated ex post facto sampling. Mathematics and Computers in Simulation, 79 (7):2106–2121, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003868.

Nie:2009:ESP

[1699] Linfei Nie, Zhidong Teng, Lin Hu, and Jigen Peng. Existence and stability of periodic solution of a predator– prey model with state-dependent impulsive effects. Mathematics and Computers in Simulation, 79(7):2122– 2134, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540800387X.

Li:2009:SNA

[1700] Hou-Biao Li, Ting-Zhu Huang, Yong Zhang, Xing-Ping Liu, and Hong Li. On some new approximate factorization methods for block tridiagonal matrices suitable for vector and parallel processors. Mathematics and Computers in Simulation, 79 (7):2135–2147, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003881.

Yang:2009:RCB

[1701] Fenglian Yang, Liang Yan, and Ting Wei. Reconstruction of the corrosion boundary for the Laplace equation by using a boundary collocation method. Mathematics and Computers in Simulation, 79(7):2148– 2156, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003893.

Jin:2009:STM

[1702] Hao Jin, Zheng Tian, and Ruibing Qin. Subsampling tests for the mean change point with heavytailed innovations. Mathematics and Computers in Simulation, 79(7):2157– 2166, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800390X.

Lee:2009:ALP

[1703] Wen-Chuan Lee, Jong-Wuu Wu, and Ching-Wen Hong. Assessing the lifetime performance index of products from progressively type II right censored data using Burr XII model. Mathematics and Computers in Simulation, 79(7):2167– 2179, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003911.

Liu:2009:ICB

[1704] Feng Liu, Zhi-Hong Guan, Hua O.

Wang, and Yuqing Li. Impulsive control of bifurcations. Mathematics and Computers in Simulation, 79 (7):2180–2191, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004060.

Gutman:2009:PIH

[1705] Semion Gutman and Junhong Ha. Parameter identifiability for heat conduction with a boundary input. Mathematics and Computers in Simulation, 79 (7):2192–2210, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408004072.

Pozna:2009:DOA

[1706] Claudiu Pozna, Fritz Troester, Radu-Emil Precup, József K. Tar, and Stefan Preitl. On the design of an obstacle avoiding trajectory: Method and simulation. Mathematics and Computers in Simulation, 79(7):2211– 2226, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408004217.

Hsu:2009:RSI

[1707] Ying-Lin Hsu, Ssu-Lang Lee, and Jau-Chuan Ke. A repairable system with imperfect coverage and reboot: Bayesian and asymptotic estimation. Mathematics and Computers in Simulation, 79(7):2227– 2239, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408004229.

Bhatti:2009:ITQ

[1708] Chad R. Bhatti. Intraday trade and quote dynamics: a Cox regression analysis. Mathematics and Computers in Simulation, 79(7):2240– 2249, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004242.

Bhatti:2009:IHI

[1709] Chad R. Bhatti. On the interday homogeneity in the intraday rate of trading. Mathematics and Computers in Simulation, 79(7):2250– 2257, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004254.

Olmos:2009:PMS

[1710] Daniel Olmos and Bernie D. Shizgal.

Pseudospectral method of solution of the Fitzhugh–Nagumo equation. Mathematics and Computers in Simulation, 79(7):2258–2278, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000020.

Anonymous:2009:NIc

[1711] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(7):2279–2280, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000573.

Anonymous:2009:ICEc

[1712] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(7):2281– 2284, March 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409000585.

Anonymous:2009:EBc

[1713] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(7):CO2, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000524.

Anonymous:2009:PM

[1714] Anonymous. Pages 1989–2290 (March 2009). Mathematics and Computers in Simulation, 79(7):??, March 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Wester:2009:FSI

[1715] Michael J. Wester, Stanly Steinberg, and Eugenio Roanes-Lozano. Foreword to the special issue on “Nonstandard applications of computer algebra”. Mathematics and Computers in Simulation, 79(8):2291– 2292, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003716.

Myllari:2009:JDN

[1716] Tatiana Mylläri. Joint distribution of the number of vertices with given different outdegrees in Galton– Watson forest. Mathematics and Computers in Simulation, 79(8):2293– 2301, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003728.

Popov:2009:UCA

[1717] Nikolaj Popov and Tudor Jebelean. Using Computer Algebra techniques for the specification, verification and synthesis of recursive programs. Mathematics and Computers in Simulation, 79(8):2302–2309, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800373X.

Vajda:2009:CLA

[1718] Robert Vajda, Tudor Jebelean, and Bruno Buchberger. Combining logical and algebraic techniques for natural style proving in elementary analysis. Mathematics and Computers in Simulation, 79(8):2310– 2316, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003741.

Roanes-Lozano:2009:ERN

[1719] Eugenio Roanes-Lozano, Luis M. Laita, Eugenio Roanes-Macías, Michael J. Wester, José Luis Ruiz-Lozano, and Carlos Roncero. Evolution of railway network flexibility: the Spanish broad gauge case. Mathematics and Computers in Simulation, 79 (8):2317–2332, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003753.

Pistella:2009:ASC

[1720] Francesca Pistella and Rosa Maria Spitaleri. Applied Scientific Computing VI: Numerical Grid Generation, Approximation and Visualization. Mathematics and Computers in Simulation, 79(8):2333, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000214.

Cerimele:2009:NME

[1721] Maria Mercede Cerimele and Rossella Cossu. A numerical modelling for the extraction of decay regions from color images of monuments. Mathematics and Computers in Simulation, 79(8):2334–2344, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000226.

Dellino:2009:KMM

[1722] G. Dellino, P. Lino, C. Meloni, and A. Rizzo. Kriging metamodel management in the design optimization of a CNG injection system. Mathematics and Computers in Simulation, 79 (8):2345–2360, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000238.

Egidi:2009:ESD

[1723] N. Egidi and P. Maponi. The efficient solution of direct medium problems by using translation techniques. Mathematics and Computers in Simulation, 79(8):2361–2372, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000275.

Kim:2009:DEA

[1724] Jong-Eun Kim, Vinay N. Rao, Roy P.

Koomullil, Doug H. Ross, Bharat K. Soni, and Alan M. Shih. Development of an efficient aerodynamic shape optimization framework. Mathematics and Computers in Simulation, 79 (8):2373–2384, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900024X.

Makhanov:2009:SFC

[1725] S. S. Makhanov. Space-filling curves in adaptive curvilinear coordinates for computer numerically controlled five-axis machining. Mathematics and Computers in Simulation, 79(8): 2385–2402, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000251.

Manzini:2009:SOT

[1726] Gianmarco Manzini. A second-order TVD implicit–explicit finite volume method for time-dependent convectionreaction equations. Mathematics and Computers in Simulation, 79(8):2403– 2428, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000263.

Pellegrino:2009:NBR

[1727] E. Pellegrino and E. Santi. Near best refinable quasi-interpolants. Mathematics and Computers in Simulation, 79(8):2429–2443, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000287.

Plaza:2009:LRB

[1728] Ángel Plaza, Alberto Márquez, Auxiliadora Moreno-González, and José P. Suárez. Local refinement based on the 7-triangle longest-edge partition. Mathematics and Computers in Simulation, 79(8):2444–2457, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000299.

Rodtook:2009:SMR

[1729] A. Rodtook and S. S. Makhanov. Selection of multiresolution rotationally invariant moments for image recognition. Mathematics and Computers in Simulation, 79(8):2458– 2475, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409000305.

Shklyar:2009:ACN

[1730] A. Shklyar and A. Arbel. Accelerated convergence of the numerical simulation of incompressible flow in general curvilinear co-ordinates by discretizations on overset grids. Mathematics and Computers in Simulation, 79(8):2476–2489, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000317.

Surla:2009:DMP

[1731] K. Surla, Z. Uzelac, and Lj. Teofanov.

The discrete minimum principle for quadratic spline discretization of a singularly perturbed problem. Mathematics and Computers in Simulation, 79 (8):2490–2505, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000329.

Villamizar:2009:GSG

[1732] Vianey Villamizar and Sebastian Acosta. Generation of smooth grids with line control for scattering from multiple obstacles. Mathematics and Computers in Simulation, 79(8):2506– 2520, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000330.

Allen:2009:MMF

[1733] David E. Allen, Jiti Gao, and Michael McAleer. Modelling and managing financial risk: an overview. Mathematics and Computers in Simulation, 79 (8):2521–2524, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004187.

Thomas:2009:MCR

[1734] Lyn C. Thomas. Modelling the credit risk for portfolios of consumer loans: Analogies with corporate loan models. Mathematics and Computers in Simulation, 79(8): 2525–2534, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004199.

Allen:2009:CAA

[1735] David Allen, Zdravetz Lazarov, Michael McAleer, and Shelton Peiris. Comparison of alternative ACD models via density and interval forecasts: Evidence from the Australian stock market. Mathematics and Computers in Simulation, 79(8):2535– 2555, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004102.

Liu:2009:PML

[1736] Shuangzhe Liu and Heinz Neudecker. On pseudo maximum likelihood estimation for multivariate time series models with conditional heteroskedasticity. Mathematics and Computers in Simulation, 79(8):2556– 2565, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004163.

Billio:2009:GDC

[1737] Monica Billio and Massimiliano Caporin. A generalized Dynamic Conditional Correlation model for portfolio risk evaluation. Mathematics and Computers in Simulation, 79(8): 2566–2578, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004138.

Asai:2009:BAS

[1738] Manabu Asai. Bayesian analysis of stochastic volatility models with mixture-of-normal distributions. Mathematics and Computers in Simulation, 79(8):2579–2596, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004114.

Kobayashi:2009:TJS

[1739] Masahito Kobayashi. Testing for jumps in the stochastic volatility models. Mathematics and Computers in Simulation, 79(8):2597– 2608, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004151.

Lai:2009:ODH

[1740] YiHao Lai, Cathy W. S. Chen, and Richard Gerlach. Optimal dynamic hedging via copula-threshold-GARCH models. Mathematics and Computers in Simulation, 79(8):2609– 2624, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540800414X.

Lam:2009:IDI

[1741] K. P. Lam and H. S. Ng. Intra-daily information of range-based volatility for MEM-GARCH. Mathematics and Computers in Simulation, 79 (8):2625–2632, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004175.

Zhu:2009:TER

[1742] Jie Zhu. Testing for expected return and market price of risk in Chinese A and B share markets: a geometric Brownian motion and multivariate GARCH model approach. Mathematics and Computers in Simulation, 79 (8):2633–2653, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408004205.

Chen:2009:ISB

[1743] Cathy W. S. Chen, Richard Gerlach, Nick Y. P. Cheng, and Y. L. Yang. The impact of structural breaks on the integration of the ASEAN-5 stock markets. Mathematics and Computers in Simulation, 79(8):2654– 2664, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004126.

Anonymous:2009:NId

[1744] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(8):2665–2666, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000858.

Anonymous:2009:ICEd

[1745] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(8):2667– 2672, April 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900086X.

Anonymous:2009:EBd

[1746] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(8):CO2, April 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000780.

Anonymous:2009:P

[1747] Anonymous. Preface. Mathematics and Computers in Simulation, 79(9):vii–viii, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001438.

Fu:2009:MES

[1748] B. Fu, L. T. H. Newham, and J. B.

Field. Modelling erosion and sediment delivery from unsealed roads in southeast Australia. Mathematics and Computers in Simulation, 79 (9):2679–2688, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003182.

Herron:2009:IIG

[1749] Natasha Herron and Barry Croke. Including the influence of groundwater exchanges in a lumped rainfallrunoff model. Mathematics and Computers in Simulation, 79(9):2689– 2700, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002929.

McPhee:2009:PES

[1750] Malcolm McPhee, Jim Oltjen, James Fadel, David Mayer, and Roberto Sainz. Parameter estimation and sensitivity analysis of fat deposition models in beef steers using acslXtreme. Mathematics and Computers in Simulation, 79(9):2701–2712, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002899.

Matsuka:2009:NSH

[1751] Maki Matsuka, Roger D. Braddock, and Igor E. Agranovski. Numerical study of hydrogen permeation flux in ytterbium doped strontium cerate and thulium doped strontium cerate (II). Mathematics and Computers in Simulation, 79(9):2713– 2723, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003820.

Aiyoshi:2009:NES

[1752] Eitaro Aiyoshi and Atsushi Maki. A Nash equilibrium solution in an oligopoly market: the search for Nash equilibrium solutions with replicator equations derived from the gradient dynamics of a simplex algorithm. Mathematics and Computers in Simulation, 79(9):2724–2732, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540800342X.

Allen:2009:MIT

[1753] David Allen, Ghialy Yap, and Riaz Shareef. Modelling interstate tourism demand in Australia: a cointegration approach. Mathematics and Computers in Simulation, 79(9):2733– 2740, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003212.

Bartolome:2009:RMI

[1754] Ana Bartolomé, Michael McAleer, Vicente Ramos, and Javier Rey-Maquieira. A risk map of international tourist regions in Spain. Mathematics and Computers in Simulation, 79 (9):2741–2758, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003418.

Bi:2009:MFR

[1755] Guang Bi and David E. Giles. Modelling the financial risk associated with U.S. movie box office earnings. Mathematics and Computers in Simulation, 79(9):2759–2766, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001754.

Chan:2009:MTV

[1756] Felix Chan. Modelling time-varying higher moments with maximum entropy density. Mathematics and Computers in Simulation, 79(9):2767– 2778, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003819.

Chan:2009:MAI

[1757] W. S. Chan, C. S. Wong, and A. H. L. Chung. Modelling Australian interest rate swap spreads by mixture autoregressive conditional heteroscedastic processes. Mathematics and Computers in Simulation, 79 (9):2779–2786, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408002796.

Chang:2009:IAG

[1758] Chia-Lin Chang and Les Oxley. Industrial agglomeration, geographic innovation and total factor productivity: the case of Taiwan. Mathematics and Computers in Simulation, 79 (9):2787–2796, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003121.

Shi:2009:TJE

[1759] Xiuhong Shi and Masahito Kobayashi. Testing for jumps in the EGARCH process. Mathematics and Computers in Simulation, 79(9):2797– 2808, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002140.

Chow:2009:RVI

[1760] Ying-Foon Chow, James T. K. Lam, and Hinson S. Yeung. Realized volatility of index constituent stocks in Hong Kong. Mathematics and Computers in Simulation, 79(9):2809– 2818, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003236.

Fukiharu:2009:AME

[1761] T. Fukiharu. Asset market equilibrium: a simulation. Mathematics and Computers in Simulation, 79 (9):2819–2829, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003832.

Hakim:2009:FCC

[1762] Abdul Hakim and Michael McAleer.

Forecasting conditional correlations in stock, bond and foreign exchange markets. Mathematics and Computers in Simulation, 79(9):2830– 2846, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002784.

Herbert:2009:SNL

[1763] Ric D. Herbert and Peter J. Stemp.

Solving a non-linear model: the importance of model specification for deriving a suitable solution. Mathematics and Computers in Simulation, 79 (9):2847–2855, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408001705.

Ho:2009:VDU

[1764] Kin-Yip Ho, Albert K. Tsui, and Zhaoyong Zhang. Volatility dynamics of the US business cycle: a multivariate asymmetric GARCH approach. Mathematics and Computers in Simulation, 79(9):2856–2868, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002887.

Konishi:2009:HTR

[1765] Yoko Konishi and Yoshihiko Nishiyama. Hypothesis testing in rank-size rule regression. Mathematics and Computers in Simulation, 79(9):2869– 2878, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408003431.

Lim:2009:AMI

[1766] Christine Lim, Michael McAleer, and Jennifer C. H. Min. ARMAX modelling of international tourism demand. Mathematics and Computers in Simulation, 79(9):2879– 2888, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002930.

Nawata:2009:ALH

[1767] Kazumitsu Nawata, Masako Ii, Aya Ishiguro, and Koichi Kawabuchi. An analysis of the length of hospital stay for cataract patients in Japan using the discrete-type proportional hazard model. Mathematics and Computers in Simulation, 79(9):2889– 2896, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002115.

Okui:2009:TSC

[1768] Ryo Okui. Testing serial correlation in fixed effects regression models based on asymptotically unbiased autocorrelation estimators. Mathematics and Computers in Simulation, 79 (9):2897–2909, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002875.

Oxley:2009:CSV

[1769] Les Oxley, Marco Reale, and Granville Tunnicliffe Wilson. Constructing structural VAR models with conditional independence graphs. Mathematics and Computers in Simulation, 79 (9):2910–2916, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003790.

Sakata:2009:IDP

[1770] Kei Sakata and C. R. McKenzie. The impact of divorce precedents on the Japanese divorce rate. Mathematics and Computers in Simulation, 79 (9):2917–2926, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003170.

Sato:2009:SMU

[1771] Kiyotaka Sato, Zhaoyong Zhang, and David Allen. The suitability of a monetary union in East Asia: What does the cointegration approach tell? Mathematics and Computers in Simulation, 79(9):2927–2937, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003224.

Ubukata:2009:EIY

[1772] M. Ubukata and M. Fukushige. Estimation and inference in the yield curve model with an instantaneous error term. Mathematics and Computers in Simulation, 79(9):2938– 2946, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475408003789.

Wang:2009:EGB

[1773] Kai Wang, Jin Hong, Dora Marinova, and Liang Zhu. Evolution and governance of the biotechnology and pharmaceutical industry of China. Mathematics and Computers in Simulation, 79(9):2947–2956, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475408003108.

Lim:2009:CIB

[1774] Lee K. Lim. Convergence and interdependence between ASEAN-5 stock markets. Mathematics and Computers in Simulation, 79(9):2957– 2966, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408004096.

Anonymous:2009:NIe

[1775] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(9):2967–2968, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001384.

Anonymous:2009:ICEe

[1776] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(9):2969– 2974, May 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001396.

Anonymous:2009:EBe

[1777] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(9):CO2, May 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001311.

Deng:2009:ASA

[1778] Shaojiang Deng, Xiaofeng Liao, and Songtao Guo. Asymptotic stability analysis of certain neutral differential equations: a descriptor system approach. Mathematics and Computers in Simulation, 79(10):2981– 2993, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000056.

Yongzhen:2009:CIH

[1779] Pei Yongzhen, Li Changguo, and Chen Lansun. Continuous and impulsive harvesting strategies in a stagestructured predator–prey model with time delay. Mathematics and Computers in Simulation, 79(10):2994– 3008, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900007X.

Sobol:2009:DBG

[1780] I. M. Sobol’ and S. Kucherenko. Derivative based global sensitivity measures and their link with global sensitivity indices. Mathematics and Computers in Simulation, 79(10):3009– 3017, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000354.

Liu:2009:SMC

[1781] Leipo Liu, Zhengzhi Han, and Xiushan Cai. Sliding mode control for polytopic differential inclusion systems. Mathematics and Computers in Simulation, 79(10):3018–3025, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000366.

Wu:2009:SOM

[1782] Ling-Yun Wu, Zhenping Li, Rui-Sheng Wang, Xiang-Sun Zhang, and Luonan Chen. Self-organizing map approaches for the haplotype assembly problem. Mathematics and Computers in Simulation, 79(10):3026– 3037, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000378.

Hou:2009:CIV

[1783] Juan Hou and Zhidong Teng. Continuous and impulsive vaccination of SEIR epidemic models with saturation incidence rates. Mathematics and Computers in Simulation, 79 (10):3038–3054, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900041X.

Orrillo:2009:MPI

[1784] Jaime Orrillo. Making promises in infinite-horizon economies with default and collateral. Mathematics and Computers in Simulation, 79 (10):3055–3068, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000433.

Jodra:2009:CFE

[1785] P. Jodrá. A closed-form expression for the quantile function of the Gompertz– Makeham distribution. Mathematics and Computers in Simulation, 79 (10):3069–3075, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000445.

Li:2009:CJD

[1786] Ming-Yuan Leon Li. Could the jump diffusion technique enhance the effectiveness of futures hedging models?: a reality test. Mathematics and Computers in Simulation, 79 (10):3076–3088, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409000482.

Kung:2009:TAS

[1787] James J. Kung. A two-asset stochastic model for long-term portfolio selection. Mathematics and Computers in Simulation, 79(10):3089– 3098, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409000652.

Koko:2009:OFS

[1788] Jonas Koko and Teddy Virin. Optimization of a fertilizer spreading process. Mathematics and Computers in Simulation, 79(10):3099– 3109, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540900069X.

Ling:2009:PIP

[1789] Li Ling, Wei Xu, and Minghai Li. Parametric inference for progressive Type-I hybrid censored data on a simple stepstress accelerated life test model. Mathematics and Computers in Simulation, 79(10):3110–3121, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000706.

Divino:2009:TCD

[1790] Jose Angelo Divino. Is there a case for domestic inflation target? Mathematics and Computers in Simulation, 79 (10):3122–3135, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000718.

Coelho:2009:IDE

[1791] Leandro dos Santos Coelho, Rodrigo Clemente Thom Souza, and Viviana Cocco Mariani. Improved differential evolution approach based on cultural algorithm and diversity measure applied to solve economic load dispatch problems. Mathematics and Computers in Simulation, 79 (10):3136–3147, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000731.

Wu:2009:ELM

[1792] Shifeng Wu and Siqing Gan. Errors of linear multistep methods for singularly perturbed Volterra delayintegro-differential equations. Mathematics and Computers in Simulation, 79(10):3148–3159, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000743.

Chungui:2009:HMW

[1793] Zhou Chungui, Zhang Xinong, Xie Shilin, Zhou Tong, and Zhu Changchun. Hybrid modeling of wire cable vibration isolation system through neural network. Mathematics and Computers in Simulation, 79(10):3160– 3173, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409000767.

Zhang:2009:SOC

[1794] Yinping Zhang. Stationary oscillation for cellular neural networks with time delays and impulses. Mathematics and Computers in Simulation, 79 (10):3174–3178, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409000974.

Angelini:2009:BAD

[1795] Natascia Angelini, Roberto Dieci, and Franco Nardini. Bifurcation analysis of a dynamic duopoly model with heterogeneous costs and behavioural rules. Mathematics and Computers in Simulation, 79(10):3179– 3196, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001025.

Anonymous:2009:NIf

[1796] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(10):3197–3198, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001554.

Anonymous:2009:ICEf

[1797] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(10):3199– 3204, June 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001566.

Anonymous:2009:EBf

[1798] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(10):CO2, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001505.

Anonymous:2009:PJb

[1799] Anonymous. Pages 2981–3210 (June 2009). Mathematics and Computers in Simulation, 79(10):??, June 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Hernando:2009:SNE

[1800] Antonio Hernando, Luis de Ledesma, and Luis M. Laita. Showing the nonexistence of solutions in systems of linear Diophantine equations. Mathematics and Computers in Simulation, 79 (11):3211–3220, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900072X.

Doha:2009:ESU

[1801] E. H. Doha and W. M. Abd-Elhameed. Efficient spectral ultraspherical-dual-Petrov–Galerkin algorithms for the direct solution of (2n + 1) th-order linear differential equations. Mathematics and Computers in Simulation, 79 (11):3221–3242, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001013.

Yeung:2009:VTS

[1802] W. W. H. Yeung. Vortex trapping on a surface with an indentation and corrugations. Mathematics and Computers in Simulation, 79 (11):3243–3257, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001190.

Feng:2009:AFR

[1803] Li-Hua Feng and Gao-Yuan Luo. Analysis on fuzzy risk of landfall typhoon in Zhejiang province of China. Mathematics and Computers in Simulation, 79(11):3258–3266, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001244.

Wong:2009:MPE

[1804] Wing-Keung Wong and Michael McAleer. Mapping the Presidential Election Cycle in US stock markets. Mathematics and Computers in Simulation, 79(11):3267– 3277, July 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001268.

Cheng:2009:EUQ

[1805] Haiyan Cheng and Adrian Sandu. Efficient uncertainty quantification with the polynomial chaos method for stiff systems. Mathematics and Computers in Simulation, 79(11):3278– 3295, July 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900127X.

Biagiola:2009:WHU

[1806] S. I. Biagiola and J. L. Figueroa.

Wiener and Hammerstein uncertain models identification. Mathematics and Computers in Simulation, 79 (11):3296–3313, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001281.

Ilison:2009:PTS

[1807] Lauri Ilison and Andrus Salupere.

Propagation of sech^{2}-type solitary waves in hierarchical KdV-type systems. Mathematics and Computers in Simulation, 79(11):3314– 3327, July 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001293.

Marchi:2009:PPR

[1808] A. Marchi, A. Liverani, and A. Del Giudice. Polynomial pseudo-random number generator via cyclic phase. Mathematics and Computers in Simulation, 79(11):3328–3338, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001463.

Delis:2009:FVM

[1809] A. I. Delis and E. N. Mathioudakis. A finite volume method parallelization for the simulation of free surface shallow water flows. Mathematics and Computers in Simulation, 79 (11):3339–3359, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001633.

Anonymous:2009:NIg

[1810] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(11):3360–3361, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002079.

Anonymous:2009:ICEg

[1811] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(11):3362– 3367, July 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002080.

Anonymous:2009:EBg

[1812] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(11):CO2, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900202X.

Anonymous:2009:PJc

[1813] Anonymous. Pages 3211–3374 (July 2009). Mathematics and Computers in Simulation, 79(11):??, July 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Amaziane:2009:P

[1814] Brahim Amaziane, Domingo Barrera, and Driss Sbibih. Preface. Mathematics and Computers in Simulation, 79(12):vii, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002274.

Alba-Fernandez:2009:BDS

[1815] V. Alba-Fernández and M. D. Jiménez-Gamero. Bootstrapping divergence statistics for testing homogeneity in multinomial populations. Mathematics and Computers in Simulation, 79 (12):3375–3384, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001037.

Bailly:2009:GCH

[1816] D. Bailly, R. Ababou, and M. Quintard. Geometric characterization, hydraulic behavior and upscaling of 3D fissured geologic media. Mathematics and Computers in Simulation, 79 (12):3385–3396, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001487.

Barbaro:2009:DCM

[1817] Alethea B. T. Barbaro, Kirk Taylor, Peterson F. Trethewey, Lamia Youseff, and Björn Birnir. Discrete and continuous models of the dynamics of pelagic fish: Application to the capelin. Mathematics and Computers in Simulation, 79(12): 3397–3414, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408003856.

Benkhaldoun:2009:AMA

[1818] Fayssal Benkhaldoun, Imad Elmahi, and Mohammed Seaïd. Application of mesh-adaptation for pollutant transport by water flow. Mathematics and Computers in Simulation, 79(12): 3415–3423, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001141.

Campillo:2009:PIM

[1819] Fabien Campillo, Rivo Rakotozafy, and Vivien Rossi. Parallel and interacting Markov chain Monte Carlo algorithm. Mathematics and Computers in Simulation, 79(12):3424– 3433, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900113X.

Delgado:2009:MAD

[1820] A. H. Delgado and L. Márquez. Modelling of an arch dam by polynomial interpolation. Mathematics and Computers in Simulation, 79(12):3434– 3443, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001128.

El-Amrani:2009:LES

[1821] Mofdi El-Amrani and Mohammed Seaïd. Large eddy simulation of turbulent heat transport in the Strait of Gibraltar. Mathematics and Computers in Simulation, 79(12):3444– 3454, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001177.

Foucher:2009:QSQ

[1822] Françoise Foucher and Paul Sablonnière.

Quadratic spline quasi-interpolants and collocation methods. Mathematics and Computers in Simulation, 79 (12):3455–3465, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001086.

Gavete:2009:ASS

[1823] L. Gavete, B. Alonso, M. L. Gavete, F. Ureña, and J. J. Benito. An adaptive solver for the spherical shallow water equations. Mathematics and Computers in Simulation, 79(12):

3466–3477, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001062.

Fortes:2009:CTI

[1824] M. A. Fortes, M. J. Ibáñez, and M. L. Rodríguez. On Chebyshevtype integral quasi-interpolation operators. Mathematics and Computers in Simulation, 79(12):3478– 3491, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001189.

Joly:2009:LDF

[1825] Agnès Joly and Alain Perrard. Linear driving force models for dynamic adsorption of volatile organic compound traces by porous adsorbent beds. Mathematics and Computers in Simulation, 79(12):3492– 3499, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001153.

Khlaifi:2009:PSI

[1826] Anis Khlaifi, Anda Ionescu, and Yves Candau. Pollution source identification using a coupled diffusion model with a genetic algorithm. Mathematics and Computers in Simulation, 79 (12):3500–3510, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001219.

Kouibia:2009:DAV

[1827] A. Kouibia and M. Pasadas. Discrete approximation by variational vector splines for noisy data. Mathematics and Computers in Simulation, 79 (12):3511–3522, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001220.

Manouzi:2009:SWS

[1828] Hassan Manouzi and Mohammed Seaïd. Solving Wick-stochastic water waves using a Galerkin finite element method. Mathematics and Computers in Simulation, 79(12):3523– 3533, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001098.

Mocenni:2009:ISS

[1829] Chiara Mocenni and Emiliano Sparacino. Identification and simulation of a spatial ecological model in a lake with fractal boundary. Mathematics and Computers in Simulation, 79(12): 3534–3546, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001116.

Amat:2009:EEB

[1830] S. Amat and M. Moncayo. Exact error bounds for the reconstruction processes using interpolating wavelets. Mathematics and Computers in Simulation, 79(12):3547– 3555, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001049.

Reinoso:2009:FFB

[1831] J. F. Reinoso, M. Moncayo, M. Pasadas, F. J. Ariza, and J. L. García. The Frenet frame beyond classical differential geometry: Application to cartographic generalization of roads. Mathematics and Computers in Simulation, 79(12):3556– 3566, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001104.

Pasadas:2009:AMD

[1832] M. Pasadas and M. L. Rodríguez. An approximation method with data selection process. Mathematics and Computers in Simulation, 79(12):3567– 3576, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001207.

Rossler:2009:NSS

[1833] Andreas Rößler, Mohammed Seaïd, and Mostafa Zahri. Numerical simulation of stochastic replicator models in catalyzed RNA-like polymers. Mathematics and Computers in Simulation, 79(12):3577– 3586, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001232.

Sbibih:2009:NAR

[1834] D. Sbibih, A. Serghini, and A. Tijini. A new algorithm for a recursive construction of the minimal interpolation space. Mathematics and Computers in Simulation, 79(12): 3587–3598, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001050.

Belkhatir:2009:CFB

[1835] Bachir Belkhatir and Ahmed Zidna.

Construction of flexible blending parametric surfaces via curves. Mathematics and Computers in Simulation, 79 (12):3599–3608, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001074.

Anonymous:2009:NIh

[1836] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 79(12):3609–3610, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475409002213.

Anonymous:2009:ICEh

[1837] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 79(12):3611– 3615, August 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002225.

Anonymous:2009:EBh

[1838] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 79(12):CO2, August 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540900216X.

Taha:2009:Fa

[1839] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 80(1):1, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001712.

Ibrahim:2009:ISN

[1840] Slim Ibrahim and Philippe Guyenne. Instability in supercritical nonlinear wave equations: Theoretical results and symplectic integration. Mathematics and Computers in Simulation, 80(1):2–9, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001724.

Ghazaryan:2009:CNI

[1841] Anna Ghazaryan. On the convective nature of the instability of a front undergoing a supercritical Turing bifurcation. Mathematics and Computers in Simulation, 80(1):10– 19, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001736.

Moore:2009:CMS

[1842] Brian E. Moore. Conformal multisymplectic integration methods for forced-damped semi-linear wave equations. Mathematics and Computers in Simulation, 80(1):20–28, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001748.

Choi:2009:NSW

[1843] Wooyoung Choi. Nonlinear surface waves interacting with a linear shear current. Mathematics and Computers in Simulation, 80(1):29– 36, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900175X.

Choudhury:2009:SWF

[1844] S. Roy Choudhury. Solitary wave families of NLPDES via reversible systems theory. Mathematics and Computers in Simulation, 80(1):37– 45, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001761.

Todorov:2009:ILC

[1845] M. D. Todorov and C. I. Christov. Impact of the large crossmodulation parameter on the collision dynamics of quasi-particles governed by vector NLSE. Mathematics and Computers in Simulation, 80 (1):46–55, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001773.

Christov:2009:ISWa

[1846] Christo I. Christov, Tchavdar T. Marinov, and Rossitza S. Marinova. Identification of solitary-wave solutions as an inverse problem: Application to shapes with oscillatory tails. Mathematics and Computers in Simulation, 80(1):56–65, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001785.

Carstea:2009:PWN

[1847] A. S. Carstea. Proteomic waves in networks of transcriptional regulators. Mathematics and Computers in Simulation, 80(1):66–72, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001797.

Mancas:2009:SSC

[1848] Stefan C. Mancas and Roy S. Choudhury. Snake solitons in the cubic–quintic Ginzburg–Landau equation. Mathematics and Computers in Simulation, 80(1):73–82, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001803.

Schober:2009:MWA

[1849] Constance Schober and Tomasz H.

Wlodarczyk. Multisymplecticity and wave action conservation. Mathematics and Computers in Simulation, 80 (1):83–90, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001815.

Christov:2009:CQP

[1850] C. I. Christov. The concept of a quasi-particle and the nonprobabilistic interpretation of wave mechanics. Mathematics and Computers in Simulation, 80(1):91–101, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001827.

Ambrose:2009:SFM

[1851] David M. Ambrose. Singularity formation in a model for the vortex sheet with surface tension. Mathematics and Computers in Simulation, 80(1): 102–111, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001839.

Gerdjikov:2009:GPC

[1852] V. S. Gerdjikov, N. A. Kostov, E. V.

Doktorov, and N. P. Matsuka. Generalized perturbed complex Toda chain for Manakov system and exact solutions of Bose–Einstein mixtures. Mathematics and Computers in Simulation, 80 (1):112–119, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001840.

Gorsky:2009:SDL

[1853] Jennifer Gorsky and David P. Nicholls. A small dispersion limit to the Camassa–Holm equation: a numerical study. Mathematics and Computers in Simulation, 80(1):120–130, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001852.

Baksmaty:2009:NEV

[1854] L. O. Baksmaty, Y. Liu, U. Landman, N. P. Bigelow, and H. Pu. Numerical exploration of vortex matter in Bose–Einstein condensates. Mathematics and Computers in Simulation, 80 (1):131–138, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001864.

Nguyen:2009:OSN

[1855] Nguyet Thanh Nguyen and Henrik Kalisch. Orbital stability of negative solitary waves. Mathematics and Computers in Simulation, 80(1): 139–150, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409001876.

Skogestad:2009:BVP

[1856] Jan Ole Skogestad and Henrik Kalisch. A boundary value problem for the KdV equation: Comparison of finite-difference and Chebyshev methods. Mathematics and Computers in Simulation, 80(1):151– 163, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001888.

Herman:2009:NRS

[1857] Russell L. Herman and Andrew Rose. Numerical realizations of solutions of the stochastic KdV equation. Mathematics and Computers in Simulation, 80(1):164–172, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S037847540900189X.

Gorsky:2009:WPK

[1858] Jennifer Gorsky and A. Alexandrou Himonas. Well-posedness of KdV with higher dispersion. Mathematics and Computers in Simulation, 80(1): 173–183, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001906.

Yu:2009:EGT

[1859] Guo-Fu Yu and Xing-Biao Hu. Extended Gram-type determinant solutions to the Kadomtsev–Petviashvili equation. Mathematics and Computers in Simulation, 80(1):184–191, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001918.

Christov:2009:ISWb

[1860] Ivan Christov. Internal solitary waves in the ocean: Analysis using the periodic, inverse scattering transform. Mathematics and Computers in Simulation, 80(1):192–201, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540900192X.

Jordan:2009:SRN

[1861] P. M. Jordan. Some remarks on nonlinear poroacoustic phenomena. Mathematics and Computers in Simulation, 80(1):202–211, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001931.

Triki:2009:CTS

[1862] Houria Triki and Thiab R. Taha. On the calculation of the timing shifts in the variable-coefficient Korteweg– de Vries equation. Mathematics and Computers in Simulation, 80(1):212– 222, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409001943.

Ludu:2009:EAM

[1863] A. Ludu and C. Cibert. Elastic axonemal model: Solitary wave shapes. Mathematics and Computers in Simulation, 80(1):223–230, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001955.

Anonymous:2009:NIi

[1864] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(1):231–232, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0378475409002535.

Anonymous:2009:ICEi

[1865] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(1):233–237, September 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002547.

Anonymous:2009:EBi

[1866] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(1):CO2, September 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847540900247X.

Pulch:2009:PCS

[1867] Roland Pulch and Cathrin van Emmerich. Polynomial chaos for simulating random volatilities. Mathematics and Computers in Simulation, 80 (2):245–255, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001645.

Ryashko:2009:CTA

[1868] L. Ryashko, I. Bashkirtseva, A. Gubkin, and P. Stikhin. Confidence tori in the analysis of stochastic 3D-cycles. Mathematics and Computers in Simulation, 80(2):256–269, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001967.

Swiercz:2009:DRA

[1869] Ewa Swiercz and Andrzej Pieniezny. Detection-recognition algorithm based on the Gabor transform for unknown signals embedded in unknown noise. Mathematics and Computers in Simulation, 80(2):270– 293, October 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001979.

Halilovic:2009:NEN

[1870] Miroslav Halilovič, Marko Vrh, and Boris Štok. NICE — an explicit numerical scheme for efficient integration of nonlinear constitutive equations. Mathematics and Computers in Simulation, 80(2):294–313, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409001992.

Yang:2009:OMI

[1871] Liu Yang, Zui-Cha Deng, Jian-Ning Yu, and Guan-Wei Luo. Optimization method for the inverse problem of reconstructing the source term in a parabolic equation. Mathematics and Computers in Simulation, 80 (2):314–326, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002006.

Yip:2009:SSM

[1872] Iris W. H. Yip and Mike K. P. So. Simplified specifications of a multivariate generalized autoregressive conditional heteroscedasticity model. Mathematics and Computers in Simulation, 80 (2):327–340, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002146.

Gao:2009:SLS

[1873] Fuzheng Gao and Hongxing Rui. A split least-squares characteristic mixed finite element method for Sobolev equations with convection term. Mathematics and Computers in Simulation, 80(2):341–351, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002316.

Qin:2009:QRT

[1874] Hai-Hua Qin and Ting Wei. Quasireversibility and truncation methods to solve a Cauchy problem for the modified Helmholtz equation. Mathematics and Computers in Simulation, 80 (2):352–366, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002328.

Shitan:2009:PSO

[1875] Mahendran Shitan and Shelton Peiris. On properties of the second order generalized autoregressive GAR(2) model with index. Mathematics and Computers in Simulation, 80(2):367– 377, October 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540900233X.

Kung:2009:OPU

[1876] James J. Kung and Lung-Sheng Lee. Option pricing under the Merton model of the short rate. Mathematics and Computers in Simulation, 80 (2):378–386, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002341. See

comment [1929].

Dookhitram:2009:NMA

[1877] K. Dookhitram, R. Boojhawon, and M. Bhuruth. A new method for accelerating Arnoldi algorithms for large scale eigenproblems. Mathematics and Computers in Simulation, 80 (2):387–401, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002353.

Chen:2009:SOS

[1878] Cheng-Wu Chen. The stability of an oceanic structure with T– S fuzzy models. Mathematics and Computers in Simulation, 80(2):402– 426, October 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002365.

Coelho:2009:SOM

[1879] Leandro dos Santos Coelho. Selforganizing migration algorithm applied to machining allocation of clutch assembly. Mathematics and Computers in Simulation, 80(2):427– 435, October 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002377.

Liu:2009:SAU

[1880] Duyu Liu, Shouming Zhong, Xinzhi Liu, and Yuanqing Huang. Stability analysis for uncertain switched neutral systems with discrete time-varying delay: a delay-dependent method. Mathematics and Computers in Simulation, 80(2):436–448, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002389.

Garcia:2009:SMD

[1881] José Paulo F. Garcia, Lizete Maria C. F. Garcia, Gisele C. Apolinário, and Fernando B. Rodrigues. Sliding mode for detection and accommodation of computation time delay fault. Mathematics and Computers in Simulation, 80(2):449–465, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002419.

Anonymous:2009:NIj

[1882] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(2):466–467, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003036.

Anonymous:2009:ICEj

[1883] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(2):468– 472, October 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003048.

Anonymous:2009:EBj

[1884] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(2):CO2, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003000.

Anonymous:2009:PO

[1885] Anonymous. Pages 245–478 (October 2009). Mathematics and Computers in Simulation, 80(2):??, October 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Gaudiano:2009:CCD

[1886] Marcos Gaudiano, Germán Ariel Torres, and Cristina Turner. On a convective condition in the diffusion of a solvent into a polymer with non-constant conductivity coefficient. Mathematics and Computers in Simulation, 80(3): 479–489, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475408002413.

Shu:2009:GAS

[1887] Huisheng Shu, Zidong Wang, and Zengwei Lü. Global asymptotic stability of uncertain stochastic bidirectional associative memory networks with discrete and distributed delays. Mathematics and Computers in Simulation, 80(3):490–505, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475408002474.

Sevastianov:2009:SFL

[1888] P. Sevastianov and L. Dymova. Synthesis of fuzzy logic and Dempster– Shafer Theory for the simulation of the decision-making process in stock trading systems. Mathematics and Computers in Simulation, 80(3):506– 521, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409001980.

Jiao:2009:DAD

[1889] Jianjun Jiao, Xiaosong Yang, Shaohong Cai, and Lansun Chen. Dynamical analysis of a delayed predatorprey model with impulsive diffusion between two patches. Mathematics and Computers in Simulation, 80(3): 522–532, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002304.

Yang:2009:MSA

[1890] Xiaoping Yang, Hao Zhang, and Xikui Ma. Modeling and stability analysis of cascade buck converters with N power stages. Mathematics and Computers in Simulation, 80(3):533– 546, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002596.

Marin:2009:SLZ

[1891] J. Marin and J. Monnier. Superposition of local zoom models and simultaneous calibration for 1D–2D shallow water flows. Mathematics and Computers in Simulation, 80(3): 547–560, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002870.

Bartolucci:2009:EBS

[1892] Al Bartolucci, Sejong Bae, Karan Singh, and H. Randall Griffith. An examination of Bayesian statistical approaches to modeling change in cognitive decline in an Alzheimer’s disease population. Mathematics and Computers in Simulation, 80(3):561– 571, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002882.

Hendrickx:2009:MVF

[1893] Wouter Hendrickx, Dirk Deschrijver, Luc Knockaert, and Tom Dhaene. Magnitude Vector Fitting to interval data. Mathematics and Computers in Simulation, 80(3):572–580, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002973.

Petrova:2009:AOS

[1894] Svetozara I. Petrova. Applications of one-shot methods in PDEs constrained shape optimization. Mathematics and Computers in Simulation, 80(3): 581–597, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002985.

Lastra:2009:SSW

[1895] Miguel Lastra, José M. Mantas, Carlos Ureña, Manuel J. Castro, and José A. García-Rodríguez. Simulation of shallow-water systems using graphics processing units. Mathematics and Computers in Simulation, 80 (3):598–618, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409003139.

Yu:2009:DAE

[1896] Hengguo Yu, Shouming Zhong, and Mao Ye. Dynamic analysis of an ecological model with impulsive control strategy and distributed time delay. Mathematics and Computers in Simulation, 80(3):619–632, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003140.

Anonymous:2009:NIk

[1897] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(3):633–634, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475409003292.

Anonymous:2009:ICEk

[1898] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(3):635–639, November 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003309.

Anonymous:2009:EBk

[1899] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(3):CO2, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003243.

Anonymous:2009:PN

[1900] Anonymous. Pages 479–646 (November 2009). Mathematics and Computers in Simulation, 80(3):??, November 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Taha:2009:Fb

[1901] Thiab R. Taha. Foreword. Mathematics and Computers in Simulation, 80(4):647, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002638.

Malomed:2009:GSF

[1902] Boris A. Malomed, V. A. Nascimento, and Sadhan K. Adhikari. Gap solitons in fermion superfluids. Mathematics and Computers in Simulation, 80 (4):648–659, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002675.

Triki:2009:CTA

[1903] Houria Triki and Thiab R. Taha. Calculation of timing and amplitude jitter in a dispersion-managed Korteweg– de Vries system. Mathematics and Computers in Simulation, 80(4):660– 665, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002687.

Linzon:2009:WIW

[1904] Yoav Linzon, Boris A. Malomed, Michael Zaezjev, R. Morandotti, Maite Volatier-Ravat, Vincent Aimez, Richard Ares, and Shimshon Bar-Ad. Wave interactions and wavestructure scattering in nonlinear patterned waveguides. Mathematics and Computers in Simulation, 80(4):666– 673, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002699.

Kamata:2009:ADN

[1905] Masaru Kamata and Atsushi Nakamula. Aspects of q-discretized Nahm equations. Mathematics and Computers in Simulation, 80(4):674–681, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002705.

Norton:2009:CHH

[1906] Guy V. Norton. Comparison of homogeneous and heterogeneous modeling of transient scattering from dispersive media directly in the time domain. Mathematics and Computers in Simulation, 80(4):682–692, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002717.

Espinola-Rocha:2009:TSC

[1907] J. Adrián Espínola-Rocha and P. G. Kevrekidis. Thresholds for soliton creation in the Ablowitz–Ladik lattice. Mathematics and Computers in Simulation, 80(4):693–706, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002729.

Karali:2009:BIM

[1908] Georgia D. Karali and Panayotis G.

Kevrekidis. Bubble interactions for the Mullins–Sekerka problem: Some case examples. Mathematics and Computers in Simulation, 80(4):707– 720, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002730.

Hoq:2009:ENW

[1909] Q. E. Hoq, J. Gagnon, P. G.

Kevrekidis, B. A. Malomed, D. J. Frantzeskakis, and R. Carretero-González. Extended nonlinear waves in multidimensional dynamical lattices. Mathematics and Computers in Simulation, 80(4):721–731, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002602.

Mishmash:2009:UAO

[1910] R. V. Mishmash and L. D. Carr. Ultracold atoms in 1D optical lattices: mean field, quantum field, computation, and soliton formation. Mathematics and Computers in Simulation, 80 (4):732–740, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002754.

Schober:2009:DGV

[1911] C. M. Schober and T. H. Wlodarczyk.

Dispersion, group velocity, and multisymplectic discretizations. Mathematics and Computers in Simulation, 80 (4):741–751, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002614.

Steinhoff:2009:LRN

[1912] John Steinhoff and Subhashini Chitta. Long range numerical simulation of short waves as nonlinear solitary waves. Mathematics and Computers in Simulation, 80(4):752–762, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002626.

Straughan:2009:NAW

[1913] B. Straughan. Nonlinear acceleration waves in porous media. Mathematics and Computers in Simulation, 80 (4):763–769, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540900264X.

Hoseini:2009:SPT

[1914] S. M. Hoseini and T. R. Marchant. Soliton perturbation theory for a higher order Hirota equation. Mathematics and Computers in Simulation, 80(4): 770–778, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002651.

Wang:2009:CSG

[1915] Hong-Yan Wang. Commutativity of source generation procedure and Bäcklund transformation: a BKP equation. Mathematics and Computers in Simulation, 80(4):779–785, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002663.

Xing:2009:FSC

[1916] Jiuxing Xing, Alan M. Davies, and Jarle Berntsen. Free surface, current profile and buoyancy effects upon internal wave energy flux profiles in sill regions. Mathematics and Computers in Simulation, 80(4): 786–793, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002766.

Chugunova:2009:UCC

[1917] Marina Chugunova and Dmitry Pelinovsky. On the uniform convergence of the Chebyshev interpolants for solitons. Mathematics and Computers in Simulation, 80(4):794–803, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002778.

Xu:2009:SGR

[1918] Runzhang Xu and Jihong Shen. Some generalized results for global wellposeness for wave equations with damping and source terms. Mathematics and Computers in Simulation, 80 (4):804–807, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S037847540900278X.

Runzhang:2009:ABB

[1919] Xu Runzhang. Asymptotic behavior and blow up of solutions for semilinear parabolic equations at critical energy level. Mathematics and Computers in Simulation, 80(4):808– 813, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409002791.

Trombettoni:2009:SDL

[1920] A. Trombettoni, H. E. Nistazakis, Z. Rapti, D. J. Frantzeskakis, and P. G. Kevrekidis. Soliton dynamics in linearly coupled discrete nonlinear Schrödinger equations. Mathematics and Computers in Simulation, 80(4): 814–824, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409002808.

Deconinck:2009:SIU

[1921] Bernard Deconinck and Michael Nivala. Symbolic integration using homotopy methods. Mathematics and Computers in Simulation, 80(4):825– 836, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S037847540900281X.

Dutykh:2009:TGD

[1922] Denys Dutykh and Frédéric Dias. Tsunami generation by dynamic displacement of sea bed due to dipslip faulting. Mathematics and Computers in Simulation, 80(4):837–848, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0378475409002821.

Triki:2009:EAS

[1923] Houria Triki and Thiab R. Taha. Exact analytic solitary wave solutions for the RKL model. Mathematics and Computers in Simulation, 80(4): 849–854, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002833. See

comment [1932] and reply [1933].

Mickens:2009:TWS

[1924] Ronald E. Mickens. Traveling-wave solutions for a discrete Burgers equation with nonlinear diffusion. Mathematics and Computers in Simulation, 80 (4):855–859, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002845.

Mitsotakis:2009:BST

[1925] D. E. Mitsotakis. Boussinesq systems in two space dimensions over a variable bottom for the generation and propagation of tsunami waves. Mathematics and Computers in Simulation, 80 (4):860–873, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475409002857.

Anonymous:2009:NIl

[1926] Anonymous. News of IMACS. Mathematics and Computers in Simulation, 80(4):874–875, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/pii/S0378475409003449.

Anonymous:2009:ICEl

[1927] Anonymous. IMACS calendar of events. Mathematics and Computers in Simulation, 80(4):876–880, December 2009. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0378475409003450.

Anonymous:2009:EBl

[1928] Anonymous. Editorial Board. Mathematics and Computers in Simulation, 80(4):CO2, December 2009. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475409003383.

Cui:2010:COP

[1929] Zhenyu Cui and Don Mcleish. Comment on “Option pricing under the Merton model of the short rate” by Kung and Lee [Math. Comput. Simul. 80 (2009) 378–386]. Mathematics and Computers in Simulation, 81(1):1–4, September 2010. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475410001953. See

[1876].

Liang:2014:CLS

[1930] TaChen Liang and Ming-Chung Yang.

Comments on “The life-span prediction of a system connected in series”. Mathematics and Computers in Simulation, 97:189–191, March 2014. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0378475413002231. See

[1666].

BerguesCabrales:2018:CMM

[1931] Luis Enrique Bergues Cabrales, Andrés Ramírez Aguilera, Rolando Placeres Jiménez, Manuel Verdecia Jarque, Héctor Manuel Camué Ciria and Juan Bory Reyes, Miguel Angel O Farril Mateus, Fabiola Suárez Palencia, and Marisela González Ávila. Corrigendum to “Mathematical modeling of tumor growth in mice following low-level direct electric current” [Math. Comput. Simulation 78 (2008) 112–120]. Mathematics and Computers in Simulation, 144:266, February 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S037847541730085X. See

[1452].

Al-Ghafri:2021:CEA

[1932] K. S. Al-Ghafri and E. V. Krishnan. Comment on: “Exact analytic solitary wave solutions for the RKL model” [Math. Comput. Simulation 80 (4) (2009) 849–854]. Mathematics and Computers in Simulation, 181:113–116, March 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/science/

                                        article/pii/S0378475420301750. See

[1923] and reply [1933].

Triki:2021:RCE

[1933] Houria Triki and Thiab R. Taha. Reply to ‘Comment on: “Exact analytic solitary wave solutions for the RKL model” [Math. Comput. Simulation 80 (4) (2009) 849–854]’. Mathematics and Computers in Simulation, 182:234, April 2021. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0378475420303256. See

[1923, 1932].